A Complete Bibliography of Publications in SIAM Journal on Mathematical Analysis for 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 2.22

Title word cross-reference

(1 + 3) [CCLM15]. (1 + 4) [LMTT15]. (N − 1) [Tan15]. (t, x) [BT12]. −x · y [Che15]. 1 [BOO18, CSW15, CN15, CM19, GM10, GMM13, JWX13, KT18, Per10, Sug16, WX11, XYY19]. 2 [AHKM15, BP19, BM12b, CWY14, DSZ19, Dro18, DYZ19b, ER19, FY13a, HHK18, Kim09, Kim13, LWY18, LW14b, SWZ15, WWX15, XY18, YZZ10, ZT17]. 2n/(n + 2) [CLW12]. 3 [BYH15, CW16, CS10b, CHS13a, FT17, FQ11, GL12, Kha13, KLL12, Lau10, LXZ13, LWY18, MZZ12, PR13]. A = B = C = 1 [XYZ16]. α^{2} [GLT10]. B_{1RevT ilted} [GCGJL18]. BV [FF12]. C^{1,α} [Che15]. C^{2} [GL12]. C^{∞} [LWZ16]. C^{ω} [LWZ16]. d [WW18]. D^{2} [GM10]. Ḃ_{2,1}^{5/2} [MOR^{+}16]. Ḣ^{3/2} [MOR^{+}16]. Ḣ^{5/2} [MOR^{+}16]. Ḣ^{s} [HF13, Mur14]. G [MMT19, XY14]. Γ [AZ12, DFP14a, DFP14b, Gin19, JS14, Moa11, Olb19, SZ12b, PSV10]. H [BRS17]. H^{m}(R^{3}) [DS14]. H^{s}(T^{n}) [Wan13]. H_{c_{1}} [Pen17]. int(K^{lc}) [RZZ18]. κ [DWZ10]. L^{1} [CNSS17, EJ14, OSW19]. L^{2} [CJ19, KM17, KPR15]. L^{2}(R^{d}) [BJLO17]. L^{3} [JS13a]. L^{∞} [FHK11, Kat19, Ohn14, Ohn16, DT15]. L^{p} [ZF12, vNVW12]. L_{2} [HNSW11]. L_{2}(R^{n}) [HR12]. L_{∞} [HNSW11]. L_{p} [Mar18, MP12]. m [KM18]. R [LWX11]. R^{2} [HKT18, NN19]. R^{3} [Dai17, DSY18, FPZ14, Wan12]. R^{d} [BM19]. R^{N} [BC14, KM18, FQ16]. S^{1} [GMM13]. N [Miz11, Tan15, YZ16, HMW11, Rei18]. p [BFDJ13, Gas19, GT16, Mei10, MRS16, SZ12a, ST19, Xu18]. p(x) [BGL16]. P^{⊤}P [GS17]. Q [ADL14, BCS15, CRWX16, Dai17, GGRB14, PZ11]. R^{N} [Moa11]. s [BV10]. S^{2} [GM10]. T_{n}^{(}k) [NNS18]. V [dlHHMV16]. φ [Tak13]. W^{1,1} [BKK18]. W_{2}^{1} [Kry10].

-Body [CCLM15, Rei18, YZ16]. -Bounds [Mur14]. -Convergence [AZ12, BRS17, DFP14a, DFP14b, Moa11, PSV10]. -currents [MMT19]. -D [CN15, CHS13a, CM19, FY13a, FQ11, Kha13, MZZ12, Per10, SWZ15, XYY19]. -Dimensional [BP19, HMW11, LWY18, WW18, ZT17, BOO18, LMTT15, Tan15]. -Dynamo [GLT10]. -Equation [XY14]. -Estimates [MP12]. -Exponential [Tak13]. -Harmonic [Gas19, GM10, GMM13]. -Initial [JS13a]. -Laplace [GT16, Xu18]. -Laplacian [MRS16, SZ12a, ST19]. -Laplacian-like [BGL16]. -Limit [Gin19, Olb19, SZ12b]. -Limits [JS14]. -Norm [CJ19]. -Peakon [HHK18]. -Points [BV10]. -Regularity [vNVW12]. -Schrödinger [BFDJ13]. -Soliton [Miz11]. -Solutions [KPR15]. -Sphere [DSZ19]. -Stability [Ohn14, Ohn16]. -States [dlHHMV16]. -System [Mei10]. -Tensor [KM18, BCS15, CRWX16, GGRB14, PZ11, ADL14, Dai17]. -type [OSW19].

. [GS17].

1 [Val15].

2 [GS18].

ABC [XYZ16]. Absence [Daf14, LvR15]. Absorbing [AKKL17, HJJ18]. Absorption [EP12, GM15, Kal12, KL18a, LS12a]. Abstract [HS10c]. Accelerated [GK10, Hal12, HS14]. Accelerating [Gar11].

Account [EW18]. Accuracy [Che14]. Acoustic [AZ17, CM14, GLZ17, IN13, Liu18a, NUW11, Ngu13, Spe14]. Acoustic-Elastic [GLZ17]. Acoustical [BV10]. Acousto [CS17]. Acousto-Optic [CS17]. Acting [CNR17]. Action [AHP13b, Bor19, Pan12b]. Active [CMWZ18, GPPP13, Gli13, HKK13]. Addendum [BA12, CK13]. Additional [LT11, Olb19]. Adhesion [DGVBW10]. Adhesive [Rou13]. Adjoint [BC11, Sen17]. Adjointness [Kre19]. Admissible [HR15]. Advection [AKKL17, EF15, Fis13, GKR18, Lam12, ASS16]. Advection-Diffusion [EF15]. Advection-Driven [Fis13]. Affine [AMW10, RZZ18]. Aggregation [BGL12, BGLV16, CKY13, Don11, Kai17]. Aharonov [AF16]. Airy [FK19a, GHH17, MS18a]. Aleksandrov [BL19]. Alfvén [RZ17]. Algebraic [Bra16, LW12, MP16]. Algebraically [vBM14]. Algebro [HGW14, HZFQ13]. Algebro-geometric [HGW14, HZFQ13]. Algorithm [ABCD^{+}18]. Algorithms [BCD^{+}11, YMYC10]. Allen [MT16, XZL10, AAD13, ABBK16, MT15, Tan15]. Almost [AMW10, CL17, KZ11, MO14, NPS13, Oh15]. Along [HHR17, PPPV16]. Alpert [GI15, GIV17]. Alzheimer [BFTT18]. Ambrosio [FI14, FL17]. Among [BLS15, Bét16]. Ampère [DF13, LR17]. Amphiphilic [DP15]. Amplitude [KDT19, RZ16, Whe13]. Analysis [ALS15, AAGP18, AAK14, ADK15, AH16, ÁCDP14, AV16, AZ12, ADHZ15, BC17a, BMSR^{+}13, BTZ15, Bos16, BP12b, BP14b, BPZ17, BL15, Cal15, CF11, CL13a, CDM13, CDR17, CFO19, CSZ19, CV12, DDGVM18, DMZ19, DSX17, ES10, EP12, FGN13, FI14, FFGHR17, FHX10, GLZ17, GI15, Gro10, HO15, IISD15, Ito18, JS13b, JMZ18, KM18, Lei13, LJ17, LZ19, MR15, NS13, NKV19, PUW18, Pra13, QWE19, Rey12, RP18, ST19, Tay18, WNRJ13, WK17, Zha19].

Analytic [HS13]. Analytical [CS14a, CT14, LM14, LX17c]. Analyticity [EMZ17, Hen10, MPN14]. Analyzing [Lau18]. Anderson [BFK16]. Angelesco [GIV17]. Angle [AAD13, Fel18, Mer18]. Angles [Wol19]. Anisotropic [AC14, BAC12, CGM16, CM11, CHN18, CP19, De 18, DDMM18, EW15a, FK19b, FHMP16, GSV19, HMN14, KLS15, LV12, Mil18, MRV12, TZ15, Zha14]. Annuli [AKKY17]. Anomalous [AKKY17, DN18, LL16b, MN18, Ngu17, Rey12]. Ansatz [Fri19, LXY19]. Antiplane [HO15]. Any [CK12]. Anyon [AN15]. Appearing [FK19b, Kia16]. Application [AHP13a, APSV19, ACZ14, AV16, BFV17, DSZ19, FL15, GP14, HZ10, LL18a, LN10, MM18a, MS18a, MS13a, MR15, PT11, ZBL19, BOS17]. Applications [AR19, BBR19, BS16b, BSW16, BMC18, BDWZ12, CLW17, CD11, Cri19, EW15a, EMZ17, FK13, Fri18, HMS14, HS16, Ign10, IT15, KPS18, Kol11, KMS15, LMR15, RTY16, Sch18b, vBM14]. Applied [ABGS10, AC14, DWZ10]. Approach [ALS15, AH18, AMP10, ALZ19, BP12a, BBS11, CCLCP13, CWE10, CS14a, CN15, DNS12, FG15, GPPP13, Gie14, GPI18, LS17, MT13, MZ18, MOS14, MSTY16, MZ13, SV19, Sou19]. Approaches [GO18]. Approaching [LTW14]. Approximate [NV12, Ngu13, NT19, dHGR14]. Approximating [CDZ13, Mil18]. Approximation [AF15, AM15, And12, AP14, BFGPE^{+}12, Beb16, BM10, BB10a, BFG^{+}13, BOO18, BLS15, BPW15, CMP13, Che14, CM12, CL13b, CT14, Cri19, DS10a, DNWW19, DKS16, FL17, FHK13, Gie14, GR13b, HNW10, HNSW11, Han18, HS13, JLX15, LS13c, MZZ12, PSSW15, WU14a, YMYC10]. Approximations [BHWY12, CT11, GS19, Hal12, Hal13, KLL12, LSW17, ZCO16]. Aquatic [HKK17]. Arbitrarily [BACP18].

Arbitrary [BHND18, LS13a, Ngu17]. Arc [And12]. Area [CL13b, Dai10, GL15, SSST15]. Area-Preserving [Dai10]. Area-Type [SSST15]. Arguments [DLZ15]. Arise [GLS10]. Arising [BG17, CGP13, DFT17, DX19, DH10, ERV17, HLWW18, JLX15, NOS12, PPPV16, Rou13, TW18b, ZLMZ18]. Artifacts [Ngu15, PUW18]. Artificial [MMP13]. Assisted [LJ17]. Associated [Bar14, BC11, FL19]. Assumption [AP15]. Assumptions [Kry13]. Asymmetric [BWW14, RW14]. Asymptotic [ADK15, AIK10, AZ12, BT16, BC17a, BFG^{+}13, Bia18, BDT12, Bos19, BP12b, Bre13, Can10a, Can10b, CFRT10, CDLLSG13, CF11, CL18, CY19, CH11a, CDK11, CEQW16, Cui13, DS14, Dek19, DS13, DSX17, Duc10, DIT15, DM14, DGVBW10, EP12, FRX19, FI14, FFGHR17, DFHM14, GM17b, Ghi13, GMT16, GW13, GH12, HMWY11, IKM17, JLL13, Mae17, MPR10, MY12, Mei10, Mor19, MR15, Ngu10, NOS12, Ohn15, PZZ19, Pol17, Pra13, Rei18, RP18, ST11, UWK12, WFL12, WK17, XY14, XYY19, Yam13, YZZ10, Yao19, YZ15, Yos17, Yos18, Yam16]. Asymptotically [MW17]. Asymptotics [AP11, BHND18, BC17b, CK11, CK13, CG10a, CQW18, Dan17, DF10, ET16, Gna15, HY14, KL18b, Len16, Per10, PY14, TZ13, dMIS10]. Atmospheres [HCHY16]. Atmospheric [LT19]. Atom [ZN19]. Atoms [CS18]. Attachment [GLL17, MS13b]. Attachment-Detachment-Limited [GLL17]. Attachment-Limited [MS13b]. Attenuated [Mon16, ST15a]. Attenuation [ILW16]. Attracting [BDFS18]. Attracting/Repelling [BDFS18]. Attraction [DFHM14, KK10]. Attractive [BLW19, CDNP16, GLW17, Kai17]. Attractor [YCW10]. Attractors [DN12, GGAS14, Pim16, ZK15]. Aubry [CGT11, DZ14]. Aurora [FMP18].

Austenite [Bev11]. Autocatalytic [LW12]. Autonomous [BBG16]. Average [LS13c]. Average-Distance [LS13c]. Averaging [BP10, Cer11, CL17, FS14]. Away [GN19]. Axial [HTW18]. Axis [Moo16, ZCO15, AAS19]. Axisymmetric [CS14c, DWX18, Kai10, LS13b, XZL10, Zha19].

B [FZ16, LZ18, SX13]. B-Spline [SX13]. B.C. [LT11]. Background [Dua11]. Backscatter [HHR09, HHR11]. Backscattering [BFRV13]. Backward [BST17b, BGT19, GS19, ST10, Ter11, YFK11]. Bakelman [BL19]. Balance [CT11, Daf14]. Balanced [DFT17]. Banach [AG17, FWW17, ZBL19]. Band [ALZ19, FW18a]. Bandlimited [Lit13]. Barotropic [Che19, DD16, FW18b, LZZ15]. Barrier [GHH17, LS15]. Barycenter [KP18a]. Barycenters [AC11, BCP19]. Based [ALM10, DSX17, HU13, Hel12, NN12, TW11b, XYD18, HL19a]. Bases [BJLO17, HR12]. Basic [Val15]. Basins [HS18]. Basis [BCD^{+}11, GI15]. Bath [BCL11]. Bayesian [TSA18]. Beam [BT19, HTW18]. Beams [DFP14a, DFP14b]. Becker [MP16]. Behavior [ADHZ15, BT16, BDT12, Bos19, Bre13, Can10a, Can10b, CDLLSG13, CTW17, CY19, CH11a, Cho16, CG11, CDK11, CEQW16, Daf13, DS14, DDF18, DS13, DM14, DFHM14, GW13, Hel12, HMW11, HMWY12, IS13, LvR15, LMW17, MY12, PZZ19, QW11, Rei18, ST10, Sou19, TW15, Ter11, WFL12, Wu14b, YZZ10, YZ15, Yos17, Yos18]. Behaviors [BHK^{+}19, HKR18, JWX13, LYZ16, Tak13]. Bellman [IS13, Zho15]. Below [BV13, BD18, KM17]. Belt [BHM17]. Benamou [BBS11]. Bending [CKZ17, XY14]. Benjamin [Oh10, HLR^{+}19, Wu16, Wu17]. Bent [GCGJL18]. Bent-Core [GCGJL18]. Besov [Tak10, XXK13]. Best [Mei10, TZ15]. Between [ABDD19, CS18, CC11, CM13, Wei12]. Beyond [AHP13a]. BGK [Yun15]. Bianisotropic [IKS12]. Biaxial [MN16]. Bifurcation [CCM12, CO12, CCLM15, DVW15, FHO16, GH10, HD17, LS13b, Lóp12]. Bifurcations [BL14, KP18b, Sch18b]. Bilayer [DHPW14, HD17]. Binary [BOS11, GM17c, Wan11]. Biofilm [DMZ19]. Biology [HLGMMM14]. Bioluminescence [BCS16]. Bipolar [HMW11, HMWY12, lLWW18, WFL12]. Bistable [HMSZ13, TV18]. Black [ILR17]. Blasius [CT14]. Bloch [ALZ19, CGLS18, GH18]. Block [LWZ16]. Block-Diagonalization [LWZ16]. Blow [BK13, BH17, BH18, BCJ20, HJ11, LMR13, Rey12, VW11]. Blow-Down [VW11]. Blow-Up [BK13, BH17, BH18, BCJ20, HJ11, LMR13]. Blowflies [LLLM14]. Blowing [MOR^{+}16]. Blowing-Up [MOR^{+}16]. Blowup [BGL12, CDW13, GM10, Sou19]. Bluetongue [GRT14]. BMO [DK11]. Bodies [CJP13, Ces11, DKN11, MR15, Rou13]. Body [ADMR14, BL11, BL14, CS15b, CCLM15, GS12b, KM13, MPS19, NT13, Rei18, SR14, YZ16]. Bohm [AF16]. Bohr [AF17]. Boltzmann [Li11, ABCL18, AN15, BCG10, BCL11, Cao19, Che18, CMWY16, DJMZ16, DYZ19a, DY10, Dua11, DL15b, GPT19, HX14, HY14, HWWY13, Jia12, JX15, JZ19, KY12a, Li09, LYZ16, lLWW18, Löb18, MT19, Str10, SY14, TAGP18, Wan11, WW15, Wan18, Wu14b]. Born [BFGPE^{+}12, CP13, MP14]. Bose [BIP16, BAC12, ER19, GM17c, GLW17, LLP16]. Bottlenecks [LMP11]. Bound [AKKL17, BV13, BCT19, Dro18, FY13a, Wal14]. Boundaries [AT10, DLZ15, FKN^{+}14, GS15a, JH18, KvMY19, KR10, Ves15]. Boundary [ADL14, ABL13, AI12, ADMR14, AGS13, AS13, AVP16, BMMP16, BOS17, BBT14, BKLU18, BSW16, BPS16, Beb16, BMY16, BdHQ13, BdHFS16, Ber12a, BBG12, BGLV16, BM12b, BK15b, BNDHV10, BKP13, BP12b, CDN10, CHN18, CDLLSG13, CCK18, CJN19, CDZ13, Che15, CH15, CKS15, CH19b, CGS17, CN15, CHS13a, Cui13, DDGVM18, DWY12, DSV15, DKT19, DN18, DL10, DMZ14, Ell12, ERV17, EHM16, FT17, FPVR13, GVWK16, GR13a, GP14, GW13, GX17, GSWZ18, Han18, HK15, HI19, Hof12, HLWW18, HL15, HMWY12, HJJ18, IY12, Iye19, JZ19, JT13, KMV18, KLO10, KT11, KMS15, KLS11, LL16a, Lee17, Len16, LX17a, LMZZ17, LMZZ18, LT11, LW14b, LXY19, LY19, Löb18, MPS19, Mel12, MMP13, MT15, MT16, MB16, Nes14, Ngu10, NS13, NP16, NOS12, Ohn15, Otw10, PPP13]. Boundary [Pra13, RSZ18, RP18, RTV17, Sal12, SZ12a, SM16, Sin10, TW18a, Ves15, WXY15b, WX19, Wan19, WNRJ13, Wol19, XY18, dHHI^{+}14, vBW11, DL13]. Bounded [BPS16, BC17a, BNDHV10, CR18, CEIV17, CIN18, EJ14, GG10, HX10, HJJ18, Kai17, KPS18, KMS15]. Boundedness [Löb18]. Bounding [HNW10]. Bounds [AHP13a, AV19, BOS11, CQ19, CT14, HPS19, HS10c, MOR^{+}16, Mur14, Spe14]. Boussinesq [DKS16, EW15a, JMWZ14, LPS13, MSZ13, SWX17, TZ18]. Branched [BBS11, COS16]. Branching [Cav12]. Bravais [Bét16]. Breaking [BFN^{+}13, CFGL17, FV18, GL19a, HD17, HT18, LPS10]. Brenier [Bon13, BBS11, CGS10]. Bresse [MM17]. Bridging [BLZ16]. Brinkman [EG19]. Brittle [BLZ16]. Brownian [GGAS14, KUV16, LSW17]. BSDEs [BEH15]. Bubble [CCHR18, CTW13, SW11b, YT11]. Bubbles [RW14]. Bubbly [AZ17, ALZ19].

Bulk

[ALS15, AET18, ERV17, FKM^{+}16, KUV16]. Bulk-Surface [ALS15, AET18, ERV17, FKM^{+}16]. Bundles [FGW13]. Burgers [AIK10, BP12a, BN14, BGN14, HI12, MZ13, ST15b]. BV [BBG16, CD11, Daf14, GY16b, PT18]. BV-Regularity [GY16b].

Cahn [GMT19, MT16, AAD13, ABK12, ABBK16, BBMN12, CGS17, DG11, Del18, DHPW14, EG19, KNR12, KP18b, MT15, OW14, OSW19, SP13, Tan15, XZL10]. Calculus [MZ13]. Calderón [HPS12, KT13, KLO10]. Calibration [BD18]. Camassa [CFGL17, DIT15, GL17, GLL18, LZ17, Tan18, dMIS10]. Can [EW18]. Canards [VW15b]. Cancer [SSW14]. Canonical [KP18a]. Cantilevered [HTW18]. Cap [BFM12]. Capillarity [BC17c, Can10a, Can10b, CH13, Gie14]. Capillarity-Viscosity [CH13]. Capillary [CP12, Hen10, MZZ12, NT18]. Capillary-Gravity [Hen10, MZZ12]. Caputo [LL18a]. Car [GM14, LMP11]. Car-Following [GM14]. Carcinogenesis [GP14]. Carrier [CDM13, CStW17]. Cascade [Lei16, MS14]. Case [ACM^{+}12, BDEM18, BG14, Ber12a, Bes16, BOO18, CL17, CQW18, DP14, GIP^{+}13, HLWW18, HL15, JT13, MS18a, NN12, RR13]. Cases [Kai10]. Catastrophes [Pao16]. Cauchy [AT14, AN15, DLZ12b, DLZ12a, HKT18, ILR17, LX19, LPS18, LS12b, MY12, SWX17, Yos17, Yos18, ZLMZ18]. Cavitation [BD18]. Cavities [CCH10, CJP13, CH11b, MC14]. Cavity [BYZ12, Lei10]. Cell [DGVBW10, EW18, GJMC12, JMNR11, NRS17, Sch18b, SSW14]. Cells [Gli13, GP14, ST11]. Centennial [Naz12]. Center [NRS17, VF13, VF15]. Centered [ZH10]. Central [BL11, Moo16]. Certain [KZ11, LT17]. CFIE [CM14]. CGO [KS14]. Chain [CD11, PWW17].

Chains [DP14, HR10]. Change [Mit10, NV12]. Changing [BV18]. Channel [GJZ15, NT13]. Chaos [BÖ19, HS13, YFK11]. Chaotic [Bes12]. Chaplygin [CQ12]. Chapman [GSW16]. Character [Igb17]. Characteristics [CCMW19]. Characterization [BGL12, Bra16, KDT19, Lei10, MPR10, ST15a]. Characterizing [Kar12]. Charge [CCM14, CDM13, Sab13]. Charge-Carrier [CDM13]. Charged [BKR16, Bos12, BPZ17, WLT16]. Cheap [MM11]. Cheeger [BFLS12]. Chemical [DFT17, GLS10, JMNR11, LW12, NT14]. Chemically [BP14b, BPZ17]. Chemostats [CHK15]. Chemotactic [AKKL17]. Chemotaxis [Bla18a, Fis13, GN15, HLWW18, Jia19, KK17a, NT14, TW11a, TW15, WWW12, Win15, ZZ14, ZLMZ18]. Chemotaxis-Fluid [Bla18a]. Chemotaxis-Haptotaxis [TW11a, TW15]. Chemotaxis-Navier [ZZ14]. Chirality [CFO19]. Circle [DL15a, JT13, Pim16, Sim16]. Circular [GW18, YT11]. Circularly [LW14b]. Clark [CLW17]. Class [AT10, BBR19, BÖ19, BST17b, BDT12, BKP13, Cal15, FZ16, FZZ18, GLS10, GLY18, GYY18, HMS14, HS16, HMN14, HHPZ17, Kar12, KZ11, LWZ18, LWX16, LT17, LX19, MPR10, MQS12, Mar10, MX19, Pas13, Sch18a, SP13, ST10, Sof18]. Classes [BL14, De 18, LR15b, WZ17]. Classical [CQX18, CC11, Cho16, DWYZ12, DLZ12a, HKN16, JLZ18, LXZ13, MP13, WZ13b, Xu11, YY10, Zhu15]. Clausius [Alm17]. Climate [ABR17]. Cloak [Ngu17]. Cloaking [AKKY17, NV12, Ngu13, Ngu17, NT19]. Close [LX17a, Che15]. Closed [DFP14b, EW15b, MJ14]. Closest [BBCD^{+}18]. Closure [JS14]. Cnoidal [ETZ13]. Coagulation [CMM10, DF10, LvR15, Sri11].

Coagulation-Fragmentation [DF10, LvR15]. Coarsening [BOS11, MS13b]. Coating [BC11]. Coefficient [AL10, DV10, EP12, LT11, PZZ19, RZ16, WZ16, ZF10]. Coefficients [AB10, AC14, ACZ14, Beb16, BCD^{+}17a, BC17c, CDL16, CMM10, DO16, DK11, DK14, FK19b, GH10, HKOP10, ILN11, ILW16, JK10, JJL10, KPZ19, Kia16, Kry10, Kry14, LYZZ14, RZ11, SSH19, Sus13, Zho15]. Coercivity [Zha10]. Coherent [Ilm16]. Cold [Otw10]. Collections [Dai10]. Collision [FPTT12, Jia12, Löb18, MR15]. Collisionless [NS13, Sch14a, Sch17]. Collisions [HY14]. Column [CCC^{+}17]. Combined [Mel12]. Combining [KK16]. Combustion [Lai14]. Coming [CMM13]. Communication [CH19a]. Commutability [JS14]. Commutative [MNS11]. Commutator [Tak10]. Compact [BLS15, CCLCP13, CC11, Lau10, Syl12, Tan15]. Compaction [HS18]. Compactly [HMZ15, KLL12, WU14a]. Compactness [AGN19, BPS16, CL13a, FZ14, FL15, IISD15, LL18b, Vis18]. Comparison [RTZ17]. Compensated [ZCO15, ZCO16]. Competition [ABR17, LMS16, LTW14, Mor19]. Competition-Diffusion [Mor19]. Competitive [DP15, FGW13]. Competitive-Cooperative [FGW13]. Complete [WK17]. Completely [LA14]. Completeness [Pan12a]. Complex [AC14, AH16, CDW13, DFT17, FK19a, GHH17, HMZ15, Kre19, LTV17, ST15b]. Complex-Valued [LTV17]. Compliance [CLLS17]. Component [Ber12a, CFGL17, IT15, Mil18, PSZ19b, JRK19]. Components [QS12]. Composite [ABL13, DFP14a, DFP14b, DX19, Lai18]. Composites [BLZ16, Bel10, Bel17, Mil18]. Compressed [Bar14, BCO17, BDWZ12]. Compressible [BP19, Bre13, CJN19, CPZ17, CMWZ18, Che12, Che14, CTW13, CHS13a, CWYZ16, DF11, DWYZ12, DLZ12b, Evj13a, FZZ18, FGN12, FKM19, FPZ14, GT10, GJZ15, Hof12, HKT18, HHPZ17, HW17, HW13a, HW13b, HLX11, HLW12, JTW16, JK10, JJL10, JLX15, JLZ18, JWX13, JLL13, Jün10, Kot12, Kwe12, KK17b, LXZ13, LZZ15, LMW17, aLW18, Li19, LLW15, LYZZ14, LW14b, LT19, LY19, LZ18, MPZ15, NP11, PWG11, Per10, PSZ19b, QW11, SS15, SS19, SW11b, Smi17, TYZZ13, TY11, VW15a, VY16, WX11, WFL12, WXY15a, WY15, WXY15b, WZ13b, WZ17, WK17, Xu11, ZF10, Zhu15]. Compressive [KNW15]. Computable [Lee10]. Computer [LJ17]. Computing [BC11, vdBW19]. Concavity [SR14]. Concentrated [BM18, CCNP17, RP18]. Concentration [DR13, PPPV16]. Concept [ACZ14]. Condensate [BAC12]. Condensates [ER19, GM17c, GLW17]. Condensation [BIP16, LLP16, LS12a]. Condensation/Evaporation [LS12a]. Condition [CDLLSG13, CDZ13, CLW17, CGS17, GW13, HI19, JZ19, Kim09, KL18a, Sin10, TZ15, dCPS16, Kim13]. Conditional [BdHFS16, LZ18]. Conditions [AAD13, AS13, BMMP16, BPS16, BK15b, BL19, BNDHV10, BKP13, CN15, Ell12, EHM16, FT17, FY13b, FPVR13, GR13a, HS10a, HK15, Hof12, HJJ18, KMV18, Lee10, LL12, LS18, Löb18, MPS19, MMP13, MT15, MT16, MB16, PPP13, SM16, WXY15b, vBW11]. Conducting [Alm17, BPZ17, MNT13, MPZ15, NP11, WZ13b]. Conductive [AH13]. Conductivities [GIP^{+}13, LTV17]. Conductivity [AHP13b, AGS13, DV10, HMN14, ILW16, LTV17, LX17a, Li19, MNT13, Ped15]. Conduit [EW15b]. Cone [AT14, HZ19]. Cone-Like [AT14]. Cones [FWW17, Lee16]. Configurations [BL11, Car14, HSV17, HO15, dlLSZ17]. Confined [CCM14]. Confinement [AMV15, BAC12, HK10, MPS17].

Conformal [BK15a, HMN14, KLO10].

Conformally [CCFdlL14, dHHI^{+}14]. Conic [HZ19]. Conical [Moo16, MO14]. Conjecture [DN18, MW17, YZ16]. Connected [BOO18, BLS15, dlHHMV16]. Connectedness [DNWW19]. Connecting [vdBMJLM11]. Connection [Igb12]. Conservation [AGN19, CR18, CW13, CD11, Daf13, DWY12, GS19, HLGMMM14, Jun14, KPS18, KMS15, MS13a, MY12, NPS18, PT18, VK18, WDL18, XYY19, Yos17, Yos18, ZLMZ18]. Conservative [Ell12, LPR12, LZ17]. Conserved [Daf14]. Conserving [ABBK16]. Consistency [Bar14, OR17]. Consistent [AHP13a, CH11a, RR15]. Constant [AFT15, BCDG16, CMM10, HNW10, HKOP10, KPR15, Li19, Zho15, vBW11]. Constituted [BGMŚG12]. Constitutive [Chu14]. Constrained [Koc16, MN16]. Constraint [Ber17, BHL18, DNWW19]. Constraints [AF15, AP14, CL13b, LW15]. Construction [CM18, HHK18, LM14, MPZ15, WU14a]. Constructive [FW18b, PD17]. Contact [BP19, BS16b, Fel18, HMS14, lLWW18, Rou13, VW15a, WX11, WY15, Wol19, ZT17]. Contacts [Pao16]. Container [GR13a]. Containing [DLZ12a]. Continua [Gas19, HWZ12]. Continuation [CGS10, LN10, SW18]. Continuity [BCDG16, CDS19, CNSS17, MS11, Sto19]. Continuous [AG17, ACJ12, DF10, GJMC12, Ohn14, TZ18]. Continuum [AG16, BCS15, BK18b, CEH14, GLL17, HD17, HY19, LX17c, TSA18]. Contraction [EJ14]. Contrast [Beb16, Bel10, Bel17, BPP15, CCC18b, PSZ19a]. Contrasting [GS15a]. Contrasts [LL16a]. Control [ABL13, ABØP19, BV18, FQS10, MM11, MNS11, PRT15, Sof18, Zhi19]. Controllability [BCS16, Lau10, dAdM18]. Controlled [MS16]. Convection [ACJ12, BP10, CCC^{+}17, CPT10, JT13, Poh15]. Convection-Diffusion [JT13, Poh15, ACJ12]. Convective [AH13]. Convergence [Ale16, AET18, AZ12, BA10, BA12, BBMN12, Bia18, BCD^{+}11, BRS17, BHWY12, CMM10, CDPS17, CPSW16, CCLCP13, CS15b, Che12, CT11, CT16, DFP14a, DFP14b, DGV16, Del18, DLZ15, FF12, FW18b, GL17, Gie14, GM13, HHR17, HY19, HMWY11, IM10, Kar12, Le10, LP19, MT15, MT16, Moa11, MPS17, MN12, NTW19, Pen15, Pen17, SW11a, Sof18, Sri11, Xu16, dCPS16, Can10a, PSV10]. Converges [HS19]. Convex [AS14, Ash13, Ber12b, BPW15, Cao19, Dai10, DS19, DK14, Feh13, Fri19, HR10, KY15, LV13, MMT19, RZZ18, Sch18a, SV18, Tan15, VK18, XYY19, DLSV12]. Convexity [AGN19, FGR15, Kry13, ZCO15, ZCO16]. Conveyor [BHM17]. Convolution [CM14, DL15a, PZ17]. Convolution-Type [PZ17]. Convolutions [ZBL19]. Cooling [AL10, ABCL18]. Cooperative [ÁCDP14, DP13, FGW13, ILN11]. Coordinates [Hyn19]. Copolymer [CP10, CP11, Gla17]. Copolymer-solvent [Gla17]. Core [GCGJL18]. Cores [BBO19]. Corner [Ben17, Kwe12, LX17b, ZH10]. Corners [Blå18b]. Corrector [BAP13]. Correctors [BFFO17, CM12, Pra13]. Correlation [Nov18]. Correspondence [AF17]. Corrigendum [BCJ20, HM13, Kim13]. Corrosion [PPPV16]. Cosmological [AFT15]. Cosserat [Gas19]. Cost [CDS19, GKR19, KW12, Che15]. Costs [Cav12]. Coulomb [BST17a, GZ18, HY14]. Counterexample [Fri19]. Counterexamples [Tak10]. Counting [BR11, LV15]. Coupled [ADL14, ALS15, AET18, BG17, CPP18, DZ14, ERV17, EKR18, KK17a, KLS11, LMP11, LN10, MSTY16, NRS17, Ohn16, PZ11, Sch18b, Tro17, XZL10]. Coupling [BT19, MLD19, RT17, ST11]. Covariance [CdGDN18]. Cowan [DN12]. Cracks [IO16]. Cradle [BFDJ13]. Crime [BWW14, CCM12]. Criteria [LL18b, WZ13a]. Criterion [CCFdlL14, CJP13, HLX11]. Critical [BK13, BD18, BCL18, CLW12, CM18, CP19, CQW18, DF15, DS10b, DR13, EEW11, GN19, JMWZ14, Kac14, LWX11, MM18b, Mur14, NNS18, Wan13, XXK13, YZ18, ZF12, dR18]. Criticality [Mos18]. Critically [KM13, KMS17]. Cross [BDPS10, CFSS18, CDNP16, DMZ19, DFP14a, DFP14b, DLM14, GV19, HJ11, HNP15, Kut15, LPR12, BCJ20]. Cross-Diffusion [BDPS10, CFSS18, DMZ19, GV19, HNP15, Kut15]. Cross-Sections [DFP14b]. Crowded [WLT16]. Crystal [ALZ19, BBO19, CRWX16, CKY18, CMWZ18, CDK11, DS14, DW13, GW13, JL19, Kom15, LR13, LLW15, LX16, LX17c, SW11a, Xu18]. Crystalline [GZ14]. Crystallization [Bét16]. Crystals [CHT18, GCGJL18, GSV19, HM12b, HW13b, INSZ14, KK18, SV19, SS15, Tay18, HM13]. Cubic [KMV18, Liu19, HTX15]. Cubic-Quintic [KMV18]. Cucker [BHK^{+}19, CFRT10, CY19, CH19a, HKK15, HKR18, Pes15, PRT15]. Cues [EW18]. Curl [BB10a, CP13]. Curl-Free [BB10a]. Current [AHP13b, BMSR^{+}13, DWZ10, HMN14]. Currents [ABGS19, MMT19]. Curse [FV18]. Curvature [BC19, BK18a, CHW16, CMP13, CM13, Dai10, DHPW14, GMT16, Hel12, Olb19, TW18b]. Curve [ESvR12, dR18]. Curved [Ara16, HZ19, NW17, PR13]. Curvelet [YY14]. Curves [AP15, CCG18, Dai10, LMS16, WX16, Wan19]. Cusp [MR15]. Cut [KK17b]. Cutaneous [FHX10]. Cutoff [CH15, TAGP18, YY18]. Cylinder [CDN10, RTV17, Sen17]. Cylinders [Lóp12, MN12]. Cylindrical [BHK^{+}19, GH14, Kal12].

D [Kim13, AHKM15, BM12b, BYH15, CWY14, CSW15, CW16, CS10b, CN15, CHS13a, CM19, DYZ19b, ER19, FY13a, FT17, FQ11, GL12, JWX13, Kha13, Kim09, KT18, KLL12, LXZ13, LW14b, MZZ12, Per10, PR13, SWZ15, Sug16, WX11, WWX15, XY18, XYY19, YZZ10]. DAEs [Ria10]. Damage [HS16, RR15]. Damping [DYZ19b, Hyn13, Kia16, LLW17, LT11, Mei10, RZ17, RR17, RZ16, VY16, WWX15]. Data [ADL14, AGS13, BFRV13, BYH15, BDG13, BV18, CCK18, CJN19, CT15, DWYZ12, DT14, DLZ12a, FZ16, FZZ18, FT13, FQ16, GL12, HKT18, JS13a, Kia16, LT17, LYZZ14, Liu19, Ngu15, NN19, Oh15, RZZ18, TYZZ13, Tsu12, WZ17, Win15]. Datum [LX19]. Dean [CSZ19]. Decay [BPS19, BW12, BHM17, Bra16, CL18, CWYZ16, DKN11, Deu13, DNK12, EW15a, EKR18, HHMM18, HL15, HW13a, KKT17, LW12, MP16, SW11b, UWK12, VZ15, Wan12, Wu14c]. Decaying [FY13b, vBM14]. Decomposition [ABCD^{+}18, Sch14b, YY14]. Defect [BR11, CCC18b, DVW15, Gli13, HW11, Rou13]. Defected [IO16]. Defined [BOO18]. Definition [LL18a]. Deflagration [GSW16]. Defocusing [BK15b, KM17, YZ18]. Deformation [BC17a, KLO10]. Deformations [MO14]. Degasperis [HZFQ13]. Degeneracy [Sug16]. Degenerate [ÁCDP14, BEH15, BKLU18, BL11, Cal15, CKY13, CLW12, DMZ19, Del18, DF10, DFV18, EJ14, Fis13, FS15, GR15a, Ghi13, Gia15, GP11, Hal13, Kai17, KLW17, KZ18, KSW13, LZZ17, LT19, MY12, NSS17, Nii12, Pan12a, RTZ17, Sho10, WWW12, WZ16, Wan19, Zho15]. Degond [ST18]. Degradating [GR13a]. Delay [AHØP13, CH19a, CN15, DAP19, GH12, HWZ12, KPR15, KKT17, LWY11, LW15, MPN14, TV18, YCW10]. Delay-Differential [MPN14]. Delayed [MOZ10, MNS11, ST11, ST15b, VF13, VF15, MOZ12]. Delays [CHK15]. Delta [CY15, CM19]. Demixing [BOS11]. Dendritic [IT15]. Denoising [Val15]. Dense [HR19, Med14, Tay18]. Densities [Gli13]. Density [CP19, DFV18, Fri18, Gla17, GR13b, HMN14, HW14, JWX13, KLO16, Li19, PT11, SV11, ZF10]. Density-Dependent [JWX13, ZF10, HW14]. Dependence [ACJ12, GS18, HKT18]. Dependent [AN15, BJ10, BMP18, CKS15, CG19, CMM13, DF10, Deu13, EHM16, Evj13a, GIP^{+}13, GS10b, GP15, GH12, HWZ12, JK10, JWX13, Kia16, KS19, LYZZ14, LTW14, MSZ19, MO15, RTZ17, RT17, SCB17, YMYC10, ZF10, HW14, NT19, SCB20]. Depending [BT12, Can10a, Can10b]. Deposition [dCPS16]. Derivation [BFG^{+}13, CM12, CSZ19, CFF19, FPP19, JM12, LR13, Tri18, WZZ15, WLT16]. Derivative [AS15, BDEM18, FS15, GHLN13]. Derivatives [LL18a, LRdS18, LP14]. Derived [SX13]. Descent [Len16, LLN19]. Describing [BP14b, BGL16, Evj11, KLW17]. Desreux [ST18]. Detachment [GLL17]. Detectors [SY17]. Determination [ADMR14, BDEM18, CKS15, DV10, LR15a, Moo16, MRV12, SZ12a, Sin10]. Determined [DMZ14]. Determining [KS19, KLS11, LLM19, LT11]. Detonation [GSW16]. Detonations [LWZ15]. Deviations [BM12b]. Dewetting [Lóp12]. Diagonalization [LWZ16]. Diameter [BHL18]. Diatomic [FW18a, Qin15]. Diblock [CP10, CP11]. Dichotomy [DL10, DL13]. Dictionaries [KNW15]. Diffeomorphism [XYD18]. Diffeomorphism-Based [XYD18]. Difference [FK19a, GK10, LL16a, MJ14]. Different [AFK^{+}18]. Differentiability [BFV17]. Differential [AP14, AMW10, AP18, BCD^{+}17a, BC11, CCC18a, CMM13, DdMH15, EKR18, Gar11, GS18, GK10, Hal12, Hal13, IM10, KT13, KRW15, LWY11, LW15, LRdS18, MPN14, Mit10, MP14, RZ11, Sch10, SSZ19, YCW10, YZ14b, BGHP18]. Differentiation [GJMC12]. Diffraction [dHHI^{+}14]. Diffuse [Abe12, ALS15, BM10, Cao19, ES10, CP11]. Diffuse-Interface [CP11]. Diffusion [ASC19, AB10, ÁCDP14, ACM^{+}12, AS13, BAP13, BMP18, BK13, BFG^{+}13, Bia18, Bla18a, BCD^{+}17a, BP10, BCJ20, BDPS10, CWY14, CCV15, CFSS18, Cer11, CL17, CKY13, CLW12, CQ19, CK11, CK13, CEQW16, CFF19, DMZ19, DF10, DLM14, DFT17, DH10, DZ15, DLZ15, EP12, EF15, EM10, FHK13, GV19, GR15a, GLS10, GLY18, GG10, GM15, GP14, GST13, HS10a, HHMM18, HJ11, HNP15, HV13, ILN11, IKM17, JT13, Kai17, KZ18, KSW13, Kut15, LS13b, LPR12, LZZ17, MN18, MOZ10, MX19, MB16, MS16, Mor19, MPZ15, NT14, PSV10, PZZ19, Poh15, RTY16, RTV17, SP13, SM16, SV14, TW11a, TW15, TV18, VW11, WWW12, WW12, WW10, YMYC10, ZZ19, ACJ12, MOZ12, ASS16]. Diffusion-Absorption [EP12]. Diffusion-Homogenization [BAP13]. Diffusions [BM15a, INRZ10]. Diffusive [CPSW16, DL10, DL13, DM14, HNP13, JX15, LWZ18, LT13, Löb18, Wan11]. Diffusivity [GS15a]. Dilute [Alm17, HM12a]. Dimension [CPSW16, CH13, CK12, CKY18, CG11, Dro18, EM10, DFHM14, FV18, HPS19, HL15, KK16, Lau10, LLW15, Nol11, Nov18, VK18, XV10]. Dimensional [ABCL18, ALST14, ABK12, BDX14, BP19, BFN^{+}13, BV10, BBMN12, Bét16, BD18, BCJ20, BMR14, CCNP17, CF14, CNS10, CCLCP13, CHKP19, CQ12, CKZ17, CCHR18, CQX18, CCC18b, CFO19, CP19, CEQW16, CQW18, Del18, DS10b, DWX18, EP12, FRX19, FPP19, FL19, GIP^{+}13, GZ13, GR13a, GS12b, GW15, GLW17, HLR^{+}19, HJ11, Hof12, HHPZ17, HLWW18, HW13a, HW13b, HMW11, HLX11, HLW12, HW14, HV13, JK10, JMWZ14, Kai10, KLW17, Lac15, Lai18, Lei16, LR11, LS12a, LWY18, aLW18, Li19, LS15, LPS13, Mar10, Mas11, MOR^{+}16, Mil18, MMP13, Nii12, NW17, Nov19, OW14, PW18, PSSW15, RV12, ST15a, SW11a, SdlL13, TYZZ13, Tan15, WY15, WXY15b, WY13, WZ17, WW18, XY14, XZ15, YY14, YZ18, ZH10, ZZ14, ZT17, BOO18, DZ15, LMTT15, LYZZ14, MPT19, Pol17]. Dimensionality [BFY15]. Dimensions [BYZ12, BFRV13, BBG16, BW12, CCLCP13, CRWX16, CDX17, CY18, Daf13, GMP13, GH18, IY12, KY15, LMR15, LX17c, Pin19, Pu13, vdBW19, LY19]. Dini [WZ16]. Dipolar [BAC12, BJ16, ER19, Tri18]. Dipole [Le19]. Dirac [AMV15, BCT19, BC17b, CCNP17, DR13, KK10, Lee16, Sab13]. Direct [KDT19, LP14, Wu17]. Direction [Kai10]. Directionality [HMZ15]. Directions [GP11]. Directors [KK16]. Dirichlet [AT14, Ash13, BFV17, BHL18, BM15b, CCMW19, GL19b, Han18, HPS12, Kal12, LT11, LS12b, OR17, RZZ18, EH16, TZ13, TZ15]. Discharge [ST18]. Discontinuities [BP19, BFRV13, WY15]. Discontinuity [VW15a, WX11]. Discontinuous [BGL16, CP12, DO16, DV10, GS19, LT17, Ohn16, RZ11]. Discrete [BF19, BFDJ13, BPP15, BK18b, CGH10, CMP13, GM13, GJMC12, HY19, LS13c, Mae17, NSS17, Pes15, SSH19, Wu16]. Discreteness [GVZ16, Syl12]. Discretization [CM14, HS13]. Discretized [Hal12]. Disease [Ai10, BFTT18, LWZ18]. Disease-Transmission [LWZ18]. Dislocation [BBO19, Gin19, HO15, NN12, SZ12b]. Dislocations [AV16, FPP19, MPS17, SV19]. Dispersal [LTW14]. Dispersion [AI12, ABR17, AMP10, BCdSN18, CHL17, CG10a, DM15, KDT19]. Dispersions [GR13a]. Dispersive [Aud12, CDL16, GLL18, HJ15, LST12, LPS13, Pu13]. Displacement [DT14, Sch18a]. Dissipation [DN18, DZ15, HLW12, JMWZ14, LR11, MPS17]. Dissipative [ABCL18, ATSR19, Bra16, BM19, CCCdlL17, CN15, Dua11, UWK12, Yam13, Yam16]. Dissipativity [Zha14]. Dissolution [HKK17]. Distance [Cav12, GR13b, KW12, LMS16, LS13c, MRT14, ZCO15]. Distances [CG10c]. Distorted [CDN10]. Distributed [GST13, Mit10]. Distribution [BCG10, BKR16, LTW14]. Distributions [CdGDN18, KZ11, Tak13]. Divergence [Ale16, BB10a, BCDG16, DKR15, DK11, DX19, Kry10, Ria10]. Divergence-Free [BB10a]. Divergence-Type [Ale16]. Divisible [FJ18]. Divisors [TY11]. Does [Ber12a]. Domain [ALS15, AH13, AET18, BHK^{+}19, BDEM18, BPS19, BMSR^{+}13, BCG10, BNDHV10, BP12b, CHL19, CDLLSG13, CM14, CPP18, GS10b, GY16b, HKK15, JX15, Lau18, MR15, PPP13, Sal12, Sch14b, WK17, Zha19]. Domains [AT10, AT14, AVP16, BDEM18, BPS16, BAH17, BDSS18, BO16, CDL16, Cao19, CHT18, CDD^{+}12, CR18, CIN18, CEQW16, CQW18, DN18, FGR15, GS10a, Har18, HX10, HK15, Hof12, HJJ18, JJN13, JH18, Kal12, KPS18, KN18, Lac15, MS14, Pla14, DT15, Sei14, WXY15b, Xu16]. Dominated [CCV15]. Dominating [BT16]. Doniach [Pen17]. Doping [LMZZ17, LMZZ18, TWW15]. Doppler [ST15a]. Döring [Lai14, MP16]. Double [DG11, DFV18, RW14]. Double-Well [DFV18]. Doubly [AS14, HL19b, dlHHMV16]. Down [BHM17, CKV18, KSW13, VW11]. Drainage [CV12]. Drift [GG10, INRZ10, KT11, MX19, WW12, WZ16]. Drift-Diffusion [GG10, MX19, WW12]. Drifts [KK15, KZ18, MPT19]. Driven [BBT14, BHR16, BCL11, Fis13, DFHM14, GGAS14, MPS17, QWE19, ZBL19]. Driving [GMT16]. Drop [FL15, Tre13]. Droplet [ABBK16, GW18]. Droplets [Hel12]. Drops [Wol19]. Dry [Pao15, Pao16]. Dual [ABDD19]. Duality [DLM14, LR15a, PY10]. Due [HCHY16]. Dumbbell [LS12b]. Dynamic [CGS17, DS19, Fel18, FGN13]. Dynamical [BGAHS17, CV16, CDZ13, LN10, YZ15]. Dynamics [CFRT10, CSW15, CDZ13, ERV17, Evj13b, FGW13, FKM^{+}16, FL12b, GRT14, HKK15, HKT18, IM18, KLW17, KT17, KvMY19, KT18, Kot12, Kwe12, Lam12, LM14, LLP16, LT19, LS17, LSW17, MM17, MJ14, MY17, MV19, Muñ12, NTW19, PSZ19b, SS15, Sch18b, SWZ15]. Dynamo [GLT10].

Eckhaus [GSWZ18]. Ecological [RSS17]. Eddy [BMSR^{+}13]. Edge [DXZ18, FPP19, PR13]. Edges [Blå18b, CNS10]. Effect [ABR17, BBO19, CDN10, HMWY12, HT17, JJ18, Jun14, LWX16, LX19, Nad10, PPP13, RTV17, WWW12]. Effective [AZ17, HM12a, LS13a, Mil18]. Effects [BFLN16, Bel10, BBG12, BDT12, BGL16, Han18, PW18, SR14]. Eigenfrequencies [NT13]. Eigenfunction [BDWZ12]. Eigenfunctions [Bec18, Blå18b, CDN10, CCC18b]. Eigenvalue [AF16, BFK16, BR11, CCG10, CG10c, Kol11, LZZ17, Nad10, PZZ19, Ria10, TZ13]. Eigenvalues [BBG17, BHL18, BL15, CGH10, CLM17, DL15a, GVZ16, HS10c, HKOP10, HKOP11, LV15, LR15a, Syl12]. Einstein [AFT15, BIP16, BAC12, ER19, Faj16, GM17c, GL15, GLW17, HM12a, LLP16, Tha19]. Elastic [ADMR14, BFGPE^{+}18, Bel10, BC17a, Bel17, CJP13, CS10b, DD18, EH13, GLZ17, HS18, KS14, Len14, LL19, LS12b, MO15, MRV12, MN16, Rou13]. Elastica [BPW15]. Elasticity [AV16, AKKY17, BFGPE^{+}12, BD18, BTZ15, CHN18, CC10, CDK11, FI14, Gie14, Han14, MC14, RZZ18, Zha10]. Elasto [BMY16, DF15]. Elasto-Plastic [BMY16]. Elasto-Plasticity [DF15]. Elastodynamics [CY18, NP16]. Elastomers [Ces11]. Elastostatic [LL16b]. Electric [AHP13b, ABGS19, HHR09, HHR11]. Electrical [HS10b, HU13, KT13, KLS11]. Electrically [BPZ17]. Electrified [FHO16]. Electrode [Sei14]. Electrokinetic [BFS14, FGJ11]. Electromagnetic [BYZ12, CLM17, CH11b, HL11, KS19, LR15a, Lei10, LWZ11, LX17b, ZN19]. Electromagnetism [CC11]. Electron [GMP13, GL19a, Sab13, ZN19]. Electron/Positron [Sab13]. Electronic [Gli13]. Electrorheological [BR17]. Electrostatic [Li09, Li11, Sim16]. Electrothermal [BGL16]. Electrowetting [FGJ11]. Ellipsoidal [Yun15]. Elliptic [AC14, Ale16, AM15, AR19, ABCD^{+}18, BMMP16, Beb16, BFFO17, BRS17, BCD^{+}17a, DP13, DKR15, DK11, DX19, FR17, HKOP11, Kar12, Kat19, KK15, Lam12, Ohn16, PZZ19, QS12, Sen17, Sus13, WX19, Wan19, Xu16, Zho15]. Elliptic-Type [BRS17]. Elliptical [YZ16]. Ellipticity [LV12, MRS16]. Embedded [KW12]. Embedding [Fri18]. Embeddings [KW11]. Emergence [CFO19]. Emergent [HKK15, HKR18]. Enclosure [KLS15]. Ends [CDN10]. Energies [AG16, BPP15, BCS15, BK18b, CCV15, CT16, De 18, FI14, NSS17, PSZ19a, RTT19]. Energy [AL10, BPS19, Bev11, BKP13, BW17, BPW15, CCMW19, Gin19, Gli13, HSV17, HHMM18, JS14, KP18b, LMTT15, LS18, Li09, Li11, LLW15, Mas11, Moa11, MN16, Pen17, SZ12b, Tri18, VZ15, dL14]. Enhanced [EW19, HI12]. Ensemble [BHK^{+}19]. Ensembles [DXZ18].

Enstrophy [Lei16]. Entropic

[CDPS17, DL18, RR15]. Entropies [JJN13]. Entropy [AIK10, AGN19, BFY15, Can10a, CHY19, CV15, DDGW18, DLM14, Gie14, LT13, Lu13, LZ18, MLD19, Pan12a, Ter11, VW15a]. Entropy-Preserving [MLD19]. Enumeration [BL14]. Environment [ÁCDP14]. Environments [IM18]. Epidemiological [CTW17]. Epitaxially [GZ14]. Equal [ILN11]. Equality [LS18]. Equation [ASC19, AB10, AIK10, AAD13, ABCL18, ACM^{+}12, AT14, ABK12, ABBK16, AN15, ADHZ15, Ash13, BT19, BCS16, BIP16, BBT14, BSW16, BBMN12, BdHQ13, BdHFS16, BCG10, Ber12a, BP12a, Bes12, Bes16, BCL18, BFDJ13, BCL11, BCdSN18, BGM19, BCT19, BGHP18, BP10, BC17b, BN14, Bre13, BGN14, CCNP17, CEH14, CMM10, Cao19, CDW13, CHKP19, CWH18, CKY13, CS10a, CDZ13, Che19, CR10, CMWY16, CH11a, CPT10, CG10a, CM18, CEQW16, CQW18, CV12, CHS13b, CKV18, DDF18, DDGW18, DG11, Dek19, Del18, DAP19, DSY18, DV10, DLVW13, DKT19, DHPW14, Don11, DM14, DP14, DGVBW10, EW19, EP12, EL17, Ess16, FT17, FG15, GS12a, GPT19, GL17, GLL18, GZ13, GP19, GS10b, Gna15, GM15, Gra19, GPI18, GZ18, GHLN13, GY16b, GJMC12, HX14]. Equation [HPS19, HL19a, HY14, HLR^{+}19, HHK18, HI19, HWWY13, HI12, HJJ18, ILR17, Igb12, Igb17, Ign10, IY12, ILP16, JX15, JZ19, JZ10, Kai17, KVM17, KVM18, KM17, Kha13, KPR15, KS19, KY12a, KY15, KKT17, KK10, KNR12, KMM11, LLW17, LvR15, Lau10, Le10, LP19, LLP16, Li09, Li11, LWX16, LX19, LLM19, LW16b, LLLM14, LPS10, LP16, Liu19, Löb18, MS18a, MN18, Med14, MZ18, MBPS13, Mit10, MMP13, MT15, MT16, MBK13, MZ13, NS12, NV12, NT18, OW14, OSW19, Ovc11, Pan12a, Poh15, PY14, PWW17, RV12, Sac18, Sou19, SV14, ST17, Str10, SY14, Tak13, Tan18, Tan15, TAGP18, TV18, Tsu12, Wan13, WW15, Wan18, Wan19, WY13, Wu14b, Wu17, XY14, Yam13, Yam16, YY18, dL14, dMIS10, vdBMJLM11]. Equations [AHØP13, AS14, AC14, ACJ12, AS15, AAS19, AMW10, AP18, BAP13, BMMP16, BAC12, BR18, BSW16, BFLN16, BFFO17, BF19, BGL12, BST17b, BGT19, BM12b, BT12, BYH15, BL19, BW12, BDG13, BDT12, BCD^{+}17a, Bos19, BM19, BO16, CDL16, CCCdlL17, CF14, Cal15, CCMW19, CCM12, CWY14, CDS10, CSW15, Cer11, CL17, Cha14, CW16, CCK18, CDX12, CKM14, CPZ17, Che18, CS10b, Che12, Che14, CS14a, CS14b, CH19b, CMM13, CSZ18, CO12, CEIV17, CHS13a, CNSS17, DZ14, DF13, DdMH15, DO16, Des14, DF11, DWYZ12, DR13, DDMM18, DK14, DW13, DYZ19b, DLZ12b, DLZ12a, EF15, EJ14, EKR18, EI11, ET16, FRX19, FZZ18, FK19a, Feh13, FY13a, FK19b, FQS10, FY13b, FG18b, FL19, FHK13, GGAS14, Gar11, Ges13, GR15a, GT16, Ghi13, Gia15, GMT16, GY16a, GN19]. Equations [GKR18, GP18, GH12, GW15, GST13, GP11, GK10, HSS17, Hal12, Hal13, HS10a, Höf18, HKK13, HHPZ17, HW17, HWZ12, HLW12, HWWY13, HJ15, Hyn13, IM10, IS13, IISD15, IKM17, JM12, JJL10, JJL16, JWX13, JMWZ14, Joh13, JLL13, JT13, Jün10, KK17a, Kat19, Kia16, Kim09, Kim13, KK15, KZ18, KT11, Kry13, Kry14, KMVW14, Lai18, Lam12, LS13a, LS16, LMTT15, LM11, Lei13, Lei16, LR15b, LS18, LXZ13, LT17, LZ17, LMW17, LWY18, aLW18, Li19, LWY11, LPS18, LW15, LYZZ14, LRdS18, LT19, LY19, MPN14, MPR10, MJ14, MQS12, Mar18, Mar10, MM18b, MS13a, MOR^{+}16, MOZ10, MOZ12, Mel10, MP13, MX19, MP14, Muñ12, MP16, NPS13, NT19, NN19, NOS12, NP11, Oh15, Ohn15, PWG11, Per10, Per15, Pim16, PD17, PW15, Pol17, QW11, RSZ18, RZ11, SS19]. Equations [Sch10, SS17, SM19, Sho10, ST10, SW18, SSZ19, SCB17, Sri11, ST17, Sto19, ST15b, Sug16, Tak10, TYZZ13, TWW15, TZ18, TW10, TY11, VY16, VW11, VF13, VF15, VZ15, WZ13a, WXY15a, WZ13b, WZ17, XZL10, Xu11, XXK13, YFK11, YCW10, YZ14b, ZF12, ZZ14, Zho15, Zhu15, ZBL19, dlHHMV16, vBW11, vNVW12, vBM14, HR15, HY19, MT19, SCB20]. Equilibria [BL14, BH11, CCLM15, Gla17, RR13, SdlL13, Tay18]. Equilibrium [AET18, BCL11, BHM17, CS15b, DFT17, FPTT12, GHMZ10, HHMM18, HS19, JLZ18, KKT17, LMW17, MP16, NTW19, OW14, SW11a, ZT17, dlLSZ17]. Equivalence [CM13]. Equivalents [WU14a]. Equivariant [Sch18b]. Equivariante [LNZ14]. Ergodicity [BHR16, ESvR12, GT16, KSW13]. Ericksen [JL19, WZZ15]. Erosion [She15]. Erratum [BH18, DL13, HHR11, Li11, MOZ12, MT16, SCB20, VF15, Yam16]. Error [BFFO17, CT14]. Ersatz [Kry13]. Escape [CF11, HCHY16]. Estimate [DMZ14, Lee17, SM19]. Estimates [ACJ12, AGN19, BSW16, Bec18, BFFO17, BdHFS16, BCL18, BB10b, Bia18, BL19, BW12, BDG13, BDWZ12, CCK18, CT15, CNR17, CG10c, DNK12, DLVW13, DFV18, DX19, EW15a, GPT19, GS15b, GX17, HS10c, HNP15, HL15, HWZ12, Ign10, IN13, KKT17, LMTT15, LM14, LX17a, MS18a, MP12, NT13, Ovc11, RZ16, SV11, Tak10, VZ15, Ves15, ZBL19]. Estimating [BD18]. Estimation [BCD^{+}17a, GR13b, PT11]. Euclidean [Com17, KW12, MRT14, Sal12, ZCO16, dHHI^{+}14]. Euler [Aud12, BDX14, BFLN16, BPW15, BM18, CDX12, CPZ17, CKZ17, Che12, Che14, CS14a, CS14b, CHS13a, CNSS17, DWX18, DKT19, DLZ12b, FRX19, FKV15, FKM19, FL19, GMP13, GL15, GS19, HWWY13, Iye19, KMVW14, Lai18, LM11, LP16, NOS12, Ohn15, PWG11, Pen15, Per15, Pu13, SWZ15, Tak10, TY11, UWK12, WFL12, Wen14, XX10, Xu11, XXK13, ZH10, dlHHMV16]. Eulerian [CNSS17, KT18]. Evans [dRDR16]. Evaporation [LW14a, LS12a]. Even [GZ18, vdBMJLM11]. Eventual [GM15, Jia19]. Evolution [AH18, ABR17, ADHZ15, BFM12, BM18, CF14, CCM16, DP15, DL15a, DGVBW10, EHM16, GGAS14, GS17, IISD15, LX17c, Mar18, MC14, Sho10, Tak13, TW18b, VW11, YFK11, vNVW12]. Evolutionary [BFLS18, IM18]. Evolutions [AV19, EMZ17, MPS17]. Exact [BHND18, DS10b, FW18a, GH14, HS10b, dHGR14]. Examples [LL12]. Excitable [NW17]. Excited [BMY16, FKN^{+}14, GP11]. Exclusion [BDPS10]. Exemplar [ALM10]. Exemplar-Based [ALM10]. Existence [ABGS19, AF15, AS15, AM15, BEH15, BT16, BKLU18, BR17, BK13, BMC14, BFLS18, BWW14, BF19, Bes16, BDSS18, Bor19, BDLM19, BN14, BP14b, BPZ17, CGH10, CP12, CSW15, CH13, CQ19, CQX18, CHL17, Chu14, CG19, CMM13, CGS17, CY18, DDM11, DGV16, DS13, DR13, DG16, DLZ12a, DGVBW10, EW19, FT17, FKM^{+}16, FG18a, FG18b, FL15, GSW16, GM14, GG10, GZ18, GY16b, HSS17, HK15, HNP15, HKK17, HHPZ17, HW13a, HNP13, JS13b, JMZ18, KK17a, KMT13, KvMY19, Kry13, Le19, Leq11, LS10, LT17, LMW17, LW15, LX16, LT19, LY19, Lu13, LMMNR17, Mae17, MQS12, MSZ13, NPS13, PZ11, Pao15, PWG11, PT18, RT17, SM16, Smi17, Tha19, XX10, Xu18, YZ14a, YZZ10, YT11, YMYC10, ZF12, dAdM18]. Exit [INRZ10]. Exothermically [CKZ17]. Expanding [WX15]. Expansion [Con12, Yam13, Yam16]. Expansions [Ben17, BFG^{+}13, GH12, IKM17, Mor19, dHGR14]. Expectations [KLO16]. Explicit [JHN12, KKT17, Spe14]. Exponent [BO16, CLW12]. Exponential [BHM17, Dan17, DYZ19b, HL15, Jia19, LLLM14, MN12, Tak13, ZBL19]. Exponential-type [Dan17]. Exponentially [GPT19]. Exponents [BGL16, GH14]. Extended [CP13]. Extensible [HTW18, LS12b]. Extension [CCFdlL14, HZ10, RR13, SX13, ST17]. Exterior [BL19, CDL16, CEQW16, CQW18, DYZ19a, Sch14b]. External [Bos19, Cao19, GW13, HSV17, LZZ15]. Extinction [BS16a]. Extrapolated [KPR15]. Extremal [Lit13]. Extreme [CCC18b].

Factoring [SZ11]. Factorization

[CHL19, dRDR16]. Factorizations [LL16a]. Family [Sto19]. Far [CHL19, HSV16]. Far-Field [HSV16]. Farfield [KS14]. Fast [Bos12, CS15a, CL17, FY13b, GR15a, MB16, RTV17, VW11]. Fat [BGHP18]. Fat-Tailed [BGHP18]. Feedback [CW13, LW16a, MNS11]. FENE [HY13]. Fermi [FW18a, Miz11]. Fermionic [KK18]. Ferrofluids [NTW19]. Ferromagnetic [Car14, KN18, PWW17]. Feynman [SSZ19]. Fiber [BHM17, KSW13]. Fibrous [LS12a]. Field [ABØP19, AHP13b, BÖ19, Bos19, CHL19, CP12, CDM16, CL13b, CP19, CQW18, DWZ10, Due16, FG18a, FGJ11, GP15, GPI18, GLT10, HSV17, HD17, HK15, HSV16, IT15, Kac14, KVM18, KK10, Lee17, LLN19, Mel12, MS18b, Mit10, ZN19, DNWW19, LNZ14]. Fields [ABGS10, BB10a, BBG16, Bos12, Cao19, CEIV17, JJ18, Kre19, KM18, MM18a, Mon16, Pen17, VF13, VF15, Zho18]. Filled [MPS19]. Filling [CV16, DMZ19, WWW12]. Film [ABGS10, CPT10, CM12, CKV18, Ess16, FG18b, Gna15, LM17, Mel10, RZ16]. Films [FHO16, GZ14, KK16]. Filtered [BMY16]. Finite [BK13, BGL12, Bos16, CF14, CCLCP13, CDW13, CHL15, CMWY16, DF15, FG15, FV18, Ges13, GM10, Gia15, GR13a, GS17, GK10, KK18, Le19, LLW15, Mos18, RR13, SW11a, Sug16, Tro17]. Finite-Time [CDW13, Mos18]. Finitely [LS12b]. Finiteness [Wu16]. FIOs [AFK^{+}18]. First [ALZ19, BA10, BA12, BÖ19, Bos16, CH19b, CP19, IM10, Sch18a, TZ13, Val15]. First-Kind [CH19b]. First-Order [IM10, Val15]. Fisher [CHS13b, GM17a, BGHP18, CS10a, EM10, GM17b, RTV17, SV14]. Fitness [LTW14]. Fitness-Dependent [LTW14]. FitzHugh [CS15a, CKM14, CCHR18, SSH19, CFF19]. Fixed [LL16a, Mer18, Moo16]. Flame [XY14]. Flat [LS13a, MW17, Ngu15, TWW15]. Floating [NT13, Tre13]. Flocking [BHK^{+}19, CFRT10, KMT13, Pes15, PRT15]. Floquet [FGW13]. Flow [Abe12, BC19, BK18a, BP12b, BH11, Bre13, BP14b, CHY19, CMP13, CL13a, CY15, CKZ17, CL13b, Dai10, DD16, DKR16, DKN11, DNS12, DT14, DYZ19b, ER12, ESvR12, Evj13b, FW18b, FHK11, GVWK16, GM14, GMM13, GMT16, GS12b, HL12, Hel12, Hof12, HTW18, HW13b, HZ19, JMNR11, JK10, JMZ18, JRK19, KPS18, Lac15, LV10, LW16a, LMP11, LS13b, LS10, LWY18, LLW15, LW14b, Ono11, PSZ19b, PWW17, QWE19, RZ16, TW18b, WXY15a, WK17, XYZ16, YZZ10, YZZ12, ZH10, ZT17]. Flows [BP19, BFN^{+}13, BH17, BKP13, BFS14, BP14a, BGMŚG12, BC11, BM18, CM11, Can10a, Can10b, CP12, CDPS17, CPSW16, CRWX16, CDX12, CKY18, CDX17, CS14c, DWX18, Deu13, DZ15, Due16, DFHM14, FMP18, GM17a, GM10, GW13, GJZ15, HHR17, HW13a, HLX11, Iye19, KV19, Kwe12, KK17b, LZZ15, MM18a, MOS14, SW11a, Vis18, WX15, WY15, WX16, WX19, Wan19, Wen14, XX10, XY14, ZF10, BH18]. Fluctuations [BFK16, GKR18]. Fluid [ALST14, BST17a, BDEM18, Bla18a, BTZ15, BP12b, BP14a, BP14b, BPZ17, Cho16, CTW13, CWYZ16, Duc10, EW15b, EW18, FW18b, HX10, Hel12, JX15, KK17a, Kha13, Len14, Leq11, Liu18a, MPS19, MR15, Ohn15, RR13, SW11b, WLT16, WK17, WX13, XXK13]. Fluid-Boundary [Ohn15]. Fluid-Elasticity [BTZ15]. Fluid-Filled [MPS19]. Fluid-Particle [ALST14]. Fluid-Rigid [BST17a]. Fluid-Structure [BDEM18, Leq11]. Fluids [AZ17, Ara16, BR17, Ber12b, BGMŚG12, BBS16, DM15, FGN12, FKN^{+}14, FPZ14, GT10, JLZ18, JH18, Jün10, KR10, Kot12, LS12b, Nes14, NP11, Smi17, Wan11, WZ13b, ZF12, DLZ12b]. Fluttering [LW16a]. Flux [ACM^{+}12, BBT14, EHM16, GS19, LR11, MY12, NPS18, RZ14, Yos18]. Flux-Limited [ACM^{+}12]. Fock [GL19a]. Focusing [LWX11]. Fokker [ASC19, GHMZ10, JZ18, KKT17, LMW17, YY10, BR18, HJJ18, MT19, Zha14]. Folded [VW15b]. Following [GM14]. Föppl [MPT18]. Force [CH11a, GW13]. Forced [Bou13, CCCdlL17, CS14c, DN12, FFGHR17]. Forces [KT18, LZZ15]. Forcing [BBS16]. Form [DK11, DX19, Neu16, Kry10]. Formal [Mor19]. Formation [BHRW16, BLW19, CPZ17, CQX18, PLPSS18, WWW12]. Forms [JHN12]. Formula [Alm17, CD11]. Formulae [ABDD19]. Formulas [Hal14, SSZ19]. Formulation [MS18b, Per15, Sho10]. Formulations [CWH18, PD17, SS17]. Forward [BST17b, BGT19, LLW17, ST10, Ter11]. Forward-Backward [BST17b, BGT19, ST10, Ter11]. Foundations [ZCO16]. Four [BL11, BL14, WW10]. Four-Body [BL11, BL14]. Fourier [JLX15, AHKM15, AM15, BV13, Con12, JJN13, LJ17, PPP13, VW15a]. Fourth [BGT19, LN10]. Fourth-Order [BGT19]. FPU [HR10]. FPU-Type [HR10]. Fractal [AT10, AIK10, CV16, CWH18]. Fraction [CP10, CP11]. Fractional [ASC19, ACJ12, ALP15, AR19, BGAHS17, CG19, EW19, FY13b, GGAS14, Gra19, HL19a, IKM17, Joh13, LP19, LR11, LL18a, LL18b, LPS18, LRdS18, LP14, SS17, SV14, ST17, VZ15, Yam13, Yam16, ASS16]. Fractional-Diffusion [ASS16]. Fractionally [BM19]. Fragmentation [DDGW18, DF10, LvR15]. Frame [DSX17, WX12]. Framelets [HMZ15, SX13]. Frames [AG17, GL12, KLL12]. Framework [Bru16, GJMC12, ZF12]. Frank [PWW17]. Fréchet [BDEM18]. Free [AL10, BB10a, BM12b, BP12b, BCD17b, CCV15, Che19, CHS13a, Cui13, DSV15, DKT19, DL10, DL13, DMZ14, DLZ15, Duc10, ERV17, Evj13b, Han14, Hen10, KR10, KP18b, Lee17, Li09, Li11, LY19, LTW14, Nes14, WX19, Wan19, WNRJ13]. Free-Boundary [CHS13a, Lee17]. Freedericksz [CKY18]. Freely [NT13]. Frenkel [Fri19]. Frequencies [CStW17]. Frequency [AZ17, ALZ19, BSW16, CMM13, CdGDN18, DDGVM18, ETZ13, Jun14, KT13, LWZ15, TW11b]. Frequency-Dependent [CMM13]. Friction [FPVR13, Kim09, Kim13, MO15, Pao15, Pao16]. Frictional [HMS14, Pao16]. Friedrichs [CHW16]. Front [ABK12, BGHP18, GLY18, LW12, Tan15]. Fronts [BCN11, HS14, ILN11, HR15]. Frustrated [CFO19]. Full [CRWX16, CDX12, DWX18, FKV15, GM17b, HW17, JLL13, Lai18, LY19, NV12, PZ11, QW11, WZ17]. Fully [ADL14, DIT15, Kry13, RTZ17, Zho15]. Function [BBCD^{+}18, FGN13, GN19, Hal14, LV15, LS17, Mil18, Moo16, TD17, ZCO15, dRDR16]. Functional [AMW10, AP18, BM10, CP10, CP11, Gla17]. Functionalized [DHPW14, KP18b]. Functionals [BB10b, BOO18, CP13, CLW17, DLSV12, DDMM18, JS14, Olb19, PT11, Sch18a]. Functions [And12, BC17a, Bét16, CK11, CK13, CNR17, Cri19, DS10a, FK19a, KMS15, KLL12, Lit13, LA14, Mon16, Pla14, PX13, Rod16, Wei12, XYD18, ZCO16]. Fundamental [CS10a, Zha14]. Fungal [Ai10].

Gain [Jia12, MNS11]. Galaxies [SW17]. Galerkin [AM15, HS13]. Game [CDM16]. Games [FG18a, GP15, MPR10, MS18b]. Gamma [Ces11, GL19b]. Gamma-Limits [Ces11, GL19b]. Gap [ALZ19, DXZ18, LZ19]. Gaps [CGLS18, Nii12]. Gas [CHY19, CSW15, CQ12, Evj11, Evj13b, EW15b, FHK11, GL19a, HL12, HKT18, LM14, MLD19, MY17, Ohn14, SR14, SWZ15, ST18, YZZ10, YZZ12]. Gas-Kick [Evj11]. Gas-Liquid [Evj11, Evj13b, EW15b]. Gaseous [Liu18b]. Gases [AL10, BJ16, CH15, Lai18, LM11, Rey12]. Gauge [GS10a]. Gauss [CHW16]. Gaussian [BBT14, LSW17, dHGR14]. Gelation [LvR15]. Gelfand [LX19]. General [AG17, BG14, BYH15, BC17c, CHY19, CJN19, CHL17, DLSV12, DSX17, GR13a, Gro10, HS10a, Hof12, JMWZ14, NPS18, PT18, SS19, WXY15b, Wan19, Xu16]. Generalized [AHP13a, Ara16, ABCD^{+}18, BKK18, BNDHV10, CCMW19, CNR17, DO16, DL15a, FQ16, GW15, HS13, INSZ14, JZ10, Le 13, Len14, LL18a, LMMNR17, MN18, Muñ12, Nes14, RTY16, RTT19, WLT16, Win15, Wun10, Zhi19]. Generated [BK15a, ZN19]. Generating [HSV17]. Generator [BEH15]. Generic [CWYZ16, JT13, LZ17, Sch18b]. Gennes [HM13, HM12b, MN16, Nov18, WZZ15]. Genuine [Evj13b]. Geodesic [LMS16, Mon16]. Geodesics [PPPV16]. Geological [She15]. Geomagnetic [Kai10]. Geometric [BFLN16, Ben17, CDS10, DP15, JS14, Val15, WW18, HGW14, HZFQ13].

Geometry [Fri19, LS13a, Lóp12, Tak13]. Geostrophic [BM19, Che19, EW15a, Nov19, ZBL19, FT13]. Gevrey [BT19, LWX16]. Ghost [HT17]. Gibbs [Oh10]. Gilbert [FT17, KMM11, Mel10]. Ginzburg [ABGS10, ABGS19, CDW13, COS16, DWZ10, GS10a, GSV19, INSZ14, Kac14]. Given [BBCD^{+}18, WX19]. Gives [Tan15]. GKDV [CM18]. Glaciology [BG17, CGP13]. Global [BK13, Bla18a, BC14, BFS14, BN14, CCMW19, CCM12, CHY19, CSW15, CRWX16, CH13, CDX12, CY15, CMWZ18, CQX18, Cho16, Chu14, CGS17, CY18, DWYZ12, DZ15, EW15b, FZ16, FZZ18, FPZ14, GL17, GMP13, GZ18, GN15, HNS17, HX10, HD17, HM12b, HM13, HNP15, HKK17, HHPZ17, HWZ12, HW13a, HW13b, HZ19, HW14, HCHY16, IK11, IN13, JLZ18, JZ19, Jia19, Jün10, KT18, KK10, KNR12, Lau10, LM11, LPR12, LLP16, LS10, LWX11, LS12a, LXZ13, LZZ15, LMW17, LWY18, Li19, LLW15, LT11, Liu18a, LT19, Lu13, LMMNR17, MOZ10, MOZ12, MP13, MN12, NPS13, NTW19, PZ11, Pan12b, PZ13, PWG11, Pen15, Pim16, RZ14, SM16, ST11, Smi17, SSW14, Str10, TYZZ13, TW18a, VY16, WWW12, Wan12, WX15, Wan18, WZ13b, WZ17, Win15, WX13, WWX15, XX10, XY18, Xu11, XXK13, XZ15, YY10]. Global [YY18, YZ18, YZZ10, YT11, YCW10, Yun15, ZF12, ZZ14, Zhi19, ZLMZ18]. Global-in-Time [XY18]. Globally [BHND18, WFL12]. Gordon [Com17, DKS16, NS12]. Governed [FRX19, Kha13]. Gowdy [GL15]. Grade [Ber12b]. Gradient [AF15, BT16, BK15a, BBG12, CM11, CDPS17, CPSW16, DNS12, DX19, Due16, DFHM14, GM17a, Gin19, GS18, GR13b, HNP15, KV19, LLN19, MOS14, NN12]. Gradients [IS13]. Granular [AL10, DP14, Igb12, PW18, Rey12]. Graph [KVM18, Med14, TSA18]. Graphs [BCT19, DD18, KVM17, Koc16, Med14]. Gravitating [RR13]. Gravitational [LMR13]. Gravity [Hen10, HCHY16, MZZ12, Ngu16]. Gray [vdBMJLM11]. Grazing [FPTT12, HY14]. Greedy [BCD^{+}11]. Green [GN19]. Greene [CCFdlL14]. Grids [Mos18]. Grisvard [DT15]. Gromov [GM13]. Gross [BAC12, CPP18, DDF18, HTX15, ILR17, KMS17, Tri18, WY13]. Ground [Bec18, CCV15, CDNP16, CO12, GLW17, Kom15, KK18, LMR13, ZN19]. Growing [CCC18a, CCM16, Kry10]. Growth [ABR17, AP18, BEH15, BKK18, BO16, CL18, Cui13, DMZ19, DLV10, DDGW18, DLSV12, DH10, EG19, Gia15, Gin19, NSS17, OR19, WNRJ13, XY14]. Growth-Fragmentation [DDGW18]. Gurtin [EI11].

Hairline [IO16]. Half

[AHP13b, FR17, RSZ18]. Half-Plane [AHP13b]. Half-Space [FR17]. Hall [CW16]. Hamilton [BT12, CEH14, CCMW19, DZ14, DDMM18, Feh13, IS13, Igb17, MBPS13, MSTY16, Pin19]. Hamiltonian [BT12, KDT19, LNZ14, YZ14b]. Hamiltonians [BFK16]. Hankel [GR15b]. Haptotaxis [TW11a, TW15]. Hard [CH15, DXZ18, TAGP18]. Hardy [Kre19]. Harmonic [AC14, Gas19, GKR19, GM10, GMM13, Rod16, Spe14, BM12a]. Hartmann [XY18]. Hartree [GL19a, Lei13]. Hartree-Type [Lei13]. Hausdorff [GM13]. Head [BCS15]. Head-to-Tail [BCS15]. Heat [AH13, BPS19, BDT12, BPZ17, FK13, GIP^{+}13, HI19, HCHY16, KVM17, KPR15, KLS11, LV10, LS12a, Li19, Med14, MPZ15, NP11, EH16, Sou19, WZ13b]. Heat-Conducting [MPZ15, WZ13b]. Hele [Ono11, TW18b, YT11]. Helfrich [BM10]. Hellinger [KV19, LMS16]. Helmholtz [BSW16, BdHFS16, CDL16, CH19b, HPS19, Mel12, MMP13, Sch14b]. Hemivariational [BS16b, HMS14, Sof18]. Herglotz [BBR19]. Heteroclinic [HR10]. Heterogeneities [PT18]. Heterogeneous [AH13, ACZ14, BCN11, Can10a, Can10b, KK16, PLPSS18, TD17]. Hierarchical [MLD19]. Hierarchy [HGW14, HTX15, HZFQ13, Thi19]. High [AL10, ALZ19, BSW16, Beb16, Bel10, Bel17, BPP15, CCC18b, DDGVM18, DWZ10, ETZ13, FWW17, HV13, HR12, Jun14, LWZ15, PSZ19a]. High- [DWZ10]. High-Contrast [Beb16, BPP15, CCC18b, PSZ19a]. High-Energy [AL10]. High-Field [DWZ10]. High-Frequency [ALZ19, BSW16, LWZ15]. High-Performance [HR12]. High-Rank [FWW17]. Higher [BFFO17, CCHR18, GVZ16, GL19b, Hal13, HLR^{+}19, Pu13, RZZ18]. Higher-Order [BFFO17, GVZ16]. Highly [AVP16, BCN11, CGM16, CP19, GS15a, GSV19]. Hilbert [AAK14, ADK15, BN14, HI12]. Hilliard [EG19, ABK12, BBMN12, CGS17, DG11, Del18, DHPW14, GMT19, KNR12, KP18b, OW14, OSW19, SP13]. Hinged [SV18]. History [MSZ19]. History-Dependent [MSZ19]. Hitting [IT15]. Hitting-Times [IT15]. Hodge [CH19b]. Hohenberg [MBK13]. Hölder [KS19, MS11, Sto19, WZ16]. Holes [BCG10, CDD^{+}12]. Holm [CFGL17, DIT15, GL17, GLL18, LZ17, Tan18, dMIS10]. Holomorphy [CSZ18, CStW17]. Homoclinics [vdBMJLM11]. Homogeneous [CHN18, CMWY16, HL19b, LX19, MM18b, Sch17, Sch18b, SY14]. Homogenization [Ale16, AMP10, AH13, ALZ19, BAP13, BLZ16, BCN11, BFFO17, Bel17, BCG10, BPP15, CGM16, CDM16, DDMM18, DG16, Feh13, FF12, GS15a, GS15b, GX17, GS18, HW11, IM10, JMNR11, JS14, KPZ19, KK16, MM11, NSS17, PSZ19a, PZ17, Pra13, Sch10, Sen17, Sus13, Xu16]. Homotopy [Kol11]. Honeycomb [Lee16]. Hopf [Naz12]. Horizon [ABL13, AHØP13]. Hörmander’s [Kry14]. Hotspot [BWW14]. Human [GP14]. Hunter [LPS10, Wun10]. Hybrid [YMYC10]. Hydrodynamic [CDM13, CH19a, GW13, HKK15, HMWY11, HMW11, HMWY12, HCHY16, JXZ16, LMZZ17, LMZZ18, WX13]. Hydrodynamics [FW18b, JLX15, WX11]. Hydrogen [CS18]. Hydrology [BG17]. Hydrostatic [KMVW14]. Hyperbolic [BIP16, Ben17, BHWY12, CHL15, CT11, CN15, Cui13, Daf14, DNK12, DS13, Ell12, FHK13, GPPP13, GYY18, GN15, HK15, HLWW18, JL19, LT11, Lu13, LNZ14, Mei10, NP16, Ohn16, Ves15, Yao19]. Hyperbolic-Elliptic [Ohn16]. Hyperbolic-hyperbolic [LNZ14]. Hyperbolic-Parabolic [DS13, GYY18, HK15]. Hypergeometric [GI15]. Hypersingular [HSV17]. Hypersurfaces [EF15]. Hypocoercivity [DJMZ16, JZ18]. Hypoelliptic [Sto19]. Hysteresis [GST13].

ICE [BG17]. ICE-Hydrology [BG17].

Ideal

[DYZ19b, GS12b, JJL16, Lac15, LTW14]. Identical [ST11]. Identifiability [Liu18a]. Identification [HSV16, Kat19, MRV19]. Identities [Lei10]. IDSA [BFG^{+}13]. II [LMZZ18, AHP13b, AFK^{+}18, Ben17, BG14, Can10b, CP11, DFP14b, GS17, HNSW11, Lam12, LW16b, MY17, Pao16, SWX17, dR18]. Ill [HHK18]. Ill-Posedness [HHK18]. Image [CM11, CM13, DSX17, SSST15]. Imaging [AFK^{+}18, AGS13, CS17, FGN13, HMN14, MNT13]. Imbeddings [CC11]. Immersed [SR14]. Immersion [CHW16]. Immiscible [CP12]. Impact [AP15, Pao15, Pao16]. Impedance [BOS17, HHR09, HHR11, HS10b, HU13, HMN14, KT13, Sin10]. Impenetrable [KvMY19]. Imperfectly [KLO10]. Implicit [AH18, Li09, Li11, CMP13]. Implicitly [BGMŚG12]. Imply [Mur14]. Implying [LS18]. Improved [Che14, KW11, Kre19]. Inclined [RZ16]. Including [CDR17, Gli13]. Inclusion [ADMR14, DLVW13, MRV12, MRV19, MP14]. Inclusions [Alm17, BLZ16, BFV17, KRW15, LX17a, MNT13]. Incoming [JZ19]. Incompatibility [AV16]. Incompressible [BFN^{+}13, BBG16, BKP13, BGMŚG12, BP14b, BPZ17, BC17c, CWY14, CDM13, Che14, CDX17, CS14b, CY18, DKN11, Deu13, DKT19, FKN^{+}14, INRZ10, JJL10, JJL16, JL19, JMWZ14, Kha13, Lac15, Len14, MM18a, SW11a, WK17, XY14, XZ15, YZZ12, ZF12, ZZ14, Abe12]. Increasing [BA10, BA12, ILW16]. Independent [Kar12, Ngu17, RT17, Rou10]. Index [BC11, LS16]. Individual [BM15a]. Induced [AHP13b, BG14, CM12]. Inelastic [BCL11, FPTT12]. Inequalities [AH18, BS16b, BCD17b, DNS12, FG18a, GH14, HMS14, Har18, HS16, Sal12, Sof18, ZBL19]. Inequality [BM15a, BMY16, DSZ19, Fri18, Kre19, Neu16, SV11, Wal14]. Inertialess [Höf18]. Inf [BCDG16]. Inf-Sup [BCDG16]. Infeld [CP13, MP14]. Infinite [ABL13, AHØP13, CGH10, CCM14, CCLCP13, KLW17, LMR15, Mos18, PSSW15, Sen17, VW11, WWW12, WX15, YZ18]. Infinite-Time [VW11, WWW12]. Infinitely [BP19, CDX12, KY15, Pas13, XX10]. Infinity [Deu13, FW18a, FY13a, LWZ15, WZ17, vBW11]. Inflow [CJN19, FRX19, HW17, QW11]. Information [GM17b]. Inhibitory [RW14]. Inhomogeneities [KMS17]. Inhomogeneous [ADL14, BK13, BC17c, CY18, MS16, YY18].

Initial

[AI12, DWYZ12, FZ16, FZZ18, FT13, FY13b, GYY18, HS10a, HKT18, HW17, JS13a, JX15, KMV18, KT11, Len16, LT17, LX19, MZ13, NP16, Oh15, TW18a, Tsu12, Ves15, dCPS16]. Initial-Boundary [AI12, Len16, TW18a]. Initial-Layers [GYY18]. Initial-Value [KMV18]. Injection [Ono11]. Injectivity [FGR15]. Inpainting [ALM10]. Inside-Outside [LR15a]. Instabilities [BAH17, ST15b]. Instability [Ben17, BGM19, Bou13, BC11, FHO16, GP19, GT10, HKN16, JZ10, Le19, ZZ19]. Instantaneous [TW11b]. Insulating [MNT13]. Integrable [BBCD^{+}18]. Integral [BA10, BA12, BSW16, BK18b, CWH18, Chu14, CH19b, JHN12, Lei10, Mel12]. Integral-Representation [BK18b]. Integrals [BHND18, DLV10, MP12]. Integration [KY15, RZZ18, VK18]. Integro [BGHP18, Gar11, GS18, Sch10]. Integro-Differential [Gar11, GS18, Sch10, BGHP18]. Integrodifferential [Igb12]. Interacting [DGV16, IT15, Len14]. Interaction [ALST14, BST17a, BTZ15, CCV15, CS10b, CT16, DKR16, Due16, Evj13a, GLZ17, HGW14, HY14, JMNR11, KT18, KK10, MPS17, Ohn15, vBM14]. Interaction-Driven [MPS17]. Interactions [AMV15, BÖ19, BK18b, BCQ12, BLW19, CS18, CDNP16, DSY18, DF11, FPZ14, FMP18, Lai18, Nii12, Sab13, WX13]. Interactive [DDGVM18]. Interconversion [LA14]. Intercritical [Mur14]. Interface [Abe12, CCHR18, CP10, CP11, Evj13b, Ono11]. Interfaces [BM10, CC11, DP15, DHPW14, Gli13, KS14]. Interferences [CdGDN18]. Interior [CCG10, CCH10, CH11b, DKT19, Fai14, KK17b, LV12, LV15, LR15a, MNT13, WZ13a, Zho15]. Intermittency [CS14a]. Internal [CT15, Duc10, DIT15, JTW16].

Interpolation [BOS17, GH10, Har18]. Interpolations [ZCO16]. Interpolatory [Wei12]. Interpretation [Sim16]. Interval [JHN12]. Intrinsic [AV16]. Intrusion [CDR17]. Invariance [BM15a, Nol11, Oh10]. Invariant [GPPP13, GL19a, GW15, NS12]. Invasion [HS14, SSW14, dAdM18]. Inverse [ACZ14, BJ10, BM12a, BFGPE^{+}18, BdHQ13, BdHFS16, BK15b, CHL19, CLM17, CHN18, CT15, CS17, DLVW13, GIP^{+}13, GY16a, HHR09, HHR11, HSV16, IY12, IN13, Kar16, Kat19, KLO10, LL16a, LTV17, LL19, LT11, Poh15, SZ12a, TSA18, Ves15, ZZ19]. Inverses [JHN12]. Inversion [ABDD19, Hal14, HL19a, Mon16]. Investigation [SCB17, SCB20]. Inviscid [BP19, BFN^{+}13, CV15, CEIV17, CIN18, CTW13, EW15a, ER12, Kim09, Kim13, LW16a, NN19, Nov19, VW15a]. Involving [KRW15, Xu18]. Ionic [Dua11]. Irregular [DK14, GS12b, LV10]. Ischemic [FHX10]. Isentropic [FZZ18, FKV15]. Ising [AG16]. Ising-Type [AG16]. Isolated [KN18]. Isometric [CHW16]. Isometry [KW11]. Isomorphism [MM18a]. Isoperimetric [BC14, GO18, KL18b, MS14]. Isothermal [CQ19, PW15, Kot12]. Isotropic [BM12a, CY18, GZ18, MRV19, BFG^{+}13]. Issues [Gas19, MX19].

Jacobi [IS13, BT12, CEH14, CCMW19, DZ14, DDMM18, Feh13, Igb17, MBPS13, MSTY16, Pin19]. Jacobians [KRW15]. Jensen [Wal14]. Jet [WX19]. Jin [Bia18]. JKO [GM17a]. Johnson [KW11]. Jost [Wu17]. Jouguet [GSW16]. Jump [BHR16, KK17b, MP12, Val15]. Junction [AMW11]. Justification [BFLN16, CCFdlL14, DP14, LXY19, MT13]. Justified [DIT15].

Kac [ADHZ15, FPTT12, SSZ19, YY18]. Kadomtsev [EL17, ILP16, JZ10, LPS18]. KAM [PX13, SdlL13, dlLSZ17].

Kantorovich [FV18, Fri19, GM17a, KV19, LMS16]. Kármán [MPT18]. Kawasaki [Kut15, CSZ19, Le10, vdBW19]. KdV [CR10, DP14, ETZ13, Joh13, Muñ12, Tsu12]. KdV-Type [Joh13]. KdV/KP [CR10]. KdV/KP-I [CR10]. Keller [BK13, BH17, BH18, BCJ20, KY12b, CLW12, HJ11, ZLMZ18]. Kelvin [DAP19, PZ13, RR17]. Kernel [FW18b, HNW10, HNSW11, Löb18, PZ17]. Kernels [BGHP18, FK13, LW16b]. Kick [Evj11]. Kind [CH19b]. Kinetic [ASS16, BCL18, CFRT10, CS15b, CY19, CHS13b, FKM^{+}16, FL12b, JM12, JXZ16, KMT13, MT13, NT18, Ovc11, Per15, PRT15, RR13, Sto19, WDL18, Zha14]. Kinetics [MO15, MS13b]. Kirchhoff [Ghi13, MRV19]. Kirkwood [CS18]. Klein [Com17, DKS16, NS12]. Knocking [CHKP19]. Knothe [Bon13, CGS10]. Known [HMN14, KLO10]. Koch [DW13]. Kohn [SCB20, SCB17]. Koiter [CS10b, Len14]. Kolmogorov [AGN19, FL19]. Kompaneets [BIP16, LLP16]. Kontorova [Fri19]. Korn [BCD17b, Fri18, GH14, Har18]. Korn-Type [GH14]. Korteweg [Aud12, BC17c, CH13, CG10a, DM15, EW19, Kot12, LP19]. Korteweg-Type [DM15]. KP-I [CR10]. KPP [BGHP18, CQ19, CHS13b, FY13b, RTV17, SV14]. KPP-Fisher [CHS13b]. Kramers [ET16, MZ18]. Kullback [PSSW15]. Kuramoto [KVM18, Tro17]. Kuznetsov [RV12].

Lagrangian [BC17c, CNSS17, DKT19, FT13, Gia15, Hyn19, MY17, MSTY16]. Lake [Dek19]. Landau [HM13, ABGS10, ABGS19, CDW13, COS16, DWZ10, FT17, GS10a, GSV19, GZ18, HM12b, INSZ14, Kac14, KMM11, LX19, Mel10, MP13, MN16, Nov18, PWW17, Wan12, WZZ15, dL14]. Landau-Type [GSV19]. Lane [BHRW16].

Langevin [MN18]. Language [Mit10]. Laplace [Ash13, BNDHV10, GT16, TZ13, Xu18]. Laplacian [BHL18, BM15b, BGL16, MRS16, PR13, Rod16, SZ12a, ST19, Yam13, Yam16]. Laplacians [Kal12]. Laptev [Kre19]. Large [AHP13b, BYZ12, BACP18, BM12b, Ces11, Cho16, DF10, DWYZ12, FZ16, FZZ18, FR17, FMP18, GO18, HHPZ17, HW17, HMW11, IS13, Kim09, Koc16, Kut15, Lam12, LXZ13, LZZ15, LMW17, LWY11, LYZZ14, PZZ19, QW11, RZ14, RZ16, SWX17, TYZZ13, TW15, WZ13b, WZ17, Whe13, Win15, Kim13]. Large-Amplitude [RZ16, Whe13]. Large-Data [Win15]. Large-Scale [FR17]. Large-Time [Cho16, QW11]. Largest [NNS18]. Larmor [Bos16]. Lattice [BGAHS17, BFK16, GZ13, HMSZ13, HO15, HV13, Lee16]. Lattices [Bét16, Miz11, Qin15]. Law [AP14, BST17a, Chu14, CW13, EKR18, GZ14, LV15, Le10, MY12, SS19, Yos17, Yos18, JM12]. Lawrence [Pen17]. Laws [AGN19, CT11, CR18, CD11, Daf13, Daf14, DWY12, GS19, Jun14, KPS18, KMS15, MS13a, NPS18, PT18, VK18, WDL18, XYY19, ZLMZ18]. Lax [CHW16, LW16b]. Lay [BHM17, KSW13]. Lay-Down [BHM17]. Layer [DDGVM18, GVWK16, HL15, JX15, JT13, Kal12, KL18a, LXY19, Pra13, WW12, XY18]. Layered [Pen17]. Layers [CV16, EEW11, GYY18, GN19, HL11, HLWW18, Iye19, LW14b, Ngu10, NOS12, Ohn15, She15]. LCD [Dai17]. Leading [AF16]. Learning [ST19]. Least [Moa11]. Lebesgue [HNW10]. Leffler [TAGP18]. Legendre [GIV17]. Leibler [PSSW15]. Lemma [FL15, Naz12]. Length [BLS15, CLLS17, Nov18, TZ13]. Leray [GP18]. Leslie [JL19, WZZ15]. Less [AC14]. Level [Bec18, CCMW19, CM13, GMT16, Lau18]. Level-Set [GMT16]. Lévy [FPTT12, KPZ19, ZBL19]. Lévy-type [FPTT12, KPZ19, ZBL19]. Life [HI12]. Lifetime [CTW13]. Lifshitz [FT17, KMM11, Mel10, MP13, PWW17, dL14]. Lifting [CKV18]. Ligand [AET18, ERV17]. Ligand-Receptor [AET18]. Lighting [Kar16]. Like [AT14, BMR14, BGL16, Muñ12]. Limit [ASS16, ASC19, AHP13b, AH16, BM12b, BCT19, BGHP18, BP14a, CEH14, CDM16, CH13, CDM13, CCHR18, CR10, CV15, CP10, CP11, CG10a, CEIV17, CIN18, CDK11, CHS13a, Daf19, DD16, DM15, FGN12, FKM19, FJ18, FPTT12, Gin19, GYY18, GPI18, HY14, Höf18, HY19, HLW12, HWWY13, IT15, JMNR11, JJL10, JX15, JLX15, JJL16, JZ19, KVM18, Kim09, Kim13, LWZ16, LST12, LP16, LLN19, MM17, MB16, NN19, Nov18, Olb19, PWG11, PR13, Pu13, RZ11, SZ12b, SS19, Str10, Tha19, VW15a, Wan11, WX11, WW12, WXY15b, YZZ12, ZZ19]. Limitations [Mer18]. Limited [ACM^{+}12, FQ16, GLL17, MS13b, Ngu15]. Limiting [Kal12, KL18a, Kut15, Lam12]. Limits [ABGS10, AG16, BAP13, BCS15, BK18b, Ces11, Che12, Due16, EG19, FKN^{+}14, GT16, Ghi13, HKN16, JS14, JXZ16, JLL13, Med14, Mel10, Ped15, TSA18, GL19b]. Lindenstrauss [KW11]. Line [DL18, GHH17, OSW19, SZ12b, WW10, ZT17]. Line-Tension [SZ12b]. Linear [AHKM15, AKKY17, AS13, AZ12, BKK18, BAP13, BFFO17, BCG10, Bos16, Dan17, DFP14a, DFP14b, DNK12, DKR15, DYZ19a, EW15a, FHK13, GP19, Gia15, GT10, GK10, Han14, IKS12, KM13, LRdS18, MO19, NKV19, PZZ19, RZ17, YFK11, Yao19, Zha10, Zha19, DS13]. Linearization [BTZ15, DP13, GS17, HS10b, JS14]. Linearized [Che18, CGLS18, DJMZ16, DKN11, FPP19, FI14, FK18, TY11, Wu14b]. Linearly [Len14, MY12]. Lines [Bos19, CM13, SWZ15]. Linkages [MO15].

Lions [Löb18]. Liouville [FR17]. Lipid [HD17]. Lipschitz [AC14, BPS16, BdHQ13, BdHFS16, CKZ17, CHT18, GS10b, HL19b, KVM18, LTV17, WZ16, Xu16]. Liquid [CRWX16, CKY18, CMWZ18, CHT18, DS14, DW13, Evj11, Evj13b, EW15b, FL15, GCGJL18, GSV19, GW13, HL12, HM12b, HM13, HW13b, INSZ14, JL19, LLW15, SS15, SW11a, Tay18, Wol19, YZZ10, YZZ12]. Liquid-Gas [HL12, YZZ10, YZZ12]. Liquids [BOS11]. Live [BFGPE^{+}12]. Lizorkin [Tak10]. Load [MO15]. Load-Dependent [MO15]. Loads [BFGPE^{+}12, BD18]. Local [AG16, AS15, Bev11, BC14, DLZ12a, Faj16, FQ11, GT16, GM10, GLW17, HNS17, Hof12, HR12, JTW16, KY12a, KMVW14, LY19, Pin19, ST18, Tsu12, WXY15a, ZH10, ZT17, Zho18, dAdM18, dHHI^{+}14, Ngu16]. Locality [BBO19]. Localization [CDN10, CCC18b, NT13, PPP13, TD17]. Localized [AKKY17, BCT19, DVW15, LMTT15, LL16b, Ngu17, PY14]. Locally [AVP16]. log [DP14]. Logarithmic [CCV15, DNS12, GP15, JHN12]. Logistic [DL10, DL13, DM14, TW11a]. Long [ADHZ15, BP19, BF19, BM18, CDX12, CPT10, Daf13, DDF18, DF11, FPZ14, FMP18, FJ18, GW13, HMWY12, IM18, LP16, LS17, MM17, MM18b, MZZ12, MSZ13, ST10, XX10, YZ14b]. Long-Range [FJ18]. Long-term [ADHZ15]. Long-Time [GW13, MM17, MSZ13, ST10, HMWY12]. Long-Wave [CPT10, MZZ12]. Longtime [CTW17]. Loss [Ria10, UWK12]. Love [MRV19]. Low [Beb16, DD16, FKM19, Gie14, HTX15, JRK19, LST12, LWX11, Liu18a, MMB11]. Low-Rank [Beb16]. Low-Temperature [MMB11]. Lower [AV19, AKKL17, BMC18, BPW15, DLV10, MOR^{+}16, Wal14]. LQ [ABL13]. LQ-Problem [ABL13]. Lucquin [ST18]. Lucquin-Desreux [ST18].

Ma [LL12]. Mach [DD16, FKM19, LST12]. Macro [JLZ18]. Macroscopic [CH15]. Magnetic [ABGS10, AHP13b, Bos12, Bos19, CWY14, DZ15, JJ18, Kre19, Lee17, Pen17, PR13, Zho18]. Magnetically [CCM14]. Magnetoelasticity [BFLS18]. Magnetohydrodynamic [DZ15, FMP18, GLT10, HHPZ17, JJL10, JJL16, JLL13, LXZ13, Zhu15]. Magnetohydrodynamical [XZ15]. Magnetohydrodynamics [Bou13, CWY14, CW16, Lee17, WZ13a]. Magnetostatic [CP13]. Main [GVZ16]. Malik [CG11, KY15]. Malliavin [MZ13]. Management [CHL17]. Manakov [BK15b]. Manev [LMR13]. Manifold [GPPP13, TW10, VF13, VF15]. Manifolds [BMMP16, BMP18, BDLM19, FGR15, HNW10, HNSW11, KMS15, Lau10, NS12, NRS17, Wei12]. Many [KY15, Pas13]. Map [Ash13, BFV17, Bon13, CGS10, HPS12]. Mapping [Mel12]. Mappings [MS11]. Maps [BB17, BK15a, Gas19, GKR19, RSS17, ZCO15]. Marcinkiewicz [GP18]. Marginal [FV18]. Marginals [Pas13]. Mass [AKKL17, ABBK16, Bor19, EHM16, LWX11, Mar10, Pan12b, RP18, SM16, YZ18, Zhi19]. Mass-Action [Bor19, Pan12b]. Mass-Conserving [ABBK16]. Mass-Critical [LWX11, YZ18]. Master [Bes16, ST17]. Matched [Kal12]. Matching [BOS17, MRT14]. Material [Bev11, DX19, LS16]. Materials [CC11]. Mathematical [BDEM18, BMSR^{+}13, BFTT18, BDFS18, CDR17, ERV17, FHX10, LZ19]. Mather [CGT11]. Matrices [MNS11]. Matrix [DXZ18, HZ10, LW16b, SZ11]. Matter [Igb12]. Maxima [AMW10]. Maximal [Jun14, ZBL19, vNVW12]. Maximizing [BBG17]. Maximum [AHØP13, BL19, CCK18, KY12b]. Maxwell [Dua11, LYZ16, YY10, AC14, BPS16, BAH17, DLZ12b, HKN16, IK11, IKS12, JS13b, KLS15, LS13a, LS16, NT19, PWG11, PD17, Tha19, UWK12, WFL12, Xu11, XXK13, Zha19]. Maxwellian [KY12a, Yun15]. Mean [ABØP19, BC19, BK18a, BÖ19, CDM16, CMP13, CNR17, Dai10, Due16, FG18a, GMT16, GP15, GPI18, GLT10, Hel12, HHR17, IT15, KVM18, KK10, LLN19, MPR10, MS18b, Mit10]. Mean-Field [ABØP19, Due16, FG18a, GP15, GLT10, IT15, Mit10]. Meander [DHPW14]. Means [Hal14]. Measurable [Kry14]. Measure [ABCL18, BGT19, BDG13, CCG18, DDGW18, DT14, EHM16, Mit10, NP16, Oh10, Rod16]. Measure-Valued [BGT19, CCG18, EHM16, Mit10]. Measurement [MNT13]. Measurements [AHKM15, ADMR14, AGS13, CKS15, HPS12, KS14, KNW15, KLS11, LV13, NKV19]. Measures [BKK18, BMC18, BK15a, BBV14, CGT11, CMWY16, Ito18, KRW15, LS13c, PSSW15, Rou13, SSST15, XV10]. Mechanics [BS16b, JMNR11]. Mechanism [BLZ16, GCGJL18]. Mechanisms [AS13, Cal15]. Media [AMP10, Alm17, ACM^{+}12, AT14, BCN11, BC17a, CGM16, Can10a, Can10b, DKR16, DL15a, DT14, Ges13, HKK17, IKS12, JRK19, LV12, LS10, LS12a, MMB11, NW17, QWE19, SdlL13, dlLSZ17]. Medial [ZCO15]. Mediating [MO15]. Medium [ACZ14, AZ17, BFGPE^{+}18, BDG13, CT15, CQW18, DGV16, FG15, HSV16, LL16a]. Meets [PR13]. Membrane [TZ15]. Membranes [RP18]. Memory [ABØP19, BDT12, PW18, dCPS16]. MEMS [GS15c]. Mesh [Kar12]. Meshless [BB10a]. Meta [LS16]. Meta-Material [LS16]. Metal [LZ19, PUW18]. Metamaterials [CC11, CMM13]. Metastability [Car14, MS13a]. Metastable [GM17b]. Method [AM15, AVP16, BS16b, BD18, BRS17, CHL19, CS18, CGS10, CPSW16, CV15, CDD^{+}12, DDGW18, FK19a, FW18b, GL17, Kal12, Kar12, Kol11, KLS15, LP19, Len16, NKV19, PY10, VW15a]. Methods [BCD^{+}11, BC17c, GH10, HPS19, Ito18, Lau18, VZ15, YMYC10, ZCO16]. Metric [BFY15, FK13, Igb17, KP13, MOS14, dHHI^{+}14]. Metrics [GM13]. MFG [CG19]. MHD [DYZ19b, FFGHR17, LXY19, TW18a, WWX15, XY18]. Micro [JLZ18]. Micro-Macro [JLZ18]. Microelectromechanical [CG10c]. Microenvironment [EW18]. Microlocal [FQ11, FGN13, KM18]. Micromagnetics [DSZ19, PY10]. Micromotions [Kha13]. Micropatterns [BW17]. Micropolar [BR17, BP12b]. Microscopic [LR13]. Microstructure [Bev11]. Microstructures [COS16]. Microswimmers [DDM11]. Midrange [BCQ12]. Migration [EW18]. Mild [KPR15, KvMY19]. Mildly [Ghi13]. Mindlin [BL15]. Mineral [HKK17]. Minimal [AP15, Ito18, JS13a, Mar10, Pla14, dL14]. Minimality [BC14, CJP13, ZK15]. Minimization [AS14, BHL18, BLS15, DS19, De 18, Li09, Li11]. Minimizer [Bev11]. Minimizers [DLSV12, FL15, GO18, HM12b, HM13, KL18b, LS13c, MS14]. Minimizing [BK18a, YZ16]. Minimum [CQ19]. Minnaert [AZ17]. Miscible [DT14]. Mittag [TAGP18]. Mittag-Leffler [TAGP18]. Mixed [BPS16, BCdSN18, BLW19, CDN10, CNR17, Gin19, HK15, Sho10, WW12]. Mixing [RZ17]. Mixture [Evj13b, JMZ18, PSZ19b]. Mixtures [BPZ17, Gla17, GM17c, JS13b, MPZ15]. mKdV [KM17]. Mobility [Del18, Fis13]. Mode [YY14]. Model [Abe12, ASS16, AAGP18, AKKL17, Ai10, ABGS19, AF17, ALST14, Aud12, BFM12, BIP16, BFLS18, BWW14, Ber17, BFTT18, Bev11, Bia18, BCQ12, BH11, BP14b, BPZ17, BDFS18, CCC18a, Can10b, CFRT10, CK12, CP13, CDM13, CTW17, CY19, Cho16, CH19a, CDR17, CDNP16, CCM16, COS16, CP19, CSZ19, CHS13b, CWYZ16, CS14c, DDGVM18, DD16, DF10, DS13, DSX17, DT14, DL10, DWZ10, DL13, DIT15, DKS16, EG19, ERV17, Evj11, Evj13a, EW15b, EW18, FZ16, Fel18, FKM^{+}16, Fis13, FGJ11, FL15, FHMP16, FHK11, FHX10, Fri19, FL12b, FPTT12, GV19, GLL17, GM14, Gie14, Gla17, Gli13, GRT14, GW13, GN15, GW18, GGRB14, HL12, HD17, Hel12, HJ11, HMWY11, HMWY12, HO15, HNP13, HT18, Igb12, INSZ14, JMNR11, JLX15, JLZ18, JL19, KVM18, KK17a]. Model [KY12b, Kom15, KK18, KT11, Kut15, Lai14, LMP11, LLP16, LR13, LMZZ17, LMZZ18, LS12b, LX16, LX17c, Liu18b, LZ18, LMMNR17, MO15, NT14, Nov18, Ohn14, Otw10, OR19, Pes15, PRT15, Sab13, SV11, SZ12b, ST18, TW11a, TW15, Tro17, WX11, WWW12, WW12, WLT16, Xu18, YZZ10, YZZ12, Yun15, ZLMZ18, dCPS16, dAdM18, BCJ20]. Modeled [Höf18]. Modeling [ATSR19, BMY16, CCM12, CMM13, Cui13, DMZ19, DH10, DGVBW10, FMP18, GR13a, KSW13, MBPS13, PPPV16, SSW14, WNRJ13, WX13]. Models [BS16a, BK13, BÖ19, BHR16, CNS10, CL13a, Chu14, CG10c, DFP14a, DFP14b, Duc10, GLS10, GZ14, GJMC12, HS16, HMSZ13, HR19, HY13, HMW11, JXZ16, KMT13, LWZ18, Lóp12, LTW14, MLD19, Mit10, PLPSS18, RR13, SP13, Wen14]. Moderate [MV19]. Modes [Bar14, BCO17, BR11, DVW15, HW11]. Modified [GL17, GLL18, GHLN13, Liu19]. Modulated [BCS16]. Modulation [MBK13]. Modulus [CCK18]. Moist [CCC^{+}17]. Molecular [Li09, Li11]. Moments [CL18, CMWY16, Pal14, Sal12, TAGP18]. Monge [DF13, DL18, Fri19, LR17, Pas11]. Mono [MO19]. Mono-Stable [MO19].

Monochromatic [IN13]. Monoid [Sch18b]. Monostable [MOZ10, MOZ12, HR15]. Monotone [BA10, BA12, FZ14, Feh13, LA14, Vis18, FWW17]. Monotonicity [HU13, HPS19, HL19a, KMVW14, LL16a]. Monotonicity-Based [HU13, HL19a]. Morphogens [CCM16, WW10]. Morrey [GP18]. Morrow [ST18]. Mossotti [Alm17]. Motion [ABK12, ABBK16, CM13, DDM11, Fel18, GGAS14, GS12b, KMM11, KMS17, Le10, LZZ15, MPS19, RZ14, SR14]. Motions [BR17, Bes12, YZ16]. Movements [BK18a, BFV17]. Moving [AF16, AET18, BMSR^{+}13, BHM17, CS15b, CS10b, EF15, FKN^{+}14, HKK15, Iye19, LMP11, Liu18a]. Muckenhoupt [AR19]. Mullins [Le10]. Multi [FV18, MRT15]. Multi-Marginal [FV18]. Multi-solitons [MRT15]. Multibody [BK18b]. Multibubble [CM18]. Multicomponent [JS13b, LS10, MPZ15]. Multidimensional [BEH15, Ber17, BGL12, BGLV16, BGN14, CP12, CLW12, Don11, FKV15, FMP18, HL12, LWY18, LX16, PP19, TW15, ZH10]. Multilane [HR19]. Multilayered [Gla17]. Multimarginal [CDS19, Fri19, GKR19, Pas11]. Multimaterial [MMT19]. Multiphase [CDS10]. Multiple [DN12, EEW11, GIV17, JZ18, KMVW14, LX17c]. Multiplicative [Cer11, Tan18]. Multiresolution [GI15]. Multiscale [FF12, FS14, SSW14, Wei12, ZCO15, Bos16]. Multispecies [DJMZ16]. Multivalued [GR15a, MS11]. Multiwave [Yos17]. Multiwavelets [GIV17, HZ10]. Multiwell [CC10, JS14]. Mumford [PSZ19a]. Muskat [LM17]. Mutually [BDFS18]. Myers [LMTT15].

Nagumo

[CS15a, CKM14, CCHR18, CFF19, SSH19]. Nano [LZ19]. Nano-Gap [LZ19]. Nanowires [Car14]. Narrow [CF11].

Natural [FT17]. Naturally [FHMP16]. Navier [Kim13, SS17, WLT16, ADL14, BT19, BV13, BFGPE^{+}12, BFLN16, BM12b, BW12, Bre13, BMR14, BC17c, CDLLSG13, CRWX16, CCK18, CJN19, CS10b, CSZ18, CEIV17, Deu13, DF11, DWYZ12, FZZ18, FPVR13, GMT19, GHMZ10, GP18, GW15, Hof12, HW17, HLW12, HW14, Hyn13, IK11, JJN13, JS13a, JLX15, JZ19, JWX13, Jün10, Kim09, Kwe12, KK17b, LLW17, Lei16, LR15b, LS18, LST12, LMW17, aLW18, Li19, LYZZ14, LY19, MPS19, MOR^{+}16, NPS13, NN19, PZ11, Per10, PW15, QW11, RZ14, SS19, TYZZ13, TWW15, VW15a, VY16, WXY15b, WZ13b, WZ17, XZL10, ZZ14, vBW11]. Navier-Slip [WXY15b]. Near [ALZ19, CH15, CQW18, JLZ18, KY12a, LMW17, MRS16, Pen17, DT15, RP18, SWZ15, YT11, Yun15, ZZ19, ZT17, AHP13b, AZ17, GHMZ10, Kac14, LP19, Str10]. Near-Circular [YT11]. Near-Equilibrium [ZT17]. Near-Field [CQW18]. Nearly [BBG16, Sch17]. Necessary [Lee10]. Negative [CHW16, HV13, LS16]. Nelson [AF17]. Nematic [CRWX16, Ces11, CKY18, CHT18, Dai17, GSV19, GW13, HM12b, HM13, LLW15, Nov18, SS15, Tay18]. Nematics [MN16]. Nernst [HL15, WLT16]. Network [GN15, LWZ18]. Networks [BW17, CCC18a, CG10b, DFT17, IT15, NRS17, PT11]. Neumann [Lai14, MT16, AKKY17, AMW11, Ash13, BFV17, GL19b, GS18, HPS12, LT11, MT15, MB16, MR15, Sus13]. Neumann-to [HPS12]. Neural [IT15, VF13, VF15]. Neutral [DM15, GPI18, GH12, WW12]. Neutrino [BFG^{+}13]. Neutron [GY16b]. Nevanlinna [BOS17]. Never [HS19]. Newton [BFDJ13, Kar12]. Newtonian [CK13, Ara16, BKP13, BP14a, BP14b, BBS16, CK11, GVWK16, JH18, Kot12, Len14, Str10]. Nicholson [LLLM14]. NLS [CM19, HF13, KMV18, LWX11, Mae17, Mur14]. No [GS15c]. Node [VW15b]. Noise [BBT14, BMY16, BG14, BHR16, BGAHS17, Cer11, CDM16, SSZ19, Tan18, WZ16]. Noise-Induced [BG14]. Non [AR19, BP14b, BBS16, BC11, CQ19, GVWK16, JH18, KVM18, KSW13, Kot12, Sen17, TWW15, TW18a, WZ16, YY18]. Non- [KSW13]. Non-cutoff [YY18]. Non-Flat [TWW15]. Non-isothermal [Kot12]. Non-KPP [CQ19]. Non-Lipschitz [KVM18, WZ16]. Non-Muckenhoupt [AR19]. Non-Newtonian [BP14b, BBS16, GVWK16, JH18, Kot12]. Non-Resistive [TW18a]. Non-Self-Adjoint [BC11, Sen17]. Nonanalyticity [MPN14]. Nonassociative [BFM12]. Nonautonomous [CHK15, CL17, FHK13, NPS18, HR15]. Noncharacteristic [Ngu10]. Nonclassical [Can10b]. Noncoercive [DDMM18]. Noncompact [BDLM19, DL15a, JH18]. Noncompatible [JT13]. Nonconservative [CWYZ16]. Nonconvex [CGT11, NSS17, Pin19, GY16b]. Noncylindrical [BDSS18]. Nondiagonal [AS13, MS18a]. Nondominated [CEH14]. Nonexistence [GKR19, GM14]. Nonexponential [RR17]. Nonfocal [FGR15]. Nonhomogeneous [Ell12, FPVR13, RSZ18, LT11]. Noninteger [AS15]. Nonintegrable [EJ14]. Nonisentropic [JJL16, LW14b, WFL12]. Nonisothermal [FHK11, LS10]. Nonlinear [AS14, ACJ12, AS15, AAS19, BBT14, BR18, BHSZ10, BF19, BGT19, BD18, Bla18a, BDT12, BCdSN18, BGM19, BCT19, BHWY12, BDPS10, BO16, CDS10, CHKP19, CGP13, CS10b, CC10, CR10, CH11a, CK11, CK13, CO12, CN15, CG10c, DSY18, DMZ14, DLZ15, FY13a, FK18, GZ13, Gie14, Gin19, GM15, Gra19, Gro10, GSWZ18, GHLN13, GP11, HKN16, HI19, HY19, HJ15, HL19b, JZN11, Jun14, Kol11, KK10, Kry13, Lau10, Len16, LS12b, MO19, MPR10, Mar10, MM18b, MS13a, Med14, MOS14, MO15, MC14, NS12, Oh15, PPPV16, Pan12a, Ped15, RSZ18, RTZ17, Sab13, Sac18, SZ12b, SP13, Sch10, Sho10, SSZ19, Sou19, SV14, TW11a, Wan13, YZ18, Yos18, Zho15, vBM14]. Nonlinearities [AS15, BCT19, CF14, Cer11, CHL17, DG11, DR13, Feh13, GP15]. Nonlinearity [BC17b, CCNP17, IS13, MM18b]. Nonlocal [ABR17, AG16, AM15, ADHZ15, BT16, BMC14, BMC18, BC14, BRS17, BGM19, BCQ12, Cha14, CMP13, CR18, CW13, CEQW16, CT16, CFF19, DSV15, DH10, DM14, DGVBW10, DFHM14, GV19, GO18, GT16, HLGMMM14, IISD15, KLW17, KPS18, Le10, LR11, MOZ10, MOZ12, MX19, MS14, PZ17, SV11, SM19, ST17, Sto19, TD17, TW18b, VZ15, WDL18, Zha14, vBM14, BMY16, BL19]. Nonmonotone [LLLM14, YZ15]. Nonnegative [CPT10, Li19]. Nonnegativity [LW15]. Nonneutral [GMP13]. Nonoscillating [Dan17]. Nonradiating [Blå18b]. Nonrelativistic [BCT19, BC17b, JLX15, RR13]. Nonshear [Iye19]. Nonsmooth [BO16, CLW17, Dan17, FK19b, HK15]. Nonstandard [CGS17]. Nonstationary [CW16, Kha13, PY14, RZ14]. Nonstrictly [Lu13]. Nonsymmetric [AZ12, Nad10]. Nontrivial [DWX18]. Nonuniform [HKT18, TW11b]. Nonuniqueness [KZ11]. Nonvanishing [AFT15, KMV18]. Nonzero [BK15b, CCK18, Deu13, vBW11]. Norm [CNR17, CJ19, HNSW11]. Normal [AHP13b, Ber12a]. Normalized [CH19a]. Norms [BFFO17]. Novikov [HHK18]. Nozzle [CHY19, FL12a, LW14a]. Nozzles [BDX14, BP19, CDX12, DWX18, WX15, XX10]. Null [dAdM18]. Number [Bou13, DD16, DLZ15, FKM19, LST12]. Numerical [ADHZ15, GO18, YMYC10]. Numerics [QWE19, vdBMJLM11].

Object [Ngu17]. Objective [MT19]. Observation [Ngu15]. Obstacle [ALP15, Cav12, DD18, Igb17, EH16]. Obstacles [NUW11]. Occurring [DKR15]. ODE [INSZ14, LMP11, SSW14]. ODEs [LWZ16, LL18a]. Off [AAS19]. Off-axis [AAS19]. Ohm [JM12]. Ohta [Le10, vdBW19]. Oil [Can10b, Evj11]. Oil-Trapping [Can10b]. Oldroyd [FZ16, LZ18]. Oldroyd-B [FZ16]. Oleinik [Naz12]. One [ABCL18, ALST14, ABK12, BBMN12, CCNP17, CPSW16, CHKP19, CH13, CQX18, CCC18b, CG11, CDK11, CEQW16, Del18, DY10, EP12, EM10, FRX19, DFHM14, GIP^{+}13, GZ13, HS10b, HLWW18, HL15, HLW12, JK10, LR11, Li19, LYZZ14, MPT19, MNT13, Nii12, Nol11, OW14, PW18, Pol17, SdlL13, TYZZ13, VK18, ZN19]. One-Dimensional [ABCL18, ALST14, ABK12, BBMN12, CCNP17, CHKP19, CQX18, CCC18b, CEQW16, Del18, EP12, FRX19, GIP^{+}13, GZ13, HLWW18, HLW12, JK10, LR11, Li19, Nii12, OW14, PW18, SdlL13, TYZZ13, LYZZ14, MPT19, Pol17]. One-Species [DY10]. One-Step [HS10b]. Only [Can10a, Can10b, CWY14]. Ono [HLR^{+}19, Oh10, Wu16, Wu17]. Onsager [DN18]. Open [DFP14a]. Operator [AH16, AV16, AKKY17, AVP16, Bar14, BBR19, BNDHV10, BC11, CP13, CH19b, FQ11, GP14, GHH17, GH18, GLT10, GHLN13, HKK13, Jia12, KP13, Nad10, PZZ19, Wu16, Xu18, Zha14]. Operators [AF16, Ale16, AZ12, AR19, AMV15, DL15a, DKR15, Dro18, FR17, GR15b, HS10c, HKOP10, HKOP11, HW11, JHN12, KPZ19, Mel12, Nii12, PZ17, Sen17, Syl12]. Operators.Part [GS18]. Opinion [PLPSS18]. Oppositely [BKR16]. Optic [CS17]. Optical [ES10, FW18a, YZ14a]. Optics [Ben17, CDS10, Ngu13, NT19, WW18]. Optima [BH11]. Optimal [ABØP19, APSV19, BPS19, BBG17, BKR16, Bon13, BFLS12, BW17, Bru16, BV18, Can10a, CGS10, CDPS17, Cav12, CLLS17, CCG18, CDS19, DKR15, FV18, Fri19, GKR19, GX17, HW13a, KW12, KN18, Lee10, LX17a, LMS16, LR17, MRT14, MS11, MRV19, OSW19, Pas11, Pas13, Sof18, VZ15, Che15]. Optimality [AHP13a, AHKM15]. Optimally [GL12, KLL12]. Optimization [BFV17, BBV14, HS16, LR17]. Optimize [EH16]. Orbit [BG14]. Orbital [CPP18, KK18]. Orbitally [BCdSN18]. Orbits [GL15, XYZ16, vdBMJLM11]. Order [AI12, BFFO17, BHND18, BÖ19, BGT19, Bos16, CJP13, DKR15, FQS10, GVZ16, Gie14, GL19b, Hal13, IM10, Kac14, Koc16, KNR12, LN10, MS18b, PZZ19, Sch18a, Val15]. Ordinary [GS12a, IM10]. Organized [JXZ16, Mos18]. Oriented [FL12b]. Orlicz [BCD17b, NP16]. Orr [GN19]. Orthogonal [GIV17, SSST15]. Orthonormal [HZ10]. Oscillating [Bos12, CGM16, EP12, GS15a, KPZ19, LW15]. Oscillations [CTW13, CV12, DM15, LXZ13, SW11b]. Oscillator [BMY16, Tro17]. Oscillatory [AVP16, CS15a, Dro18, DVW15]. Oseen [PWW17]. Ostrovsky [GP19, HSS17, LPS10]. Ostwald [Yos17]. Other [BKR16, CG10c, FGJ11, VZ15]. Outflow [CJN19]. Output [CW13]. Overlap [AAK14, ADK15]. Oxidation [MMB11].

P1 [JLX15]. Packets [dHGR14]. Painlevé [LW16b, dMIS10]. Painlevé-Type [dMIS10]. Pair [Dek19, Sab13]. Pairs [ABDD19, SZ11]. Palais [CLW17]. Panel [LW16a]. Parabolic [AI12, BBT14, BT16, BST17b, BGT19, BO16, Cal15, Cui13, Dan17, DdMH15, DV10, DS13, DK11, DK14, EJ14, EMZ17, FS15, Gia15, GYY18, Hal13, HK15, HS13, KLW17, Kry13, Kry14, LMR15, MPR10, MS13a, NKV19, Pan12a, Pim16, PY14, Pol17, RTZ17, ST10, ST17, Ter11, ZLMZ18]. Parabolic-Hyperbolic [Cui13]. Paradox [LL11]. Parallel [Bre13, Kar16]. Parameter [Kac14]. Parameters [KLS11, Liu18a, Pen15]. Parametric [ABCD^{+}18, HS13]. Paraorthogonal [Sim16]. Parseval [GL12]. Part [GS17, AHP13b, DFP14a, DFP14b, Pao15, Pao16, Val15]. Partial [AGS13, AAS19, BCD^{+}17a, CCFdlL14, CW16, CMM13, DdMH15, DLSV12, DZ15, GK10, Hal12, Hal13, Hyn13, Igb12, Kia16, LRdS18, OR19, SSZ19, YZ14b]. Partially [ATSR19, DK11, MY12]. Particle [ALST14, BCL11, CPSW16, FW18b, GL17, Höf18, Hyn19, IT15, KvMY19]. Particles [Bos12, Ces11, DGV16, FL12b, HKR18, SY14, WLT16]. Partitions [BK18a, OR17]. Passage [BG14, FK18]. Past [CY15, HZ19, CKZ17]. Pasta [FW18a, Miz11]. Patch [BGLV16]. Patches [SWZ15]. Path [BB10b, RTZ17]. Path-Dependent [RTZ17]. Paths [LSW17]. Pathwise [Tan18]. Patlak [BH18, BK13, BH17, KY12b]. Pattern [HSV16, Yos17]. Patterns [BLW19, dRDR16, dR18]. PCM [WX12]. PDE [ALS15, CWH18, CK12, Dan17, HL19b, LMP11, LMR15, RR15, SSW14]. PDEs [ABL13, BEH15, Bos16, HS13, LL18b, RTZ17]. Peak [BOS17]. Peaked [GP19]. Peakon [HHK18]. Pearling [DHPW14, KP18b]. Pekar [LR13]. Penalization [BCP19, CLLS17, Olb19]. Penalty [AF15]. Pencil [Ria10]. Pencils [BBR19, Kol11]. Penetrable [HL11, LX17b, NUW11]. Perfect [BKR16, DS19, LX17a, MR15]. Perfectly [Alm17, Kal12, MNT13]. Perforated [PPP13]. Performance [HR12, WX12]. Peridynamics [BMC14, BMC18, SM19]. Perimeter [DNWW19]. Periodic [Ale16, AVP16, BM15a, BMP18, BG14, BCG10, BGM19, Bre13, BR11, CL17, CHKP19, CCC18b, CKS15, CDD^{+}12, Daf13, DL15a, DVW15, EF15, GP19, GP14, HSS17, Hen10, HW11, HWZ12, HJ15, IM10, IM18, JX15, JZ10, Joh13, KZ11, KL18a, KDT19, LP19, LR15b, LZZ17, LW15, Mae17, Nad10, Nii12, Oh15, PZ17, Pra13, RZ16, SSH19, Sch10, Sen17, Sus13, TZ18, TY11, Tsu12, Wan13, XY14, XYZ16, XYY19, YCW10, dRDR16, dR18, dlLSZ17, vBW11, GZ13, HGW14, SdlL13]. Periodic-Coefficient [RZ16]. Periodically [CCCdlL17, DN12]. Periods [HWZ12]. Permanence [BS16a, RSS17]. Permittivity [GS15c]. Perona [CG11, KY15]. Perpendicular [AHP13b]. Perry [LMTT15]. Persistence [GRT14, Le 13, Pan12b]. Persistent [CKV18]. Persuasion [PLPSS18]. Perturbation [ETZ13, Ess16, HW17, Le 13]. Perturbations [BFN^{+}13, DVW15, GVZ16, JT13, Lau18, Mar10, Smi17, XYY19]. Perturbed [AMW11, Cer11, CDZ13, GL19b, ILR17, Mor19]. Peterlin [LMMNR17]. Petviashvili [EL17, ILP16, JZ10, LP19, LPS18]. Phase [AG17, ALP15, AP11, BKLU18, BBG12, BCQ12, Can10a, Can10b, CP12, Cho16, CDNP16, CL13b, DNWW19, Evj13a, Evj13b, FL12a, FGJ11, FL12b, GCGJL18, GM17c, GR13b, HNS17, HL12, HD17, HK15, HMSZ13, JMZ18, KLO16, LWZ16, QWE19, RZ17, RR15, Ter11, YZZ10, YZZ12, Abe12, JRK19]. Phase-Field [CL13b, FGJ11, HD17, DNWW19]. Phase-Space [AP11]. Phaselocked [Tro17]. Phases [BKR16, Mil18, ST11]. Phenomena [BDFS18, FQS10, FGJ11, FMP18, Gia15]. Phenomenon [RTY16]. Phenotypically [IM18]. Photoacoustic [SY17]. Physical [EW18, Lei16]. Physics [NOS12]. Phytoplankton [DH10]. Pick [BOS17]. Piecewise [CHN18, Fri18, Sei14, Tej17]. Pinning [WY13]. Pipes [Ara16]. Pipkin [EI11]. Pitaevskii [BAC12, CPP18, DDF18, HTX15, ILR17, KMS17, Tri18, WY13]. Place [TZ15]. Placement [EH16]. Planar [BL11, BL14, BOO18, BM18, CKM14, CJ19, GLY18, aLW18, Pol17, SY17, dlHHMV16]. Planck [WLT16, ASC19, BR18, GHMZ10, HL15, HJJ18, JZ18, KKT17, LMW17, MT19, YY10, Zha14]. Plane [AHP13b, BK15a, CW16, Dai10, ESvR12, HPS12, LTV17, MPS17, RSZ18, Sei14, Tej17]. Planetary [HCHY16]. Plans [Fri19]. Plasma [CCM14, HK10, NOS12, Ohn15, Otw10, Sch14a, Sch17, WFL12]. Plasmas [NS13]. Plasmon [LZ19]. Plasmonic [LZ19]. Plastic [BMY16]. Plasticity [BFM12, CDK11, DF15, DS19, Gin19, GS17, SZ12b]. Plate [BGM19, MRV12, MRV19, SV18]. Plateau [AAD13]. Plates [CNS10, MPT18]. Poincaré [AKKY17, BB17, DSZ19, DNS12]. Poincaré-Type [DSZ19]. Point [AMV15, ALM10, GM15, HSV17, HPS12, Lee16, Nii12, PT11, Sei14, vBM14]. Points [BOS17, BV10, CKV18, GS15c, LWZ15, Ono11]. Pointwise [DKN11, GPT19, Wu14b]. Poisson [DYZ19a, DY10, DL15b, JZ18, Li11, lLWW18, Wan12, BDX14, BAC12, BACP18, Bos19, CL18, CJ19, DGV16, Des14, GM11, GMP13, GPI18, HL15, Kom15, KK18, Li09, LP16, LS17, Mas11, NOS12, Ohn15, Pal14, Pen15, Pu13, TYZZ13, TWW15, WLT16, Wen14]. Polarized [GG10]. Polarons [LR13]. Pole [AF16, DXZ18]. Polycrystals [FPP19]. Polygon [Ash13]. Polygonal [BDEM18, BFV17, DT15]. Polygonal-Shaped [BDEM18]. Polyharmonic [BB10a, Han18]. Polyhedrons [Ber12b]. Polymeric [CL13a, JLZ18, LS12b]. Polynomial [Cer11, HS13, RR17]. Polynomials [And12, GIV17, SSST15, Sim16]. Polytopes [DK14]. Polytropic [FHK11]. Population [APSV19, GV19, Lam12, Mit10, MV19]. Populations [IM18]. Poro [HS18].

Poro-Visco-Elastic [HS18]. Poroelastic [LP14]. Porous [AMP10, ACM^{+}12, AT14, BDG13, Can10a, Can10b, CQW18, DKR16, DT14, FG15, Ges13, HKK17, JRK19, LS10, LS12a, MMB11, QWE19]. Posed [CM11, CWH18, HKK13]. Posedness [ADL14, AET18, ALST14, AN15, BG17, BFTT18, BTZ15, BFS14, CGP13, CY19, CHS13a, Faj16, GV19, GM10, Gna15, HNS17, HL12, HX10, HY14, HHK18, IKS12, JTW16, JL19, KM17, LS13b, LPR12, Li19, LPS13, Mar18, MSZ19, NN12, RV12, TW18a, Tsu12, WXY15a, WLT16, Wan18, Wu14c, XXK13, YZ18, ZZ14, ZT17, Abe12, DZ15, GW18, HTW18, KMVW14]. Positive [Bor19, CF14, DDF18, FG18b, LV15, HL19a]. Positivity [SV18]. Positron [Sab13]. Posteriors [TSA18]. Postflutter [HTW18]. Potential [ADHZ15, BM15a, BDG13, CH15, CKS15, CM19, DXZ18, DSY18, GZ13, GZ18, JS13a, Kai17, KS19, LW16a, LZZ15, LWY18, LLM19, LS15, Mae17, SM19]. Potentials [BBV14, CK11, CK13, DFV18, FY13a, DFHM14, HL19a, Kia16, TAGP18, Tej17]. Power [AF15, AS15, AGS13, AP14, EKR18, DFHM14]. Power-Law [AP14]. Prandtl [Iye19, KMVW14, LWX16, LXY19, WXY15a, XY18]. Precipitation [HKK17]. Precompression [DP14]. Predator [BS16a, HS19]. Predator-Prey [HS19]. Prescribed [BOS17, CJ19, KRW15, KL18b, TZ13]. Presence [ABGS19, BW12, KM13, MNT13, WBS13]. Preservation [dCPS16]. Preserving [BKP13, Dai10, MLD19]. Pressure [CP12, Evj13a, GS12a, JMZ18, SS19, WX19]. Pressure-Dependent [Evj13a]. Pressureless [Ber17, CSW15]. Preventing [BCJ20, HJ11]. Prey [BS16a, HS19]. Primitive [LT17, LT19, TW10]. Principal [DL15a, LZZ17, Nad10, PZZ19]. Principle [AHØP13, AF17, BM15a, Cer11, CL17, Kal12, KY12b, KL18a, KR10, Nol11, SX13].

Principles [BL19, FR17]. Priori [LM14]. Probabilistic [BR18]. Probability [CMWY16, DXZ18, PSSW15]. Problem [ABL13, AHP13a, AAGP18, AI12, AH13, ALP15, AM15, AT14, AMW11, AN15, BL11, BL14, BDEM18, BdHQ13, BdHFS16, BMSR^{+}13, Ber12b, BG17, BC14, BV18, CCG10, CCH10, CGM16, CHL19, CCMW19, CDN10, CHN18, Cav12, CLLS17, CF11, CGP13, CP10, CP11, CT15, CCLM15, CH11b, Cui13, Dai17, DD18, DKR16, DLVW13, DL18, DSV15, DG16, DYZ19a, DH10, DLZ12b, Fai14, FRX19, Fel18, GO18, GSV19, GS15c, HNS17, Han18, HHR09, HHR11, Hof12, HKT18, HW17, HLGMMM14, HCHY16, HPS12, ILR17, IY12, JTW16, JJ18, Kai10, KM13, Kar16, Kat19, KLW17, KMV18, KT13, KW12, KLO10, KT11, Lai14, LV12, LTV17, LM17, LX17a, LX19, LPS18, LR17, LT11, LS12b, LY19, LS13c, MMT19, MY12, MRT14, MBPS13, MS14, MR15, NKV19, PPPV16, Pas11, PPP13]. Problem [Pla14, Poh15, QW11, Rei18, RTT19, RZZ18, SZ12a, SWX17, ST17, Sus13, TW18a, TY11, Ter11, TW18b, Ves15, WW12, WX19, Wan19, WNRJ13, Wu17, Yos17, Yos18, YZ16, ZZ19, ZT17, ZLMZ18, vdBW19, Che15]. Problems [Ale16, AF15, ABCD^{+}18, BKK18, BMP18, BFGPE^{+}18, Bar14, BT16, Beb16, BMC14, BMC18, Ben17, BMY16, BRS17, BLS15, BR11, BBV14, CQ12, CM14, CCC18b, CG10c, DKR15, DMZ14, DLZ15, ES10, Ell12, ERV17, EHM16, FQ16, GIP^{+}13, GY16a, GL19b, HMS14, HSV16, HWWY13, Kar12, KDT19, LV10, Len16, LP14, MPT18, MM11, Moa11, NP16, Otw10, Pao15, Pao16, Pas13, Pin19, RSZ18, RR17, Sei14, Tre13, TSA18, Xu16]. Procesi [HZFQ13]. Process [AH18, BHM17, HS14, KSW13]. Processes [ABØP19, MSZ19, MP12, PT11, RT17, Rou10, YMYC10, ZK15, ZBL19]. Processing [CM13]. Product [DS10a, GS12a, HMZ15, Liu19].

Production [Sab13]. Profile [ABCL18, GS15c, LMZZ17, LMZZ18, Mei10, TWW15]. Profiles [CCHR18, GP18, GLW17, Lam12, LMN^{+}10, dCPS16]. Programming [LV13]. Projection [KP13]. Projector [HNSW11]. Proof [CSW15, CM13, HM12a, RR13]. Propagating [Pol17]. Propagation [BÖ19, DO16, DIT15, FY13b, FG15, Ges13, Gia15, HX14, ILP16, Pal14, Pin19, Qin15, RTV17]. Propagations [MO19]. Propelled [DDM11]. Proper [ABCD^{+}18]. Properties [AIK10, BCDG16, BST17b, BKR16, DGVBW10, FQ11, GH18, KY12b, KZ18, LS13c, Mel12, Rod16, WU14a, ZK15]. Property [BPS16, CKY13, Dua11, Jia12, KW11, LN10, Ngu16, SW18]. Proximity [Gro10]. Pseudo [BB10a, Ngu16]. Pseudo-local [Ngu16]. Pseudo-Polyharmonic [BB10a]. Pseudoidentity [SZ11]. Pucci [BL19]. Pull [CG10c]. Pull-In [CG10c]. Pullback [ZK15]. Pulse [HSS17, dRDR16, dR18]. Pulses [CS15a]. Pure [BHR16, dHGR14].

Quadratic [GH10, IS13, Kar12, LV13, Thi19]. Quadrature [CM14, Lit13]. Qualitative [ÁCDP14, CKY13, Lei13, Ter11]. Quantitative [BdHFS16, CJP13, Neu16, PUW18]. Quantities [Daf14]. Quantization [WX12]. Quantum [BJ16, CKS15, Ilm16, Jün10, KLO16, LR13, LP16, VY16]. Quasi [Abe12, And12, BFM12, BK15a, BM19, CCCdlL17, CV16, Che19, DNK12, DS13, DLSV12, DM15, EW15a, GZ13, GS17, GPI18, HGW14, Kal12, Kre19, LRdS18, Mae17, NKV19, Nov19, PX13, SdlL13, Tsu12, WW12, ZBL19, dlLSZ17]. Quasi-Conformal [BK15a]. Quasi-convex [DLSV12]. Quasi-cylindrical [Kal12]. Quasi-Filling [CV16]. Quasi-Geostrophic [BM19, Che19, EW15a, Nov19, ZBL19]. Quasi-incompressible [Abe12].

Quasi-Linear

[DNK12, LRdS18, NKV19, DS13].

Quasi-Neutral [DM15, GPI18, WW12]. Quasi-Periodic [Mae17, Tsu12, dlLSZ17, GZ13, HGW14, SdlL13]. Quasi-Periodically [CCCdlL17]. Quasi-Reversibility [NKV19]. Quasi-Self-Adjointness [Kre19]. Quasi-Smooth [And12]. Quasi-Static [GS17, BFM12]. Quasi-Töplitz [PX13]. Quasilinear [BYH15, DdMH15, Ria10, Sug16]. Quasineutral [HKN16]. Quasistatic [AH18, MC14]. Quintic [KMV18].

Radial [BLW19, CK11, CK13, CNR17]. Radially [ACM^{+}12, BGL12, LM14]. Radiating [Ohn14]. Radiation [JLX15, KL18a, WX11]. Radiational [Liu18b]. Radiative [AH13, BCS16, DD16, FRX19, LMN^{+}10, SW11b, ZZ19]. Radius [Bos16]. Radon [ABDD19, CNR17, FQ16, Moo16, Ngu15, PW15, Rod16]. Ramified [AT10]. Random [Alm17, BB17, BP10, CDZ13, DXZ18, FR17, GGAS14, GKR18, HS19, HS13, KVM17, Koc16, LMR15, LLM19, LL19, LW16b, MJ14, MPT19, MZ13, MV19, Nol11, Smi17]. Randomly [EKR18, FFGHR17]. Range [FJ18, MM18b, ST15a]. Rank [Beb16, FWW17]. Rao [GM17a]. Rapidly [Dro18, EP12]. Rarefaction [DL15b, FRX19, HLW12, JWX13, aLW18, lLWW18, LW17, Per10, XYY19, ZH10]. Ratchets [KUV16]. Rate [AHKM15, ABCL18, BHWY12, CMM10, CT11, Dai10, IM10, RT17, Rou10]. Rate-Dependent [RT17]. Rate-Independent [RT17, Rou10]. Rates [AP18, BCD^{+}11, BOS11, Che12, CWYZ16, Ess16, HW13a, MS13b, OSW19, Sri11, XY14, Xu16, dCPS16]. Ratio [KL18b]. Ray [KM18, Mon16, Zho18, PUW18]. Rayleigh [GH10, GT10, JJ18, WW18]. Reacting [BP14b, BPZ17, CKZ17]. Reaction [AKKL17, ÁCDP14, BMP18, CFSS18, Cer11, CL17, CFF19, DFT17, DH10, FHK13, GLS10, GLY18, GST13, HS10a, HHMM18, LW12, LZZ17, LMS16, MOZ10, MOZ12, MO15, MB16, MS16, PSV10, SM16, TV18, WW10]. Reaction-Diffusion [BMP18, CFF19, DFT17, GLY18, LZZ17, MOZ12]. Reaction-Hyperbolic [FHK13]. Reactions [HKK17, JMNR11]. Reactive [AMP10, HKK17]. Real [DL18, DKR15, WW10]. Real-World [DKR15]. Rearrangement [Bon13, Nad10]. Receiver [AFK^{+}18]. Receptor [AET18, ERV17]. Receptor-Ligand [ERV17]. Reciprocals [And12]. Reconstruction [BFRV13, HS10b, HU13, KS14, KLO10, NUW11, Tej17, dHHI^{+}14]. Reconstructions [AHP13a, AHKM15]. Recovering [Hal14]. Recovery [FK19b, Kia16, LV13, Mer18, SSST15, TW11b]. Rectangular [BPS19]. Rectifiable [De 18, MMT19]. Recurrent [FGW13, PY14]. Recursive [XYD18]. Reduced [ABGS19, BCD^{+}11, GP19, GZ14]. Reduction [BFY15, CCLCP13, KK16, Nov18, SW11a]. Redundant [KNW15]. Refined [GLW17, Sal12, Sou19]. Refinement [MS18a]. Reflecting [NS13]. Reflection [Cao19, Muñ12]. Reflections [DG11]. Reflectivity [FGN13]. Refraction [Muñ12]. Refractive [Kar16]. Regime [BS16a, BJ16, BCL11, Bos16, DIT15, GLL17, JRK19, LLN19, XY18]. Regime-Switching [BS16a]. Regimes [HTW18]. Region [HMSZ13, HCHY16]. Regions [CCH10, CH11b, KMVW14, TZ13]. Regression [XYD18]. Regular [Cha14, FT17, GL15, PZ13]. Regularity [AC14, AET18, BT19, BKLU18, BGLV16, BCL18, BT12, Cao19, CW16, CLLS17, Che18, Dai17, DF13, DdMH15, DLSV12, DKT19, DKR15, DW13, FR17, FS15, GS12a, Gas19, GMT19, GS15b, GGRB14, GY16b, HX14, HS13, HY13, HTX15, Hyn13, HL19b, ILP16, JZ18, KZ18, KT18, KW12, Kwe12, KK17b, Lee10, LJ17, LWX11, LZ17, Liu18a, LZ18, MRS16, MX19, OR19, PZ11, Pal14, Rod16, RZZ18, SWZ15, ST17, UWK12, WWW12, WZ13a, WXY15b, WX13, ZCO15, Zho15, vNVW12, Che15]. Regularity-Loss [UWK12]. Regularization [BGT19, DL18, ES10, GLL18, Kat19, KP18a, ST19, Val15]. Regularized [CSZ19, CT16, GH12, HSS17, Ngu13]. Regularizing [AAS19]. Reissner [BL15]. Reiterated [FF12]. Related [CL13a, INSZ14, LW16b, MPR10, MP12, MRT15, MP14, Tak10, ZF12]. Relation [KDT19]. Relations [MJ14]. Relationships [LA14]. Relative [BL14, CV15, CK11, CK13, CCLM15, DDGW18, Gie14, JJN13, LT13, LZ18, VW15a]. Relativistic [BAH17, HX14, RR13, Str10, SY14, Wan18, ZN19, Zha19]. Relaxation [BKK18, BMC18, Daf14, Daf19, Ess16, GYY18, LT13, MMT19, OW14, OSW19, PWG11]. Relaxed [LPR12]. Releases [APSV19]. Relevant [Evj11]. Reloaded [FS14]. Remarks [CEIV17]. Remodeled [MJ14]. Renormalization [BBG16, Sab13]. Renormalized [AF17]. Reparametrizations [Bru16]. Repelling [BDFS18]. Replacement [APSV19]. Replication [CK12]. Replicator [KLW17]. Representation [BR18, BK18b, Wei12]. Representations [GL12, WX12]. Repulsion [DFHM14]. Repulsive [BLW19, CDS19, GKR19, Rei18]. Resampling [AHP13a]. Rescaled [MT19]. Rescaling [DDMM18]. Reservoir [Evj13a]. Resistance [AP15, Pla14]. Resistive [KT11, TW18a]. Resolution [LL11, PP19]. Resolved [Gli13]. Resolving [AGS13]. Resonance [AKKY17, FQS10, LL16b, LZ19, MMB11, Ngu17]. Resonances [LS15]. Resonant [AZ17, HGW14, PD17, YZ18, dlLSZ17]. Resonators [Lei10]. Respect [BDEM18, FWW17, Sal12]. Response [CCCdlL17]. Restitution [AL10]. Restoration [DSX17]. Restricted [BL11, BL14, KW11, KM18]. Restriction [Sch14a]. Result [BRS17, BK18b, BM15b, CP12, DDM11, DdMH15, DKT19, LMMNR17, MRS16, Pao15, PT18, Poh15, VF13, VF15]. Results [AB10, BC14, CGP13, Chu14, DP13, FKM^{+}16, Fri18, GS12a, INSZ14, JS14, KZ11, RV12, RT17, Sof18]. Réthy [CDX17]. Retrieval [AG17]. Reversibility [NKV19]. Reversible [Bor19]. Revisited [CS18]. Revisiting [DF15]. Reynolds [Bou13]. Ribbons [FHMP16]. Riccati [dRDR16]. Riemann [AAGP18, CQ12, FKV15, HWWY13, Lai14]. Riemannian [BMMP16, BMP18, BC19]. Riesz [Due16, HSV17, HR12]. Rigid [BST17a, CDK11, GS12b, MPS19, MRV12, MRV19, MR15]. Rigidity [CC10, CFO19]. Rigorous [BFLN16, CCC^{+}17, CM12, CT14, CFF19, NTW19, WZZ15, vdBMJLM11]. Rigorously [vdBW19]. Rimming [BC11]. Ring [ST11, WY13]. Rings [LS13a]. Ripples [FW18a]. Ritz [GH10]. Robin [AS13, BMMP16, Sin10]. Robust [RSS17]. Rods [BFLS12]. Roles [TW11a]. Roll [JZN11, Le 13, RZ16]. Roll-Waves [Le 13]. Root [Bru16]. Rosensweig [NTW19]. Rotating [Che12, CQX18, DKN11, FGN12, SW17]. Rotation [CFGL17, EL17]. Rotation-Two-Component [CFGL17]. Rotational [KY12a, MO14]. Rotationally [GM10]. Rotator [GPPP13]. Rothe [BS16b]. Rough [AB10, BP12b, CDLLSG13, CWE10, DG16, EH13, GVWK16, HL11, LWZ11, She15, Wol19]. Roughness [CM12]. Roughness-Induced [CM12]. Rule [CD11]. Running [AMW10].

Saari [YZ16]. Saddle [VW15b]. Saddle-Node [VW15b]. Saddles [BHND18]. Sampled [ZCO16]. Samples [TW11b]. Sampling [AHP13a, AHKM15, BFY15]. Sandpile [FJ18]. Sandpiles [CCC18a]. SAR [AFK^{+}18, FGN13]. Satisfying [LL12]. Saturated [Mar10]. Saturation [Cal15]. Saxton [Wun10]. SBD [Cri19, Fri18]. SBV [BT12, DLV10]. Scalar [AGN19, CGM16, CV15, CKS15, CW13, Daf13, HKK13, IM10, Jun14, KPS18, KDT19, Lai14, LMN^{+}10, LW17, MS13a, MY12, VK18, WDL18, XYY19, Yos17, Yos18]. Scale [FR17, GW15]. Scale-Invariant [GW15]. Scaled [ET16, KMS17]. Scales [JZ18, Lei16]. Scaling [Bia18, BW17, FJ18, GZ14, GH14, LLN19, dCPS16]. Scatterers [Lee16]. Scattering [ACZ14, BYZ12, BDEM18, BV10, BK15b, CGM16, CHL19, CLM17, CWE10, CWH18, CM14, EH13, GHLN13, HL11, HSV16, IN13, KL18a, LL16a, LWZ11, LL19, LS15, LX17b, Liu19, MM18b, Mer18, Mur14, NUW11, Spe14, Wu16, Wu17, YZ18]. Scheme [CHW16, CG11, FL17, GM17a, Ngu13]. Schemes [BHSZ10, CDPS17, CFSS18, Gro10, GK10]. Schrödinger [AV19, AH16, AS15, AAS19, Bar14, BdHQ13, BF19, BFDJ13, BL19, BCdSN18, BDLM19, BLW19, CDS10, CHKP19, CR10, CJ19, CO12, Com17, DSY18, Dro18, FY13a, GZ13, GM11, Gra19, GHLN13, GP11, HL19a, HS10c, HW11, HY19, Ign10, IY12, KS19, Kom15, KK18, Lau10, LLM19, Liu19, LS17, Mar10, Mas11, MM18b, MS16, Nii12, Oh10, Oh15, RSZ18, Wan13, YZ18]. Schrödinger-Type [BdHQ13]. Schwarz [Nad10]. Scott [vdBMJLM11]. Screening [BKR16]. Screens [CWH18]. Screw [BBO19, HO15]. SDE [WZ16]. SDEs [MPT19]. Sea [CS15b]. Seawater [CDR17]. Second [AI12, BFFO17, Ber12b, CJP13, DKR15, FQS10, Har18, Kac14, LN10, MS18b, PZZ19]. Second-Fourth [LN10]. Second-Order [BFFO17, DKR15, FQS10, MS18b]. Sections [DFP14a, DFP14b]. Sedimentary [HS18]. Sedimentation [Höf18]. Segel [BH18, BCJ20, BK13, BH17, CLW12, HJ11, KY12b, ZLMZ18]. Segregation [CFSS18, GM17c]. Sekerka [Le10]. Selection [MO19]. Self [AT10, ABCL18, BC11, CMM10, CH11a, CDNP16, CKV18, DDM11, ER12, Gna15, GM15, GP18, JXZ16, Kre19, LLW17, LvR15, LM17, LMR13, LNZ14, Mos18, RR13, ZN19, Sen17]. Self-Consistent [CH11a]. Self-Generated [ZN19]. Self-Gravitating [RR13]. Self-Organized [JXZ16]. Self-Organized-Criticality [Mos18]. Self-Propelled [DDM11]. Self-Similar [AT10, ABCL18, CKV18, ER12, Gna15, GP18, LLW17, LvR15, LMR13, LNZ14]. Self-Similarity [CMM10, LM17, GM15]. Semi [CMP13, FT13, Vis18]. Semi-geostrophic [FT13]. Semi-implicit [CMP13]. Semi-Monotone [Vis18]. Semiclassical [AH16, AP11, DR13, HS10c, LWZ16, MZ18, PR13]. Semiconductors [HMWY11, HMW11, HMWY12, LMZZ17, LMZZ18, WW12]. Semicontinuity [BMC18, DLV10]. Semicontinuous [BPW15]. Semidiscrete [BHSZ10, CG11, GM14]. Semiflows [FZ14, FWW17]. Semigroup [EW15a, HTW18, MT13]. Semihyperbolic [SWZ15]. Semilinear [BMMP16, GKR18, HS16, KVM17, Lam12, MQS12, Mar18, Moa11, Pim16, PY14, Pol17, QS12]. Semipermeable [GHH17]. Semisupervised [ST19]. Semitrivial [CO12]. Sensing [BDWZ12, KNW15]. Sensitivities [Win15]. Sensitivity [BL15]. Separating [Evj13b]. Sequence [LJ17]. Sequences [KVM18]. Serrin [Fel18, HLX11]. Serrin-Type [HLX11]. Set [BL11, Bec18, CGH10, Fri19, GMT16, Lau18, Pan12a, Tan15, Val15, ZK15]. Set-Valued [ZK15]. Sets [BOO18, BLS15, BFLS12, DZ14]. Setting [AG17, CGT11, NP16]. Several [Daf13]. Shah [PSZ19a]. Shallow [AAGP18, Che12, CQX18, DLVW13, DLZ12a, Duc10, HT18]. Sham [SCB20, SCB17]. Shape [BFV17, BL15, BV18, CSZ18, CTW13, HS10b, HU13, HS16, HSV16, KN18, LS13a, XYD18]. Shape-Reconstruction [HS10b]. Shaped [BDEM18]. Shapes [BBG17]. Sharp [AHP13a, BSW16, Bra16, CNS10, CCHR18, CQ19, CP10, CN15, DSZ19, DMZ14, EW15a, LV15, ZH10, ZCO15]. Sharp-Interface [CP10]. Sharpening [CM11]. Shaw [Ono11, TW18b, YT11]. Shear [BH17, BH18]. Shearlet [KLL12]. Shearlets [GL12]. Sheets [MO14]. Shell [AMV15, BHR16, CS10b, DLVW13, Tha19, Len14]. Shells [GH14, Yao19]. Shift [ABR17, DT15]. Shigesada [Kut15]. Shilov [LX19]. Shock [HZ19, LMN^{+}10, LWY18, Ohn14, WW15, XYY19, Yos18]. Shocks [BHSZ10, Can10b, CV15, GM14, Sch14a]. Short [CV15, DF11, FPZ14, FMP18, HSS17]. Short-Time [CV15]. Shortening [CM13, Dai10, ESvR12]. Shrinking [Fis13, Kut15]. Shubin [LX19]. Side [BHRW16]. Side-Stepping [BHRW16]. Sided [BHSZ10]. Sign [BV18, LL16a]. Signal [LV13, TW15]. Signed [LV15]. Similar [AT10, ABCL18, CKV18, ER12, Gna15, GP18, LLW17, LvR15, LMR13, LNZ14]. Similarity [CMM10, CT14, Don11, LM17, GM15]. Simple [CSW15, CK12, DK14, Fri19, Mon16, RR13]. Simplicity [Wu16]. Simplified [Sou19]. Single [AP15, CCC^{+}17, DHPW14, HSV16, JMZ18, SV19]. Single-Curvature [DHPW14]. Single-Phase [JMZ18].

Singular [ABØP19, AFK^{+}18, BÖ19, BP10, BBS16, Cha14, Che12, DS10a, DMZ19, DG11, DG16, EG19, FGN12, FT13, GIP^{+}13, GR15a, GS18, JT13, KK15, KT18, Lac15, Lau18, MM17, MP12, MBPS13, Pes15, WX11]. Singular-Degenerate [GR15a]. Singular/Regular [Cha14]. Singularities [CHL15, DXZ18, DO16, GL12, JS13a, Mer18, Pin19, VW11]. Singularity [CH15, CPZ17, CQX18, Kwe12, WWW12]. Singularly [AMW11, CDZ13, GL19b, ILR17, Mor19]. SIS [CTW17]. Sixth [KNR12]. Size [BDPS10, DF10, DLVW13, Tro17]. Size-Dependent [DF10]. Slater [CS18]. Sliding [GW18]. Slightly [CDLLSG13]. Slip [CDK11, MPS17, WXY15b]. Slip-Plane [MPS17]. Slonczewski [MP13]. Slow [Ale16, CL17, CG11, NT14, She15, TW15, TW10]. Slowly [FY13b, LW15, Muñ12]. Smale [BHK^{+}19, CFRT10, CLW17, CY19, CH19a, HKK15, HKR18, Pes15, PRT15]. Small [BYH15, Bou13, Ces11, CDM16, CP10, CP11, CG10a, CM19, GS10a, HMSZ13, IK11, Joh13, KDT19, KL18b, Liu19, Pen15, PZZ19, Rou10, TY11, WWX15, XZ15]. Small-Amplitude [KDT19]. Small/Large [PZZ19]. Smearing [Löb18]. Smectic [SW11a]. Smectic-A [SW11a]. Smoluchowski [CMM10, ET16, MZ18, Sri11]. Smooth [And12, Bes16, FPZ14, GN15, HNP15, HI12, Lau18, LM11, PWG11, Pen15, Sei14, Tej17, WFL12, WX15, WX16]. Smoothed [FW18b]. Smoothing [Aud12, BCL18, Com17, IO16, Jia12, Jun14, LWX16, LX19]. Smoothly [IT15]. Smoothness [Jia19]. Sobolev [AR19, BM15a, CM11, CWE10, DF13, DKR15, DNS12, NP16, RZZ18, SV11, WU14a]. Soft [BLZ16]. Solar [Gli13]. Solenoidal [MM18a]. Soler [BC17b]. Soler-Type [BC17b]. Solid [dAdM18].

Solids [Rou10]. Solitary [BC17b, CFGL17, EL17, FW18a, KK10, Le19, Whe13]. Soliton [DSY18, GR15b, Mar10, Miz11, Muñ12, NS12]. Soliton-like [Muñ12]. Soliton-Potential [DSY18]. Solitonic [CG10a]. Solitons [BDLM19, CHL17, CM19, ILR17, MRT15]. Solubility [JRK19]. Solution [BC19, Can10a, CH13, CS10a, CMM13, CT14, Cui13, DWYZ12, Ess16, EHM16, GLL17, HKK17, HW13b, HW14, HT17, LS10, LS12a, Liu19, Löb18, Mas11, MPZ15, Ter11, Wan12, WWX15, YY18, ZF12, Zha14]. Solution-Dependent [EHM16]. Solutions [AV19, AIK10, ABGS19, ABCL18, AFT15, ACM^{+}12, AMW11, AMW10, BDX14, BMMP16, BMP18, BR18, BR17, BACP18, BFLS18, BP12a, BST17b, BGT19, Bes16, BYH15, Bla18a, BW12, BDSS18, BGM19, Bra16, BN14, Bre13, BMR14, CCCdlL17, CCM12, CHY19, CRWX16, CKY13, CY15, CTW17, CMWZ18, Che18, CQX18, CS14b, Cho16, CPT10, CG19, CM18, CNSS17, Daf13, Daf14, DS14, DGV16, DDGW18, DO16, DNK12, DS13, DR13, Don11, DWZ10, DW13, DLZ12a, DGVBW10, EW19, EF15, EJ14, EMZ17, Evj11, EW15b, FZ16, FZZ18, FKV15, FT17, FT13, FG18a, FG18b, FHK11, FPZ14, Gar11, GZ13, GMP13, GMT16, GS10a, GG10, GM15, GZ18, GN15, GP18, GW18, GW15, HGW14, HK15, HHK18, HNP15, Hof12, HY13, HHPZ17, HW17, HZFQ13, HL15, HWZ12, HMW11, HMWY12, HCHY16, HNP13]. Solutions [HI12, Hyn13, IK11, IS13, ILP16, IKM17, JJN13, JLZ18, JZ19, Jia19, Jün10, KK17a, KS14, KMT13, KPR15, KY12b, KY15, KK15, KvMY19, KNR12, Kut15, LLW17, LM11, LM14, LR15b, LMR13, Len14, Leq11, LS18, LWX11, LW12, LXZ13, LT17, LZ17, LMW17, LL18b, LLW15, LW15, LT19, LY19, Lu13, Mae17, MPN14, MQS12, MY12, MT19, MOR^{+}16, MP13, MPT19, MSZ13, MRT15, MY17, Moa11, Muñ12, NPS13, NP16, NTW19, Nov19, NP11, Otw10, Pan12a, Pas11, PZ13, PWG11, Pen15, PY14, Pol17, QW11, Rei18, RTZ17, RR15, RZZ18, SM19, SM16, ST10, Smi17, SSW14, TYZZ13, Tan18, TZ18, TY11, Tha19, Tre13, Tro17, UWK12, VY16, WFL12, WZ13a, WW10, WDL18, WZ13b, WZ17, Win15, WX13, Wu17, XZL10, Xu11, XZ15, Yam13, Yam16, YY10, YFK11, YZZ10]. Solutions [YCW10, YMYC10, Yos17, Yos18, YZ16, Zhu15, dAdM18, vBW11]. Solvability [AI12, DKR16, HI19, ST18]. Solvation [Li09, Li11]. Solvent [Li09, Li11, Gla17]. Some [AV19, BBS16, Chu14, CMM13, Cri19, Due16, GY16a, HS10a, Jia19, Koc16, Lau10, LR15b, LW12, LL18b, Lu13, MM11, MPT19, MS18b, MP12, Tre13, WZ17, TD17]. Sommerfeld [GN19, LL11]. Sonic [LMZZ17, LMZZ18, SWZ15, WX16, Wan19]. Sorting [BDFS18, CEH14]. Source [BFG^{+}13, BP10, GMT16, GM15, Kat19, LLM19, LL19, QWE19, TW11a]. Sources [BM12a, Blå18b]. Space [AC11, AN15, AP11, BC17a, BCP19, Can10a, Can10b, CD11, Daf13, DW13, FWW17, FR17, Gra19, GK10, KK17a, KLO16, LMTT15, LR15b, LX17c, Liu19, LS17, MM18a, Mas11, MW17, NP16, Pal14, SS17, SM19, SSZ19, VK18, WXY15a, WY13]. Space-Dependent [AN15]. Space-Time [DW13, Gra19, LS17, SS17, SSZ19]. Space-Times [MW17, LMTT15]. Spaces [AG17, AR19, BB10b, BCD17b, CWE10, CL13a, Com17, FV18, GP18, Kry10, LJ17, Mar18, PSSW15, Tak10, TD17, Wan13, XXK13, ZCO16, ZBL19]. Spacetimes [GL15]. Span [HI12]. Sparse [DS10a, GL12, HS13, KVM17, KLL12, LV13, PP19]. Spatial [Deu13, Hal12, Hal13, HL15, LW12, Löb18, PT18, PT11]. Spatially [GST13, LX19, Mit10, SSH19, SY14, YY18, dRDR16, dR18]. Spatiotemporally [ÁCDP14]. SPDEs [Kry10]. Species [ABR17, BDFS18, DY10, LTW14, NT14].

Specific [FW18b]. SPECT [FQ11]. Spectra [CNS10, EI11, dRDR16, dR18]. Spectral [AAK14, AH16, Ash13, BB17, BBV14, CGLS18, Fai14, GH18, KDT19, Nii12, PPP13, dR18]. Spectrum [AKKY17, AMV15, BM15b, GLT10, LYZ16, LWY11, Wu16, vBM14]. Specularly [NS13]. Speed [CQ19, DS10b, DMZ14, DLZ15, FG15, Ges13, Gia15, GMT16, Ito18, MO19, SY17]. Speeds [AFK^{+}18, DL15a, HS10a, XY14, YZ15, HR15]. Sphere [DSZ19, GM11, KW12]. Spheres [HM12a]. Spherical [CNR17, Hal14, KV19, Ngu15]. Spherically [DWYZ12, FL12a, LW14a, RR13]. Spike [GLW17]. Spikes [BWW14]. Spiky [WW10]. Spin [AG16, BCS15, CFO19, GG10, GL19a, KMM11, LNZ14, PWW17]. Spin-field [LNZ14]. Spin-Polarized [GG10]. Spin-Transfer [KMM11]. Spinodal [HMSZ13]. Spline [CCG18, Han18, SX13]. Splines [WU14a]. Splitting [CFSS18, GM17a, HKK13]. Spot [CK12]. Spreading [DL15a, DL10, DMZ14, HS10a, YZ15, DL13]. Spreading-Vanishing [DL10, DL13]. SQG [CIN18]. Square [BBCD^{+}18, Bru16]. Squared [KW12, ZCO15]. Squared-Distance [ZCO15]. Stability [ATSR19, BJ10, BMMP16, BMP18, BYZ12, BFN^{+}13, BHSZ10, Ben17, BdHQ13, BdHFS16, BF19, BDLM19, BFS14, CV15, CT15, CPP18, CM19, Cui13, DYZ19b, DY10, DL15b, EL17, ER19, FRX19, FKV15, Fel18, FHO16, GLS10, GLY18, GY16a, GW18, GSWZ18, HX14, HNS17, HSS17, HW17, HLWW18, HL15, HJ15, IN13, ILW16, JZN11, Joh13, KT17, KDT19, KK18, KP13, LWZ15, LMN^{+}10, LS13b, LW12, aLW18, lLWW18, LLLM14, LW15, LT11, LW14b, Mae17, MOZ10, MOZ12, Moa11, MRV19, Ngu10, NS13, NOS12, Ohn14, Ohn15, Ono11, Pan12b, Ria10, RR17, RZ16, Sin10, TWW15, Tej17, TW10, Ves15, Vis18, WY15, WW15, WW10, WX13, XY18, XYY19, YZ14b, ZH10, Zha19, dRDR16, dR18, vBM14, Ohn16]. Stabilization [CW13, Jia19, LW16a, Lau10]. Stabilizing [GW13, JJ18]. Stable [AHKM15, ADMR14, BCdSN18, CKS15, DV10, GP11, HO15, LMR13, MO19, Mar10, MRV12]. Stably [KS19]. Stacked [ILN11]. Stage [HS14]. Standing [BCdSN18, CKM14, CO12, FL12a, LNZ14]. Stark [Sac18]. Stars [Liu18b, SW17]. State [AHP13b, Bec18, BCL11, FV18, GG10, GH12, HWZ12, Kom15, Poh15, Sch18b, YMYC10, ZN19]. State-Dependent [GH12, HWZ12, YMYC10]. States [Bor19, BCT19, CF14, CCV15, CDNP16, CO12, Dro18, FY13a, GLW17, GP11, Kai17, KK18, LMR13, Miz11, MS16, Sch17, SW17, TWW15, dlHHMV16, vdBW19]. Static [AFT15, GS17, Lei13, Sab13, Tha19, BFM12]. Stationary [BP12a, Che18, CS14b, CJ19, CSZ18, Cui13, DKN11, Ess16, FG18a, GW15, GR13b, HW17, HMWY11, Kai17, Kut15, LWZ16, Löb18, MS18b, MS16, Rod16, ZZ19, vdBW19]. Statistical [BMR14]. Steady [BDX14, Ber12a, BW12, Bor19, CF14, CDX12, CKZ17, Dek19, EEW11, ER12, GG10, GJZ15, HZ19, Iye19, JH18, LMZZ17, LMZZ18, NP11, Poh15, Sch18b, SW17, TWW15, WBS13, WY15, Wen14, XX10, XY14, ZH10]. Steepest [Len16]. Stefan [BCQ12, HNS17, JS13b]. Stein [LLN19]. Stekloff [CLM17]. Steklov [BBG17, PPPV16]. Step [HS10b]. Stepping [BHRW16]. Sticky [GL17, Hyn19]. Stochastic [ABK12, ABBK16, AMW10, BBT14, BFFO17, BMY16, BM12b, BHR16, BGAHS17, BDT12, BGN14, BM19, CL17, CDZ13, DG11, DdMH15, DDMM18, ESvR12, Feh13, FQS10, FG15, FG18b, FL19, GS12a, GGAS14, Ges13, GR15a, GT16, GK10, Hal12, Hal13, Igb12, IT15, LMR15, LRdS18, MQS12, Mar18, MBK13, MZ13, MB16, NSS17, RZ11, SSZ19, YFK11, ZBL19, vNVW12]. Stokes [BV13, BC17c, GMT19, GHMZ10, IK11, JJN13, JLX15, Kim13, LMW17, TYZZ13, TWW15, VW15a, XZL10, ADL14, BT19, BFLN16, BM12b, BW12, Bre13, BMR14, CDLLSG13, CRWX16, CCK18, CJN19, CGP13, CS10b, CSZ18, Con12, CEIV17, Deu13, DF11, DWYZ12, EW18, FZZ18, FPVR13, GS15b, GX17, GP18, GW15, Höf18, HW17, HLW12, HW14, Hyn13, JS13a, JZ19, JWX13, Jün10, Kha13, Kim09, Kwe12, KK17b, LLW17, Lei16, LR15b, LS18, LST12, aLW18, Li19, LYZZ14, LY19, MOR^{+}16, NPS13, NN19, PZ11, Per10, PW15, QW11, RZ14, SS19, SS17, VY16, WXY15b, WLT16, WZ13b, WZ17, ZZ14, ZT17, vBW11]. Stokes-Like [BMR14]. Stokes [?]Abels:2014:WPF. Stokes/Allen [XZL10]. Stopping [ABØP19]. Storativity [CDR17]. Straight [Bos19]. Strain [Gin19, Yao19]. Strain-Gradient [Gin19]. Strained [GZ14]. Strains [Rou10]. Strategies [APSV19]. Strategy [Kat19]. Stratified [BC17a, Bel17, DYZ19b]. Strauss [MW17]. Streamlines [Hen10]. Stress [SV18]. Strichartz [Ign10, MS18a, Ovc11]. Striped [Lóp12]. Strong [Abe12, ABGS10, ABGS19, AP11, Bos19, CRWX16, CH13, DAP19, HY19, HW14, HT17, JJN13, KPR15, KP18b, Leq11, LMW17, LY19, LZ18, MPT19, Neu16, NPS18, PSZ19b, RR17, SS15, WX11, WX13, XY14, ZF12]. Strongly [BA10, BA12, EP12, GS15a, LPS13]. Structural [Vis18, WY15, ZH10]. Structure [BST17a, BDEM18, BGL16, GLZ17, Kut15, Leq11, LYZ16, Pin19, UWK12]. Structured [IM18, KNW15, MO15]. Structures [BC19, DVW15, LWZ18]. Study [Ber17]. Sturm [Pim16]. Style [Sch14b]. Sub [BC19, Igb17]. Sub-Hamilton [Igb17]. Sub-Riemannian [BC19]. Subdifferential [NN12]. Subdiffusion [VZ15]. Subdivision [Gro10]. Subject [KvMY19, LZZ15]. Subjected [Wol19]. Sublinear [AP18, SW18]. Submerged [KM13]. Submitted [ABR17]. Submonolayer [dCPS16]. Subsonic [BDX14, CDX12, Han14, HMSZ13, LW16a, LMZZ17, WX16, WX19, Wan19, Wen14, XX10]. Subsonic-Sonic [WX16, Wan19]. Suitable [WZ13a]. Sup [BCDG16]. Super [PP19]. Super-Resolution [PP19]. Superconducting [ABGS19, GS10a]. Superconductivity [AHP13b]. Superconductors [COS16, CP19, Pen17]. Supercritical [CCMW19, NPS13]. Superlinear [BEH15]. Supernova [BFG^{+}13]. Superposition [lLWW18, WW15]. Supersolutions [Zhi19]. Supersonic [CY15, CKZ17, HZ19, LMZZ18, WX15, WY15, ZH10]. Support [BCO17, CL18, Fis13]. Supported [HMZ15, KLL12, WU14a]. Suppression [BH17, BH18]. Sure [AMW10, NPS13]. Surface [ALS15, AET18, Bev11, CHW16, Che19, CTW13, CHS13a, DKT19, Duc10, EW15a, ERV17, FKM^{+}16, GLL17, GVWK16, GL12, Han18, Han14, HD17, Hen10, JTW16, KT17, LS13b, LZ19, LPS13, LX16, LX17c, Nes14, OR19, RTT19, SWX17, Wu14c, Xu18, YT11]. Surface-Internal [JTW16]. Surface-Waves-Type [SWX17]. Surfaces [CWE10, DG16, EH13, GP14, Kar16, LWZ11, Mon16, Ngu15, Yao19]. Surfactant [FKM^{+}16]. Surfactants [KT17]. Surgery [BM15b]. Surrounding [WX19]. Suspension [HM12a]. SVD [ADK15]. Sweat [LS12a]. Sweeping [AH18, MSZ19]. Swift [MBK13]. Swimmer [Kha13]. Swirl [DWX18]. Switching [BS16a, GCGJL18, HS19, LMR15, YMYC10]. Symmetric [ACM^{+}12, ABCD^{+}18, BWW14, BGL12, BCS15, DWYZ12, Ell12, FL12a, GM10, HZ10, KM18, LM14, LW14a, LW14b, RR13, RW14, vdBMJLM11, vdBW19]. Symmetries [BBO19]. Symmetrization [Des14]. Symmetry [BFN^{+}13, BM18, DP13, GL19a, GL15, HZ10, HD17, HM12b, MO14, Pol17, QS12, Tre13, HM13]. Symmetry-Breaking [HD17]. Symplectic [CCFdlL14, CdGDN18]. Synchronization [ST11]. Synchronized [Mos18]. Synchrosqueezed [YY14]. Synchrosqueezing [TW11b]. Synchrosqueezing-Based [TW11b]. System [ADL14, ALS15, AI12, AET18, ÁCDP14, ATSR19, AFT15, BT19, BDX14, BV13, BST17a, BPS19, BAH17, BACP18, BK15b, Bla18a, Bou13, BL15, BC17c, CS15a, CDLLSG13, CRWX16, CDM16, CJN19, CH13, CLW12, CFGL17, CL18, CMWZ18, CCHR18, CQX18, CFO19, CJ19, CO12, CGS17, CDK11, CFF19, DS14, Dai17, DJMZ16, DMZ19, DGV16, DN12, DYZ19a, DY10, Dua11, DL15b, Duc10, Faj16, FPP19, FKV15, FKM19, FT13, FPVR13, GM11, GMP13, GS10a, GMT19, GHMZ10, HKK15, HX10, HK15, HS19, HNP15, HLWW18, HW14, HT17, Hyn19, IK11, IT15, IKS12, JMNR11, JJN13, JZ18, KLS15, LMR13, LPR12, Leq11, LST12, LS12a, LL16b, lLWW18, LN10, LNZ14, Mas11, Mei10, MRT15, Ngu16, NTW19, Nov19, Oh10, Ohn16, PZ11, Pal14, PZ13, PW18, Pu13, RW14, RR15, SWZ15, SSW14, Tha19]. System [UWK12, VW15a, Wan11, WFL12, Wan12, WW10, Win15, WK17, WX13, WWX15, Wun10, XZ15, YY10, YZ18, YCW10, Zha19, ZLMZ18]. Systems [ABL13, AG16, AS13, BEH15, BA10, BA12, BGAHS17, Bor19, BCS15, Bra16, BHWY12, BMR14, BGL16, BLW19, CCFdlL14, CFSS18, CCLCP13, Cer11, CL17, CHL15, CQ19, Com17, Con12, CPP18, CN15, CG10c, DP13, DZ14, DFT17, DNK12, DL15a, DK11, DX19, EW15a, EKR18, FGW13, Feh13, FS15, FPZ14, GLS10, GLY18, GPPP13, GYY18, GG10, GS15b, GX17, HKN16, HHMM18, HL15, HV13, ILN11, Jia19, JS13b, LW12, LYZ16, LZZ17, LPS13, Lu13, MM17, MS18b, MSZ13, MSTY16, MB16, MS16, Mor19, Pan12b, Pen15, QS12, SWX17, SSH19, Sch18b, SM19, SM16, Sus13, TW18a, YZ15, YMYC10, Zhi19, LW16b]. Szego [Thi19].

Tail [BCS15, MBPS13]. Tailed [BGHP18]. Tails [AL10, CS15a, MV19]. Tangent [GZ13]. Tangential [Pao16]. Tartar [BRS17]. Tataru [DW13]. Taxis [Zhi19]. Taylor [GT10, JJ18]. Technique [Val15]. Teeth [CHKP19]. Temperature [BV13, DDF18, JK10, LYZZ14, MMB11]. Tempered [ACM^{+}12]. Tending [FY13a]. Tension [CTW13, CHS13a, DKT19, Nes14, SZ12b, YT11]. Tensor [BCS15, CRWX16, DS10a, GGRB14, HMZ15, HS13, KM18, Mil18, PZ11, Win15, Zho18, ADL14, Dai17]. Tensor-Valued [Win15]. Tensors [Mil18, Yao19]. Teramoto [Kut15]. Term [AF16, AI12, Bev11, BP10, Jia12, Olb19, WWX15, ADHZ15]. Terminal [NKV19]. Terms [BT16, GVZ16, GMT16, QWE19]. Ternary [RW14]. Terraces [Pol17]. Textile [LS10]. their [ZK15]. Theorem [CLW17, Kry13, Kry14, DT15, PX13]. Theorems [AT10, CL13a, LX16, RZ11, TD17, Xu18]. Theoretical [Ber17, GO18, SCB17, SCB20, ZCO16]. Theories [Yao19]. Theory [ABL13, AC14, AMP10, AS15, AZ17, BAP13, BB17, BBG12, BBS16, CGLS18, DS10b, Fai14, FR17, FW18b, GR15b, GX17, HM12b, HM13, JT13, Lee10, LW16b, MT13, NT18, DT15, ST17, SdlL13, WZZ15, dlLSZ17]. Thermal [Con12]. Thermo [RR17, SY17]. Thermo- [SY17]. Thermodynamic [HKR18]. Thermodynamically [RR15]. Thermodynamics [Rou10]. Thermomechanical [DKR16]. Thermoviscoelastic [PZ13]. Theta [Bét16]. Thick [LS15]. Thin [ABGS10, AVP16, BFLS12, BGHP18, BP12b, CNS10, CDN10, CDLLSG13, CPT10, CM12, CKV18, DFP14a, DFP14b, Ess16, FG18b, FHO16, Gna15, Har18, KK16, LM17, Mel10, MS14, MO14, RP18, RZ16, Tha19, Yao19]. Thin-Film [CKV18, Ess16, FG18b, Gna15, Mel10]. Thin-Walled [DFP14a, DFP14b]. Third [AI12]. Three [BFN^{+}13, BV10, BD18, BMR14, CP19, CY18, DS10b, DWX18, GMP13, GR13a, HGW14, Hof12, HHPZ17, HW13a, HW13b, HLX11, Kai10, Lei16, LS12a, LLW15, Mar10, MOR^{+}16, Nov19, RV12, WY15, WXY15b, WY13, WZ17, XZ15, ZH10, vdBW19, LY19]. Three-Dimensional [BFN^{+}13, BV10, BMR14, CP19, DS10b, DWX18, GR13a, Hof12, HHPZ17, HW13a, HW13b, HLX11, Kai10, Lei16, LS12a, Mar10, MOR^{+}16, Nov19, RV12, WY15, WXY15b, WY13, WZ17, XZ15, ZH10]. Three-Wave [HGW14]. Threshold [IT15]. Thresholds [CHL17]. Thrombus [WNRJ13]. Tight [HMZ15, WX12]. Time [AC14, BJ10, BM12a, BMP18, BK13, BGL12, Bre13, BM18, CHL19, CDW13, CMP13, CM14, CHL15, CV15, Cho16, CH19a, CKS15, CG19, CG11, CdGDN18, CN15, Daf13, DDF18, DF10, Deu13, DK14, DW13, EW19, EF15, FGW13, FS14, FS15, GIP^{+}13, GM10, Gia15, GS10b, GP15, Gra19, GW13, GGRB14, Hal12, HMW11, IS13, IM18, Kia16, KS19, Kry14, LR15b, LMW17, LL18b, LRdS18, LS17, MM17, MOZ10, MOZ12, MSZ13, Mos18, NT19, QW11, SWX17, SS17, ST10, SSZ19, Spe14, SCB17, SCB20, ST15b, Sug16, ST18, TW15, TY11, UWK12, VW11, VZ15, WWW12, Wan12, XY18, YZ14b, vBW11, HMWY12]. Time-Delay [CN15]. Time-Delayed [MOZ10, ST15b, MOZ12]. Time-Dependent [BJ10, BMP18, CKS15, CG19, Deu13, GIP^{+}13, GS10b, GP15, Kia16, KS19, SCB17, NT19, SCB20]. Time-Discrete [CMP13]. Time-Fractional [LRdS18, VZ15].

Time-Frequency [CdGDN18]. Time-Harmonic [Spe14, BM12a]. Time-Local [ST18]. Time-Periodic [Bre13, EF15, IM18, LR15b, TY11]. Time-Recurrent [FGW13]. Times [BF19, FG15, IT15, INRZ10, MW17, dHHI^{+}14, LMTT15]. Timoshenko [ATSR19]. Tissue [CCM16, dAdM18]. Tissues [JMNR11]. Tokamak [HK10]. Tomography [BCS16, ES10, HHR09, HHR11, HS10b, HU13, Ilm16, KT13, Liu18a, PUW18, SY17]. Tool [IISD15]. Töplitz [PX13]. Torque [KMM11]. Torrey [GH18]. Torsion [Bel10, BFLS12]. Torsional [BGM19]. Tortorelli [FI14, FL17]. Torus [Wu14b]. Touch [CKV18]. Touch-Down [CKV18]. Touchdown [GS15c]. Trace [AT10, BCD17b, LT11, EH16, TD17]. Trace-Free [BCD17b]. Traces [KMS15, NPS18]. Tracking [KR10]. Traffic [BH11, CG10b, GM14, HR19, KPS18, LMP11]. Trajectory [BMR14]. Transfer [AH13, KLS11, KMM11]. Transform [AAK14, ADK15, ABDD19, BK15b, CNR17, FQ16, GR13b, KM18, Mon16, Moo16, Ngu15, PW15, ST15a, YY14, Zho18]. Transformation [Ngu13, NT19, dRDR16]. Transient [BMSR^{+}13, GLZ17, MO15]. Transition [CKY18, DLZ15, FL12b, HMSZ13, LSW17, HR15]. Transitions [BKLU18, BBG12, BCQ12, CFO19, FL12a, GPPP13, RR15]. Translating [DKN11, GW18, YT11]. Translation [GL19a]. Translation-Invariant [GL19a]. Transmission [Blå18b, CGH10, CCG10, CCH10, CH11b, Fai14, GVZ16, HKOP10, HKOP11, LV12, LV15, LWZ18, LR15a, RR17, Syl12]. Transmitter [AFK^{+}18]. Transonic [HCHY16, LW14a, WX16]. Transport [AMP10, BJ10, BM12a, BAP13, BCS16, Ber12a, BFG^{+}13, BCL18, Bon13, Bos12, BW17, BBS11, CGS10, CDPS17, Cav12, Cha14, CDM13, CCG18, CS17, CDS19, CHS13b, FV18, Fri19, GKR19, GS10b, GY16a, GY16b, GJMC12, HKK17, JK10, LS12a, LMS16, LR17, LYZZ14, MMT19, Ovc11, SM16, XV10, ZZ19, Che15]. Transportation [GM13, KW12, Lee10, Pas11, Pas13]. Transverse [JZ10]. Trapped [NT13]. Trapping [Can10b, CM19]. Travel [dHHI^{+}14]. Traveling [Ai10, BF19, CQ19, CHS13b, FZ14, GLS10, HSS17, HV13, HJ15, Ito18, JZ10, LWZ18, LW12, LLLM14, MOZ10, MOZ12, MN12, NW17, Nol11, Ohn16, SSH19, Tan15, TV18, YZ15, dL14]. Travelling [Hen10, HR10]. Treatment [CCC^{+}17]. Tree [Ign10]. Trees [FK13]. Trend [DFT17, KKT17]. Triangular [Syl12]. Tridiagonal [FGW13]. Triebel [Tak10]. Triple [AMW11]. Trudinger [LL12]. Truncated [AAK14, ADK15]. Tsingou [FW18a]. Tug [MPR10]. Tug-of-War [MPR10]. Tumor [Cui13, EG19]. Tumors [dAdM18]. Turbulence [CS14a, KT11, NT18]. Turbulent [XY14]. Turing [AS13]. Turning [LWZ15]. Twisted [FHMP16]. Two [Abe12, AAGP18, ALP15, BDX14, BYZ12, BFRV13, BFN^{+}13, BC19, BHSZ10, Bét16, BBG16, BW12, BCJ20, BDFS18, CNS10, Can10a, Can10b, CP12, CS18, CRWX16, CQ12, CFGL17, CKZ17, CKY18, CDX17, Cho16, CNR17, CDNP16, CFO19, CQW18, CWYZ16, DG11, DLZ12b, Duc10, Evj13a, Evj13b, EW18, FPP19, FL19, GS12b, GH18, GW15, GLW17, HNS17, HL12, HJ11, HS14, HW14, IY12, JMWZ14, JRK19, Kai10, Lac15, Lai18, aLW18, lLWW18, LW16b, LS15, LPS13, LTW14, Mas11, Mil18, MMP13, NT14, NW17, Ono11, PSZ19b, Pin19, QWE19, ST15a, Ter11, WXY15a, WW15, XY14, XXK13, YY14, YZZ10, YZZ12, ZZ14, DZ15]. Two-Component [PSZ19b, JRK19]. Two-Dimensional [BDX14, BFN^{+}13, Bét16, BCJ20, CNS10, CQ12, CKZ17, CFO19, CQW18, FPP19, GS12b, GW15, GLW17, HJ11, JMWZ14, Kai10, Lac15, Lai18, aLW18, LS15, LPS13, Mas11, Mil18, MMP13, NW17, ST15a, XY14, YY14, ZZ14, DZ15]. Two-Fluid [CWYZ16, Duc10, EW18, XXK13]. Two-fluids [DLZ12b]. Two-Phase [ALP15, Can10a, Can10b, CP12, Cho16, Evj13a, Evj13b, HNS17, HL12, QWE19, Ter11, YZZ10, YZZ12, Abe12, JRK19]. Two-Sided [BHSZ10]. Two-Species [LTW14]. Two-Stage [HS14]. Two-Weight [CNR17]. Type [Ale16, AG16, AM15, AS13, BL19, BDG13, BRS17, BNDHV10, BC17b, BM19, CDM13, COS16, DF13, DdMH15, DSZ19, DM15, EI11, FMP18, GH14, HR10, HLX11, Joh13, KNR12, Kot12, Lei13, Len14, LZ17, PZ13, PZ17, SSST15, SWX17, SSH19, SM16, UWK12, VW15b, Yos17, ZLMZ18, dMIS10, Dan17, KPZ19, OSW19, BdHQ13, GSV19, FPTT12, ZBL19]. Type-I [COS16].

Ulam [FW18a, Miz11]. Ultrasound [BCS16]. Unbounded [BAH17, CWE10, CTW13, EH13, GLZ17, JJN13, LWZ11, MPT19, Pim16]. Uncertainty [JZ18]. Unconditional [HF13, HTX15]. Underlying [MV19]. Unequal [EW15b]. Unfolding [AVP16, CDD^{+}12, GP14]. Uniaxial [HM12b, HM13]. Unified [BBS16]. Uniform [AHP13a, AF15, ATSR19, AGN19, Bec18, Bia18, CL18, Dua11, GRT14, GS15b, HX14, JZ18, Lee17, MRS16, WXY15b]. Uniformly [CEIV17, Pen15]. Unifying [FG15]. Unilateral [Ber17]. Unipolar [HMWY11]. Unique [LN10, Otw10, SW18]. Uniqueness [BEH15, BFGPE^{+}18, BT16, BC19, BCL11, CHN18, CG19, CMM13, Daf19, DDM11, FKM^{+}16, GP19, GS10a, GMT19, GGRB14, GLW17, HF13, HKK17, HTX15, INSZ14, JJN13, KZ11, KvMY19, Lac15, LTV17, LT17, LW15, LT11, LY19, LZ18, Pas11, Poh15, YMYC10]. Unit [Sim16]. Unitarization [ABDD19]. Unitary [SX13]. Universal [Hal14, HS10c, Zha10]. Unknown [LLM19, Sin10, Ves15]. Unsaturated [DKR16]. Unstable [BJ16, BG14, CPT10]. Unsteady [BR17, BGMŚG12, DDGVM18, LWY18]. Unwinding [CStW17]. Upper [BOS11, Syl12]. Using [CM14, GL12, GR13b, KS14, KR10, KLS11, Ngu13, VW15a].

Vacuum

[DWYZ12, DLZ12a, HHPZ17, HLW12, HW14, JWX13, Lee17, LXZ13, LZZ15, MY17, Per10, Str10, WZ13b, WZ17, Zhu15]. Validation [LX17c]. Validity [DKS16, Mor19]. Value [AI12, Beb16, BMY16, BdHQ13, BdHFS16, CDN10, CHN18, IY12, KMV18, KT11, Len16, MPR10, NNS18, NP16, NKV19, Otw10, RSZ18, TW18a, Ves15]. Valued [AAD13, BGT19, CCG18, EHM16, LTV17, Mit10, Win15, YZ18, ZK15]. Values [GMM13]. Vanish [Ber12a, Blå18b]. Vanishing [BHWY12, CH13, CT11, CG10b, DL10, GS19, JJL10, Nov18, Rou13, SS19, WXY15b, DL13]. Variable [AL10, BGL16, BC17c, BO16, CDL16, CHK15, DLV10, Kry14, SY17]. Variables [CH15, NV12, WXY15a]. Variably [DK11]. Variant [CLW17]. Variation [AF16, KMS15, Mit10]. Variational [AH18, AF15, ALM10, Bar14, BS16b, BMC14, BMC18, BMY16, BP12a, BDSS18, BOO18, CWE10, CFO19, FG18a, FHMP16, HMS14, HS16, JHN12, KR10, LLN19, MPT18, MS18b, MN12, NP11, SV11, SV19, SS17]. Variational-Hemivariational [BS16b, HMS14]. Variations [AAS19, KY12b]. Varifolds [De 18]. Varying [LJ17, Muñ12, WK17]. Vasculogenesis [DS13]. Vector [AAD13, BB10a, BBG16, MM18a, Mon16, YZ18]. Vector-Valued [AAD13, YZ18]. Vehicular [HR19]. Velocities [AAGP18, Cha14, EHM16, EW15b, MZ13]. Velocity [Ber12a, Bru16, CEIV17, CW13, Deu13, KUV16, WWX15]. Ventcel [BNDHV10]. Version [Bes12, HLR^{+}19]. versus [LWZ16, Muñ12, WWW12]. Vertex [Koc16]. Very [Bla18a, HR12]. Vesicle [WX13]. Vesicles [HD17]. Via [BH17, BH18, BFFO17, BMC18, AZ12, BM10, BRS17, BLW19, CM11, CP19, CN15, DP13, FF12, GS18, HSV17, Ito18, KY12b, KP18a, KDT19, KW11, LV13, LR17, LS13c, MMT19, MOS14, NV12, Ngu17, NT19, PSV10, SV11, Syl12, VZ15, dRDR16]. Vibrating [DGV16, RP18]. Vibro [Pao15, Pao16]. Vibro-Impact [Pao15, Pao16]. Vicinal [GLL17]. View [ALM10]. Viewpoint [CCG18]. Vineyard [Ai10]. Violent [Ben17]. Visco [HS18, Rou13]. Visco-elastic [Rou13]. Viscoelastic [ATSR19, BP14a, Chu14, DAP19, HX10, HW13a, Hyn13, LLW17, LMMNR17, Nes14, ZF12]. Viscoelasticity [DNK12, FK18, MOS14, RR17]. Viscoplasticity [NN12]. Viscosities [LT19]. Viscosity [BC19, BM12b, BHWY12, BC17c, CH13, CT11, CG10b, Gie14, GS19, HM12a, JJL10, JWX13, RTZ17, Rou13, SS19, WXY15b, Yos17, Yos18, ZF10]. Viscous [Abe12, BFN^{+}13, BOS11, CV15, CGS17, DWY12, DKN11, EW15b, Feh13, FPZ14, GT10, GJZ15, HL12, HLWW18, HLX11, JTW16, JZN11, Kwe12, KK17b, Le 13, Lee17, LZZ15, LT17, lLWW18, LW17, MS13a, MY12, Ngu10, QWE19, RTT19, Rou10, Smi17, TW18a, WW15, WZ13b, Wu14c, YZZ10, YZZ12, Yos18]. Vlasov [GHMZ10, LMW17, Tha19, ASC19, AFT15, BAH17, BACP18, Bes12, Bos19, CL18, CH11a, DGV16, Des14, DYZ19a, DY10, Dua11, DL15b, Faj16, GPI18, HKN16, Höf18, JZ18, KKT17, LMR13, LYZ16, lLWW18, Pal14, SR14, Wan11, Wan12, YY10, Zha19]. Voigt [DAP19, PZ13, RR17]. Volterra [AP18]. Volume [CP10, CP11, CL13b, DMZ19, GO18, WWW12]. Volume-Filling [DMZ19, WWW12]. Volume-Fraction [CP11]. Vortex [CP19, Dek19, DWZ10, KMM11, KMS17, NN19, WY13]. Vortex-Wave [NN19]. Vortices [YZ14a]. Vorticity [CNSS17, Hen10, Whe13]. Vries [CG10a, EW19, LP19]. VSC [HNP13].

Waals [CS18, JMZ18, Lai18, LM11]. Waele [Yos17]. Waele-Type [Yos17]. Waiting [FG15, Gia15]. Walk [MV19]. Wall [Car14, BW12]. Walled [DFP14a, DFP14b]. Walls [CKZ17, CPP18]. Wang [LL12]. Wannier [Sac18]. War [MPR10]. Wasserstein [AC11, BB10b, BCP19, KK17a, KP18a, MPS17, Tak13]. Water [AAGP18, Che12, CQX18, CGLS18, DS10b, DLZ12a, Duc10, EEW11, Hen10, HT18, KM13, Le19, MRT15, Ngu16, WBS13, Whe13]. Water-waves [MRT15]. Wave [BPS19, BYH15, CCCdlL17, CDZ13, CQ19, CPT10, DO16, DAP19, DF11, DL15b, FRX19, FK19b, FPZ14, FMP18, GP19, HL11, HGW14, HZ19, HLW12, HT18, Ito18, JTW16, KM13, Kia16, KT11, LMTT15, aLW18, lLWW18, LPS13, LPS10, MO19, MQS12, MZZ12, NV12, NT18, NN19, Qin15, RTT19, Sug16, Wu14c, ZH10, dHGR14]. Wave-Heat [BPS19]. Wave-Long [DF11, FPZ14, FMP18]. Wavefronts [CKM14]. Waveguide [CKS15, MMP13]. Wavelength [LP16]. Wavelet [AHKM15, DS10a, DSX17, HR12]. Wavenumber [Spe14]. Wavenumber-Explicit [Spe14]. Waves [Ai10, AZ17, ALZ19, BF19, BCdSN18, BC17b, CFGL17, CJ19, CO12, CGLS18, CHS13b, DS10b, Duc10, DIT15, EEW11, ETZ13, EL17, FL12a, FZ14, FW18a, GSW16, GLS10, HSS17, Han14, Hen10, HR10, HMSZ13, HMWY11, HV13, HJ15, JWX13, JZ10, JZN11, Joh13, Jun14, KT17, KDT19, KK10, Lai18, LWZ18, LP19, Le19, Le 13, LWY18, lLWW18, LLLM14, LW14a, LW17, LP14, LNZ14, MMB11, MOZ10, MOZ12, MZZ12, MN12, NT13, Ngu16, NT18, NW17, Nol11, Ohn14, Ohn16, Per10, RZ17, RZ16, SWX17, SSH19, TV18, WBS13, WW15, Whe13, XYY19, YZ15, Yos18, dL14, MRT15]. Wavetrains [WW18]. Weak [BPS16, BR17, BFFO17, BFLS18, Ben17, Bla18a, BN14, CMWZ18, CS14b, EL17, Evj11, FZ14, FG18a, GLL17, GZ18, GGRB14, HK15, Hyn13, JJN13, Jia19, Jün10, KK17a, KMT13, KY15, KK15, KNR12, Len14, LS18, LS10, LS12a, LT17, LL18b, LLW15, LT19, Lu13, LZ18, MQS12, MP13, MPZ15, NPS13, NP16, NTW19, Nov19, NP11, Pan12a, Ped15, Sab13, Smi17, SSW14, TZ18, VY16, WZ13a, YZZ10]. Weak* [MY17]. Weak-Strong [JJN13, LZ18]. Weakly [BCL11, Bor19, BP14a, CDS10, DZ14, GL15, MSTY16, Pes15]. Weidl [Kre19]. Weight [CNR17, Pes15]. Weighted [BDWZ12, CWE10, CL13a, FL17, GPT19]. Weights [AR19, CH19a, Kry10]. Welds [IO16]. Well [Abe12, ADL14, AET18, ALST14, AN15, BG17, BFTT18, BTZ15, BFS14, CM11, CWH18, CGP13, CY19, CHS13a, DFV18, DZ15, Evj13a, Faj16, GV19, GM10, Gna15, GW18, HNS17, HL12, HX10, HY14, HKK13, HTW18, IKS12, JTW16, JL19, KM17, KMVW14, LS13b, LPR12, Li19, LS15, LPS13, Mar18, MSZ19, NN12, RV12, TW18a, Tsu12, WXY15a, WLT16, Wan18, Wu14c, XXK13, YZ18, ZZ14, ZT17]. Well-Posed [CM11, CWH18, HKK13]. Well-Posedness [ADL14, ALST14, AN15, BFTT18, BTZ15, BFS14, CGP13, CY19, CHS13a, Faj16, GV19, GM10, Gna15, HNS17, HL12, HX10, HY14, IKS12, JTW16, JL19, KM17, LS13b, LPR12, Li19, LPS13, Mar18, MSZ19, NN12, RV12, TW18a, Tsu12, WXY15a, WLT16, Wan18, Wu14c, XXK13, YZ18, ZZ14, Abe12, DZ15, GW18, HTW18, KMVW14]. Well-Reservoir [Evj13a]. Wells [Evj11]. Wetting [Lóp12]. Weyl [LV15]. Where [Ber12a, TZ15]. Whitham [DKS16]. Whole [GK10, LR15b, Mas11, Moa11]. Whose [Rod16]. Wiener [YFK11]. Wigner [BBCD^{+}18, LS17]. Willmore [CL13b, KL18b, Olb19]. Wilson [BJLO17, DN12]. Wind [WBS13]. Wing [CY15]. Wireless [PT11]. Without [AGN19, Kry10, TAGP18, CLW17, Kry13, Nes14]. WKB [FK19a]. Wolbachia [APSV19]. World [DKR15]. Wounds [FHX10]. Wright [CS10a, EM10]. Wulff [Neu16].

X [Mon16, PUW18]. X-Ray [Mon16, PUW18]. Xin [Bia18].

Young [BKK18, BMC18, BK15a, Ito18, KRW15, NP16]. Yudovich [DT15].

Zakharov [Com17, NT18, RV12]. Zaremba [Naz12]. Zealotry [PLPSS18]. Zeldovich [Lai14]. Zero [BM12b, CHS13a, Daf19, DLZ15, FY13a, GS15c, HLW12, Lee17]. Zeros [Sim16].

References

Alikakos:2013:PAC

[AAD13] Nicholas D. Alikakos, Panagiotis Antonopoulos, and Apostolos Damialis. Plateau angle conditions for the vectorvalued Allen–Cahn equation. SIAM Journal on Mathematical Analysis, 45(6):3823– 3837, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aguillon:2018:ARP

[AAGP18] Nina Aguillon, Emmanuel Audusse, Edwige Godlewski, and Martin Parisot. Analysis of the Riemann problem for a shallow water model with two velocities. SIAM Journal on Mathematical Analysis, 50(5): 4861–4888, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Al-Aifari:2014:SAT

[AAK14] Reema Al-Aifari and Alexander Katsevich. Spectral analysis of the truncated Hilbert transform with overlap. SIAM Journal on Mathematical Analysis, 46(1):192– 213, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antonelli:2019:RNS

[AAS19] Paolo Antonelli, Jack Arbunich, and Christof Sparber. Regularizing nonlinear Schrödinger equations through partial off-axis variations. SIAM Journal on Mathematical Analysis, 51(1): 110–130, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aksoylu:2010:RDE

[AB10] Burak Aksoylu and Horst R. Beyer. Results on the diffusion equation with rough coefficients. SIAM Journal on Mathematical Analysis, 42(1): 406–426, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antonopoulou:2016:MDS

[ABBK16] D. C. Antonopoulou, P. W. Bates, D. Blömker, and G. D. Karali. Motion of a droplet for the stochastic mass-conserving Allen–Cahn equation. SIAM Journal on Mathematical Analysis, 48(1): 670–708, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Azaiez:2018:NAP

[ABCD^{+}18] M. Azaïez, F. Ben Belgacem, J. Casado-Díaz, T. Chacón Rebollo, and F. Murat. A new algorithm of proper generalized decomposition for parametric symmetric elliptic problems. SIAM Journal on Mathematical Analysis, 50(5): 5426–5445, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alonso:2018:ODD

[ABCL18] Ricardo Alonso, Véronique Bagland, Yingda Cheng, and Bertrand Lods. Onedimensional dissipative Boltzmann equation: Measure solutions, cooling rate, and selfsimilar profile. SIAM Journal on Mathematical Analysis, 50(1):1278–1321, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alberti:2019:UIF

[ABDD19] Giovanni S. Alberti, Francesca Bartolucci, Filippo De Mari, and Ernesto De Vito. Unitarization and inversion formulae for the Radon transform between dual pairs. SIAM Journal on Mathematical Analysis, 51(6):4356– 4381, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abels:2012:SWP

[Abe12] Helmut Abels. Strong wellposedness of a diffuse interface model for a viscous, quasi-incompressible two-phase flow. SIAM Journal on Mathematical Analysis, 44 (1):316–340, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p316_s1.

Alama:2010:TFL

[ABGS10] Stan Alama, Lia Bronsard, and Bernardo Galvão-Sousa. Thin film limits for Ginzburg– Landau with strong applied magnetic fields. SIAM Journal on Mathematical Analysis, 42(1):97–124, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almog:2019:ESS

[ABGS19] Yaniv Almog, Leonid Berlyand, Dmitry Golovaty, and Itai Shafrir. Existence of superconducting solutions for a reduced Ginzburg–Landau model in the presence of strong electric currents. SIAM Journal on Mathematical Analysis, 51(2):873–912, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antonopoulou:2012:FMO

[ABK12] D. C. Antonopoulou, D. Blömker, and G. D. Karali. Front motion in the one-dimensional stochastic Cahn–Hilliard equation. SIAM Journal on Mathematical Analysis, 44(5): 3242–3280, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Acquistapace:2013:TIH

[ABL13] Paolo Acquistapace, Francesca Bucci, and Irena Lasiecka. A theory of the infinite horizon LQ-problem for composite systems of PDEs with boundary control. SIAM Journal on Mathematical Analysis, 45(3): 1825–1870, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Agram:2019:SCO

[ABØP19] Nacira Agram, Achref Bachouch, Bernt Øksendal, and Frank Proske. Singular control optimal stopping of memory mean-field processes. SIAM Journal on Mathematical Analysis, 51(1):450– 468, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alfaro:2017:ECS

[ABR17] Matthieu Alfaro, Henri Berestycki, and Gaël Raoul. The effect of climate shift on a species submitted to dispersion, evolution, growth, and nonlocal competition. SIAM Journal on Mathematical Analysis, 49(1):562– 596, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Agueh:2011:BWS

[AC11] Martial Agueh and Guillaume Carlier. Barycenters in the Wasserstein space. SIAM Journal on Mathematical Analysis, 43(2):904– 924, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p904_s1.

Alberti:2014:ERT

[AC14] Giovanni S. Alberti and Yves Capdeboscq. Elliptic regularity theory applied to time harmonic anisotropic Maxwell’s equations with less than Lipschitz complex coefficients. SIAM Journal on Mathematical Analysis, 46(1):998– 1016, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alvarez-Caudevilla:2014:QAC

[ÁCDP14] Pablo Álvarez-Caudevilla, Yihong Du, and Rui Peng. Qualitative analysis of a cooperative reaction–diffusion system in a spatiotemporally degenerate environment. SIAM Journal on Mathematical Analysis, 46(1):499–531, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alibaud:2012:CDE

[ACJ12] Nathaël Alibaud, Simone Cifani, and Espen R. Jakobsen. Continuous dependence estimates for nonlinear fractional convection-diffusion equations. SIAM Journal on Mathematical Analysis, 44(2): 603–632, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i2/p603_s1.

Andreu:2012:RSS

[ACM^{+}12] F. Andreu, V. Caselles, J. M.

Mazón, J. Soler, and M. Verbeni. Radially symmetric solutions of a tempered diffusion equation. A porous media, flux-limited case. SIAM Journal on Mathematical Analysis, 44(2):1019–1049, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2014:CHS

[ACZ14] Habib Ammari, Yat Tin Chow, and Jun Zou. The concept of heterogeneous scattering coefficients and its application in inverse medium scattering. SIAM Journal on Mathematical Analysis, 46(4): 2905–2935, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Armstrong:2015:LTB

[ADHZ15] Seth Armstrong, Sarah Duffin, Jianlong Han, and Chunlei Zhang. Long-term behavior and numerical analysis of a nonlocal evolution equation with Kac potential. SIAM Journal on Mathematical Analysis, 47(2):1234– 1252, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alaifari:2015:AAS

[ADK15] Rima Alaifari, Michel Defrise, and Alexander Katsevich. Asymptotic analysis of the SVD for the truncated Hilbert transform with overlap. SIAM Journal on Mathematical Analysis, 47(1): 797–824, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abels:2014:WPF

[ADL14] Helmut Abels, Georg Dolzmann, and YuNing Liu. Wellposedness of a fully coupled Navier–Stokes/Q-tensor system with inhomogeneous boundary data. SIAM Journal on Mathematical Analysis, 46(4):3050–3077, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alessandrini:2014:SDI

[ADMR14] Giovanni Alessandrini, Michele Di Cristo, Antonino Morassi, and Edi Rosset. Stable determination of an inclusion in an elastic body by boundary measurements. SIAM Journal on Mathematical Analysis, 46(4):2692–2729, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alphonse:2018:CLR

[AET18] Amal Alphonse, Charles M. Elliott, and Joana Terra. A coupled ligand-receptor bulksurface system on a moving domain: Well posedness, regularity, and convergence to equilibrium. SIAM Journal on Mathematical Analysis, 50(2): 1544–1592, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alvarez:2015:EAV

[AF15] Felipe Alvarez and Salvador Flores. Existence and approximation for variational problems under uniform constraints on the gradient by power penalty. SIAM Journal on Mathematical Analysis, 47(5):3466–3487, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abatangelo:2016:LTE

[AF16] Laura Abatangelo and Veronica Felli. On the leading term of the eigenvalue variation for Aharonov–Bohm operators with a moving pole. SIAM Journal on Mathematical Analysis, 48(4):2843– 2868, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2017:BCP

[AF17] Zied Ammari and Marco Falconi. Bohr’s correspondence principle for the renormalized Nelson model. SIAM Journal on Mathematical Analysis, 49(6):5031–5095, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ambartsoumian:2018:SFS

[AFK^{+}18] G. Ambartsoumian, R. Felea, V. P. Krishnan, C. J. Nolan, and E. T. Quinto. Singular FIOs in SAR imaging, II: Transmitter and receiver at different speeds. SIAM Journal on Mathematical Analysis, 50(1):591–621, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andreasson:2015:SSE

[AFT15] Håkan Andréasson, David Fajman, and Maximilian Thaller. Static solutions to the Einstein–Vlasov system with a nonvanishing cosmological constant. SIAM Journal on Mathematical Analysis, 47(4):2657–2688, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alicandro:2016:LNC

[AG16] Roberto Alicandro and Maria Stella Gelli. Local and nonlocal continuum limits of Isingtype energies for spin systems. SIAM Journal on Mathematical Analysis, 48(2): 895–931, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alaifari:2017:PRG

[AG17] Rima Alaifari and Philipp Grohs. Phase retrieval in the general setting of continuous frames for Banach spaces. SIAM Journal on Mathematical Analysis, 49(3): 1895–1911, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ancona:2019:KEC

[AGN19] Fabio Ancona, Olivier Glass, and Khai T. Nguyen. On Kolmogorov entropy compactness estimates for scalar conservation laws without uniform convexity. SIAM Journal on Mathematical Analysis, 51(4): 3020–3051, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2013:PDR

[AGS13] Habib Ammari, Josselin Garnier, and Knut Sølna. Partial data resolving power of conductivity imaging from boundary measurements. SIAM Journal on Mathematical Analysis, 45(3):1704–1722, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Allaire:2013:HCC

[AH13] Grégoire Allaire and Zakaria Habibi. Homogenization of a conductive, convective, and radiative heat transfer problem in a heterogeneous domain. SIAM Journal on Mathematical Analysis, 45(3): 1136–1178, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almog:2016:SAC

[AH16] Yaniv Almog and Raphaël Henry. Spectral analysis of a complex Schrödinger operator in the semiclassical limit. SIAM Journal on Mathematical Analysis, 48(4): 2962–2993, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adly:2018:ISP

[AH18] S. Adly and T. Haddad. An implicit sweeping process approach to quasistatic evolution variational inequalities. SIAM Journal on Mathematical Analysis, 50(1):761– 778, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adcock:2015:LSS

[AHKM15] Ben Adcock, Anders C. Hansen, Gitta Kutyniok, and Jackie Ma. Linear stable sampling rate: Optimality of 2D wavelet reconstructions from Fourier measurements. SIAM Journal on Mathematical Analysis, 47(2): 1196–1233, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Agram:2013:MPI

[AHØP13] N. Agram, S. Haadem, B. Øksendal, and F. Proske. A maximum principle for infinite horizon delay equations. SIAM Journal on Mathematical Analysis, 45(4): 2499–2522, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adcock:2013:BCR

[AHP13a] Ben Adcock, Anders C.

Hansen, and Clarice Poon. Beyond consistent reconstructions: Optimality and sharp bounds for generalized sampling, and application to the uniform resampling problem. SIAM Journal on Mathematical Analysis, 45(5):3132– 3167, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almog:2013:SNN

[AHP13b] Yaniv Almog, Bernard Helffer, and Xing-Bin Pan. Superconductivity near the normal state in a half-plane under the action of a perpendicular electric current and an induced magnetic field, Part II: The large conductivity limit. SIAM Journal on Mathematical Analysis, 44(6): 3671–3733, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ai:2010:TWM

[Ai10]  Shangbing Ai.  Traveling

waves for a model of a fungal disease over a vineyard. SIAM Journal on Mathematical Analysis, 42(2): 833–856, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aiki:2012:SIB

[AI12] Masashi Aiki and Tatsuo Iguchi. Solvability of an initial-boundary value problem for a second order parabolic system with a third order dispersion term. SIAM Journal on Mathematical Analysis, 44(5):3388– 3411, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alibaud:2010:APE

[AIK10] Nathael Alibaud, Cyril Imbert, and Grzegorz Karch. Asymptotic properties of entropy solutions to fractal Burgers equation. SIAM Journal on Mathematical Analysis, 42(1):354–376, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahn:2017:LBM

[AKKL17] Jaewook Ahn, Kyungkeun Kang, Junha Kim, and Jihoon Lee. Lower bound of mass in a chemotactic model with advection and absorbing reaction. SIAM Journal on Mathematical Analysis, 49 (2):723–755, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ando:2017:SNP

[AKKY17] Kazunori Ando, Hyeonbae Kang, Kyoungsun Kim, and Sanghyeon Yu. Spectrum of Neumann–Poincaré operator on annuli and cloaking by anomalous localized resonance for linear elasticity. SIAM Journal on Mathematical Analysis, 49(5):4232– 4250, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alonso:2010:FCH

[AL10] Ricardo J. Alonso and Bertrand Lods. Free cooling and highenergy tails of granular gases with variable restitution coefficient. SIAM Journal on Mathematical Analysis, 42(6): 2499–2538, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aleksanyan:2016:SCP

[Ale16]  Hayk Aleksanyan.  Slow

convergence in periodic homogenization problems for divergence-type elliptic operators. SIAM Journal on Mathematical Analysis, 48(5): 3345–3382, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aujol:2010:EBI

[ALM10] Jean-François Aujol, Saïd Ladjal, and Simon Masnou. Exemplar-based inpainting from a variational point of view. SIAM Journal on Mathematical Analysis, 42(3): 1246–1285, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almog:2017:CMF

[Alm17] Y. Almog. The Clausius– Mossotti formula for dilute random media of perfectly conducting inclusions. SIAM Journal on Mathematical Analysis, 49(4):2885– 2919, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Allen:2015:TPF

[ALP15] Mark Allen, Erik Lindgren, and Arshak Petrosyan. The two-phase fractional obstacle problem. SIAM Journal on Mathematical Analysis, 47(3): 1879–1905, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abels:2015:ADD

[ALS15] Helmut Abels, Kei Fong Lam, and Björn Stinner. Analysis of the diffuse domain approach for a bulk-surface coupled PDE system. SIAM Journal on Mathematical Analysis, 47(5):3687–3725, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andreianov:2014:WPO

[ALST14] Boris Andreianov, Frédéric Lagoutière, Nicolas Seguin, and Takéo Takahashi. Wellposedness for a one-dimensional fluid-particle interaction model. SIAM Journal on Mathematical Analysis, 46(2):1030– 1052, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2018:SPR

[aLW18] Lin an Li and Yi Wang. Stability of planar rarefaction wave to two-dimensional compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 50(5): 4937–4963, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2019:BWB

[ALZ19] Habib Ammari, Hyundae Lee, and Hai Zhang. Bloch waves in bubbly crystal near the first band gap: a highfrequency homogenization approach. SIAM Journal on Mathematical Analysis, 51(1): 45–59, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andres:2015:TNE

[AM15] Fuensanta Andrés and Julio Munñoz. A type of nonlocal elliptic problem: Existence and approximation through a Galerkin–Fourier method. SIAM Journal on Mathematical Analysis, 47(1): 498–525, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Allaire:2010:HAD

[AMP10] Grégoire Allaire, Andro Mikelić, and Andrey Piatnitski. Homogenization approach to the dispersion theory for reactive transport through porous media. SIAM Journal on Mathematical Analysis, 42(1): 125–144, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arrizabalaga:2015:SID

[AMV15] Naiara Arrizabalaga, Albert Mas, and Luis Vega. Shell interactions for Dirac operators: On the point spectrum and the confinement. SIAM Journal on Mathematical Analysis, 47(2):1044–1069, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Appleby:2010:ASR

[AMW10] John A. D. Appleby, Xuerong Mao, and Huizhong Wu. On the almost sure running maxima of solutions of affine stochastic functional differential equations. SIAM Journal on Mathematical Analysis, 42 (2):646–678, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ao:2011:TJS

[AMW11] Weiwei Ao, Monica Musso, and Juncheng Wei. Triple junction solutions for a singularly perturbed Neumann problem. SIAM Journal on Mathematical Analysis, 43(6): 2519–2541, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2519_s1.

Arkeryd:2015:WPC

[AN15] Leif Arkeryd and Anne Nouri. Well-posedness of the Cauchy problem for a spacedependent anyon Boltzmann equation. SIAM Journal on Mathematical Analysis, 47(6): 4720–4742, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrievskii:2012:AFR

[And12] Vladimir Andrievskii. Approximation of functions by reciprocals of polynomials on a quasi-smooth arc. SIAM Journal on Mathematical Analysis, 44(4):2329– 2343, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Athanassoulis:2011:SPS

[AP11] Agissilaos Athanassoulis and Thierry Paul. Strong phasespace semiclassical asymptotics. SIAM Journal on Mathematical Analysis, 43(5): 2116–2149, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2116_s1.

Ansini:2014:PLA

[AP14] Nadia Ansini and Francesca Prinari. Power-law approximation under differential constraints. SIAM Journal on Mathematical Analysis, 46(2): 1085–1115, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Akopyan:2015:MRC

[AP15] Arseniy Akopyan and Alexander Plakhov. Minimal resistance of curves under the single impact assumption. SIAM Journal on Mathematical Analysis, 47(4):2754– 2769, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Appleby:2018:GRS

[AP18] John A. D. Appleby and Denis D. Patterson. Growth rates of sublinear functional and Volterra differential equations. SIAM Journal on Mathematical Analysis, 50(2): 2086–2110, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Almeida:2019:ORP

[APSV19] Luis Almeida, Yannick Privat, Martin Strugarek, and Nicolas Vauchelet. Optimal releases for population replacement strategies: Application to Wolbachia. SIAM Journal on Mathematical Analysis, 51(4):3170–3194, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antil:2019:SSN

[AR19] Harbir Antil and Carlos N.

Rautenberg. Sobolev spaces with non-muckenhoupt weights, fractional elliptic operators, and applications. SIAM Journal on Mathematical Analysis, 51(3):2479–2503, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arada:2016:GNF

[Ara16] Nadir Arada. On generalized Newtonian fluids in curved pipes. SIAM Journal on Mathematical Analysis, 48(2): 1210–1249, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anma:2013:TTM

[AS13] Atsushi Anma and Kunimochi Sakamoto. Turing type mechanisms for linear diffusion systems under nondiagonal Robin boundary conditions. SIAM Journal on Mathematical Analysis, 45(6): 3611–3628, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Akagi:2014:DNE

[AS14] Goro Akagi and Ulisse Stefanelli. Doubly nonlinear equations as convex minimization. SIAM Journal on Mathematical Analysis, 46(3): 1922–1945, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ambrose:2015:LET

[AS15] David M. Ambrose and Gideon Simpson. Local existence theory for derivative nonlinear Schrödinger equations with noninteger power nonlinearities. SIAM Journal on Mathematical Analysis, 47(3):2241–2264, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aceves-Sanchez:2019:FDL

[ASC19] Pedro Aceves-Sanchez and Ludovic Cesbron. Fractional diffusion limit for a fractional Vlasov–Fokker–Planck equation. SIAM Journal on Mathematical Analysis, 51(1): 469–488, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ashton:2013:SDN

[Ash13] A. C. L. Ashton. The spectral Dirichlet–Neumann map for Laplace’s equation in a convex polygon. SIAM Journal on Mathematical Analysis, 45(6): 3575–3591, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aceves-Sanchez:2016:FDA

[ASS16] Pedro Aceves-Sánchez and Christian Schmeiser. Fractional-Diffusion–Advection limit of a kinetic model. SIAM Journal on Mathematical Analysis, 48(4):2806–2818, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Achdou:2010:TTC

[AT10] Yves Achdou and Nicoletta Tchou. Trace theorems for a class of ramified domains with self-similar fractal boundaries. SIAM Journal on Mathematical Analysis, 42(4):1449– 1482, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andreucci:2014:CDP

[AT14] Daniele Andreucci and Anatoli F. Tedeev. The Cauchy– Dirichlet problem for the porous media equation in cone-like domains. SIAM Journal on Mathematical Analysis, 46(2):1427–1455, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alves:2019:MUS

[ATSR19] Michele O. Alves, Eduardo H. Gomes Tavares, Marcio A. Jorge Silva, and José H. Rodrigues. On modeling and uniform stability of a partially dissipative viscoelastic Timoshenko system. SIAM Journal on Mathematical Analysis, 51(6):4520–4543, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Audiard:2012:DSE

[Aud12] Corentin Audiard. Dispersive smoothing for the Euler– Korteweg model. SIAM Journal on Mathematical Analysis, 44(4):3018–3040, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amstutz:2016:AIO

[AV16] Samuel Amstutz and Nicolas Van Goethem. Analysis of the incompatibility operator and application in intrinsic elasticity with dislocations. SIAM Journal on Mathematical Analysis, 48(1): 320–348, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Agirre:2019:SLB

[AV19] Mikel Agirre and Luis Vega.

Some lower bounds for solutions of Schrödinger evolutions. SIAM Journal on Mathematical Analysis, 51(4): 3324–3336, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arrieta:2016:UOM

[AVP16] José M. Arrieta and Manuel Villanueva-Pesqueira. Unfolding operator method for thin domains with a locally periodic highly oscillatory boundary. SIAM Journal on Mathematical Analysis, 48(3):1634– 1671, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ansini:2012:AAN

[AZ12] Nadia Ansini and Caterina Ida Zeppieri. Asymptotic analysis of nonsymmetric linear operators via Γconvergence. SIAM Journal on Mathematical Analysis, 44 (3):1617–1635, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:2017:EMT

[AZ17] Habib Ammari and Hai Zhang. Effective medium theory for acoustic waves in bubbly fluids near Minnaert resonant frequency. SIAM Journal on Mathematical Analysis, 49(4):3252–3276, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Banaji:2010:CSM

[BA10] Murad Banaji and David Angeli. Convergence in strongly monotone systems with an increasing first integral. SIAM Journal on Mathematical Analysis, 42(1):334–353, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See addendum [BA12].

Banaji:2012:ACS

[BA12] Murad Banaji and David Angeli. Addendum to “Convergence in Strongly Monotone Systems with an Increasing First Integral”. SIAM Journal on Mathematical Analysis, 44 (1):536–537, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/

1/sjmaah/v44/i1/p536_s1.

See [BA10].

Bao:2012:GPP

[BAC12] Weizhu Bao, Naoufel Ben Abdallah, and Yongyong Cai. Gross–Pitaevskii– Poisson equations for dipolar Bose–Einstein condensate with anisotropic confinement. SIAM Journal on Mathematical Analysis, 44(3):1713– 1741, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ben-Artzi:2018:ALS

[BACP18] Jonathan Ben-Artzi, Simone Calogero, and Stephen Pankavich. Arbitrarily large solutions of the Vlasov– Poisson system. SIAM Journal on Mathematical Analysis, 50(4):4311–4326, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ben-Artzi:2017:IRV

[BAH17] Jonathan Ben-Artzi and Thomas Holding. Instabilities of the relativistic Vlasov–Maxwell system on unbounded domains. SIAM Journal on Mathematical Analysis, 49(5):4024–4063, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2013:CTD

[BAP13] Guillaume Bal, Naoufel Ben Abdallah, and Marjolaine Puel. A corrector theory for diffusion-homogenization limits of linear transport equations. SIAM Journal on Mathematical Analysis, 44(6): 3848–3873, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barekat:2014:CCM

[Bar14] Farzin Barekat. On the consistency of compressed modes for variational problems associated with the Schrödinger operator. SIAM Journal on Mathematical Analysis, 46(5): 3568–3577, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benbourhim:2010:MPP

[BB10a] M. N. Benbourhim and A. Bouhamidi. Meshless pseudo-polyharmonic divergencefree and curl-free vector fields approximation. SIAM Journal on Mathematical Analysis, 42(3):1218–1245, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bianchini:2010:EPF

[BB10b] S. Bianchini and A. Brancolini. Estimates on path functionals over Wasserstein spaces. SIAM Journal on Mathematical Analysis, 42(3): 1179–1217, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baudel:2017:STR

[BB17] Manon Baudel and Nils Berglund. Spectral theory for random Poincaré maps. SIAM Journal on Mathematical Analysis, 49(6):4319– 4375, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ben-Benjamin:2018:WWF

[BBCD^{+}18] J. S. Ben-Benjamin, L. Cohen, N. C. Dias, P. Loughlin, and J. N. Prata. What is the Wigner function closest to a given square integrable function? SIAM Journal on Mathematical Analysis, 50(5): 5161–5197, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertini:2012:BEG

[BBG12] Lorenzo Bertini, Paolo Buttà, and Adriana Garroni. Boundary effects in the gradient theory of phase transitions. SIAM Journal on Mathematical Analysis, 44(2):926– 945, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bianchini:2016:RAN

[BBG16] Stefano Bianchini, Paolo Bonicatto, and Nikolay A. Gusev. Renormalization for autonomous nearly incompressible BV vector fields in two dimensions. SIAM Journal on Mathematical Analysis, 48 (1):1–33, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bogosel:2017:OSM

[BBG17] B. Bogosel, D. Bucur, and A. Giacomini. Optimal shapes maximizing the Steklov eigenvalues. SIAM Journal on Mathematical Analysis, 49(2): 1645–1680, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellettini:2012:COD

[BBMN12] Giovanni Bellettini, Lorenzo Bertini, Mauro Mariani, and Matteo Novaga. Convergence of the one-dimensional Cahn– Hilliard equation. SIAM Journal on Mathematical Analysis, 44(5):3458–3480, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braun:2019:ECS

[BBO19] Julian Braun, Maciej Buze, and Christoph Ortner. The effect of crystal symmetries on the locality of screw dislocation cores. SIAM Journal on Mathematical Analysis, 51(2): 1108–1136, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barreiro:2019:ACH

[BBR19] Andrea K. Barreiro, Jared C.

Bronski, and Zoi Rapti. Applications of a class of Herglotz operator pencils. SIAM Journal on Mathematical Analysis, 51(1):256–275, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brasco:2011:BBA

[BBS11] L. Brasco, G. Buttazzo, and F. Santambrogio. A Benamou–Brenier approach to branched transport. SIAM Journal on Mathematical Analysis, 43(2):1023–1040, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p1023_s1.

Bulicek:2016:UTS

[BBS16] M. Bulícek, J. Burczak, and S. Schwarzacher. A unified theory for some non-Newtonian fluids under singular forcing. SIAM Journal on Mathematical Analysis, 48(6): 4241–4267, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbu:2014:SPE

[BBT14] Viorel Barbu, Stefano Bonaccorsi, and Luciano Tubaro. A stochastic parabolic equation with nonlinear flux on the boundary driven by a Gaussian noise. SIAM Journal on Mathematical Analysis, 46(1): 780–802, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bucur:2014:SOP

[BBV14] Dorin Bucur, Giuseppe Buttazzo, and Bozhidar Velichkov. Spectral optimization problems for potentials and measures. SIAM Journal on Mathematical Analysis, 46(4): 2956–2986, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burchard:2011:CII

[BC11] Almut Burchard and Marina Chugunova. On computing the instability index of a non-self-adjoint differential operator associated with coating and rimming flows. SIAM Journal on Mathematical Analysis, 43(1):367– 388, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i1/p367_s1.

Bonacini:2014:LGM

[BC14] M. Bonacini and R. Cristoferi.

Local and global minimality results for a nonlocal isoperimetric problem on R^{N}. SIAM Journal on Mathematical Analysis, 46(4):2310– 2349, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellieud:2017:AAS

[BC17a] Michel Bellieud and Shane Cooper. Asymptotic analysis of stratified elastic media in the space of functions with bounded deformation. SIAM Journal on Mathematical Analysis, 49(5): 4275–4317, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boussaid:2017:NAS

[BC17b] Nabile Boussaïd and Andrew Comech. Nonrelativistic asymptotics of solitary waves in the Dirac equation with soler-type nonlinearity. SIAM Journal on Mathematical Analysis, 49(4): 2527–2572, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burtea:2017:LMG

[BC17c] Cosmin Burtea and Frédéric Charve. Lagrangian methods for a general inhomogeneous incompressible Navier– Stokes–Korteweg system with variable capillarity and viscosity coefficients. SIAM Journal on Mathematical Analysis, 49(5):3476–3495, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baspinar:2019:UVM

[BC19] E. Baspinar and G. Citti. Uniqueness of viscosity mean curvature flow solution in two sub-Riemannian structures. SIAM Journal on Mathematical Analysis, 51(3): 2633–2659, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Binev:2011:CRG

[BCD^{+}11] Peter Binev, Albert Cohen, Wolfgang Dahmen, Ronald DeVore, Guergana Petrova, and Przemyslaw Wojtaszczyk. Convergence rates for greedy algorithms in reduced basis methods. SIAM Journal on Mathematical Analysis, 43(3): 1457–1472, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1457_s1.

Bonito:2017:DCE

[BCD^{+}17a] Andrea Bonito, Albert Cohen, Ronald DeVore, Guergana Petrova, and Gerrit Welper. Diffusion coefficients estimation for elliptic partial differential equations. SIAM Journal on Mathematical Analysis, 49(2): 1570–1592, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Breit:2017:TFK

[BCD17b] D. Breit, A. Cianchi, and L. Diening. Trace-free Korn inequalities in Orlicz spaces. SIAM Journal on Mathematical Analysis, 49(4):2496– 2526, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernardi:2016:CPI

[BCDG16] Christine Bernardi, Martin Costabel, Monique Dauge, and Vivette Girault. Continuity properties of the infsup constant for the divergence. SIAM Journal on Mathematical Analysis, 48(2): 1250–1271, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonheure:2018:OSS

[BCdSN18] Denis Bonheure, Jean-Baptiste Casteras, Ederson Moreira dos Santos, and Robson Nascimento. Orbitally stable standing waves of a mixed dispersion nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 50(5): 5027–5071, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernard:2010:HLB

[BCG10] Etienne Bernard, Emanuele Caglioti, and François Golse. Homogenization of the linear Boltzmann equation in a domain with a periodic distribution of holes. SIAM Journal on Mathematical Analysis, 42(5):2082–2113, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braukhoff:2020:CCD

[BCJ20] Marcel Braukhoff, Xiuqing Chen, and Ansgar Jüngel. Corrigendum: Cross Diffusion Preventing Blow-Up in the Two-Dimensional Keller– Segel Model. SIAM Journal on Mathematical Analysis, 52(2):2198–2200, ???? 2020. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bisi:2011:UWI

[BCL11] Marzia Bisi, José A. Cañizo, and Bertrand Lods. Uniqueness in the weakly inelastic regime of the equilibrium state to the Boltzmann equation driven by a particle bath. SIAM Journal on Mathematical Analysis, 43(6): 2640–2674, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.

siam.org/sima/resource/
1/sjmaah/v43/i6/p2640_s1.

Bez:2018:SEK

[BCL18] Neal Bez, Jayson Cunanan, and Sanghyuk Lee. Smoothing estimates for the kinetic transport equation at the critical regularity. SIAM Journal on Mathematical Analysis, 50(2):2280–2294, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barles:2011:HFH

[BCN11] Guy Barles, Annalisa Cesaroni, and Matteo Novaga. Homogenization of fronts in highly heterogeneous media. SIAM Journal on Mathematical Analysis, 43(1):212– 227, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p212_s1.

Barekat:2017:SCM

[BCO17] Farzin Barekat, Russel Caflisch, and Stanley Osher. On the support of compressed modes. SIAM Journal on Mathematical Analysis, 49(4): 2573–2590, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bigot:2019:PBW

[BCP19] Jérémie Bigot, Elsa Cazelles, and Nicolas Papadakis. Penalization of barycenters in the Wasserstein space. SIAM Journal on Mathematical Analysis, 51(3):2261– 2285, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brandle:2012:PTM

[BCQ12] Cristina Brändle, Emmanuel Chasseigne, and Fernando Quirós. Phase transitions with midrange interactions: a nonlocal Stefan model. SIAM Journal on Mathematical Analysis, 44(4):3071– 3100, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braides:2015:TCE

[BCS15] Andrea Braides, Marco Cicalese, and Francesco Solombrino. Q-tensor continuum energies as limits of headto-tail symmetric spin systems. SIAM Journal on Mathematical Analysis, 47(4): 2832–2867, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2016:UMB

[BCS16] Guillaume Bal, Francis J. Chung, and John C. Schotland. Ultrasound modulated bioluminescence tomography and controllability of the radiative transport equation. SIAM Journal on Mathematical Analysis, 48(2): 1332–1347, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Borrelli:2019:NDE

[BCT19] William Borrelli, Raffaele Carlone, and Lorenzo Tentarelli. Nonlinear Dirac equation on graphs with localized nonlinearities: Bound states and nonrelativistic limit. SIAM Journal on Mathematical Analysis, 51(2):1046– 1081, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bevan:2018:CME

[BD18] Jonathan J. Bevan and Jonathan H. B. Deane. A calibration method for estimating critical cavitation loads from below in three dimensional nonlinear elasticity. SIAM Journal on Mathematical Analysis, 50(3):2566– 2587, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barucq:2018:MDF

[BDEM18] Hélène Barucq, Rabia Djellouli, Elodie Estecahandy, and Mohand Moussaoui. Mathematical determination of the Fréchet derivative with respect to the domain for a fluid-structure scattering problem: Case of polygonalshaped domains. SIAM Journal on Mathematical Analysis, 50(1):1010–1036, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burger:2018:SPM

[BDFS18] Martin Burger, Marco Di Francesco, Simone Fagioli, and Angela Stevens. Sorting phenomena in a mathematical model for two mutually Attracting/Repelling species. SIAM Journal on Mathematical Analysis, 50(3): 3210–3250, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bogelein:2013:PMT

[BDG13] Verena Bögelein, Frank Duzaar, and Ugo Gianazza. Porous medium type equations with measure data and potential estimates. SIAM Journal on Mathematical Analysis, 45(6): 3283–3330, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:2016:IBV

[BdHFS16] Elena Beretta, Maarten V. de Hoop, Florian Faucher, and Otmar Scherzer. Inverse boundary value problem for the Helmholtz equation: Quantitative conditional Lipschitz stability estimates. SIAM Journal on Mathematical Analysis, 48(6):

3962–3983, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:2013:LSI

[BdHQ13] Elena Beretta, Maarten V.

de Hoop, and Lingyun Qiu. Lipschitz stability of an inverse boundary value problem for a Schrödinger-Type equation. SIAM Journal on Mathematical Analysis, 45(2): 679–699, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Borthwick:2019:ESS

[BDLM19] David Borthwick, Roland Donninger, Enno Lenzmann, and Jeremy L. Marzuola. Existence and stability of Schrödinger solitons on noncompact manifolds. SIAM Journal on Mathematical Analysis, 51(5):3854–3901, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burger:2010:NCD

[BDPS10] Martin Burger, Marco Di Francesco, Jan-Frederik Pietschmann, and Bärbel Schlake. Nonlinear cross-diffusion with size exclusion. SIAM Journal on Mathematical Analysis, 42(6): 2842–2871, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v42/i6/p2842_s1.

Bogelein:2018:EVS

[BDSS18] Verena Bögelein, Frank Duzaar, Christoph Scheven, and Thomas Singer. Existence of variational solutions in noncylindrical domains. SIAM Journal on Mathematical Analysis, 50(3):3007–3057, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonaccorsi:2012:ABC

[BDT12] Stefano Bonaccorsi, Giuseppe Da Prato, and Luciano Tubaro. Asymptotic behavior of a class of nonlinear stochastic heat equations with memory effects. SIAM Journal on Mathematical Analysis, 44(3): 1562–1587, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burq:2012:WEE

[BDWZ12] Nicolas Burq, Semyon Dyatlov, Rachel Ward, and Maciej Zworski. Weighted eigenfunction estimates with applications to compressed sensing. SIAM Journal on Mathematical Analysis, 44(5):3481– 3501, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bae:2014:SSS

[BDX14] Myoungjean Bae, Ben Duan, and Chunjing Xie. Subsonic solutions for steady Euler–Poisson system in twodimensional nozzles. SIAM Journal on Mathematical Analysis, 46(5):3455–3480, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bebendorf:2016:LRA

[Beb16]  Mario Bebendorf.  Low-

rank approximation of elliptic boundary value problems with high-contrast coefficients. SIAM Journal on Mathematical Analysis, 48(2): 932–949, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beck:2018:ULS

[Bec18] Thomas Beck. Uniform level set estimates for ground state eigenfunctions. SIAM Journal on Mathematical Analysis, 50(4):4483–4502, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bahlali:2015:EUM

[BEH15] K. Bahlali, E. Essaky, and M. Hassani. Existence and uniqueness of multidimensional BSDEs and of systems of degenerate PDEs with superlinear growth generator. SIAM Journal on Mathematical Analysis, 47(6): 4251–4288, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellieud:2010:TEE

[Bel10]  Michel Bellieud.  Torsion

effects in elastic composites with high contrast. SIAM Journal on Mathematical Analysis, 41(6):2514– 2553, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellieud:2017:HSE

[Bel17] Michel Bellieud. Homogenization of stratified elastic composites with high contrast. SIAM Journal on Mathematical Analysis, 49(4): 2615–2665, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benoit:2017:GOE

[Ben17] Antoine Benoit. Geometric optics expansions for hyperbolic corner problems II: From weak stability to violent instability. SIAM Journal on Mathematical Analysis, 49(5): 3335–3395, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernard:2012:STE

[Ber12a] J.-M. Bernard. Steady transport equation in the case where the normal component of the velocity does not vanish on the boundary. SIAM Journal on Mathematical Analysis, 44(2):993–1018, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernard:2012:PSG

[Ber12b] Jean-Marie Bernard. Problem of second grade fluids in convex polyhedrons. SIAM Journal on Mathematical Analysis, 44(3):2018–2038, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berthelin:2017:TSM

[Ber17]  F. Berthelin.  Theoret-

ical study of a multidimensional pressureless model with unilateral constraint. SIAM Journal on Mathematical Analysis, 49(3):2287– 2320, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bessi:2012:CMV

[Bes12] Ugo Bessi. Chaotic motions for a version of the Vlasov equation. SIAM Journal on Mathematical Analysis, 44(4): 2496–2525, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bessi:2016:ESM

[Bes16] Ugo Bessi. Existence of solutions of the master equation in the smooth case. SIAM Journal on Mathematical Analysis, 48(1):204–228, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Betermin:2016:TDT

[Bét16]  Laurent Bétermin.  Two-

dimensional theta functions and crystallization among Bravais lattices. SIAM Journal on Mathematical Analysis, 48(5):3236–3269, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bevan:2011:ALM

[Bev11] J. J. Bevan. Austenite as a local minimizer in a model of material microstructure with a surface energy term. SIAM Journal on Mathematical Analysis, 43(2):1041– 1073, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p1041_s1.

Bernier:2019:EST

[BF19] Joackim Bernier and Erwan Faou. Existence and stability of traveling waves for discrete nonlinear Schrödinger equations over long times. SIAM Journal on Mathematical Analysis, 51(3):1607– 1656, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bidegaray-Fesquet:2013:NCD

[BFDJ13] Brigitte Bidégaray-Fesquet, Eric Dumas, and Guillaume James. From Newton’s cradle to the discrete p-Schrödinger equation. SIAM Journal on Mathematical Analysis, 45(6): 3404–3430, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bella:2017:SHL

[BFFO17] Peter Bella, Benjamin Fehrman, Julian Fischer, and Felix Otto. Stochastic homogenization of linear elliptic equations: Higher-order error estimates in weak norms via second-order correctors. SIAM Journal on Mathematical Analysis, 49(6):4658– 4703, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berninger:2013:DID

[BFG^{+}13] H. Berninger, E. Frénod, M. Gander, M. Liebendörfer, and J. Michaud. Derivation of the Isotropic Diffusion Source Approximation (IDSA) for supernova neutrino transport by asymptotic expansions. SIAM Journal on Mathematical Analysis, 45(6):

3229–3265, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barcelo:2012:BAL

[BFGPE^{+}12] Juan Antonio Barceló, Magali Folch-Gabayet, Salvador Pérez-Esteva, Alberto Ruiz, and Mari Cruz Vilela. A Born approximation for live loads in Navier elasticity. SIAM Journal on Mathematical Analysis, 44(4):2824–2846, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barcelo:2018:UIE

[BFGPE^{+}18] J. A. Barceló, M. Folch-Gabayet, S. Pérez-Esteva, A. Ruiz, and M. C. Vilela. Uniqueness for inverse elastic medium problems. SIAM Journal on Mathematical Analysis, 50(4):3939–3962, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biskup:2016:EFL

[BFK16] Marek Biskup, Ryoki Fukushima, and Wolfgang König. Eigenvalue fluctuations for lattice Anderson Hamiltonians. SIAM Journal on Mathematical Analysis, 48(4):2674– 2700, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bella:2016:RJE

[BFLN16] Peter Bella, Eduard Feireisl, Marta Lewicka, and Antonín Novotný. A rigorous justification of the Euler and Navier– Stokes equations with geometric effects. SIAM Journal on Mathematical Analysis, 48(6): 3907–3930, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouchitte:2012:OTT

[BFLS12] Guy Bouchitté, Ilaria Fragalà, Ilaria Lucardesi, and Pierre Seppecher. Optimal thin torsion rods and Cheeger sets. SIAM Journal on Mathematical Analysis, 44(1):483– 512, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p483_s1.

Benesova:2018:EWS

[BFLS18] Barbora Benesová, Johannes Forster, Chun Liu, and Anja Schlömerkemper. Existence of weak solutions to an evolutionary model for magnetoelasticity. SIAM Journal on Mathematical Analysis, 50(1): 1200–1236, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Babadjian:2012:QSE

[BFM12] J.-F. Babadjian, G. A. Francfort, and M. G. Mora. Quasistatic evolution in nonassociative plasticity: The cap model. SIAM Journal on Mathematical Analysis, 44(1): 245–292, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p245_s1.

Bardos:2013:STD

[BFN^{+}13] C. Bardos, M. C. Lopes Filho, Dongjuan Niu, H. J. Nussenzveig Lopes, and E. S. Titi. Stability of two-dimensional viscous incompressible flows under three-dimensional perturbations and inviscid symmetry breaking. SIAM Journal on Mathematical Analysis, 45(3):1871–1885, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barcelo:2013:RDB

[BFRV13] J. A. Barceló, Daniel Faraco, Alberto Ruiz, and Ana Vargas. Reconstruction of discontinuities from backscattering data in two dimensions. SIAM Journal on Mathematical Analysis, 45(6): 3494–3513, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bothe:2014:GWP

[BFS14] Dieter Bothe, André Fischer, and Jürgen Saal. Global well-posedness and stability of electrokinetic flows. SIAM Journal on Mathematical Analysis, 46(2):1263– 1316, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:2018:WPM

[BFTT18] Michiel Bertsch, Bruno Franchi, Maria Carla Tesi, and Andrea Tosin. Well-posedness of a mathematical model for Alzheimer’s disease. SIAM Journal on Mathematical Analysis, 50(3):2362–2388, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:2017:DDN

[BFV17] Elena Beretta, Elisa Francini, and Sergio Vessella. Differentiability of the Dirichlet to Neumann map under movements of polygonal inclusions with an application to shape optimization. SIAM Journal on Mathematical Analysis, 49 (2):756–776, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Batenkov:2015:SME

[BFY15] Dmitry Batenkov, Omer Friedland, and Yosef Yomdin. Sampling, metric entropy, and dimensionality reduction. SIAM Journal on Mathematical Analysis, 47(1):786–

796, ????  2015.  CO-

DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berglund:2014:NIP

[BG14] Nils Berglund and Barbara Gentz. On the noise-induced passage through an unstable periodic orbit II: General case. SIAM Journal on Mathematical Analysis, 46(1): 310–352, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertagna:2017:WPC

[BG17] Luca Bertagna and Max Gunzburger. Well posedness of a coupled ICE-hydrology problem arising in glaciology. SIAM Journal on Mathematical Analysis, 49(2):699– 722, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bessaih:2017:SLD

[BGAHS17] Hakima Bessaih, María J. Garrido-Atienza, Xiaoying Han, and Björn Schmalfuss. Stochastic lattice dynamical systems with fractional noise. SIAM Journal on Mathematical Analysis, 49(2): 1495–1518, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouin:2018:TFL

[BGHP18] Emeric Bouin, Jimmy Garnier, Christopher Henderson, and Florian Patout. Thin front limit of an integrodifferential Fisher–KPP equation with fat-tailed kernels. SIAM Journal on Mathematical Analysis, 50(3):3365– 3394, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertozzi:2012:CRS

[BGL12] Andrea L. Bertozzi, John B.

Garnett, and Thomas Laurent. Characterization of radially symmetric finite time blowup in multidimensional aggregation equations. SIAM Journal on Mathematical Analysis, 44(2):651– 681, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i2/p651_s1.

Bulicek:2016:SDE

[BGL16] Miroslav Bulícek, Annegret Glitzky, and Matthias Liero. Systems describing electrothermal effects with p(x)-Laplacian-like structure for discontinuous variable exponents. SIAM Journal on Mathematical Analysis, 48(5):3496– 3514, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertozzi:2016:RBM

[BGLV16] A. Bertozzi, J. Garnett, T. Laurent, and J. Verdera. The regularity of the boundary of a multidimensional aggregation patch. SIAM Journal on Mathematical Analysis, 48(6):3789–3819, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonheure:2019:PST

[BGM19] Denis Bonheure, Filippo Gazzola, and Ederson Moreira Dos Santos. Periodic solutions and torsional instability in a nonlinear nonlocal plate equation. SIAM Journal on Mathematical Analysis, 51(4): 3052–3091, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bulicek:2012:UFI

[BGMŚG12] Miroslav Bulícek, Piotr Gwiazda, Josef Málek, and Agnieszka Świerczewska-Gwiazda. On unsteady flows of implicitly constituted incompressible fluids. SIAM Journal on Mathematical Analysis, 44(4): 2756–2801, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brzezniak:2014:MSB

[BGN14] Zdzisław Brzeźniak, Ben Goldys, and Misha Neklyudov. Multidimensional stochastic Burgers equation. SIAM Journal on Mathematical Analysis, 46(1):871– 889, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:2019:MVS

[BGT19] Michiel Bertsch, Lorenzo Giacomelli, and Alberto Tesei. Measure-valued solutions to a nonlinear fourth-order regularization of forwardbackward parabolic equations. SIAM Journal on Mathematical Analysis, 51(1): 374–402, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2011:OEM

[BH11] Alberto Bressan and Ke Han.

Optima and equilibria for a model of traffic flow. SIAM Journal on Mathematical Analysis, 43(5):2384– 2417, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2384_s1.

Bedrossian:2017:SBP

[BH17] Jacob Bedrossian and Siming He. Suppression of blowup in Patlak–Keller–Segel via shear flows. SIAM Journal on Mathematical Analysis, 49(6): 4722–4766, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [BH18].

Bedrossian:2018:ESB

[BH18] Jacob Bedrossian and Siming He. Erratum: Suppression of Blow-Up in Patlak– Keller–Segel Via Shear Flows. SIAM Journal on Mathematical Analysis, 50(6):6365– 6372, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [BH17].

Bae:2019:FBC

[BHK^{+}19] Hyeong-Ohk Bae, Seung-Yeal Ha, Jeongho Kim, Dongnam Ko, and Sung-Ik Sohn. Flocking behaviors of a Cucker– Smale ensemble in a cylindrical domain. SIAM Journal on Mathematical Analysis, 51(3): 2390–2424, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bogosel:2018:MED

[BHL18] B. Bogosel, A. Henrot, and I. Lucardesi. Minimization of the eigenvalues of the Dirichlet–Laplacian with a diameter constraint. SIAM Journal on Mathematical Analysis, 50(5):5337– 5361, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouin:2017:EDE

[BHM17] Emeric Bouin, Franca Hoffmann, and Clément Mouhot. Exponential decay to equilibrium for a fiber lay-down process on a moving conveyor belt. SIAM Journal on Mathematical Analysis, 49(4): 3233–3251, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bennett:2018:GEA

[BHND18] T. Bennett, C. J. Howls, G. Nemes, and A. B. Olde Daalhuis. Globally exact asymptotics for integrals with arbitrary order saddles. SIAM Journal on Mathematical Analysis, 50(2):2144– 2177, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bessaih:2016:ESS

[BHR16] Hakima Bessaih, Erika Hausenblas, and Paul A. Razafimandimby. Ergodicity of stochastic shell models driven by pure jump noise. SIAM Journal on Mathematical Analysis, 48(2):1423–1458, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burger:2016:LFS

[BHRW16] Martin Burger, Sabine Hittmeir, Helene Ranetbauer, and Marie-Therese Wolfram. Lane formation by side-stepping. SIAM Journal on Mathematical Analysis, 48(2):981– 1005, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beck:2010:NSS

[BHSZ10] Margaret Beck, Hermen Jan Hupkes, Björn Sandstede, and Kevin Zumbrun. Nonlinear stability of semidiscrete shocks for two-sided schemes. SIAM Journal on Mathematical Analysis, 42(2): 857–903, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2012:CRV

[BHWY12] Alberto Bressan, Feimin Huang, Yong Wang, and Tong Yang. On the convergence rate of vanishing viscosity approximations for nonlinear hyperbolic systems. SIAM Journal on Mathematical Analysis, 44(5):3537–3563, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bianchini:2018:UAC

[Bia18]  Roberta Bianchini.  Uni-

form asymptotic and convergence estimates for the Jin– Xin model under the diffusion scaling. SIAM Journal on Mathematical Analysis, 50(2): 1877–1899, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ballew:2016:BEC

[BIP16] Joshua Ballew, Gautam Iyer, and Robert L. Pego. Bose– Einstein condensation in a hyperbolic model for the Kompaneets equation. SIAM Journal on Mathematical Analysis, 48(6):3840–3859, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2010:STD

[BJ10] Guillaume Bal and Alexandre Jollivet. Stability for time-dependent inverse transport. SIAM Journal on Mathematical Analysis, 42(2):679– 700, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellazzini:2016:DQG

[BJ16] Jacopo Bellazzini and Louis Jeanjean. On dipolar quantum gases in the unstable regime. SIAM Journal on Mathematical Analysis, 48(3): 2028–2058, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bownik:2017:WB

[BJLO17] Marcin Bownik, Mads S.

Jakobsen, Jakob Lemvig, and Kasso A. Okoudjou. On Wilson bases in L^{2}(R^{d}).

SIAM Journal on Mathematical Analysis, 49(5):3999– 4023, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bedrossian:2013:GEF

[BK13] Jacob Bedrossian and Inwon C. Kim. Global existence and finite time blowup for critical Patlak–Keller– Segel models with inhomogeneous diffusion. SIAM Journal on Mathematical Analysis, 45(3):934–964, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benesova:2015:GYM

[BK15a] Barbora Benesová and Malte Kampschulte. Gradient Young measures generated by quasi-conformal maps in the plane. SIAM Journal on Mathematical Analysis, 47(6): 4404–4435, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biondini:2015:IST

[BK15b] Gino Biondini and Daniel Kraus. Inverse scattering transform for the defocusing Manakov system with nonzero boundary conditions. SIAM Journal on Mathematical Analysis, 47(1):706– 757, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellettini:2018:MMM

[BK18a] Giovanni Bellettini and Shokhrukh Yu.

Kholmatov. Minimizing movements for mean curvature flow of partitions. SIAM Journal on Mathematical Analysis, 50(4):4117– 4148, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braides:2018:IRR

[BK18b] Andrea Braides and Leonard Kreutz. An integral-representation result for continuum limits of discrete energies with multibody interactions. SIAM Journal on Mathematical Analysis, 50(2):1485–1520, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baia:2018:GYM

[BKK18] Margarida Baía, Stefan Krömer, and Martin Kruzík. Generalized W^{1,1}–Young measures and relaxation of problems with linear growth. SIAM Journal on Mathematical Analysis, 50(1):1076– 1119, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baroni:2018:EBR

[BKLU18] Paolo Baroni, Tuomo Kuusi, Casimir Lindfors, and José Miguel Urbano. Existence and boundary regularity for degenerate phase transitions. SIAM Journal on Mathematical Analysis, 50(1): 456–490, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bothe:2013:CEP

[BKP13] Dieter Bothe, Matthias Köhne, and Jan Prüss. On a class of energy preserving boundary conditions for incompressible Newtonian flows. SIAM Journal on Mathematical Analysis, 45(6):3768–3822, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonacini:2016:ODO

[BKR16] Marco Bonacini, Hans Knüpfer, and Matthias Röger. Optimal distribution of oppositely charged phases: Perfect screening and other properties. SIAM Journal on Mathematical Analysis, 48(2): 1128–1154, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barros:2011:SDC

[BL11] Jean F. Barros and Eduardo S. G. Leandro. The set of degenerate central configurations in the planar restricted four-body problem. SIAM Journal on Mathematical Analysis, 43(2):634– 661, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p634_s1.

Barros:2014:BEC

[BL14] Jean F. Barros and Eduardo S. G. Leandro. Bifurcations and enumeration of classes of relative equilibria in the planar restricted four-body problem. SIAM Journal on Mathematical Analysis, 46(2):1185– 1203, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buoso:2015:SSA

[BL15] Davide Buoso and Pier Domenico Lamberti. Shape sensitivity analysis of the eigenvalues of the Reissner–Mindlin system. SIAM Journal on Mathematical Analysis, 47(1): 407–426, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biswas:2019:MPA

[BL19] Anup Biswas and József Lörinczi. Maximum principles and Aleksandrov–Bakelman– Pucci type estimates for NonLocal Schrödinger equations with exterior conditions. SIAM Journal on Mathematical Analysis, 51(3): 1543–1581, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Black:2018:GVW

[Bla18a] Tobias Black. Global very weak solutions to a chemotaxisfluid system with nonlinear diffusion. SIAM Journal on Mathematical Analysis, 50(4): 4087–4116, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blaasten:2018:NST

[Blå18b] Eemeli Blåsten. Nonradiating sources and transmission eigenfunctions vanish at corners and edges. SIAM Journal on Mathematical Analysis, 50(6):6255–6270, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnivard:2015:ALM

[BLS15] Matthieu Bonnivard, Antoine Lemenant, and Filippo Santambrogio. Approximation of length minimization problems among compact connected sets. SIAM Journal on Mathematical Analysis, 47(2): 1489–1529, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Byeon:2019:FRP

[BLW19] Jaeyoung Byeon, Youngae Lee, and Zhi-Qiang Wang. Formation of radial patterns via mixed attractive and repulsive interactions for Schrödinger systems. SIAM Journal on Mathematical Analysis, 51(2): 1514–1542, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barchiesi:2016:BMH

[BLZ16] Marco Barchiesi, Giuliano Lazzaroni, and Caterina Ida Zeppieri. A bridging mechanism in the homogenization of brittle composites with soft inclusions. SIAM Journal on Mathematical Analysis, 48(2): 1178–1209, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellettini:2010:AHF

[BM10] Giovanni Bellettini and Luca Mugnai. Approximation of Helfrich’s functional via diffuse interfaces. SIAM Journal on Mathematical Analysis, 42(6):2402–2433, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:2012:ITI

[BM12a] Guillaume Bal and François Monard. Inverse transport with isotropic time-harmonic sources. SIAM Journal on Mathematical Analysis, 44(1): 134–161, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p134_s1.

Bessaih:2012:LDZ

[BM12b] Hakima Bessaih and Annie Millet. Large deviations and the zero viscosity limit for 2D stochastic Navier–Stokes equations with free boundary. SIAM Journal on Mathematical Analysis, 44(3):1861– 1893, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ba:2015:SII

[BM15a] Moustapha Ba and Pierre Mathieu. A Sobolev inequality and the individual invariance principle for diffusions in a periodic potential. SIAM Journal on Mathematical Analysis, 47(3):2022– 2043, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bucur:2015:SRS

[BM15b] Dorin Bucur and Dario Mazzoleni. A surgery result for the spectrum of the Dirichlet Laplacian. SIAM Journal on Mathematical Analysis, 47(6): 4451–4466, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butta:2018:LTE

[BM18] Paolo Buttà and Carlo Marchioro. Long time evolution of concentrated Euler flows with planar symmetry. SIAM Journal on Mathematical Analysis, 50(1):735–760, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brzezniak:2019:FDS

[BM19] Zdzisław Brzeźniak and Elzbieta Motyl. Fractionally dissipative stochastic quasigeostrophic type equations on R^{d}. SIAM Journal on Mathematical Analysis, 51(3):2306– 2358, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellido:2014:ENV

[BMC14] José C. Bellido and Carlos Mora-Corral. Existence for nonlocal variational problems in peridynamics. SIAM Journal on Mathematical Analysis, 46(1):890–916, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellido:2018:LSR

[BMC18] José C. Bellido and Carlos Mora-Corral. Lower semicontinuity and relaxation via Young measures for nonlocal variational problems and applications to peridynamics. SIAM Journal on Mathematical Analysis, 50(1):779– 809, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:2016:SSS

[BMMP16] C. Bandle, P. Mastrolia, D. D. Monticelli, and F. Punzo. On the stability of solutions of semilinear elliptic equations with Robin boundary conditions on Riemannian manifolds. SIAM Journal on Mathematical Analysis, 48(1): 122–151, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:2018:RDP

[BMP18] C. Bandle, D. D. Monticelli, and F. Punzo. Reactiondiffusion problems on timedependent Riemannian manifolds: Stability of periodic solutions. SIAM Journal on Mathematical Analysis, 50(6): 6082–6099, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bronzi:2014:TSS

[BMR14] Anne C. Bronzi, Cecilia F. Mondaini, and Ricardo M. S. Rosa. Trajectory statistical solutions for threedimensional Navier–Stokeslike systems. SIAM Journal on Mathematical Analysis, 46(3):1893–1921, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bermudez:2013:TEC

[BMSR^{+}13] Alfredo Bermúdez, Rafael Munñoz-Sola, Carlos Reales, Rodolfo Rodríguez, and Pilar Salgado. A transient eddy current problem on a moving domain. mathematical analysis. SIAM Journal on Mathematical Analysis, 45(6): 3629–3650, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bensoussan:2016:NBV

[BMY16] A. Bensoussan, L. Mertz, and S. C. P. Yam. NonLocal boundary value problems of a stochastic variational inequality modeling an elasto-plastic oscillator excited by a filtered noise. SIAM Journal on Mathematical Analysis, 48(4): 2783–2805, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:2014:GEW

[BN14] Alberto Bressan and Khai T.

Nguyen. Global existence of weak solutions for the Burgers–Hilbert equation. SIAM Journal on Mathematical Analysis, 46(4):2884– 2904, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnaillie-Noel:2010:GVT

[BNDHV10] V. Bonnaillie-Noël, M. Dambrine, F. Hérau, and G. Vial. On generalized Ventcel’s type boundary conditions for Laplace operator in a bounded domain. SIAM Journal on Mathematical Analysis, 42(2): 931–945, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Byun:2016:NPE

[BO16] Sun-Sig Byun and Jihoon Ok. Nonlinear parabolic equations with variable exponent growth in nonsmooth domains. SIAM Journal on Mathematical Analysis, 48(5): 3148–3190, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berman:2019:PCC

[BÖ19] Robert J. Berman and Magnus Önnheim. Propagation of chaos for a class of first order models with singular mean field interactions. SIAM Journal on Mathematical Analysis, 51(1):159–196, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnotte:2013:KRB

[Bon13]  Nicolas Bonnotte.  From

Knothe’s rearrangement to Brenier’s optimal transport map. SIAM Journal on Mathematical Analysis, 45(1): 64–87, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonafini:2018:VAF

[BOO18] Mauro Bonafini, Giandomenico Orlandi, and Édouard Oudet. Variational approximation of functionals defined on 1dimensional connected sets: The planar case. SIAM Journal on Mathematical Analysis, 50(6):6307–6332, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boros:2019:EPS

[Bor19]  Balázs Boros.  Existence

of positive steady states for weakly reversible mass-action systems. SIAM Journal on Mathematical Analysis, 51(1): 435–449, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brenier:2011:UBC

[BOS11] Yann Brenier, Felix Otto, and Christian Seis. Upper bounds on coarsening rates in demixing binary viscous liquids. SIAM Journal on Mathematical Analysis, 43(1): 114–134, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i1/p114_s1.

Bostan:2012:TCP

[Bos12] Mihai Bostan. Transport of charged particles under fast oscillating magnetic fields. SIAM Journal on Mathematical Analysis, 44(3):1415– 1447, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bostan:2016:MAL

[Bos16] Mihai Bostan. MultiScale analysis for linear first order PDEs. The finite Larmor radius regime. SIAM Journal on Mathematical Analysis, 48(3): 2133–2188, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baratchart:2017:BNP

[BOS17] Laurent Baratchart, Martine Olivi, and Fabien Seyfert. Boundary Nevanlinna–Pick interpolation with prescribed peak points. Application to impedance matching. SIAM Journal on Mathematical Analysis, 49(2):1131–1165, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bostan:2019:ABV

[Bos19] Mihai Bostan. Asymptotic behavior for the Vlasov–Poisson equations with strong external magnetic field. straight magnetic field lines. SIAM Journal on Mathematical Analysis, 51(3):2713–2747, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouya:2013:IFM

[Bou13] Ismaël Bouya. Instability of the forced magnetohydrodynamics system at small Reynolds number. SIAM Journal on Mathematical Analysis, 45(1):307–323, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bourgeat:2010:ASR

[BP10] Alain Bourgeat and Andrey L.

Piatnitski. Averaging of a singular random source term in a diffusion convection equation. SIAM Journal on Mathematical Analysis, 42(6): 2626–2651, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertini:2012:VAS

[BP12a] Lorenzo Bertini and Marcello Ponsiglione. A variational approach to the stationary solutions of the Burgers equation. SIAM Journal on Mathematical Analysis, 44 (2):682–698, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i2/p682_s1.

Boukrouche:2012:AAM

[BP12b] Mahdi Boukrouche and Laetitia Paoli. Asymptotic analysis of a micropolar fluid flow in a thin domain with a free and rough boundary. SIAM Journal on Mathematical Analysis, 44(2):1211–1256, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bresch:2014:NLW

[BP14a] Didier Bresch and Christophe Prange. Newtonian limit for weakly viscoelastic fluid flows. SIAM Journal on Mathematical Analysis, 46(2): 1116–1159, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bulicek:2014:EAM

[BP14b] Miroslav Bulícek and Petra Pustejovská. Existence analysis for a model describing flow of an incompressible chemically reacting non-Newtonian fluid. SIAM Journal on Mathematical Analysis, 46(5): 3223–3240, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bae:2019:CDD

[BP19] Myoungjean Bae and Hyangdong Park. Contact discontinuities for 2-dimensional inviscid compressible flows in infinitely long nozzles. SIAM Journal on Mathematical Analysis, 51(3):1730– 1760, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braides:2015:HDH

[BPP15] Andrea Braides, Valeria Chiadò Piat, and Andrey Piatnitski. Homogenization of discrete high-contrast energies. SIAM Journal on Mathematical Analysis, 47(4):3064– 3091, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bauer:2016:MCP

[BPS16] Sebastian Bauer, Dirk Pauly, and Michael Schomburg. The Maxwell compactness property in bounded weak Lipschitz domains with mixed boundary conditions. SIAM Journal on Mathematical Analysis, 48(4):2912–2943, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Batty:2019:OED

[BPS19] Charles Batty, Lassi Paunonen, and David Seifert. Optimal energy decay for the waveheat system on a rectangular domain. SIAM Journal on Mathematical Analysis, 51 (2):808–819, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bredies:2015:CLS

[BPW15] Kristian Bredies, Thomas Pock, and Benedikt Wirth. A convex, lower semicontinuous approximation of Euler’s elastica energy. SIAM Journal on Mathematical Analysis, 47 (1):566–613, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bulicek:2017:EAI

[BPZ17] Miroslav Bulícek, Milan Pokorný, and Nicola Zamponi. Existence analysis for incompressible fluid model of electrically charged chemically reacting and heat conducting mixtures. SIAM Journal on Mathematical Analysis, 49(5):3776–3830, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bronski:2011:CDM

[BR11] Jared C. Bronski and Zoi Rapti. Counting defect modes in periodic eigenvalue problems. SIAM Journal on Mathematical Analysis, 43(2): 803–827, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p803_s1.

Baumle:2017:EWS

[BR17] E. Bäumle and M. Ruzicka. Existence of weak solutions for unsteady motions of micropolar electrorheological fluids. SIAM Journal on Mathematical Analysis, 49(1): 115–141, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbu:2018:PRS

[BR18] Viorel Barbu and Michael Röckner. Probabilistic representation for solutions to nonlinear Fokker–Planck equations. SIAM Journal on Mathematical Analysis, 50(4): 4246–4260, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brandolese:2016:CSD

[Bra16] Lorenzo Brandolese. Characterization of solutions to dissipative systems with sharp algebraic decay. SIAM Journal on Mathematical Analysis, 48(3):1616–1633, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brezina:2013:ABS

[Bre13] Jan Brezina. Asymptotic behavior of solutions to the compressible Navier–Stokes equation around a time-periodic parallel flow. SIAM Journal on Mathematical Analysis, 45(6):3514–3574, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonder:2017:CRN

[BRS17] Julián Fernández Bonder, Antonella Ritorto, and Ariel Martin Salort. H-convergence result for nonlocal elliptictype problems via Tartar’s method. SIAM Journal on Mathematical Analysis, 49(4): 2387–2408, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bruveris:2016:ORS

[Bru16]  Martins Bruveris.  Opti-

mal reparametrizations in the square root velocity framework. SIAM Journal on Mathematical Analysis, 48(6): 4335–4354, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bao:2016:PER

[BS16a] Jianhai Bao and Jinghai Shao. Permanence and extinction of regime-switching predator– prey models. SIAM Journal on Mathematical Analysis, 48 (1):725–739, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bartosz:2016:RMV

[BS16b] Krzysztof Bartosz and Mircea Sofonea. The Rothe method for variational-hemivariational inequalities with applications to contact mechanics. SIAM Journal on Mathematical Analysis, 48(2):861–883, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Balilescu:2017:FRS

[BST17a] Loredana Balilescu, Jorge San Martín, and Takéo Takahashi. Fluid-rigid structure interaction system with Coulomb’s law. SIAM Journal on Mathematical Analysis, 49(6):4625– 4657, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:2017:CFB

[BST17b] Michiel Bertsch, Flavia Smarrazzo, and Alberto Tesei. On a class of forward-backward parabolic equations: Properties of solutions. SIAM Journal on Mathematical Analysis, 49(3):2037–2060, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baskin:2016:SHF

[BSW16] Dean Baskin, Euan A.

Spence, and Jared Wunsch. Sharp high-frequency estimates for the Helmholtz equation and applications to boundary integral equations. SIAM Journal on Mathematical Analysis, 48(1):229–

267, ????  2016.  CO-

DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bianchini:2012:SRH

[BT12] Stefano Bianchini and Daniela Tonon. SBV regularity for Hamilton–Jacobi equations with Hamiltonian depending on (t, x). SIAM Journal on Mathematical Analysis, 44(3): 2179–2203, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barles:2016:EUA

[BT16] Guy Barles and Erwin Topp. Existence, uniqueness, and asymptotic behavior for nonlocal parabolic problems with dominating gradient terms. SIAM Journal on Mathematical Analysis, 48(2):1512– 1547, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Badra:2019:GRS

[BT19] Mehdi Badra and Takéo Takahashi. Gevrey regularity for a system coupling the Navier– Stokes system with a beam equation. SIAM Journal on Mathematical Analysis, 51(6): 4776–4814, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bociu:2015:WPA

[BTZ15] Lorena Bociu, Daniel Toundykov, and Jean-Paul Zolésio. Wellposedness analysis for a linearization of a fluid-elasticity interaction. SIAM Journal on Mathematical Analysis, 47(3): 1958–2000, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Belishev:2010:PTD

[BV10] M. I. Belishev and A. F.

Vakulenko. s-points in threedimensional acoustical scattering. SIAM Journal on Mathematical Analysis, 42(6): 2703–2720, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baer:2013:BBT

[BV13] Eric Baer and Alexis Vasseur.

A bound from below on the temperature for the Navier–Stokes–Fourier system. SIAM Journal on Mathematical Analysis, 45(4):2046– 2063, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buttazzo:2018:SOC

[BV18] Giuseppe Buttazzo and Bozhidar Velichkov. A shape optimal control problem with changing sign data. SIAM Journal on Mathematical Analysis, 50(3):2608– 2627, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boeckle:2012:DES

[BW12] Christoph Boeckle and Peter Wittwer. Decay estimates for steady solutions of the Navier–Stokes equations in two dimensions in the presence of a Wall. SIAM Journal on Mathematical Analysis, 44(5):3346–3368, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brancolini:2017:OES

[BW17] Alessio Brancolini and Benedikt Wirth. Optimal energy scaling for micropatterns in transport networks. SIAM Journal on Mathematical Analysis, 49 (1):311–359, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berestycki:2014:ESA

[BWW14] Henri Berestycki, Juncheng Wei, and Matthias Winter. Existence of symmetric and asymmetric spikes for a crime hotspot model. SIAM Journal on Mathematical Analysis, 46 (1):691–719, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bingbing:2015:SDS

[BYH15] Ding Bingbing, Liu Yingbo, and Yin Huicheng. The small data solutions of general 3D quasilinear wave equations. I. SIAM Journal on Mathematical Analysis, 47(6): 4192–4228, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bao:2012:SSL

[BYZ12] Gang Bao, KiHyun Yun, and Zhengfang Zhou. Stability of the scattering from a large electromagnetic cavity in two dimensions. SIAM Journal on Mathematical Analysis, 44(1): 383–404, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p383_s1.

Calvo:2015:ACD

[Cal15] Juan Calvo. Analysis of a class of degenerate parabolic equations with saturation mechanisms. SIAM Journal on Mathematical Analysis, 47 (4):2917–2951, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cances:2010:ABTa

[Can10a] Clément Cancès. Asymptotic behavior of two-phase flows in heterogeneous porous media for capillarity depending only on space. I. Convergence to the optimal entropy solution. SIAM Journal on Mathematical Analysis, 42(2): 946–971, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cances:2010:ABTb

[Can10b] Clément Cancès. Asymptotic behavior of two-phase flows in heterogeneous porous media for capillarity depending only on space. II. Nonclassical shocks to model oiltrapping. SIAM Journal on Mathematical Analysis, 42(2): 972–995, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2019:RBE

[Cao19] Yunbai Cao. Regularity of Boltzmann equation with external fields in convex domains of diffuse reflection. SIAM Journal on Mathematical Analysis, 51(4):3195– 3275, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carbou:2014:MWC

[Car14]  Gilles Carbou.  Metasta-

bility of wall configurations in ferromagnetic nanowires. SIAM Journal on Mathematical Analysis, 46(1):45– 95, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cavalletti:2012:OTB

[Cav12] Fabio Cavalletti. Optimal transport with branching distance costs and the obstacle problem. SIAM Journal on Mathematical Analysis, 44(1): 454–482, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p454_s1.

Chermisi:2010:MRN

[CC10] Milena Chermisi and Sergio Conti. Multiwell rigidity in nonlinear elasticity. SIAM Journal on Mathematical Analysis, 42(5):1986– 2012, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chesnel:2011:CIE

[CC11] Lucas Chesnel and Patrick Ciarlet, Jr. Compact imbeddings in electromagnetism with interfaces between classical materials and metamaterials. SIAM Journal on Mathematical Analysis, 43(5): 2150–2169, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2150_s1.

Cheng:2017:RTM

[CCC^{+}17] Bin Cheng, Jingrui Cheng, Michael Cullen, John Norbury, and Matthew Turner. A rigorous treatment of moist convection in a single column. SIAM Journal on Mathematical Analysis, 49(5): 3854–3892, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cacace:2018:DMG

[CCC18a] Simone Cacace, Fabio Camilli, and Lucilla Corrias. A differential model for growing sandpiles on networks. SIAM Journal on Mathematical Analysis, 50(3):2509– 2535, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cherdantsev:2018:ELE

[CCC18b] Mikhail Cherdantsev, Kirill Cherednichenko, and Shane Cooper. Extreme localization of eigenfunctions to onedimensional high-contrast periodic problems with a defect. SIAM Journal on Mathematical Analysis, 50(6):5825– 5856, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calleja:2017:RSQ

[CCCdlL17] Renato C. Calleja, Alessandra Celletti, Livia Corsi, and Rafael de la Llave. Response solutions for quasiperiodically forced, dissipative wave equations. SIAM Journal on Mathematical Analysis, 49(4):3161–3207, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calleja:2014:EGC

[CCFdlL14] Renato C. Calleja, Alessandra Celletti, Corrado Falcolini, and Rafael de la Llave. An extension of Greene’s criterion for conformally symplectic systems and a partial justification. SIAM Journal on Mathematical Analysis, 46(4): 2350–2384, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cakoni:2010:ITE

[CCG10] Fioralba Cakoni, David Colton, and Drossos Gintides. The interior transmission eigenvalue problem. SIAM Journal on Mathematical Analysis, 42(6): 2912–2921, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p2912_s1.

Chen:2018:MVS

[CCG18] Yongxin Chen, Giovanni Conforti, and Tryphon T. Georgiou. Measure-valued spline curves: an optimal transport viewpoint. SIAM Journal on Mathematical Analysis, 50(6): 5947–5968, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cakoni:2010:ITP

[CCH10] Fioralba Cakoni, David Colton, and Houssem Haddar. The interior transmission problem for regions with cavities. SIAM Journal on Mathematical Analysis, 42(1): 145–162, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2018:HDB

[CCHR18] Chao-Nien Chen, Yung-Sze Choi, Yeyao Hu, and Xiaofeng Ren. Higher dimensional bubble profiles in a sharp interface limit of the FitzHugh– Nagumo system. SIAM Journal on Mathematical Analysis, 50(5):5072–5095, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chang:2018:MME

[CCK18] TongKeun Chang, Hi Jun Choe, and Kyungkeun Kang. On maximum modulus estimates of the Navier–Stokes equations with nonzero boundary data. SIAM Journal on Mathematical Analysis, 50(3): 3147–3171, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carvalho:2013:RID

[CCLCP13] A. N. Carvalho, J. W.

Cholewa, G. Lozada-Cruz, and M. R. T. Primo. Reduction of infinite dimensional systems to finite dimensions: Compact convergence approach. SIAM Journal on Mathematical Analysis, 45(2): 600–638, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Corbera:2015:BRE

[CCLM15] Montserrat Corbera, Josep Cors, Jaume Llibre, and Richard Moeckel. Bifurcation of relative equilibria of the (1 + 3)-body problem. SIAM Journal on Mathematical Analysis, 47(2):1377– 1404, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cantrell:2012:GBS

[CCM12] Robert Stephen Cantrell, Chris Cosner, and Raúl Manásevich. Global bifurcation of solutions for crime modeling equations. SIAM Journal on Mathematical Analysis, 44(3):1340– 1358, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caprino:2014:MCP

[CCM14] Silvia Caprino, Guido Cavallaro, and Carlo Marchioro. On a magnetically confined plasma with infinite charge. SIAM Journal on Mathematical Analysis, 46(1):133– 164, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coclite:2016:MEM

[CCM16] G. M. Coclite, M. M. Coclite, and S. Mishra. On a model for the evolution of morphogens in a growing tissue. SIAM Journal on Mathematical Analysis, 48(3):1575– 1615, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannarsa:2019:GGC

[CCMW19] Piermarco Cannarsa, Wei Cheng, Marco Mazzola, and Kaizhi Wang. Global generalized characteristics for the Dirichlet problem for Hamilton–Jacobi equations at a supercritical energy level. SIAM Journal on Mathematical Analysis, 51(5):4213– 4244, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cacciapuoti:2017:ODD

[CCNP17] Claudio Cacciapuoti, Raffaele Carlone, Diego Noja, and Andrea Posilicano. The onedimensional Dirac equation with concentrated nonlinearity. SIAM Journal on Mathematical Analysis, 49(3):2246– 2268, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carrillo:2015:GSD

[CCV15] José Antonio Carrillo, Daniele Castorina, and Bruno Volzone. Ground states for diffusion dominated free energies with logarithmic interaction. SIAM Journal on Mathematical Analysis, 47(1):1–25, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crasta:2011:CRF

[CD11] Graziano Crasta and Virginia De Cicco. A chain rule formula in the space BV and applications to conservation laws. SIAM Journal on Mathematical Analysis, 43(1): 430–456, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p430_s1.

Cioranescu:2012:PUM

[CDD^{+}12] D. Cioranescu, A. Damlamian, P. Donato, G. Griso, and R. Zaki. The periodic unfolding method in domains with holes. SIAM Journal on Mathematical Analysis, 44(2):

718–760, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cordero:2018:SCI

[CdGDN18] Elena Cordero, Maurice de Gosson, Monika Dörfler, and Fabio Nicola. On the symplectic covariance and interferences of time-frequency distributions. SIAM Journal on Mathematical Analysis, 50(2): 2178–2193, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conti:2011:ABC

[CDK11] Sergio Conti, Georg Dolzmann, and Carolin Kreisbeck. Asymptotic behavior of crystal plasticity with one slip system in the limit of rigid elasticity. SIAM Journal on Mathematical Analysis, 43(5): 2337–2353, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i5/p2337_s1.

Cacciafesta:2016:HDE

[CDL16] Federico Cacciafesta, Piero D’Ancona, and Renato Lucà. Helmholtz and dispersive equations with variable coefficients on exterior domains. SIAM Journal on Mathematical Analysis, 48(3):1798– 1832, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Casado-Diaz:2013:ABN

[CDLLSG13] J. Casado-Díaz, M. Luna-Laynez, and F. J. Suárez-Grau. Asymptotic behavior of the Navier–Stokes system in a thin domain with Navier condition on a slightly rough boundary. SIAM Journal on Mathematical Analysis, 45(3): 1641–1674, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2013:ITL

[CDM13] Li Chen, Donatella Donatelli, and Pierangelo Marcati. Incompressible type limit analysis of a hydrodynamic model for charge-carrier transport. SIAM Journal on Mathematical Analysis, 45(3):915– 933, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cesaroni:2016:HMF

[CDM16] Annalisa Cesaroni, Nicolas Dirr, and Claudio Marchi. Homogenization of a mean field game system in the small noise limit. SIAM Journal on Mathematical Analysis, 48(4): 2701–2729, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cardone:2010:LEE

[CDN10] G. Cardone, T. Durante, and S. A. Nazarov. The localization effect for eigenfunctions of the mixed boundary value problem in a thin cylinder with distorted ends. SIAM Journal on Mathematical Analysis, 42(6):2581– 2609, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cicalese:2016:GST

[CDNP16] M. Cicalese, L. De Luca, M. Novaga, and M. Ponsiglione. Ground states of a two phase model with cross and self attractive interactions. SIAM Journal on Mathematical Analysis, 48(5): 3412–3443, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlier:2017:CES

[CDPS17] Guillaume Carlier, Vincent Duval, Gabriel Peyré, and Bernhard Schmitzer. Convergence of entropic schemes for optimal transport and gradient flows. SIAM Journal on Mathematical Analysis, 49(2): 1385–1418, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choquet:2017:MAS

[CDR17] C. Choquet, M. M. Diédhiou, and C. Rosier. Mathematical analysis of a seawater intrusion model including storativity. SIAM Journal on Mathematical Analysis, 49(1): 29–63, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carles:2010:MWN

[CDS10] Rémi Carles, Eric Dumas, and Christof Sparber. Multiphase weakly nonlinear geometric optics for Schrödinger equations. SIAM Journal on Mathematical Analysis, 42(1): 489–518, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombo:2019:CMO

[CDS19] Maria Colombo, Simone Di Marino, and Federico Stra. Continuity of multimarginal optimal transport with repulsive cost. SIAM Journal on Mathematical Analysis, 51(4): 2903–2926, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cazenave:2013:FTB

[CDW13] Thierry Cazenave, Flávio Dickstein, and Fred B. Weissler. Finite-time blowup for a complex Ginzburg– Landau equation. SIAM Journal on Mathematical Analysis, 45(1):244–266, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2012:GSS

[CDX12] Gui-Qiang G. Chen, Xuemei Deng, and Wei Xiang. Global steady subsonic flows through infinitely long nozzles for the full Euler equations. SIAM Journal on Mathematical Analysis, 44(4): 2888–2919, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2017:IRF

[CDX17] Jianfeng Cheng, Lili Du, and Wei Xiang. Incompressible Réthy flows in two dimensions. SIAM Journal on Mathematical Analysis, 49(5): 3427–3475, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2013:ADS

[CDZ13] Guanggan Chen, Jinqiao Duan, and Jian Zhang. Approximating dynamics of a singularly perturbed stochastic wave equation with a random dynamical boundary condition. SIAM Journal on Mathematical Analysis, 45(5): 2790–2814, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calder:2014:HJE

[CEH14] Jeff Calder, Selim Esedoáğlu, and Alfred O. Hero. A Hamilton–Jacobi equation for the continuum limit of nondominated sorting. SIAM Journal on Mathematical Analysis, 46(1):603–638, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Constantin:2017:RIL

[CEIV17] Peter Constantin, Tarek Elgindi, Mihaela Ignatova, and Vlad Vicol. Remarks on the inviscid limit for the Navier–Stokes equations for uniformly bounded velocity fields. SIAM Journal on Mathematical Analysis, 49(3): 1932–1946, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cortazar:2016:ABO

[CEQW16] Carmen Cortázar, Manuel Elgueta, Fernando Quirós, and Noemí Wolanski. Asymptotic behavior for a onedimensional nonlocal diffusion equation in exterior domains. SIAM Journal on Mathematical Analysis, 48(3): 1549–1574, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cerrai:2011:APS

[Cer11]  Sandra Cerrai.  Averag-

ing principle for systems of reaction–diffusion equations with polynomial nonlinearities perturbed by multiplicative noise. SIAM Journal on Mathematical Analysis, 43(6): 2482–2518, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2482_s1.

Cesana:2011:NEG

[Ces11] Pierluigi Cesana. Nematic elastomers: Gamma-limits for large bodies and small particles. SIAM Journal on Mathematical Analysis, 43(5): 2354–2383, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2354_s1.

Chen:2011:AAN

[CF11] Xinfu Chen and Avner Friedman. Asymptotic analysis for the narrow escape problem. SIAM Journal on Mathematical Analysis, 43(6):2542– 2563, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2542_s1.

Calsina:2014:PSS

[CF14] Àngel Calsina and József Z. Farkas. Positive steady states of evolution equations with finite dimensional nonlinearities. SIAM Journal on Mathematical Analysis, 46(2):1406– 1426, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crevat:2019:RDN

[CFF19] Joachim Crevat, Grégory Faye, and Francis Filbet. Rigorous derivation of the nonlocal reaction-diffusion Fitzhugh–Nagumo system. SIAM Journal on Mathematical Analysis, 51(1):346– 373, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2017:BWS

[CFGL17] Robin Ming Chen, Lili Fan, Hongjun Gao, and Yue Liu. Breaking waves and solitary waves to the rotation-two-component Camassa–Holm system. SIAM Journal on Mathematical Analysis, 49(5): 3573–3602, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cicalese:2019:VAT

[CFO19] Marco Cicalese, Marwin Forster, and Gianluca Orlando. Variational analysis of a two-dimensional frustrated spin system: Emergence and rigidity of chirality transitions. SIAM Journal on Mathematical Analysis, 51(6): 4848–4893, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carrillo:2010:AFD

[CFRT10] J. A. Carrillo, M. Fornasier, J. Rosado, and G. Toscani. Asymptotic flocking dynamics for the kinetic Cucker– Smale model. SIAM Journal on Mathematical Analysis, 42 (1):218–236, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carrillo:2018:SSS

[CFSS18] J. A. Carrillo, S. Fagioli, F. Santambrogio, and M. Schmidtchen. Splitting schemes and segregation in reaction cross-diffusion systems. SIAM Journal on Mathematical Analysis, 50(5): 5695–5718, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Claeys:2010:SAK

[CG10a] Tom Claeys and Tamara Grava. Solitonic asymptotics for the Korteweg–de Vries equation in the small dispersion limit. SIAM Journal on Mathematical Analysis, 42(5): 2132–2154, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coclite:2010:VVT

[CG10b] G. M. Coclite and M. Garavello. Vanishing viscosity for traffic on networks. SIAM Journal on Mathematical Analysis, 42(4):1761– 1783, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cowan:2010:EPD

[CG10c] Craig Cowan and Nassif Ghoussoub. Estimates on pull-in distances in microelectromechanical systems models and other nonlinear eigenvalue problems. SIAM Journal on Mathematical Analysis, 42(5):1949–1966, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombo:2011:STB

[CG11] Maria Colombo and Massimo Gobbino. Slow time behavior of the semidiscrete Perona–Malik scheme in one dimension. SIAM Journal on Mathematical Analysis, 43(6): 2564–2600, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2564_s1.

Cirant:2019:EUS

[CG19] Marco Cirant and Alessandro Goffi. On the existence and uniqueness of solutions to time-dependent fractional MFG. SIAM Journal on Mathematical Analysis, 51(2): 913–954, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cakoni:2010:EID

[CGH10] Fioralba Cakoni, Drossos Gintides, and Houssem Haddar. The existence of an infinite discrete set of transmission eigenvalues. SIAM Journal on Mathematical Analysis, 42(1): 237–255, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Craig:2018:BTS

[CGLS18] Walter Craig, Maxime Gazeau, Christophe Lacave, and Catherine Sulem. Bloch theory and spectral gaps for linearized water waves. SIAM Journal on Mathematical Analysis, 50(5):5477–5501, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cakoni:2016:HSS

[CGM16] Fioralba Cakoni, Bojan B. Guzina, and Shari Moskow. On the homogenization of a scalar scattering problem for highly oscillating anisotropic media. SIAM Journal on Mathematical Analysis, 48(4): 2532–2560, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2013:WPR

[CGP13] Qingshan Chen, Max Gunzburger, and Mauro Perego. Well-posedness results for a nonlinear Stokes problem arising in glaciology. SIAM Journal on Mathematical Analysis, 45(5):2710–2733, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlier:2010:KTB

[CGS10] G. Carlier, A. Galichon, and F. Santambrogio. From Knothe’s transport to Brenier’s map and a continuation method for optimal transport. SIAM Journal on Mathematical Analysis, 41(6): 2554–2576, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colli:2017:GEN

[CGS17] Pierluigi Colli, Gianni Gilardi, and Jürgen Sprekels. Global existence for a nonstandard viscous Cahn–Hilliard system with dynamic boundary condition. SIAM Journal on Mathematical Analysis, 49(3): 1732–1760, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cagnetti:2011:AMM

[CGT11] F. Cagnetti, D. Gomes, and H. V. Tran. Aubry–Mather measures in the nonconvex setting. SIAM Journal on Mathematical Analysis, 43(6): 2601–2629, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2601_s1.

Choi:2011:ABN

[CH11a] Sun-Ho Choi and Seung-Yeal Ha. Asymptotic behavior of the nonlinear Vlasov equation with a self-consistent force. SIAM Journal on Mathematical Analysis, 43(5): 2050–2077, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2050_s1.

Cossonniere:2011:EIT

[CH11b] Anne Cossonnière and Houssem Haddar. The electromagnetic interior transmission problem for regions with cavities. SIAM Journal on Mathematical Analysis, 43(4):1698– 1715, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1698_s1.

Charve:2013:EGS

[CH13] Frédéric Charve and Boris Haspot. Existence of a global strong solution and vanishing capillarity-viscosity limit in one dimension for the Korteweg system. SIAM Journal on Mathematical Analysis, 45 (2):469–494, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2015:SMV

[CH15] I-Kun Chen and Chun-Hsiung Hsia. Singularity of macroscopic variables near boundary for gases with cutoff hard potential. SIAM Journal on Mathematical Analysis, 47(6): 4332–4349, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2019:HCS

[CH19a] Young-Pil Choi and Jan Haskovec. Hydrodynamic Cucker–Smale model with normalized communication weights and time delay. SIAM Journal on Mathematical Analysis, 51(3):2660– 2685, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Claeys:2019:FKB

[CH19b] Xavier Claeys and Ralf Hiptmair. First-kind boundary integral equations for the Hodge–Helmholtz operator. SIAM Journal on Mathematical Analysis, 51(1): 197–227, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chae:2014:TES

[Cha14] Dongho Chae. On the transport equations with singular/regular nonlocal velocities. SIAM Journal on Mathematical Analysis, 46(2):1017– 1029, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2012:SLC

[Che12] Bin Cheng. Singular limits and convergence rates of compressible Euler and rotating shallow water equations. SIAM Journal on Mathematical Analysis, 44(2): 1050–1076, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2014:IAI

[Che14] Bin Cheng. Improved accuracy of incompressible approximation of compressible Euler equations. SIAM Journal on Mathematical Analysis, 46(6): 3838–3864, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2015:BRO

[Che15] Shibing Chen. Boundary C^{1,α} regularity of an optimal transport problem with cost close to −x · y. SIAM Journal on Mathematical Analysis, 47(4): 2689–2698, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2018:RSS

[Che18] I-Kun Chen. Regularity of stationary solutions to the linearized Boltzmann equations. SIAM Journal on Mathematical Analysis, 50(1): 138–161, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2019:BQG

[Che19] Qingshan Chen. The barotropic quasi-geostrophic equation under a free surface. SIAM Journal on Mathematical Analysis, 51(3):1836–1867, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caraballo:2015:NCV

[CHK15] Tomás Caraballo, Xiaoying Han, and Peter E. Kloeden. Nonautonomous chemostats with variable delays. SIAM Journal on Mathematical Analysis, 47(3):2178– 2199, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chaichenets:2019:KTO

[CHKP19] Leonid Chaichenets, Dirk Hundertmark, Peer Kunstmann, and Nikolaos Pattakos. Knocking out teeth in one-dimensional periodic nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 51(5): 3714–3749, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2015:FTS

[CHL15] Geng Chen, Tao Huang, and Chun Liu. Finite time singularities for hyperbolic systems. SIAM Journal on Mathematical Analysis, 47(1): 758–785, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2017:TED

[CHL17] Mi-Ran Choi, Dirk Hundertmark, and Young-Ran Lee. Thresholds for existence of dispersion management solitons for general nonlinearities. SIAM Journal on Mathematical Analysis, 49(2):1519– 1569, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cakoni:2019:FMF

[CHL19] Fioralba Cakoni, Houssem Haddar, and Armin Lechleiter. On the factorization method for a far field inverse scattering problem in the time domain. SIAM Journal on Mathematical Analysis, 51(2): 854–872, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carstea:2018:UIB

[CHN18] Catalin I. Cârstea, Naofumi Honda, and Gen Nakamura. Uniqueness in the inverse boundary value problem for piecewise homogeneous anisotropic elasticity. SIAM Journal on Mathematical Analysis, 50(3):3291– 3302, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2016:GCS

[Cho16] Young-Pil Choi. Global classical solutions and large-time behavior of the two-phase fluid model. SIAM Journal on Mathematical Analysis, 48(5): 3090–3122, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coutand:2013:WPF

[CHS13a] Daniel Coutand, Jason Hole, and Steve Shkoller. Wellposedness of the free-boundary compressible 3-D Euler equations with surface tension and the zero surface tension limit. SIAM Journal on Mathematical Analysis, 45(6): 3690–3767, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuesta:2013:TWK

[CHS13b] Carlota M. Cuesta, Sabine Hittmeir, and Christian Schmeiser. Traveling waves of a kinetic transport model for the KPPfisher equation. SIAM Journal on Mathematical Analysis, 44(6):4128–4146, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choudhury:2018:NLC

[CHT18] Anupam Pal Choudhury, Amru Hussein, and Patrick Tolksdorf. Nematic liquid crystals in Lipschitz domains. SIAM Journal on Mathematical Analysis, 50(4):4282– 4310, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chupin:2014:GER

[Chu14]  Laurent Chupin.  Global

existence results for some viscoelastic models with an integral constitutive law. SIAM Journal on Mathematical Analysis, 46(3):1859– 1873, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2016:IIS

[CHW16] Wentao Cao, Feimin Huang, and Dehua Wang. Isometric immersion of surface with negative Gauss curvature and the Lax–Friedrichs scheme. SIAM Journal on Mathematical Analysis, 48(3): 2227–2249, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2019:GES

[CHY19] Wentao Cao, Feimin Huang, and Difan Yuan. Global entropy solutions to the gas flow in general nozzle. SIAM Journal on Mathematical Analysis, 51(4):3276–3297, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Constantin:2018:ILS

[CIN18] Peter Constantin, Mihaela Ignatova, and Huy Q. Nguyen. Inviscid limit for SQG in bounded domains. SIAM Journal on Mathematical Analysis, 50(6):6196–6207, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cingolani:2019:SWP

[CJ19] Silvia Cingolani and Louis Jeanjean. Stationary waves with prescribed L^{2}-norm for the planar Schrödinger– Poisson system. SIAM Journal on Mathematical Analysis, 51(4):3533–3568, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chang:2019:CNS

[CJN19] T. Chang, B. J. Jin, and A. Novotný. Compressible Navier–Stokes system with general inflow–outflow boundary data. SIAM Journal on Mathematical Analysis, 51(2): 1238–1278, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Capriani:2013:QSO

[CJP13] Giuseppe Maria Capriani, Vesa Julin, and Giovanni Pisante. A quantitative second order minimality criterion for cavities in elastic bodies. SIAM Journal on Mathematical Analysis, 45(3): 1952–1991, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chung:2011:RNP

[CK11] Jaywan Chung and Yong Jung Kim. Relative Newtonian potentials of radial functions and asymptotics in nonlinear diffusion. SIAM Journal on Mathematical Analysis, 43(4): 1975–1994, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1975_s1.

See addendum [CK13].

Chen:2012:SPM

[CK12] Chiun-Chuan Chen and Theodore Kolokolnikov. Simple PDE model of spot replication in any dimension. SIAM Journal on Mathematical Analysis, 44(5):3564–3593, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chung:2013:ARN

[CK13] Jaywan Chung and Yong Jung Kim. Addendum to “Relative Newtonian Potentials of Radial Functions and Asymptotics in Nonlinear Diffusion”. SIAM Journal on Mathematical Analysis, 45(2): 728–731, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [CK11].

Chen:2014:PSW

[CKM14] Chao-Nien Chen, Shih-Yin Kung, and Yoshihisa Morita. Planar standing wavefronts in the FitzHugh–Nagumo equations. SIAM Journal on Mathematical Analysis, 46(1): 657–690, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choulli:2015:SDT

[CKS15] Mourad Choulli, Yavar Kian, and Eric Soccorsi. Stable determination of timedependent scalar potential from boundary measurements in a periodic quantum waveguide. SIAM Journal on Mathematical Analysis, 47(6): 4536–4558, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuesta:2018:SSL

[CKV18] Carlota M. Cuesta, Hans Knüpfer, and Juan J. L. Velázquez. Self-similar lifting and persistent touch-down points in the thin-film equation. SIAM Journal on Mathematical Analysis, 50(2): 1900–1917, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chayes:2013:AED

[CKY13] Lincoln Chayes, Inwon Kim, and Yao Yao. An aggregation equation with degenerate diffusion: Qualitative property of solutions. SIAM Journal on Mathematical Analysis, 45(5): 2995–3018, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2018:FTN

[CKY18] Yuan Chen, Soojung Kim, and Yong Yu. Freedericksz transition in nematic liquid crystal flows in dimension two. SIAM Journal on Mathematical Analysis, 50(5): 4838–4860, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2017:TDS

[CKZ17] Gui-Qiang G. Chen, Jie Kuang, and Yongqian Zhang. Two-dimensional steady supersonic exothermically reacting Euler flow past Lipschitz bending walls. SIAM Journal on Mathematical Analysis, 49 (2):818–873, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2013:APF

[CL13a] Xiuqing Chen and Jian-Guo Liu. Analysis of polymeric flow models and related compactness theorems in weighted spaces. SIAM Journal on Mathematical Analysis, 45(3): 1179–1215, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colli:2013:PFA

[CL13b] Pierluigi Colli and Philippe Laurençot. A phase-field approximation of the Willmore flow with volume and area constraints. SIAM Journal on Mathematical Analysis, 44(6): 3734–3754, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cerrai:2017:APN

[CL17] Sandra Cerrai and Alessandra Lunardi. Averaging principle for nonautonomous slow–fast systems of stochastic reaction–diffusion equations: The almost periodic case. SIAM Journal on Mathematical Analysis, 49(4): 2843–2884, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2018:AGS

[CL18] Zili Chen and Xiuting Li. Asymptotic growth of support and uniform decay of moments for the Vlasov–Poisson system. SIAM Journal on Mathematical Analysis, 50(4): 4180–4202, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chambolle:2017:ROC

[CLLS17] Antonin Chambolle, Jimmy Lamboley, Antoine Lemenant, and Eugene Stepanov. Regularity for the optimal compliance problem with length penalization. SIAM Journal on Mathematical Analysis, 49(2): 1166–1224, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Camano:2017:ESE

[CLM17] Jessika Camaño, Christopher Lackner, and Peter Monk. Electromagnetic Stekloff eigenvalues in inverse scattering. SIAM Journal on Mathematical Analysis, 49(6): 4376–4401, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2012:MDK

[CLW12] Li Chen, Jian-Guo Liu, and Jinhuan Wang. Multidimensional degenerate Keller– Segel system with critical diffusion exponent 2n/(n + 2). SIAM Journal on Mathematical Analysis, 44(2):1077– 1102, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2017:VCT

[CLW17] Shaowei Chen, Zhaoli Liu, and Zhi-Qiang Wang. A variant of Clark’s theorem and its applications for nonsmooth functionals without the Palais–Smale condition. SIAM Journal on Mathematical Analysis, 49(1):446– 470, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Calder:2011:AIS

[CM11] J. Calder and A. Mansouri.

Anisotropic image sharpening via well-posed Sobolev gradient flows. SIAM Journal on Mathematical Analysis, 43(4): 1536–1556, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1536_s1.

Chupin:2012:RDT

[CM12] Laurent Chupin and Sébastien Martin. Rigorous derivation of the thin film approximation with roughness-induced correctors. SIAM Journal on Mathematical Analysis, 44(4): 3041–3070, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ciomaga:2013:PEB

[CM13] Adina Ciomaga and Jean-Michel Morel. A proof of equivalence between level lines shortening and curvature motion in image processing. SIAM Journal on Mathematical Analysis, 45(3): 1047–1067, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2014:DTD

[CM14] Qiang Chen and Peter Monk. Discretization of the time domain CFIE for acoustic scattering problems using convolution quadrature. SIAM Journal on Mathematical Analysis, 46(5):3107– 3130, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Combet:2018:CMS

[CM18] Vianney Combet and Yvan Martel. Construction of multibubble solutions for the critical GKDV equation. SIAM Journal on Mathematical Analysis, 50(4):3715– 3790, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuccagna:2019:SSS

[CM19] Scipio Cuccagna and Masaya Maeda. On stability of small solitons of the 1-D NLS with a trapping delta potential. SIAM Journal on Mathematical Analysis, 51(6):4311– 4331, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Canizo:2010:RCS

[CMM10] José A. Cañizo, Stéphane Mischler, and Clément Mouhot. Rate of convergence to selfsimilarity for Smoluchowski’s coagulation equation with constant coefficients. SIAM Journal on Mathematical Analysis, 41(6):2283–2314, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cocquet:2013:EUS

[CMM13] Pierre-Henri Cocquet, Pierre-Alain Mazet, and Vincent Mouysset. On the existence and uniqueness of a solution for some frequency-dependent partial differential equations coming from the modeling of metamaterials. SIAM Journal on Mathematical Analysis, 44(6):3806–3833, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chambolle:2013:NMC

[CMP13] Antonin Chambolle, Massimiliano Morini, and Marcello Ponsiglione. A nonlocal mean curvature flow and its semiimplicit time-discrete approximation. SIAM Journal on Mathematical Analysis, 44(6): 4048–4077, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cho:2016:PMF

[CMWY16] Yong-Kum Cho, Yoshinori Morimoto, Shuaikun Wang, and Tong Yang. Probability measures with finite moments and the homogeneous Boltzmann equation. SIAM Journal on Mathematical Analysis, 48(4):2399–2413, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2018:GWS

[CMWZ18] Gui-Qiang G. Chen, Apala Majumdar, Dehua Wang, and Rongfang Zhang. Global weak solutions for the compressible active liquid crystal system. SIAM Journal on Mathematical Analysis, 50(4):3632– 3675, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coron:2015:DBC

[CN15] Jean-Michel Coron and Hoai-Minh Nguyen. Dissipative boundary conditions for nonlinear 1-D hyperbolic systems: Sharp conditions through an approach via time-delay systems. SIAM Journal on Mathematical Analysis, 47(3): 2220–2240, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ciaurri:2017:TWM

[CNR17] Óscar Ciaurri, Adam Nowak, and Luz Roncal. Two-weight mixed norm estimates for a generalized spherical mean Radon transform acting on radial functions. SIAM Journal on Mathematical Analysis, 49(6):4402–4439, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campbell:2010:STD

[CNS10] A. Campbell, S. A. Nazarov, and G. H. Sweers. Spectra of two-dimensional models for thin plates with sharp edges. SIAM Journal on Mathematical Analysis, 42(6): 3020–3044, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3020_s1.

Crippa:2017:ELS

[CNSS17] Gianluca Crippa, Camilla Nobili, Christian Seis, and Stefano Spirito. Eulerian and Lagrangian solutions to the continuity and Euler equations with L^{1} vorticity. SIAM Journal on Mathematical Analysis, 49(5):3973–3998, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colin:2012:BSS

[CO12] Mathieu Colin and Masahito Ohta. Bifurcation from semitrivial standing waves and ground states for a system of nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 44(1): 206–223, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p206_s1.

Compaan:2017:SZK

[Com17]  E. Compaan.  Smoothing

for the Zakharov and Klein– Gordon–Schrödinger systems on Euclidean spaces. SIAM Journal on Mathematical Analysis, 49(5):4206–4231, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Consiglieri:2012:TES

[Con12] Luisa Consiglieri. Thermal expansion on Stokes–Fourier systems. SIAM Journal on Mathematical Analysis, 44(3): 1831–1860, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conti:2016:BMG

[COS16] Sergio Conti, Felix Otto, and Sylvia Serfaty. Branched microstructures in the Ginzburg– Landau model of type-I superconductors. SIAM Journal on Mathematical Analysis, 48(4): 2994–3034, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choksi:2010:SVF

[CP10] Rustum Choksi and Mark A.

Peletier. Small volume fraction limit of the diblock copolymer problem: I. sharp-interface functional. SIAM Journal on Mathematical Analysis, 42(3):1334– 1370, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choksi:2011:SVF

[CP11] Rustum Choksi and Mark A. Peletier. Small volumefraction limit of the diblock copolymer problem: II. Diffuse-interface functional. SIAM Journal on Mathematical Analysis, 43(2):739– 763, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p739_s1.

Cances:2012:ERM

[CP12] Clément Cancès and Michel Pierre. An existence result for multidimensional immiscible two-phase flows with discontinuous capillary pressure field. SIAM Journal on Mathematical Analysis, 44(2): 966–992, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2013:FOC

[CP13] Jun Chen and Xing-Bin Pan. Functionals with operator curl in an extended magnetostatic Born–Infeld model. SIAM Journal on Mathematical Analysis, 45(4):2253– 2284, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Contreras:2019:FCF

[CP19] Andres Contreras and Guanying Peng. First critical field of highly anisotropic threedimensional superconductors via a vortex density model. SIAM Journal on Mathematical Analysis, 51(6):4490– 4519, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Contreras:2018:OSD

[CPP18] Andres Contreras, Dmitry E.

Pelinovsky, and Michael Plum. Orbital stability of domain walls in coupled Gross– Pitaevskii systems. SIAM Journal on Mathematical Analysis, 50(1):810–833, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carrillo:2016:CPM

[CPSW16] J. A. Carrillo, F. S. Patacchini, P. Sternberg, and G. Wolansky. Convergence of a particle method for diffusive gradient flows in one dimension. SIAM Journal on Mathematical Analysis, 48(6): 3708–3741, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chugunova:2010:NSL

[CPT10] Marina Chugunova, M. C.

Pugh, and R. M. Taranets. Nonnegative solutions for a long-wave unstable thin film equation with convection. SIAM Journal on Mathematical Analysis, 42(4):1826– 1853, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2017:SFC

[CPZ17] Geng Chen, Ronghua Pan, and Shengguo Zhu. Singularity formation for the compressible Euler equations. SIAM Journal on Mathematical Analysis, 49(4):2591– 2614, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2012:TDR

[CQ12] Shuxing Chen and Aifang Qu. Two-dimensional Riemann problems for Chaplygin gas. SIAM Journal on Mathematical Analysis, 44(3): 2146–2178, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2019:TWN

[CQ19] Xinfu Chen and Yuanwei Qi. Traveling wave to non-KPP isothermal diffusion systems: Existence of minimum speed and sharp bounds. SIAM Journal on Mathematical Analysis, 51(2):1436– 1453, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cortazar:2018:NFA

[CQW18] Carmen Cortázar, Fernando Quirós, and Noemí Wolanski. Near-field asymptotics for the porous medium equation in exterior domains. the critical two-dimensional case. SIAM Journal on Mathematical Analysis, 50(3):2664– 2680, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2018:SFG

[CQX18] Bin Cheng, Peng Qu, and Chunjing Xie. Singularity formation and global existence of classical solutions for onedimensional rotating shallow water system. SIAM Journal on Mathematical Analysis, 50(3):2486–2508, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chiron:2010:KKL

[CR10] D. Chiron and F. Rousset.

The KdV/KP-I limit of the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 42(1): 64–96, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombo:2018:NCL

[CR18] Rinaldo M. Colombo and Elena Rossi. Nonlocal conservation laws in bounded domains. SIAM Journal on Mathematical Analysis, 50(4): 4041–4065, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crismale:2019:ASF

[Cri19] Vito Crismale. On the approximation of SBD functions and some applications. SIAM Journal on Mathematical Analysis, 51(6):5011– 5048, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cavaterra:2016:GSS

[CRWX16] Cecilia Cavaterra, Elisabetta Rocca, Hao Wu, and Xiang Xu. Global strong solutions of the full Navier–Stokes and Q-tensor system for nematic liquid crystal flows in two dimensions. SIAM Journal on Mathematical Analysis, 48(2): 1368–1399, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2010:FSW

[CS10a] Linan Chen and Daniel W. Stroock. The fundamental solution to the Wright–Fisher equation. SIAM Journal on Mathematical Analysis, 42(2): 539–567, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:2010:INS

[CS10b] C. H. Arthur Cheng and Steve Shkoller. The interaction of the 3D Navier– Stokes equations with a moving nonlinear Koiter elastic shell. SIAM Journal on Mathematical Analysis, 42(3): 1094–1155, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheskidov:2014:EET

[CS14a] A. Cheskidov and R. Shvydkoy. Euler equations and turbulence: Analytical approach to intermittency. SIAM Journal on Mathematical Analysis, 46(1):353–374, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choffrut:2014:WSS

[CS14b] A. Choffrut and L. Székelyhidi, Jr. Weak solutions to the stationary incompressible Euler equations. SIAM Journal on Mathematical Analysis, 46(6): 4060–4074, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cullen:2014:MFA

[CS14c] Mike Cullen and Marc Sedjro. On a model of forced axisymmetric flows. SIAM Journal on Mathematical Analysis, 46(6):3983–4013, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carter:2015:FPO

[CS15a] Paul Carter and Björn Sandstede. Fast pulses with oscillatory tails in the FitzHugh– Nagumo system. SIAM Journal on Mathematical Analysis, 47(5):3393–3441, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2015:CEB

[CS15b] Xuwen Chen and Walter A. Strauss. Convergence to equilibrium of a body moving in a kinetic sea. SIAM Journal on Mathematical Analysis, 47(6): 4630–4651, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chung:2017:ITA

[CS17] Francis J. Chung and John C. Schotland. Inverse transport and acousto-optic imaging. SIAM Journal on Mathematical Analysis, 49(6):4704– 4721, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cances:2018:VWI

[CS18] Eric Cancès and L. Ridgway Scott. Van der Waals interactions between two hydrogen atoms: The Slater–Kirkwood method revisited. SIAM Journal on Mathematical Analysis, 50(1):381–410, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coifman:2017:CFH

[CStW17] Ronald R. Coifman, Stefan Steinerberger, and Hau tieng Wu. Carrier frequencies, holomorphy, and unwinding. SIAM Journal on Mathematical Analysis, 49(6): 4838–4864, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cavalletti:2015:SPG

[CSW15] Fabio Cavalletti, Marc Sedjro, and Michael Westdickenberg. A simple proof of global existence for the 1D pressureless gas dynamics equations. SIAM Journal on Mathematical Analysis, 47(1): 66–79, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:2018:SHS

[CSZ18] Albert Cohen, Christoph Schwab, and Jakob Zech. Shape holomorphy of the stationary Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 50(2): 1720–1752, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cornalba:2019:RDK

[CSZ19] Federico Cornalba, Tony Shardlow, and Johannes Zimmer. A regularized Dean– Kawasaki model: Derivation and analysis. SIAM Journal on Mathematical Analysis, 51(2):1137–1187, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Christoforou:2011:RCV

[CT11] Cleopatra Christoforou and Konstantina Trivisa. Rate of convergence for vanishing viscosity approximations to hyperbolic balance laws. SIAM Journal on Mathematical Analysis, 43(5):2307– 2336, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i5/p2307_s1.

Costin:2014:AAB

[CT14] O. Costin and S. Tanveer. Analytical approximation of the Blasius similarity solution with rigorous error bounds. SIAM Journal on Mathematical Analysis, 46(6):3782– 3813, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choulli:2015:NSE

[CT15] Mourad Choulli and Faouzi Triki. New stability estimates for the inverse medium problem with internal data. SIAM Journal on Mathematical Analysis, 47(3):1778– 1799, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Craig:2016:CRN

[CT16] Katy Craig and Ihsan Topaloglu.

Convergence of regularized nonlocal interaction energies. SIAM Journal on Mathematical Analysis, 48(1):34– 60, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Costin:2013:LSO

[CTW13] O. Costin, S. Tanveer, and M. I. Weinstein. The lifetime of shape oscillations of a bubble in an unbounded, inviscid, and compressible fluid with surface tension. SIAM Journal on Mathematical Analysis, 45(5):2924–2936, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2017:LBS

[CTW17] Xinfu Chen, Je-Chiang Tsai, and Yaping Wu. Longtime behavior of solutions of a SIS epidemiological model. SIAM Journal on Mathematical Analysis, 49(5):3925– 3950, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cui:2013:ASS

[Cui13]  Shangbin Cui.  Asymp-

totic stability of the stationary solution for a parabolichyperbolic free boundary problem modeling tumor growth. SIAM Journal on Mathematical Analysis, 45(5): 2870–2893, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuesta:2012:AOD

[CV12] C. M. Cuesta and J. J. L. Velázquez. Analysis of oscillations in a drainage equation. SIAM Journal on Mathematical Analysis, 44(3): 1588–1616, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choi:2015:STS

[CV15] Kyudong Choi and Alexis F. Vasseur. Short-time stability of scalar viscous shocks in the inviscid limit by the relative entropy method. SIAM Journal on Mathematical Analysis, 47(2):1405–1418, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Capitanelli:2016:DQF

[CV16] Raffaela Capitanelli and Maria Agostina Vivaldi. Dynamical quasi-filling fractal layers. SIAM Journal on Mathematical Analysis, 48(6):

3931–3961, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coron:2013:OFS

[CW13] Jean-Michel Coron and Zhiqiang Wang. Output feedback stabilization for a scalar conservation law with a nonlocal velocity. SIAM Journal on Mathematical Analysis, 45(5): 2646–2665, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chae:2016:PRN

[CW16] Dongho Chae and Jörg Wolf.

On partial regularity for the 3D nonstationary Hall magnetohydrodynamics equations on the plane. SIAM Journal on Mathematical Analysis, 48 (1):443–469, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chandler-Wilde:2010:VAW

[CWE10] Simon N. Chandler-Wilde and Johannes Elschner. Variational approach in weighted Sobolev spaces to scattering by unbounded rough surfaces. SIAM Journal on Mathematical Analysis, 42(6): 2554–2580, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chandler-Wilde:2018:WPP

[CWH18] Simon N. Chandler-Wilde and David P. Hewett. Well-posed PDE and integral equation formulations for scattering by fractal screens. SIAM Journal on Mathematical Analysis, 50 (1):677–717, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:2014:IME

[CWY14] Chongsheng Cao, Jiahong Wu, and Baoquan Yuan. The 2D incompressible magnetohydrodynamics equations with only magnetic diffusion. SIAM Journal on Mathematical Analysis, 46(1): 588–602, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cui:2016:DRN

[CWYZ16] Haibo Cui, Wenjun Wang, Lei Yao, and Changjiang Zhu. Decay rates for a nonconservative compressible generic twofluid model. SIAM Journal on Mathematical Analysis, 48 (1):470–512, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2015:GSS

[CY15] Shuxing Chen and Chao Yi. Global solutions for supersonic flow past a delta wing. SIAM Journal on Mathematical Analysis, 47(1):80– 126, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cui:2018:GEI

[CY18] Xiufang Cui and Silu Yin.

Global existence of inhomogeneous incompressible isotropic elastodynamics in three dimensions. SIAM Journal on Mathematical Analysis, 50(5): 4721–4751, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:2019:KCS

[CY19] Zili Chen and Xiuxia Yin. The kinetic Cucker–Smale model: Well-posedness and asymptotic behavior. SIAM Journal on Mathematical Analysis, 51(5):3819–3853, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deAraujo:2018:ESL

[dAdM18] Anderson L. A. de Araujo and Paulo M. D. de Magalhães. Existence of solutions and local null controllability for a model of tissue invasion by solid tumors. SIAM Journal on Mathematical Analysis, 50(4):3598–3631, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:2013:LTB

[Daf13] Constantine M. Dafermos. Long time behavior of periodic solutions to scalar conservation laws in several space dimensions. SIAM Journal on Mathematical Analysis, 45(4): 2064–2070, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:2014:BSH

[Daf14] Constantine M. Dafermos. BV solutions of hyperbolic balance laws with relaxation in the absence of conserved quantities. SIAM Journal on Mathematical Analysis, 46(6): 4014–4034, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:2019:UZR

[Daf19] Constantine M. Dafermos. Uniqueness of zero relaxation limit. SIAM Journal on Mathematical Analysis, 51(3): 1999–2018, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2010:SRC

[Dai10] Shibin Dai. On the shortening rate of collections of plane convex curves by the area-preserving mean curvature flow. SIAM Journal on Mathematical Analysis, 42(1): 323–333, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2017:RPN

[Dai17] Mimi Dai. Regularity problem for the nematic LCD system with Q-tensor in R^{3}. SIAM Journal on Mathematical Analysis, 49(6):5007– 5030, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Danilov:2017:NNE

[Dan17] V. G. Danilov. Nonsmooth nonoscillating exponentialtype asymptotics for linear parabolic PDE. SIAM Journal on Mathematical Analysis, 49(5):3550–3572, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Demchenko:2019:KVV

[DAP19] Hanna Demchenko, Andrii Anikushyn, and Michael Pokojovy. On a Kelvin– Voigt viscoelastic wave equation with strong delay. SIAM Journal on Mathematical Analysis, 51(6):4382–4412, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

daCosta:2016:RCS

[dCPS16] Fernando P. da Costa, João T.

Pinto, and Rafael Sasportes.

Rates of convergence to scaling profiles in a submonolayer deposition model and the preservation of memory of the initial condition. SIAM Journal on Mathematical Analysis, 48(2):1109–1127, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Danchin:2016:LMN

[DD16] Raphaël Danchin and Bernard Ducomet. The low Mach number limit for a barotropic model of radiative flow. SIAM Journal on Mathematical Analysis, 48(2):1025– 1053, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DallAcqua:2018:OPE

[DD18] Anna Dall’Acqua and Klaus Deckelnick. An obstacle problem for elastic graphs. SIAM Journal on Mathematical Analysis, 50(1):119– 137, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeBouard:2018:LTB

[DDF18] Anne De Bouard, Arnaud Debussche, and Reika Fukuizumi. Long time behavior of Gross– Pitaevskii equation at positive temperature. SIAM Journal on Mathematical Analysis, 50(6):5887–5920, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dalibard:2018:HFA

[DDGVM18] Anne-Laure Dalibard, Helge Dietert, David Gérard-Varet, and Frédéric Marbach. High frequency analysis of the unsteady interactive boundary layer model. SIAM Journal on Mathematical Analysis, 50(4): 4203–4245, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Debiec:2018:REM

[DDGW18] Tomasz Debiec, Marie Doumic, Piotr Gwiazda, and Emil Wiedemann. Relative entropy method for measure solutions of the growth-fragmentation equation. SIAM Journal on Mathematical Analysis, 50(6): 5811–5824, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DalMaso:2011:EUR

[DDM11] Gianni Dal Maso, Antonio DeSimone, and Marco Morandotti. An existence and uniqueness result for the motion of selfpropelled microswimmers. SIAM Journal on Mathematical Analysis, 43(3):1345– 1368, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1345_s1.

Debussche:2015:RRQ

[DdMH15] Arnaud Debussche, Sylvain de Moor, and Martina Hofmanová. A regularity result for quasilinear stochastic partial differential equations of parabolic type. SIAM Journal on Mathematical Analysis, 47(2):1590–1614, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dirr:2018:SHF

[DDMM18] Nicolas Dirr, Federica Dragoni, Paola Mannucci, and Claudio Marchi. Stochastic homogenization for functionals with anisotropic rescaling and noncoercive Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 50(5):5198–5242, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeRosa:2018:MAE

[De 18] Antonio De Rosa. Minimization of anisotropic energies in classes of rectifiable varifolds. SIAM Journal on Mathematical Analysis, 50(1): 162–181, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dekeyser:2019:ASV

[Dek19] Justin Dekeyser. Asymptotic of steady vortex pair in the lake equation. SIAM Journal on Mathematical Analysis, 51(2):1209–1237, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Delgadino:2018:COD

[Del18] Matias G. Delgadino. Convergence of a one-dimensional Cahn–Hilliard equation with degenerate mobility. SIAM Journal on Mathematical Analysis, 50(4):4457–4482, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Despres:2014:SVP

[Des14] Bruno Després. Symmetrization of Vlasov–Poisson equations. SIAM Journal on Mathematical Analysis, 46(4): 2554–2580, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deuring:2013:SDT

[Deu13] Paul Deuring. Spatial decay of time-dependent incompressible Navier–Stokes flows with nonzero velocity at infinity. SIAM Journal on Mathematical Analysis, 45(3):1388– 1421, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desvillettes:2010:LTA

[DF10] Laurent Desvillettes and Klemens Fellner. Large time asymptotics for a continuous coagulation-fragmentation model with degenerate size-dependent diffusion. SIAM Journal on Mathematical Analysis, 41(6): 2315–2334, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dias:2011:SWL

[DF11] João Paulo Dias and Hermano Frid. Short wave-long wave interactions for compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 43(2): 764–787, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p764_s1.

DePhilippis:2013:SRM

[DF13] Guido De Philippis and Alessio Figalli. Sobolev regularity for Monge–Ampère type equations. SIAM Journal on Mathematical Analysis, 45(3): 1812–1824, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davoli:2015:CRF

[DF15] Elisa Davoli and Gilles A. Francfort. A critical revisiting of finite elasto-plasticity.

SIAM Journal on Mathematical Analysis, 47(1):526– 565, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Francesco:2014:ABG

[DFHM14] Marco Di Francesco, Massimo Fornasier, Jan-Christian Hütter, and Daniel Matthes. Asymptotic behavior of gradient flows driven by nonlocal power repulsion and attraction potentials in one dimension. SIAM Journal on Mathematical Analysis, 46(6): 3814–3837, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davini:2014:LMCa

[DFP14a] C. Davini, L. Freddi, and R. Paroni. Linear models for composite thin-walled beams by Γ-convergence. Part I: Open cross sections. SIAM Journal on Mathematical Analysis, 46(5):3296– 3331, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davini:2014:LMCb

[DFP14b] C. Davini, L. Freddi, and R. Paroni. Linear models for composite thin-walled beams by Γ-convergence. Part II: Closed cross-sections. SIAM Journal on Mathematical Analysis, 46(5):3332– 3360, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desvillettes:2017:TER

[DFT17] Laurent Desvillettes, Klemens Fellner, and Bao Quoc Tang. Trend to equilibrium for reaction-diffusion systems arising from complex balanced chemical reaction networks. SIAM Journal on Mathematical Analysis, 49(4): 2666–2709, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dipierro:2018:DED

[DFV18] Serena Dipierro, Alberto Farina, and Enrico Valdinoci. Density estimates for degenerate double-well potentials. SIAM Journal on Mathematical Analysis, 50(6):6333– 6347, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Debussche:2011:SCH

[DG11] Arnaud Debussche and Ludovic Goudenège. Stochastic Cahn–Hilliard equation with double singular nonlinearities and two reflections. SIAM Journal on Mathematical Analysis, 43(3):1473– 1494, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.

siam.org/sima/resource/
1/sjmaah/v43/i3/p1473_s1.

Donato:2016:EHS

[DG16] Patrizia Donato and Daniela Giachetti. Existence and homogenization for a singular problem through rough surfaces. SIAM Journal on Mathematical Analysis, 48(6): 4047–4086, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeBievre:2016:PIV

[DGV16] Stephan De Bièvre, Thierry Goudon, and Arthur Vavasseur. Particles interacting with a vibrating medium: Existence of solutions and convergence to the Vlasov–Poisson system. SIAM Journal on Mathematical Analysis, 48(6):3984– 4020, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dyson:2010:EAP

[DGVBW10] Janet Dyson, Stephen A.

Gourley, Rosanna Villella-Bressan, and Glenn F. Webb. Existence and asymptotic properties of solutions of a nonlocal evolution equation modeling cell–cell adhesion. SIAM Journal on Mathematical Analysis, 42(4):1784– 1804, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2010:NRD

[DH10] Yihong Du and Sze-Bi Hsu. On a nonlocal reaction– diffusion problem arising from the modeling of phytoplankton growth. SIAM Journal on Mathematical Analysis, 42(3): 1305–1333, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deHoop:2014:EAE

[dHGR14] Maarten V. de Hoop, Karlheinz Gröchenig, and José Luis Romero. Exact and approximate expansions with pure Gaussian wave packets. SIAM Journal on Mathematical Analysis, 46(3):2229– 2253, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deHoop:2014:RCE

[dHHI^{+}14] Maarten V. de Hoop, Sean F. Holman, Einar Iversen, Matti Lassas, and Bjørn Ursin. Reconstruction of a conformally Euclidean metric from local boundary diffraction travel times. SIAM Journal on Mathematical Analysis, 46(6): 3705–3726, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Doelman:2014:MPS

[DHPW14] Arjen Doelman, Gurgen Hayrapetyan, Keith Promislow, and Brian Wetton. Meander and pearling of singlecurvature bilayer interfaces in the functionalized Cahn– Hilliard equation. SIAM Journal on Mathematical Analysis, 46(6):3640–3677, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duchene:2015:NFJ

[DIT15] Vincent Duchêne, Samer Israwi, and Raafat Talhouk. A new fully justified asymptotic model for the propagation of internal waves in the Camassa–Holm regime. SIAM Journal on Mathematical Analysis, 47(1):240– 290, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daus:2016:HLM

[DJMZ16] Esther S. Daus, Ansgar Jüngel, Clément Mouhot, and Nicola Zamponi. Hypocoercivity for a linearized multispecies Boltzmann system. SIAM Journal on Mathematical Analysis, 48(1):538– 568, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2011:PES

[DK11] Hongjie Dong and Doyoon Kim. Parabolic and elliptic systems in divergence form with variably partially BMO coefficients. SIAM Journal on Mathematical Analysis, 43(3): 1075–1098, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1075_s1.

Dong:2014:PES

[DK14] Hongjie Dong and Doyoon Kim. Parabolic equations in simple convex polytopes with time irregular coefficients. SIAM Journal on Mathematical Analysis, 46(3): 1789–1819, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deuring:2011:PDL

[DKN11] Paul Deuring, Stanislav Kracmar, and Sárka Necasová. On pointwise decay of linearized stationary incompressible viscous flow around rotating and translating bodies. SIAM Journal on Mathematical Analysis, 43(2):705– 738, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p705_s1.

Disser:2015:OSR

[DKR15] Karoline Disser, Hans-Christoph Kaiser, and Joachim Rehberg. Optimal Sobolev regularity for linear second-order divergence elliptic operators occurring in real-world problems. SIAM Journal on Mathematical Analysis, 47(3): 1719–1746, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Detmann:2016:SUP

[DKR16] Bettina Detmann, Pavel Krejcí, and Elisabetta Rocca. Solvability of an unsaturated porous media flow problem with thermomechanical interaction. SIAM Journal on Mathematical Analysis, 48(6): 4175–4201, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dull:2016:VWA

[DKS16] Wolf-Patrick Düll, Kourosh Sanei Kashani, and Guido Schneider. The validity of Whitham’s approximation for a Klein– Gordon–Boussinesq model. SIAM Journal on Mathematical Analysis, 48(6):4311– 4334, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Disconzi:2019:LIR

[DKT19] Marcelo M. Disconzi, Igor Kukavica, and Amjad Tuffaha. A Lagrangian interior regularity result for the incompressible free boundary Euler equation with surface tension. SIAM Journal on Mathematical Analysis, 51(5):

3982–4022, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2010:SVD

[DL10] Yihong Du and Zhigui Lin. Spreading-vanishing dichotomy in the diffusive logistic model with a free boundary. SIAM Journal on Mathematical Analysis, 42(1):377–405, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [DL13].

Du:2013:ESV

[DL13] Yihong Du and Zhigui Lin. Erratum: Spreading-Vanishing Dichotomy in the Diffusive Logistic Model with a Free Boundary. SIAM Journal on Mathematical Analysis, 45(3): 1995–1996, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [DL10].

deLaire:2014:MET

[dL14] André de Laire. Minimal energy for the traveling waves of the Landau–Lifshitz equation. SIAM Journal on Mathematical Analysis, 46(1):96– 132, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:2015:PEG

[DL15a] Weiwei Ding and Xing Liang. Principal eigenvalues of generalized convolution operators on the circle and spreading speeds of noncompact evolution systems in periodic media. SIAM Journal on Mathematical Analysis, 47(1):855– 896, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2015:SRW

[DL15b] Renjun Duan and Shuangqian Liu. Stability of the rarefaction wave of the Vlasov– Poisson–Boltzmann system. SIAM Journal on Mathematical Analysis, 47(5):3585– 3647, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DiMarino:2018:ERM

[DL18] Simone Di Marino and Jean Louet. The entropic regularization of the Monge problem on the real line. SIAM Journal on Mathematical Analysis, 50(4):3451–3477, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

delaHoz:2016:DCS

[dlHHMV16] Francisco de la Hoz, Taoufik Hmidi, Joan Mateu, and Joan Verdera. Doubly connected V -states for the planar Euler equations. SIAM Journal on Mathematical Analysis, 48(3): 1892–1928, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

delaLlave:2017:REC

[dlLSZ17] Rafael de la Llave, Xifeng Su, and Lei Zhang. Resonant equilibrium configurations in quasi-periodic media: KAM theory. SIAM Journal on Mathematical Analysis, 49 (1):597–625, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desvillettes:2014:EDC

[DLM14] L. Desvillettes, Th. Lepoutre, and A. Moussa. Entropy, duality, and cross diffusion. SIAM Journal on Mathematical Analysis, 46(1): 820–853, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diening:2012:PRM

[DLSV12] L. Diening, D. Lengeler, B. Stroffolini, and A. Verde. Partial regularity for minimizers of quasi-convex functionals with general growth. SIAM Journal on Mathematical Analysis, 44(5):3594– 3616, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeCicco:2010:LSS

[DLV10] Virginia De Cicco, Chiara Leone, and Anna Verde. Lower semicontinuity in SBV for integrals with variable growth. SIAM Journal on Mathematical Analysis, 42(6): 3112–3128, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3112_s1.

DiCristo:2013:SEI

[DLVW13] M. Di Cristo, C.-L. Lin, S. Vessella, and J.-N. Wang. Size estimates of the inverse inclusion problem for the shallow shell equation. SIAM Journal on Mathematical Analysis, 45(1):88– 100, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2012:LEC

[DLZ12a] Ben Duan, Zhen Luo, and Yuxi Zheng. Local existence of classical solutions to shallow water equations with Cauchy data containing vacuum. SIAM Journal on Mathematical Analysis, 44(2): 541–567, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i2/p541_s1.

Duan:2012:CPC

[DLZ12b] Renjun Duan, Qingqing Liu, and Changjiang Zhu. The Cauchy problem on the compressible two-fluids Euler–Maxwell equations. SIAM Journal on Mathematical Analysis, 44(1):102– 133, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p102_s1.

Du:2015:NDP

[DLZ15] Yihong Du, Bendong Lou, and Maolin Zhou. Nonlinear diffusion problems with free boundaries: Convergence, transition speed, and zero number arguments. SIAM Journal on Mathematical Analysis, 47(5):3555– 3584, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ducrot:2014:ABN

[DM14] Arnaud Ducrot and Pierre Magal. Asymptotic behavior of a nonlocal diffusive logistic equation. SIAM Journal on Mathematical Analysis, 46(3): 1731–1753, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donatelli:2015:QNL

[DM15] Donatella Donatelli and Pierangelo Marcati. Quasi-neutral limit, dispersion, and oscillations for Korteweg-type fluids. SIAM Journal on Mathematical Analysis, 47(3):2265– 2282, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deMonvel:2010:PTA

[dMIS10] Anne Boutet de Monvel, Alexander Its, and Dmitry Shepelsky. Painlevé-type asymptotics for the Camassa– Holm equation. SIAM Journal on Mathematical Analysis, 42(4):1854–1873, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2014:SES

[DMZ14] Yihong Du, Hiroshi Matsuzawa, and Maolin Zhou. Sharp estimate of the spreading speed determined by nonlinear free boundary problems. SIAM Journal on Mathematical Analysis, 46(1): 375–396, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daus:2019:ADS

[DMZ19] Esther S. Daus, Pina Milisić, and Nicola Zamponi. Analysis of a degenerate and singular volume-filling crossdiffusion system modeling biofilm growth. SIAM Journal on Mathematical Analysis, 51(4):3569–3605, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Decker:2012:PFW

[DN12] R. Decker and V. W. Noonburg. A periodically forced Wilson–Cowan system with multiple attractors. SIAM Journal on Mathematical Analysis, 44(2):887–905, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Drivas:2018:OCA

[DN18] Theodore D. Drivas and Huy Q. Nguyen. Onsager’s conjecture and anomalous dissipation on domains with boundary. SIAM Journal on Mathematical Analysis, 50(5): 4785–4811, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dharmawardane:2012:DES

[DNK12] Priyanjana M. N. Dharmawardane, Tohru Nakamura, and Shuichi Kawashima. Decay estimates of solutions for quasi-linear hyperbolic systems of viscoelasticity. SIAM Journal on Mathematical Analysis, 44(3):1976– 2001, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolbeault:2012:PLS

[DNS12] Jean Dolbeault, Bruno Nazaret, and Giuseppe Savaré. From Poincaré to logarithmic Sobolev inequalities: a gradient flow approach. SIAM Journal on Mathematical Analysis, 44(5): 3186–3216, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dondl:2019:PFA

[DNWW19] Patrick W. Dondl, Matteo Novaga, Benedikt Wirth, and Stephan Wojtowytsch. A phase-field approximation of the perimeter under a connectedness constraint. SIAM Journal on Mathematical Analysis, 51(5):3902– 3920, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deguchi:2016:PSG

[DO16] Hideo Deguchi and Michael Oberguggenberger. Propagation of singularities for generalized solutions to wave equations with discontinuous coefficients. SIAM Journal on Mathematical Analysis, 48(1): 397–442, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2011:SSM

[Don11] Hongjie Dong. On similarity solutions to the multidimensional aggregation equation. SIAM Journal on Mathematical Analysis, 43(4): 1995–2008, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1995_s1.

Damascelli:2013:SRC

[DP13] Lucio Damascelli and Filomena Pacella. Symmetry results for cooperative elliptic systems via linearization. SIAM Journal on Mathematical Analysis, 45(3):1003– 1026, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dumas:2014:JLK

[DP14] Eric Dumas and Dmitry Pelinovsky. Justification of the log–KdV equation in granular chains: The case of precompression. SIAM Journal on Mathematical Analysis, 46(6): 4075–4103, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2015:CGE

[DP15] Shibin Dai and Keith Promislow. Competitive geometric evolution of amphiphilic interfaces. SIAM Journal on Mathematical Analysis, 47(1): 347–380, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:2013:ECS

[DR13] Yanheng Ding and Bernhard Ruf. Existence and concentration of semiclassical solutions for Dirac equations with critical nonlinearities. SIAM Journal on Mathematical Analysis, 44(6):3755–3785, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deRijk:2018:SSS

[dR18] Björn de Rijk. Spectra and stability of spatially periodic pulse patterns II: The critical spectral curve. SIAM Journal on Mathematical Analysis, 50(2):1958–2019, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deRijk:2016:SSS

[dRDR16] Björn de Rijk, Arjen Doelman, and Jens Rademacher. Spectra and stability of spatially periodic pulse patterns: Evans function factorization via Riccati transformation. SIAM Journal on Mathematical Analysis, 48(1): 61–121, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Drouot:2018:BSR

[Dro18]  Alexis Drouot.  Bound

states for rapidly oscillatory Schrödinger operators in dimension 2. SIAM Journal on Mathematical Analysis, 50(2): 1471–1484, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dauge:2010:STP

[DS10a] Monique Dauge and Rob Stevenson. Sparse tensor product wavelet approximation of singular functions. SIAM Journal on Mathematical Analysis, 42(5):2203– 2228, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2010:ETT

[DS10b] Shengfu Deng and Shu-Ming Sun. Exact theory of threedimensional water waves at the critical speed. SIAM Journal on Mathematical Analysis, 42(6):2721–2761, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DiRusso:2013:EAB

[DS13] Cristiana Di Russo and Alice Sepe. Existence and asymptotic behavior of solutions to a quasi-linear hyperbolicparabolic model of vasculogenesis. SIAM Journal on Mathematical Analysis, 45(2): 748–776, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2014:ABS

[DS14] Mimi Dai and Maria Schonbek. Asymptotic behavior of solutions to the liquid crystal system in H^{m}(R^{3}). SIAM Journal on Mathematical Analysis, 46(5):3131– 3150, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davoli:2019:DPP

[DS19] Elisa Davoli and Ulisse Stefanelli. Dynamic perfect plasticity as convex minimization. SIAM Journal on Mathematical Analysis, 51(2):672– 730, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dipierro:2015:NFB

[DSV15] Serena Dipierro, Ovidiu Savin, and Enrico Valdinoci. A nonlocal free boundary problem. SIAM Journal on Mathematical Analysis, 47(6): 4559–4605, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2017:IRG

[DSX17] Bin Dong, Zuowei Shen, and Peichu Xie. Image restoration: a general wavelet frame based model and its asymptotic analysis. SIAM Journal on Mathematical Analysis, 49 (1):421–445, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2018:SPI

[DSY18] Qingquan Deng, Avy Soffer, and Xiaohua Yao. Solitonpotential interactions for nonlinear Schrödinger equation in R^{3}. SIAM Journal on Mathematical Analysis, 50(5):5243– 5292, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DiFratta:2019:SPT

[DSZ19] Giovanni Di Fratta, Valeriy Slastikov, and Arghir Zarnescu. On a sharp Poincaré-type inequality on the 2-sphere and its application in micromagnetics. SIAM Journal on Mathematical Analysis, 51(4):3373– 3387, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Droniou:2014:MDM

[DT14] Jérôme Droniou and Kyle S. Talbot. On a miscible displacement model in porous media flow with measure data. SIAM Journal on Mathematical Analysis, 46(5):3158– 3175, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Plinio:2015:GST

[DT15] Francesco Di Plinio and Roger Temam. Grisvard’s shift theorem near L^{∞} and Yudovich theory on polygonal domains. SIAM Journal on Mathematical Analysis, 47(1): 159–178, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2011:DPV

[Dua11]  Renjun Duan.  Dissipa-

tive property of the Vlasov– Maxwell–Boltzmann system with a uniform ionic background. SIAM Journal on Mathematical Analysis, 43(6): 2732–2757, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2732_s1.

Duchene:2010:ASW

[Duc10] Vincent Duchêne. Asymptotic shallow water models for internal waves in a twofluid system with a free surface. SIAM Journal on Mathematical Analysis, 42(5): 2229–2260, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duerinckx:2016:MFL

[Due16]  Mitia Duerinckx.  Mean-

field limits for some Riesz interaction gradient flows. SIAM Journal on Mathematical Analysis, 48(3):2269– 2300, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DiCristo:2010:SDD

[DV10] Michele Di Cristo and Sergio Vessella. Stable determination of the discontinuous conductivity coefficient of a parabolic equation. SIAM Journal on Mathematical Analysis, 42(1): 183–217, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duchene:2015:OLP

[DVW15] V. Duchêne, I. Vukićević, and M. I. Weinstein. Oscillatory and localized perturbations of periodic structures and the bifurcation of defect modes. SIAM Journal on Mathematical Analysis, 47(5): 3832–3883, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2013:STR

[DW13] Yi Du and Keyan Wang. Space-time regularity of the Koch and Tataru solutions to the liquid crystal equations. SIAM Journal on Mathematical Analysis, 45(6): 3838–3853, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2018:TDF

[DWX18] Xuemei Deng, Tian-Yi Wang, and Wei Xiang. Threedimensional full Euler flows with nontrivial swirl in axisymmetric nozzles. SIAM Journal on Mathematical Analysis, 50(3):2740–2772, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:2012:VCL

[DWY12] Shijin Deng, Weike Wang, and Shih-Hsien Yu. Viscous conservation laws with boundary. SIAM Journal on Mathematical Analysis, 44(4):2695– 2755, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:2012:GSS

[DWYZ12] Shijin Ding, Huanyao Wen, Lei Yao, and Changjiang Zhu. Global spherically symmetric classical solution to compressible Navier–Stokes equations with large initial data and vacuum. SIAM Journal on Mathematical Analysis, 44(2): 1257–1278, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2010:VSH

[DWZ10] Qiang Du, Juncheng Wei, and Chunyi Zhao. Vortex solutions of the highκ high-field Ginzburg–Landau model with an applied current. SIAM Journal on Mathematical Analysis, 42(6): 2368–2401, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2019:GED

[DX19] Hongjie Dong and Longjuan Xu. Gradient estimates for divergence form elliptic systems arising from composite material. SIAM Journal on Mathematical Analysis, 51(3): 2444–2478, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dai:2018:GPH

[DXZ18] Dan Dai, Shuai-Xia Xu, and Lun Zhang. Gap probability at the hard edge for random matrix ensembles with pole singularities in the potential. SIAM Journal on Mathematical Analysis, 50(2): 2233–2279, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duan:2010:SOS

[DY10] Renjun Duan and Tong Yang. Stability of the one-species Vlasov–Poisson–Boltzmann system. SIAM Journal on Mathematical Analysis, 41(6): 2353–2387, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dong:2019:EPL

[DYZ19a] Hongjie Dong, Tong Yang, and Mingying Zhong. Exterior problem of the linear Vlasov– Poisson–Boltzmann system. SIAM Journal on Mathematical Analysis, 51(3):1792– 1823, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2019:ESS

[DYZ19b] Yi Du, Wang Yang, and Yi Zhou. On the exponential stability of a stratified flow to the 2D ideal MHD equations with damping. SIAM Journal on Mathematical Analysis, 51(6):5077–5102, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davini:2014:ASW

[DZ14] Andrea Davini and Maxime Zavidovique. Aubry sets for weakly coupled systems of Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 46(5):3361– 3389, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:2015:GWP

[DZ15] Lili Du and Deqin Zhou.

Global well-posedness of twodimensional magnetohydrodynamic flows with partial dissipation and magnetic diffusion. SIAM Journal on Mathematical Analysis, 47(2): 1562–1589, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ehrnstrom:2011:SWW

[EEW11] Mats Ehrnström, Joachim Escher, and Erik Wahlén. Steady water waves with multiple critical layers. SIAM Journal on Mathematical Analysis, 43(3):1436–1456, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1436_s1.

Elliott:2015:TPS

[EF15] Charles M. Elliott and Hans Fritz. Time-periodic solutions of advection-diffusion equations on moving hypersurfaces. SIAM Journal on Mathematical Analysis, 47(3): 1693–1718, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ebenbeck:2019:CHB

[EG19] Matthias Ebenbeck and Harald Garcke. On a Cahn– Hilliard–Brinkman model for tumor growth and its singular limits. SIAM Journal on Mathematical Analysis, 51(3): 1868–1912, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elschner:2013:ESU

[EH13] Johannes Elschner and Guanghui Hu. Elastic scattering by unbounded rough surfaces. SIAM Journal on Mathematical Analysis, 44(6):4101– 4127, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soufi:2016:POO

[EH16] Ahmad El Soufi and Evans M.

Harrell II. On the placement of an obstacle so as to optimize the Dirichlet heat trace. SIAM Journal on Mathematical Analysis, 48(2): 884–894, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evers:2016:MVM

[EHM16] Joep H. M. Evers, Sander C.

Hille, and Adrian Muntean. Measure-valued mass evolution problems with flux boundary conditions and solution-dependent velocities. SIAM Journal on Mathematical Analysis, 48(3):1929– 1953, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eremenko:2011:SGP

[EI11] Alexandre Eremenko and Sergei Ivanov. Spectra of the Gurtin–Pipkin type equations. SIAM Journal on Mathematical Analysis, 43(5): 2296–2306, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2296_s1.

Endal:2014:CBN

[EJ14] J. Endal and E. R. Jakobsen. L^{1} contraction for bounded (nonintegrable) solutions of degenerate parabolic equations. SIAM Journal on Mathematical Analysis, 46(6): 3957–3982, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erdos:2018:PLD

[EKR18] László Erdös, Torben Krüger, and David Renfrew. Power law decay for systems of randomly coupled differential equations. SIAM Journal on Mathematical Analysis, 50(3): 3271–3290, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esfahani:2017:SSW

[EL17] Amin Esfahani and Steve Levandosky. Stability of solitary waves of the Kadomtsev– Petviashvili equation with a weak rotation. SIAM Journal on Mathematical Analysis, 49(6):5096–5133, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eller:2012:SHB

[Ell12] Matthias Eller. On symmetric hyperbolic boundary problems with nonhomogeneous conservative boundary conditions. SIAM Journal on Mathematical Analysis, 44(3): 1925–1949, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Epstein:2010:WFD

[EM10] Charles L. Epstein and Rafe Mazzeo. Wright–Fisher diffusion in one dimension. SIAM Journal on Mathematical Analysis, 42(2):568– 608, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Escauriaza:2017:ASP

[EMZ17] Luis Escauriaza, Santiago Montaner, and Can Zhang. Analyticity of solutions to parabolic evolutions and applications. SIAM Journal on Mathematical Analysis, 49(5): 4064–4092, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:2012:AAO

[EP12] Alexander Elbert and Grigory Panasenko. Asymptotic analysis of the one-dimensional diffusion-absorption equation with rapidly and strongly oscillating absorption coefficient. SIAM Journal on Mathematical Analysis, 44(3): 2099–2119, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elling:2012:SSS

[ER12] Volker Elling and Joseph Roberts. Steady and selfsimilar inviscid flow. SIAM Journal on Mathematical Analysis, 44(4):2344–2371, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eychenne:2019:SDB

[ER19] Arnaud Eychenne and Nicolas Rougerie. On the stability of 2D dipolar Bose–Einstein condensates. SIAM Journal on Mathematical Analysis, 51(2): 1371–1386, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:2017:CBS

[ERV17] Charles M. Elliott, Thomas Ranner, and Chandrasekhar Venkataraman. Coupled bulksurface free boundary problems arising from a mathematical model of receptorligand dynamics. SIAM Journal on Mathematical Analysis, 49(1):360–397, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Egger:2010:ARP

[ES10] Herbert Egger and Matthias Schlottbom. Analysis and regularization of problems in diffuse optical tomography. SIAM Journal on Mathematical Analysis, 42(5):1934– 1948, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esselborn:2016:RRP

[Ess16]  Elias Esselborn.  Relax-

ation rates for a perturbation of a stationary solution to the thin-film equation. SIAM Journal on Mathematical Analysis, 48(1): 349–396, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Es-Sarhir:2012:ESC

[ESvR12] Abdelhadi Es-Sarhir and Max-K. von Renesse. Ergodicity of stochastic curve shortening flow in the plane. SIAM Journal on Mathematical Analysis, 44(1):224– 244, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p224_s1.

Evans:2016:ASK

[ET16] Lawrence C. Evans and Peyam R. Tabrizian. Asymptotics for scaled Kramers– Smoluchowski equations. SIAM Journal on Mathematical Analysis, 48(4):2944–2961, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erdogan:2013:HFP

[ETZ13] M. B. Erdogan, N. Tzirakis, and V. Zharnitsky. High frequency perturbation of cnoidal waves in KdV. SIAM Journal on Mathematical Analysis, 44(6):4147– 4164, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evje:2011:WSG

[Evj11] Steinar Evje. Weak solutions for a gas-liquid model relevant for describing gas-kick in oil wells. SIAM Journal on Mathematical Analysis, 43(4): 1887–1922, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1887_s1.

Evje:2013:CTP

[Evj13a]  Steinar Evje.  A com-

pressible two-phase model with pressure-dependent wellreservoir interaction. SIAM Journal on Mathematical Analysis, 45(2):518–546, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evje:2013:GTP

[Evj13b] Steinar Evje. Genuine twophase flow dynamics with a free interface separating gas-liquid mixture from gas. SIAM Journal on Mathematical Analysis, 45(5):2894– 2923, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elgindi:2015:SDE

[EW15a] Tarek M. Elgindi and Klaus Widmayer. Sharp decay estimates for an anisotropic linear semigroup and applications to the surface quasi-geostrophic and inviscid Boussinesq systems. SIAM Journal on Mathematical Analysis, 47(6): 4672–4684, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evje:2015:GSV

[EW15b] Steinar Evje and Huanyao Wen. Global solutions of a viscous gas-liquid model with unequal fluid velocities in a closed conduit. SIAM Journal on Mathematical Analysis, 47 (1):381–406, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evje:2018:STF

[EW18] Steinar Evje and Huanyao Wen. A Stokes two-fluid model for cell migration that can account for physical cues in the microenvironment. SIAM Journal on Mathematical Analysis, 50(1): 86–118, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ehrnstrom:2019:EET

[EW19] Mats Ehrnström and Yuexun Wang. Enhanced existence time of solutions to the fractional Korteweg–de Vries equation. SIAM Journal on Mathematical Analysis, 51(4): 3298–3323, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Faierman:2014:ITP

[Fai14] Melvin Faierman. The interior transmission problem: Spectral theory. SIAM Journal on Mathematical Analysis, 46(1): 803–819, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fajman:2016:LWP

[Faj16] David Fajman. Local wellposedness for the Einstein– Vlasov system. SIAM Journal on Mathematical Analysis, 48(5):3270–3321, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fehrman:2013:SHM

[Feh13] Benjamin J. Fehrman. Stochastic homogenization of monotone systems of viscous Hamilton–Jacobi equations with convex nonlinearities. SIAM Journal on Mathematical Analysis, 45(4):2441– 2476, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feldman:2018:SSP

[Fel18] William M. Feldman. Stability of Serrin’s problem and dynamic stability of a model for contact angle motion. SIAM Journal on Mathematical Analysis, 50(3): 3303–3326, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ferreira:2012:RHM

[FF12] Rita Ferreira and Irene Fonseca. Reiterated homogenization in BV via multiscale convergence. SIAM Journal on Mathematical Analysis, 44(3): 2053–2098, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Foldes:2017:AAR

[FFGHR17] Juraj Földes, Susan Friedlander, Nathan Glatt-Holtz, and Geordie Richards. Asymptotic analysis for randomly forced MHD. SIAM Journal on Mathematical Analysis, 49(6):4440–4469, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fischer:2015:FSP

[FG15] Julian Fischer and Günther Grün. Finite speed of propagation and waiting times for the stochastic porous medium equation: a unifying approach. SIAM Journal on Mathematical Analysis, 47(1): 825–854, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ferreira:2018:EWS

[FG18a] Rita Ferreira and Diogo Gomes. Existence of weak solutions to stationary meanfield games through variational inequalities. SIAM Journal on Mathematical Analysis, 50(6):5969–6006, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fischer:2018:EPS

[FG18b] Julian Fischer and Günther Grün. Existence of positive solutions to stochastic thinfilm equations. SIAM Journal on Mathematical Analysis, 50 (1):411–455, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fontelos:2011:PFM

[FGJ11] M. A. Fontelos, G. Grün, and S. Jörres. On a phasefield model for electrowetting and other electrokinetic phenomena. SIAM Journal on Mathematical Analysis, 43(1): 527–563, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p527_s1.

Feireisl:2012:SLC

[FGN12] Eduard Feireisl, Isabelle Gallagher, and Antonín Novotný. A singular limit for compressible rotating fluids. SIAM Journal on Mathematical Analysis, 44(1):192–205, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sima/resource/
1/sjmaah/v44/i1/p192_s1.

Felea:2013:MAS

[FGN13] Raluca Felea, Romina Gaburro, and Clifford J. Nolan. Microlocal analysis of SAR imaging of a dynamic reflectivity function. SIAM Journal on Mathematical Analysis, 45(5): 2767–2789, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Figalli:2015:CID

[FGR15] Alessio Figalli, Thomas O.

Gallouët, and Ludovic Rifford. On the convexity of injectivity domains on nonfocal manifolds. SIAM Journal on Mathematical Analysis, 47(2): 969–1000, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2013:FBT

[FGW13] Chun Fang, Mats Gyllenberg, and Yi Wang. Floquet bundles for tridiagonal competitive-cooperative systems and the dynamics of time-recurrent systems. SIAM Journal on Mathematical Analysis, 45(4):2477– 2498, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frid:2011:SMN

[FHK11] Hermano Frid, Helge Holden, and Kenneth H. Karlsen. L^{∞} solutions for a model of nonisothermal polytropic gas flow. SIAM Journal on Mathematical Analysis, 43(5): 2253–2274, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2253_s1.

Friedman:2013:DAL

[FHK13] Avner Friedman, Bei Hu, and James P. Keener. The diffusion approximation for linear nonautonomous reactionhyperbolic equations. SIAM Journal on Mathematical Analysis, 45(4):2285–2298, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Freddi:2016:VMA

[FHMP16] Lorenzo Freddi, Peter Hornung, Maria Giovanna Mora, and Roberto Paroni. A variational model for anisotropic and naturally twisted ribbons. SIAM Journal on Mathematical Analysis, 48(6): 3883–3906, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fontelos:2016:SIB

[FHO16] Marco A. Fontelos, Hyung Ju Hwang, and Youngmin Oh. Stability, instability, and bifurcation in electrified thin films. SIAM Journal on Mathematical Analysis, 48(4): 2730–2782, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:2010:AMM

[FHX10] Avner Friedman, Bei Hu, and Chuan Xue. Analysis of a mathematical model of ischemic cutaneous wounds. SIAM Journal on Mathematical Analysis, 42(5):2013– 2040, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Focardi:2014:AAA

[FI14] M. Focardi and F. Iurlano. Asymptotic analysis of Ambrosio– Tortorelli energies in linearized elasticity. SIAM Journal on Mathematical Analysis, 46(4):2936–2955, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fischer:2013:ADS

[Fis13] Julian Fischer. Advectiondriven support shrinking in a chemotaxis model with degenerate mobility. SIAM Journal on Mathematical Analysis, 45(3):1585–1615, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frometa:2018:SLL

[FJ18] Susana Frómeta and Milton Jara. Scaling limit for a long-range divisible sandpile. SIAM Journal on Mathematical Analysis, 50(3):2317– 2361, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frank:2013:HKM

[FK13] Rupert L. Frank and Hynek Kovarík. Heat kernels of metric trees and applications. SIAM Journal on Mathematical Analysis, 45(3):1027– 1046, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedrich:2018:PNL

[FK18] Manuel Friedrich and Martin Kruzík. On the passage from nonlinear to linearized viscoelasticity. SIAM Journal on Mathematical Analysis, 50(4): 4426–4456, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fedotov:2019:CWM

[FK19a] Alexander Fedotov and Frédéric Klopp. The complex WKB method for difference equations and Airy functions. SIAM Journal on Mathematical Analysis, 51(6):4413– 4447, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feizmohammadi:2019:RNC

[FK19b] Ali Feizmohammadi and Yavar Kian. Recovery of nonsmooth coefficients appearing in anisotropic wave equations. SIAM Journal on Mathematical Analysis, 51(6): 4953–4976, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fernandez:2016:EUR

[FKM^{+}16] J. R. Fernández, P. Kalita, S. Migórski, M. C. Muñiz, and C. Núñez. Existence and uniqueness results for a kinetic model in bulksurface surfactant dynamics.

SIAM Journal on Mathematical Analysis, 48(5):3065– 3089, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feireisl:2019:LMN

[FKM19] Eduard Feireisl, Christian Klingenberg, and Simon Markfelder. On the low Mach number limit for the compressible Euler system. SIAM Journal on Mathematical Analysis, 51(2):1496– 1513, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feireisl:2014:ILF

[FKN^{+}14] Eduard Feireisl, Ondrej Kreml, Sárka Necasová, Jirí Neustupa, and Jan Stebel. Incompressible limits of fluids excited by moving boundaries. SIAM Journal on Mathematical Analysis, 46(2):1456– 1471, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feireisl:2015:SIR

[FKV15] E. Feireisl, O. Kreml, and A. Vasseur. Stability of the isentropic Riemann solutions of the full multidimensional Euler system. SIAM Journal on Mathematical Analysis, 47(3):2416–2425, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fan:2012:SWP

[FL12a] Haitao Fan and Xiao-Biao Lin. Standing waves for phase transitions in a spherically symmetric nozzle. SIAM Journal on Mathematical Analysis, 44(1):405–436, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http:

//epubs.siam.org/sima/resource/
1/sjmaah/v44/i1/p405_s1.

Frouvelle:2012:DKM

[FL12b] Amic Frouvelle and Jian-Guo Liu. Dynamics in a kinetic model of oriented particles with phase transition. SIAM Journal on Mathematical Analysis, 44(2):791– 826, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frank:2015:CLA

[FL15] Rupert L. Frank and Elliott H. Lieb. A compactness lemma and its application to the existence of minimizers for the liquid drop model. SIAM Journal on Mathematical Analysis, 47(6): 4436–4450, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:2017:WAT

[FL17] Irene Fonseca and Pan Liu. The weighted Ambrosio– Tortorelli approximation scheme. SIAM Journal on Mathematical Analysis, 49(6):4491– 4520, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Flandoli:2019:KEA

[FL19] Franco Flandoli and Dejun Luo. Kolmogorov equations associated to the stochastic two dimensional Euler equations. SIAM Journal on Mathematical Analysis, 51(3): 1761–1791, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frid:2018:MAT

[FMP18] Hermano Frid, Daniel R. Marroquin, and Ronghua Pan. Modeling Aurora type phenomena by short wave-long wave interactions in multidimensional large magnetohydrodynamic flows. SIAM Journal on Mathematical Analysis, 50(6):6156– 6195, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fanzon:2019:DLP

[FPP19] Silvio Fanzon, Mariapia Palombaro, and Marcello Ponsiglione. Derivation of linearized polycrystals from a two-dimensional system of edge dislocations. SIAM Journal on Mathematical Analysis, 51(5):3956–3981, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Furioli:2012:GCL

[FPTT12] G. Furioli, A. Pulvirenti, E. Terraneo, and G. Toscani. The grazing collision limit of the inelastic Kac model around a Lévy-type equilibrium. SIAM Journal on Mathematical Analysis, 44(2): 827–850, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ferreira:2013:NNS

[FPVR13] Lucas C. F. Ferreira, Gabriela Planas, and Elder J. Villamizar-Roa. On the nonhomogeneous Navier–Stokes system with Navier friction boundary conditions. SIAM Journal on Mathematical Analysis, 45(4): 2576–2595, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frid:2014:GSS

[FPZ14] Hermano Frid, Ronghua Pan, and Weizhe Zhang. Global smooth solutions in R^{3} to short wave-long wave interactions systems for viscous compressible fluids. SIAM Journal on Mathematical Analysis, 46(3):1946–1968, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Felea:2011:MPL

[FQ11] Raluca Felea and Eric Todd Quinto. The microlocal properties of the local 3-D SPECT operator. SIAM Journal on Mathematical Analysis, 43(3): 1145–1157, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1145_s1.

Frikel:2016:LDP

[FQ16] Jürgen Frikel and Eric Todd Quinto. Limited data problems for the generalized Radon transform in R^{n}. SIAM Journal on Mathematical Analysis, 48(4):2301– 2318, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Felmer:2010:RPS

[FQS10] Patricio Felmer, Alexander Quaas, and Boyan Sirakov. Resonance phenomena for second-order stochastic control equations. SIAM Journal on Mathematical Analysis, 42 (3):997–1024, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fischer:2017:LPL

[FR17] Julian Fischer and Claudia Raithel. Liouville principles and a large-scale regularity theory for random elliptic operators on the halfspace. SIAM Journal on Mathematical Analysis, 49(1): 82–114, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedrich:2018:PKI

[Fri18] Manuel Friedrich. A piecewise Korn inequality in SBD and applications to embedding and density results. SIAM Journal on Mathematical Analysis, 50(4):3842– 3918, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friesecke:2019:SCM

[Fri19]  Gero Friesecke.  A sim-

ple counterexample to the Monge ansatz in multimarginal optimal transport, convex geometry of the set of Kantorovich plans, and the Frenkel–Kontorova model. SIAM Journal on Mathematical Analysis, 51(6):4332– 4355, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fan:2019:ASR

[FRX19] Lili Fan, Lizhi Ruan, and Wei Xiang. Asymptotic stability of rarefaction wave for the inflow problem governed by the one-dimensional radiative Euler equations. SIAM Journal on Mathematical Analysis, 51 (1):595–625, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fishman:2014:MTA

[FS14] Shmuel Fishman and Avy Soffer. Multiscale time averaging, reloaded. SIAM Journal on Mathematical Analysis, 46(2): 1385–1405, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frehse:2015:RTD

[FS15] Jens Frehse and Sebastian Schwarzacher. On regularity of the time derivative for degenerate parabolic systems. SIAM Journal on Mathematical Analysis, 47(5): 3917–3943, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feldman:2013:LSS

[FT13] Mikhail Feldman and Adrian Tudorascu. On Lagrangian solutions for the semi-geostrophic system with singular initial data. SIAM Journal on Mathematical Analysis, 45(3): 1616–1640, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feischl:2017:ERS

[FT17] Michael Feischl and Thanh Tran. Existence of regular solutions of the Landau– Lifshitz–Gilbert equation in 3D with natural boundary conditions. SIAM Journal on Mathematical Analysis, 49(6): 4470–4490, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friesecke:2018:BCD

[FV18] Gero Friesecke and Daniela Vögler. Breaking the curse of dimension in multi-marginal Kantorovich optimal transport on finite state spaces. SIAM Journal on Mathematical Analysis, 50(4):3996– 4019, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Faver:2018:EDF

[FW18a] Timothy E. Faver and J. Douglas Wright. Exact diatomic Fermi–Pasta–Ulam– Tsingou solitary waves with optical band ripples at infinity. SIAM Journal on Mathematical Analysis, 50(1): 182–250, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Franz:2018:CSP

[FW18b] Tino Franz and Holger Wendland. Convergence of the smoothed particle hydrodynamics method for a specific barotropic fluid flow: Constructive kernel theory. SIAM Journal on Mathematical Analysis, 50(5):4752– 4784, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feng:2017:S

[FWW17] Lirui Feng, Yi Wang, and Jianhong Wu. Semiflows “Monotone with respect to high-rank cones” on a Banach space. SIAM Journal on Mathematical Analysis, 49 (1):142–161, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fei:2013:BSD

[FY13a] Mingwen Fei and Huicheng Yin. Bound states of 2-D nonlinear Schrödinger equations with potentials tending to zero at infinity. SIAM Journal on Mathematical Analysis, 45(4):2299–2331, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Felmer:2013:FPF

[FY13b] Patricio Felmer and Miguel Yangari. Fast propagation for fractional KPP equations with slowly decaying initial conditions. SIAM Journal on Mathematical Analysis, 45(2): 662–678, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2014:TWM

[FZ14] Jian Fang and Xiao-Qiang Zhao. Traveling waves for monotone semiflows with weak compactness. SIAM Journal on Mathematical Analysis, 46(6):3678–3704, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2016:GSO

[FZ16] Daoyuan Fang and Ruizhao Zi. Global solutions to the Oldroyd-B model with a class of large initial data. SIAM Journal on Mathematical Analysis, 48(2):1054– 1084, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fang:2018:GSI

[FZZ18] Daoyuan Fang, Ting Zhang, and Ruizhao Zi. Global solutions to the isentropic compressible Navier–Stokes equations with a class of large initial data. SIAM Journal on Mathematical Analysis, 50(5): 4983–5026, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garnier:2011:ASI

[Gar11]  Jimmy Garnier.  Acceler-

ating solutions in integrodifferential equations. SIAM Journal on Mathematical Analysis, 43(4):1955–1974, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1955_s1.

Gastel:2019:RIC

[Gas19] Andreas Gastel. Regularity issues for Cosserat continua and p-harmonic maps. SIAM Journal on Mathematical Analysis, 51(6):4287– 4310, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

GarcIa-Cervera:2018:SMP

[GCGJL18] Carlos J. GarcÍa-Cervera, Tiziana Giorgi, Sookyung Joo, and Xin Yang Lu. Switching mechanism in the B_{1RevT ilted} phase of bent-core liquid crystals. SIAM Journal on Mathematical Analysis, 50(5):4889–4913, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gess:2013:FSP

[Ges13] Benjamin Gess. Finite speed of propagation for stochastic porous media equations. SIAM Journal on Mathematical Analysis, 45(5):2734– 2766, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glitzky:2010:EBS

[GG10] Annegret Glitzky and Klaus Gärtner. Existence of bounded steady state solutions to spin-polarized driftdiffusion systems. SIAM Journal on Mathematical Analysis, 41(6):2489–2513, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2014:RAS

[GGAS14] Hongjun Gao, María J.

Garrido-Atienza, and Björn Schmalfuss. Random attractors for stochastic evolution equations driven by fractional Brownian motion. SIAM Journal on Mathematical Analysis, 46(4):2281– 2309, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guillen-Gonzalez:2014:WTR

[GGRB14] Francisco Guillén-González and MarÍa Ángeles RodrÍguez-Bellido. Weak time regularity and uniqueness for a Q-tensor model. SIAM Journal on Mathematical Analysis, 46(5): 3540–3567, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenlee:2010:QIR

[GH10] W. M. Greenlee and L. Hermi. Quadratic interpolation and Rayleigh–Ritz methods for bifurcation coefficients. SIAM Journal on Mathematical Analysis, 42(6):2987–3019, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v42/i6/p2987_s1.

Guglielmi:2012:AER

[GH12] Nicola Guglielmi and Ernst Hairer. Asymptotic expansions for regularized statedependent neutral delay equations. SIAM Journal on Mathematical Analysis, 44(4): 2428–2458, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grabovsky:2014:ESE

[GH14] Yury Grabovsky and Davit Harutyunyan. Exact scaling exponents in Korn and Korntype inequalities for cylindrical shells. SIAM Journal on Mathematical Analysis, 46(5): 3277–3295, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grebenkov:2018:SPB

[GH18] Denis S. Grebenkov and Bernard Helffer. On spectral properties of the Bloch– Torrey operator in two dimensions. SIAM Journal on Mathematical Analysis, 50(1): 622–676, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grebenkov:2017:CAO

[GHH17] Denis S. Grebenkov, Bernard Helffer, and Raphael Henry. The complex Airy operator on the line with a semipermeable barrier. SIAM Journal on Mathematical Analysis, 49(3): 1844–1894, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghisi:2013:ALM

[Ghi13]  Marina Ghisi.  Asymp-

totic limits for mildly degenerate Kirchhoff equations. SIAM Journal on Mathematical Analysis, 45(3):1886– 1906, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2013:MSO

[GHLN13] Zihua Guo, Nakao Hayashi, Yiquan Lin, and Pavel I. Naumkin. Modified scattering operator for the derivative nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 45(6): 3854–3871, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goudon:2010:NSV

[GHMZ10] Thierry Goudon, Lingbing He, Ayman Moussa, and Ping Zhang. The Navier–Stokes– Vlasov–Fokker–Planck system near equilibrium. SIAM Journal on Mathematical Analysis, 42(5):2177–2202, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Geronimo:2015:HBA

[GI15] Jeffrey S. Geronimo and Plamen Iliev. A hypergeometric basis for the Alpert multiresolution analysis. SIAM Journal on Mathematical Analysis, 47 (1):654–668, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomelli:2015:FSP

[Gia15] Lorenzo Giacomelli. Finite speed of propagation and waiting time phenomena for degenerate parabolic equations with linear growth Lagrangian. SIAM Journal on Mathematical Analysis, 47(3): 2426–2441, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giesselmann:2014:REA

[Gie14] Jan Giesselmann. A relative entropy approach to convergence of a low order approximation to a nonlinear elasticity model with viscosity and capillarity. SIAM Journal on Mathematical Analysis, 46(5): 3518–3539, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ginster:2019:SGP

[Gin19]  Janusz Ginster.  Strain-

gradient plasticity as the Γlimit of a nonlinear dislocation energy with mixed growth. SIAM Journal on Mathematical Analysis, 51(4): 3424–3464, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gaitan:2013:IPT

[GIP^{+}13] P. Gaitan, H. Isozaki, O. Poisson, S. Siltanen, and J. P. Tamminen. Inverse problems for time-dependent singular heat conductivities — onedimensional case. SIAM Journal on Mathematical Analysis, 45(3):1675–1690, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Geronimo:2017:AML

[GIV17] Jeffrey S. Geronimo, Plamen Iliev, and Walter Van Assche. Alpert multiwavelets and Legendre–Angelesco multiple orthogonal polynomials. SIAM Journal on Mathematical Analysis, 49(1):626– 645, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gwiazda:2012:MDC

[GJMC12] Piotr Gwiazda, Grzegorz Jamróz, and Anna Marciniak-Czochra. Models of discrete and continuous cell differentiation in the framework of transport equation. SIAM Journal on Mathematical Analysis, 44(2):1103– 1133, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2015:SVC

[GJZ15] Yan Guo, Song Jiang, and Chunhui Zhou. Steady viscous compressible channel flows. SIAM Journal on Mathematical Analysis, 47(5): 3648–3670, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gyongy:2010:AFD

[GK10] István Gyöngy and Nicolai Krylov. Accelerated finite difference schemes for linear stochastic partial differential equations in the whole space. SIAM Journal on Mathematical Analysis, 42(5): 2275–2296, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gu:2018:FRS

[GKR18] Yu Gu, Tomasz Komorowski, and Lenya Ryzhik. Fluctuations of random semilinear advection equations. SIAM Journal on Mathematical Analysis, 50(5):5293– 5336, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gerolin:2019:NOT

[GKR19] Augusto Gerolin, Anna Kausamo, and Tapio Rajala. Nonexistence of optimal transport maps for the multimarginal repulsive harmonic cost. SIAM Journal on Mathematical Analysis, 51(3): 2359–2371, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2012:OSR

[GL12] Kanghui Guo and Demetrio Labate. Optimally sparse representations of 3D data with C^{2} surface singularities using Parseval frames of shearlets. SIAM Journal on Mathematical Analysis, 44(2):851– 886, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grubic:2015:ASO

[GL15] Nastasia Grubic and Philippe G. LeFloch. On the area of the symmetry orbits in weakly regular Einstein–Euler spacetimes with Gowdy symmetry. SIAM Journal on Mathematical Analysis, 47(1):669– 683, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2017:GCS

[GL17] Yu Gao and Jian-Guo Liu. Global convergence of a sticky particle method for the modified Camassa–Holm equation. SIAM Journal on Mathematical Analysis, 49(2):1267– 1294, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gontier:2019:SSB

[GL19a] David Gontier and Mathieu Lewin. Spin symmetry breaking in the translationinvariant Hartree–Fock electron gas. SIAM Journal on Mathematical Analysis, 51(4): 3388–3423, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gravina:2019:HOG

[GL19b] Giovanni Gravina and Giovanni Leoni. Higher order gamma-limits for singularly perturbed Dirichlet– Neumann problems. SIAM Journal on Mathematical Analysis, 51(4):3337–3372, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glasner:2017:MED

[Gla17] Karl Glasner. Multilayered equilibria in a density functional model of copolymersolvent mixtures. SIAM Journal on Mathematical Analysis, 49(2):1593–1620, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glitzky:2013:EMS

[Gli13] Annegret Glitzky. An electronic model for solar cells including active interfaces and energy resolved defect densities. SIAM Journal on Mathematical Analysis, 44(6):3874– 3900, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2017:WSC

[GLL17] Yuan Gao, Jian-Guo Liu, and Jianfeng Lu. Weak solution of a continuum model for vicinal surface in the attachmentdetachment-limited regime. SIAM Journal on Mathematical Analysis, 49(3):1705– 1731, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2018:DRM

[GLL18] Yu Gao, Lei Li, and Jian-Guo Liu. A dispersive regularization for the modified Camassa–Holm equation. SIAM Journal on Mathematical Analysis, 50(3):2807– 2838, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghazaryan:2010:STW

[GLS10] Anna Ghazaryan, Yuri Latushkin, and Stephen Schecter. Stability of traveling waves for a class of reaction–diffusion systems that arise in chemical reaction models. SIAM Journal on Mathematical Analysis, 42(6):2434–2472, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gunther:2010:SMM

[GLT10] Uwe Günther, Heinz Langer, and Christiane Tretter. On the spectrum of the magnetohydrodynamic mean-field α^{2}-dynamo operator. SIAM Journal on Mathematical Analysis, 42(3):1413–1447, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2017:LUR

[GLW17] Yujin Guo, Changshou Lin, and Juncheng Wei. Local uniqueness and refined spike profiles of ground states for two-dimensional attractive Bose–Einstein condensates. SIAM Journal on Mathematical Analysis, 49(5): 3671–3715, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghazaryan:2018:SPF

[GLY18] A. Ghazaryan, Y. Latushkin, and X. Yang. Stability of a planar front in a class of reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 50(5):5569– 5615, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:2017:ATA

[GLZ17] Yixian Gao, Peijun Li, and Bo Zhang. Analysis of transient acoustic-elastic interaction in an unbounded structure. SIAM Journal on Mathematical Analysis, 49(5): 3951–3972, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomelli:2010:RSH

[GM10] Lorenzo Giacomelli and Salvador Moll. Rotationally symmetric 1-harmonic flows from D^{2} to S^{2}: Local wellposedness and finite time blowup. SIAM Journal on Mathematical Analysis, 42(6): 2791–2817, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gerard:2011:SPS

[GM11] Patrick Gérard and Florian Méhats. The Schrödinger– Poisson system on the sphere. SIAM Journal on Mathematical Analysis, 43(3):1232– 1268, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1232_s1.

Gigli:2013:GHC

[GM13] Nicola Gigli and Jan Maas. Gromov–Hausdorff convergence of discrete transportation metrics. SIAM Journal on Mathematical Analysis, 45 (2):879–899, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghorbel:2014:ENS

[GM14] A. Ghorbel and R. Monneau. Existence and nonexistence of semidiscrete shocks for a car-following model in traffic flow. SIAM Journal on Mathematical Analysis, 46(6): 3612–3639, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gordon:2015:ESS

[GM15] Peter V. Gordon and Cyrill B. Muratov. Eventual selfsimilarity of solutions for the diffusion equation with nonlinear absorption and a point source. SIAM Journal on Mathematical Analysis, 47(4): 2903–2916, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gallouet:2017:JSS

[GM17a] Thomas O. Gallouët and Léonard Monsaingeon. A JKO splitting scheme for Kantorovich–Fisher–Rao gradient flows. SIAM Journal on Mathematical Analysis, 49(2): 1100–1130, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gesu:2017:FMA

[GM17b] Giacomo Di Gesù and Mauro Mariani. Full metastable asymptotic of the Fisher information. SIAM Journal on Mathematical Analysis, 49(4): 3048–3072, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldman:2017:PSB

[GM17c] M. Goldman and B. Merlet.

Phase segregation for binary mixtures of Bose–Einstein condensates. SIAM Journal on Mathematical Analysis, 49 (3):1947–1981, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomelli:2013:HFV

[GMM13] Lorenzo Giacomelli, José M. Mazón, and Salvador Moll. The 1-harmonic flow with values into S^{1}. SIAM Journal on Mathematical Analysis, 45(3): 1723–1740, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Germain:2013:NGS

[GMP13] Pierre Germain, Nader Masmoudi, and Benoit Pausader. Nonneutral global solutions for the electron Euler–Poisson system in three dimensions. SIAM Journal on Mathematical Analysis, 45(1):267– 278, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giga:2016:ASS

[GMT16] Yoshikazu Giga, Hiroyoshi Mitake, and Hung V. Tran. On asymptotic speed of solutions to level-set mean curvature flow equations with driving and source terms. SIAM Journal on Mathematical Analysis, 48(5):3515– 3546, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giorgini:2019:URN

[GMT19] Andrea Giorgini, Alain Miranville, and Roger Temam. Uniqueness and regularity for the Navier–Stokes–Cahn– Hilliard system. SIAM Journal on Mathematical Analysis, 51(3):2535–2574, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guarguaglini:2015:GSS

[GN15] F. R. Guarguaglini and R. Natalini. Global smooth solutions for a hyperbolic chemotaxis model on a network. SIAM Journal on Mathematical Analysis, 47(6):4652– 4671, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grenier:2019:GFO

[GN19] Emmanuel Grenier and Toan T.

Nguyen. Green function of Orr–Sommerfeld equations away from critical layers. SIAM Journal on Mathematical Analysis, 51(2):1279– 1296, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gnann:2015:WPS

[Gna15] Manuel V. Gnann. Wellposedness and self-similar asymptotics for a thin-film equation. SIAM Journal on Mathematical Analysis, 47(4): 2868–2902, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Generau:2018:LVM

[GO18] François Générau and Edouard Oudet. Large volume minimizers of a nonlocal isoperimetric problem: Theoretical and numerical approaches. SIAM Journal on Mathematical Analysis, 50(3):3427– 3450, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:2011:SDD

[GP11] Stephen Gustafson and Tuoc Van Phan. Stable directions for degenerate excited states of nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 43(4): 1716–1758, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1716_s1.

Graf:2014:DSB

[GP14] Isabell Graf and Malte A. Peter. Diffusion on surfaces and the boundary periodic unfolding operator with an application to carcinogenesis in human cells. SIAM Journal on Mathematical Analysis, 46(4): 3025–3049, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gomes:2015:TDM

[GP15] Diogo A. Gomes and Edgard Pimentel. Time-dependent mean-field games with logarithmic nonlinearities. SIAM Journal on Mathematical Analysis, 47(5):3798–3812, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guevara:2018:LSS

[GP18] Cristi Guevara and Nguyen Cong Phuc. Leray’s self-similar solutions to the Navier–Stokes equations with profiles in Marcinkiewicz and Morrey spaces. SIAM Journal on Mathematical Analysis, 50(1): 541–556, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Geyer:2019:LIU

[GP19] Anna Geyer and D. Pelinovsky. Linear instability and uniqueness of the peaked periodic wave in the reduced Ostrovsky equation. SIAM Journal on Mathematical Analysis, 51(2):1188–1208, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Griffin-Pickering:2018:MFA

[GPI18] Megan Griffin-Pickering and Mikaela Iacobelli. A mean field approach to the quasineutral limit for the Vlasov– Poisson equation. SIAM Journal on Mathematical Analysis, 50(5):5502–5536, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giacomin:2013:TAR

[GPPP13] Giambattista Giacomin, Khashayar Pakdaman, Xavier Pellegrin, and Christophe Poquet. Transitions in active rotator systems: Invariant hyperbolic manifold approach. SIAM Journal on Mathematical Analysis, 44(6): 4165–4194, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gamba:2019:PEW

[GPT19] Irene M. Gamba, Natasa Pavlović, and Maja Tasković. On pointwise exponentially weighted estimates for the Boltzmann equation. SIAM Journal on Mathematical Analysis, 51(5):3921–3955, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gianni:2013:MDD

[GR13a] Roberto Gianni and Fabio Rosso. Modeling degradating dispersions in a threedimensional finite container under general boundary conditions. SIAM Journal on Mathematical Analysis, 45(4): 2332–2353, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gurumoorthy:2013:DTG

[GR13b] Karthik S. Gurumoorthy and Anand Rangarajan. Distance transform gradient density estimation using the stationary phase approximation. SIAM Journal on Mathematical Analysis, 44(6):4250– 4273, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gess:2015:SDM

[GR15a] Benjamin Gess and Michael Röckner. Singular-degenerate multivalued stochastic fast diffusion equations. SIAM Journal on Mathematical Analysis, 47(5):4058–4090, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grudsky:2015:STH

[GR15b] Sergei Grudsky and Alexei Rybkin. Soliton theory and Hankel operators. SIAM Journal on Mathematical Analysis, 47(3):2283–2323, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grande:2019:STF

[Gra19]  Ricardo Grande.  Space-

time fractional nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 51(5):4172–4212,

????  2019.  CODEN

SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grohs:2010:GPA

[Gro10] Philipp Grohs. A general proximity analysis of nonlinear subdivision schemes. SIAM Journal on Mathematical Analysis, 42(2):729– 750, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gourley:2014:UPM

[GRT14] Stephen A. Gourley, Gergely Röst, and Horst R. Thieme. Uniform persistence in a model for bluetongue dynamics. SIAM Journal on Mathematical Analysis, 46(2):1160– 1184, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giorgi:2010:GUS

[GS10a] Tiziana Giorgi and Robert G. Smits. Gauge uniqueness of solutions to the Ginzburg– Landau system for small superconducting domains. SIAM Journal on Mathematical Analysis, 42(1):163– 182, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Girault:2010:TDT

[GS10b] V. Girault and L. Ridgway Scott. On a time-dependent transport equation in a Lipschitz domain. SIAM Journal on Mathematical Analysis, 42(4):1721–1731, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galvis:2012:RRO

[GS12a] Juan Galvis and Marcus Sarkis. Regularity results for the ordinary product stochastic pressure equation. SIAM Journal on Mathematical Analysis, 44(4):2637– 2665, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glass:2012:MRB

[GS12b] Olivier Glass and Franck Sueur. On the motion of a rigid body in a twodimensional irregular ideal flow. SIAM Journal on Mathematical Analysis, 44(5): 3101–3126, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gaudiello:2015:HHO

[GS15a] Antonio Gaudiello and Ali Sili. Homogenization of highly oscillating boundaries with strongly contrasting diffusivity. SIAM Journal on Mathematical Analysis, 47(3):1671– 1692, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gu:2015:HSS

[GS15b] Shu Gu and Zhongwei Shen. Homogenization of Stokes systems and uniform regularity estimates. SIAM Journal on Mathematical Analysis, 47(5): 4025–4057, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2015:NTZ

[GS15c] Jong-Shenq Guo and Philippe Souplet. No touchdown at zero points of the permittivity profile for the MEMS problem. SIAM Journal on Mathematical Analysis, 47(1): 614–625, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grandi:2017:FPP

[GS17] D. Grandi and U. Stefanelli. Finite plasticity in P^{⊤}P. part II: Quasi-static evolution and linearization. SIAM Journal on Mathematical Analysis, 49 (2):1356–1384, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guillen:2018:NHI

[GS18] Nestor Guillen and Russell W. Schwab. Neumann homogenization via integrodifferential Operators.Part 2: Singular gradient dependence. SIAM Journal on Mathematical Analysis, 50(2):1679– 1719, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guerra:2019:VVB

[GS19] Graziano Guerra and Wen Shen. Vanishing viscosity and backward Euler approximations for conservation laws with discontinuous flux. SIAM Journal on Mathematical Analysis, 51(4): 3112–3144, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gurevich:2013:RDE

[GST13] Pavel Gurevich, Roman Shamin, and Sergey Tikhomirov. Reaction– diffusion equations with spatially distributed hysteresis. SIAM Journal on Mathematical Analysis, 45(3):1328– 1355, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golovaty:2019:GLT

[GSV19] Dmitry Golovaty, Peter Sternberg, and Raghavendra Venkatraman. A Ginzburg–Landau-Type problem for highly anisotropic nematic liquid crystals. SIAM Journal on Mathematical Analysis, 51(1):

276–320, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasser:2016:ECJ

[GSW16] Ingenuin Gasser, Peter Szmolyan, and Johannes Wächtler. Existence of Chapman–Jouguet detonation and deflagration waves. SIAM Journal on Mathematical Analysis, 48(2): 1400–1422, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guillod:2018:NSE

[GSWZ18] Julien Guillod, Guido Schneider, Peter Wittwer, and Dominik Zimmermann. Nonlinear stability at the Eckhaus boundary. SIAM Journal on Mathematical Analysis, 50(5): 4699–4720, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2010:LRT

[GT10] Yan Guo and Ian Tice. Linear Rayleigh–Taylor instability for viscous, compressible fluids. SIAM Journal on Mathematical Analysis, 42(4): 1688–1720, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gess:2016:ELL

[GT16] Benjamin Gess and Jonas M. Tölle. Ergodicity and local limits for stochastic local and nonlocal p-Laplace equations. SIAM Journal on Mathematical Analysis, 48(6): 4094–4125, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galiano:2019:WPC

[GV19] Gonzalo Galiano and Julián Velasco. Well-posedness of a cross-diffusion population model with nonlocal diffusion. SIAM Journal on Mathematical Analysis, 51(4): 2884–2902, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gerard-Varet:2016:BLN

[GVWK16] David Gérard-Varet and Aneta Wróblewska-Kamińska. Boundary layer for a non-Newtonian flow over a rough surface. SIAM Journal on Mathematical Analysis, 48(5): 3123–3147, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garcia:2016:DTE

[GVZ16] Andoni García, Esa V.

Vesalainen, and Miren Zubeldia. Discreteness of transmission eigenvalues for higherorder main terms and perturbations. SIAM Journal on Mathematical Analysis, 48(4): 2382–2398, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grasselli:2013:LTB

[GW13] Maurizio Grasselli and Hao Wu. Long-time behavior for a hydrodynamic model on nematic liquid crystal flows with asymptotic stabilizing boundary condition and external force. SIAM Journal on Mathematical Analysis, 45(3):965– 1002, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guillod:2015:GSI

[GW15] Julien Guillod and Peter Wittwer. Generalized scale-invariant solutions to the two-dimensional stationary Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 47(1):955– 968, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guidotti:2018:MSD

[GW18] Patrick Guidotti and Christoph Walker. On a model for a sliding droplet: Wellposedness and stability of translating circular solutions. SIAM Journal on Mathematical Analysis, 50(2):1656– 1678, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gu:2017:OBE

[GX17] Shu Gu and Qiang Xu. Optimal boundary estimates for Stokes systems in homogenization theory. SIAM Journal on Mathematical Analysis, 49(5):3831–3853, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golgeleyen:2016:SSI

[GY16a] Fikret Gölgeleyen and Masahiro Yamamoto. Stability for some inverse problems for transport equations. SIAM Journal on Mathematical Analysis, 48(4): 2319–2344, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:2016:EBR

[GY16b] Yan Guo and Xiongfeng Yang.

Existence and BV-regularity for neutron transport equation in NonConvex domain. SIAM Journal on Mathematical Analysis, 48(5):3467– 3495, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giovangigli:2018:RLI

[GYY18] Vincent Giovangigli, Zai-Bao Yang, and Wen-An Yong. Relaxation limit and initial-layers for a class of hyperbolic-parabolic systems. SIAM Journal on Mathematical Analysis, 50(4):4655– 4697, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Geng:2013:QPS

[GZ13] Jiansheng Geng and Zhiyan Zhao. Quasi-periodic solutions for one-dimensional nonlinear lattice Schrödinger equation with tangent potential. SIAM Journal on Mathematical Analysis, 45(6):3651– 3689, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldman:2014:SLR

[GZ14] M. Goldman and B. Zwicknagl. Scaling law and reduced models for epitaxially strained crystalline films. SIAM Journal on Mathematical Analysis, 46(1):1–24, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gualdani:2018:GEW

[GZ18] Maria Pia Gualdani and Nicola Zamponi. Global existence of weak even solutions for an isotropic Landau equation with Coulomb potential. SIAM Journal on Mathematical Analysis, 50(4): 3676–3714, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hall:2012:ASA

[Hal12] Eric Joseph Hall. Accelerated spatial approximations for time discretized stochastic partial differential equations. SIAM Journal on Mathematical Analysis, 44(5): 3162–3185, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hall:2013:HOS

[Hal13] Eric Joseph Hall. Higher order spatial approximations for degenerate parabolic stochastic partial differential equations. SIAM Journal on Mathematical Analysis, 45(4): 2071–2098, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haltmeier:2014:UIF

[Hal14] Markus Haltmeier. Universal inversion formulas for recovering a function from spherical means. SIAM Journal on Mathematical Analysis, 46(1): 214–232, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hansen:2014:SFS

[Han14] Sönke Hansen. Subsonic free surface waves in linear elasticity. SIAM Journal on Mathematical Analysis, 46(4):2501– 2524, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hangelbroek:2018:PDP

[Han18] Thomas C. Hangelbroek. On a polyharmonic Dirichlet problem and boundary effects in surface spline approximation. SIAM Journal on Mathematical Analysis, 50(4):4616– 4654, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harutyunyan:2018:KIS

[Har18] D. Harutyunyan. On the Korn interpolation and second inequalities in thin domains. SIAM Journal on Mathematical Analysis, 50(5): 4964–4982, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2016:GTS

[HCHY16] Bo-Chih Huang, Shih-Wei Chou, John M. Hong, and Chien-Chang Yen. Global transonic solutions of planetary atmospheres in a hydrodynamic region — hydrodynamic escape problem due to gravity and heat. SIAM Journal on Mathematical Analysis, 48(6):4268–4310, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Healey:2017:SBG

[HD17] Timothy J. Healey and Sanjay Dharmavaram. Symmetrybreaking global bifurcation in a surface continuum phasefield model for lipid bilayer vesicles. SIAM Journal on Mathematical Analysis, 49(2): 1027–1059, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Helmensdorfer:2012:MBF

[Hel12] Sebastian Helmensdorfer. A model for the behavior of fluid droplets based on mean curvature flow. SIAM Journal on Mathematical Analysis, 44(3): 1359–1371, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henry:2010:ASP

[Hen10] David Henry. Analyticity of the streamlines for periodic travelling free surface capillary-gravity water waves with vorticity. SIAM Journal on Mathematical Analysis, 42(6):3103– 3111, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3103_s1.

Han:2013:UUN

[HF13] Zheng Han and Daoyuan Fang. On the unconditional uniqueness for NLS in Ḣ^{s}. SIAM Journal on Mathematical Analysis, 45(3):1505– 1526, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2014:AGQ

[HGW14] Guoliang He, Xianguo Geng, and Lihua Wu. Algebrogeometric quasi-periodic solutions to the three-wave resonant interaction hierarchy. SIAM Journal on Mathematical Analysis, 46(2):1348– 1384, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Himonas:2018:CPS

[HHK18] A. Alexandrou Himonas, Curtis Holliman, and Carlos Kenig. Construction of 2peakon solutions and illposedness for the Novikov equation. SIAM Journal on Mathematical Analysis, 50(3): 2968–3006, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haskovec:2018:DEE

[HHMM18] Jan Haskovec, Sabine Hittmeir, Peter Markowich, and Alexander Mielke. Decay to equilibrium for energy–reaction– diffusion systems. SIAM Journal on Mathematical Analysis, 50(1):1037–1075,

????  2018.  CODEN

SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2017:GEC

[HHPZ17] Guangyi Hong, Xiaofeng Hou, Hongyun Peng, and Changjiang Zhu. Global existence for a class of large solutions to three-dimensional compressible magnetohydrodynamic equations with vacuum. SIAM Journal on Mathematical Analysis, 49(4): 2409–2441, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hanke:2009:IBP

[HHR09] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. An inverse backscatter problem for electric impedance tomography. SIAM Journal on Mathematical Analysis, 41(5): 1948–1966, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [HHR11].

Hanke:2011:EIB

[HHR11] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. Erratum: An inverse backscatter problem for electric impedance tomography. SIAM Journal on Mathematical Analysis, 43(3):1495–1497, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1495_s1.

See [HHR09].

Holding:2017:CAM

[HHR17] Thomas Holding, Harsha Hutridurga, and Jeffrey Rauch. Convergence along mean flows. SIAM Journal on Mathematical Analysis, 49(1): 222–271, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hunter:2012:ELS

[HI12] John K. Hunter and Mihaela Ifrim. Enhanced life span of smooth solutions of a Burgers–Hilbert equation. SIAM Journal on Mathematical Analysis, 44(3):2039– 2052, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hisa:2019:SHE

[HI19] Kotaro Hisa and Kazuhiro Ishige. Solvability of the heat equation with a nonlinear boundary condition. SIAM Journal on Mathematical Analysis, 51(1):565– 594, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hittmeir:2011:CDP

[HJ11] Sabine Hittmeir and Ansgar Jüngel. Cross diffusion preventing blow-up in the two-dimensional Keller–Segel model. SIAM Journal on Mathematical Analysis, 43(2): 997–1022, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p997_s1.

See corrigendum [BCJ20].

Hur:2015:SPT

[HJ15] Vera Mikyoung Hur and Mathew A. Johnson. Stability of periodic traveling waves for nonlinear dispersive equations. SIAM Journal on Mathematical Analysis, 47(5): 3528–3554, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hwang:2018:FPE

[HJJ18] Hyung Ju Hwang, Juhi Jang, and Jaewoo Jung. The Fokker–Planck equation with absorbing boundary conditions in bounded domains. SIAM Journal on Mathematical Analysis, 50(2):2194– 2232, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han-Kwan:2010:CTP

[HK10] Daniel Han-Kwan. On the confinement of a Tokamak plasma. SIAM Journal on Mathematical Analysis, 42(6): 2337–2367, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heinemann:2015:EWS

[HK15] Christian Heinemann and Christiane Kraus. Existence of weak solutions for a hyperbolic-parabolic phase field system with mixed boundary conditions on nonsmooth domains. SIAM Journal on Mathematical Analysis, 47(3):2044–2073, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holden:2013:OSW

[HKK13] Helge Holden, Kenneth H. Karlsen, and Trygve K. Karper. Operator splitting for well-posed active scalar equations. SIAM Journal on Mathematical Analysis, 45(1): 152–180, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ha:2015:EDH

[HKK15] Seung-Yeal Ha, Moon-Jin Kang, and Bongsuk Kwon. Emergent dynamics for the hydrodynamic Cucker–Smale system in a moving domain. SIAM Journal on Mathematical Analysis, 47(5):3813– 3831, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoffmann:2017:EUG

[HKK17] Joachim Hoffmann, Serge Kräutle, and Peter Knabner. Existence and uniqueness of a global solution for reactive transport with mineral precipitation–dissolution and aquatic reactions in porous media. SIAM Journal on Mathematical Analysis, 49(6): 4812–4837, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han-Kwan:2016:NIV

[HKN16] Daniel Han-Kwan and Toan T.

Nguyen. Nonlinear instability of Vlasov–Maxwell systems in the classical and quasineutral limits. SIAM Journal on Mathematical Analysis, 48(5): 3444–3466, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hitrik:2010:TEO

[HKOP10] Michael Hitrik, Katsiaryna Krupchyk, Petri Ola, and Lassi Päivärinta. Transmission eigenvalues for operators with constant coefficients. SIAM Journal on Mathematical Analysis, 42(6): 2965–2986, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p2965_s1.

Hitrik:2011:TEE

[HKOP11] Michael Hitrik, Katsiaryna Krupchyk, Petri Ola, and Lassi Päivärinta. Transmission eigenvalues for elliptic operators. SIAM Journal on Mathematical Analysis, 43(6): 2630–2639, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i6/p2630_s1.

Ha:2018:EBT

[HKR18] Seung-Yeal Ha, Jeongho Kim, and Tommaso Ruggeri. Emergent behaviors of thermodynamic Cucker–Smale particles. SIAM Journal on Mathematical Analysis, 50(3):3092– 3121, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holmes:2018:NDI

[HKT18] John Holmes, Barbara Keyfitz, and Feride Tiglay. Nonuniform dependence on initial data for compressible gas dynamics: The Cauchy problem on R^{2}. SIAM Journal on Mathematical Analysis, 50(1):1237–1254, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haddar:2011:EWS

[HL11] Houssem Haddar and Armin Lechleiter. Electromagnetic wave scattering from rough penetrable layers. SIAM Journal on Mathematical Analysis, 43(5):2418–2443, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2418_s1.

Hao:2012:WPM

[HL12] Chengchun Hao and Hai-Liang Li. Well-posedness for a multidimensional viscous liquid-gas two-phase flow model. SIAM Journal on Mathematical Analysis, 44(3): 1304–1332, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsieh:2015:EDE

[HL15] Chia-Yu Hsieh and Tai-Chia Lin. Exponential decay estimates for the stability of boundary layer solutions to Poisson–Nernst–Planck systems: One spatial dimension case. SIAM Journal on Mathematical Analysis, 47(5): 3442–3465, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrach:2019:MBI

[HL19a] Bastian Harrach and Yi-Hsuan Lin. Monotonicitybased inversion of the fractional Schrödinger equation I. Positive potentials.

SIAM Journal on Mathematical Analysis, 51(4):3092– 3111, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hynd:2019:LRH

[HL19b] Ryan Hynd and Erik Lindgren. Lipschitz regularity for a homogeneous doubly nonlinear PDE. SIAM Journal on Mathematical Analysis, 51(4): 3606–3624, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsu:2014:NPC

[HLGMMM14] S. B. Hsu, J. López-Gómez, L. Mei, and M. Molina-Meyer. A nonlocal problem from conservation biology. SIAM Journal on Mathematical Analysis, 46(6):4035–4059, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hickman:2019:HDV

[HLR^{+}19] Jonathan Hickman, Felipe Linares, Oscar G. Riaño, Keith M. Rogers, and James Wright. On a higher dimensional version of the Benjamin–Ono equation. SIAM Journal on Mathematical Analysis, 51(6):4544–4569, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2012:ZDL

[HLW12] Feimin Huang, Mingjie Li, and Yi Wang. Zero dissipation limit to rarefaction wave with vacuum for one-dimensional compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 44(3): 1742–1759, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hou:2018:SBL

[HLWW18] Qianqian Hou, Cheng-Jie Liu, Ya-Guang Wang, and Zhian Wang. Stability of boundary layers for a viscous hyperbolic system arising from chemotaxis: One-dimensional case. SIAM Journal on Mathematical Analysis, 50(3): 3058–3091, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2011:STC

[HLX11] Xiangdi Huang, Jing Li, and Zhouping Xin. Serrintype criterion for the threedimensional viscous compressible flows. SIAM Journal on Mathematical Analysis, 43(4): 1872–1886, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1872_s1.

Haines:2012:PEE

[HM12a] Brian M. Haines and Anna L. Mazzucato. A proof of Einstein’s effective viscosity for a dilute suspension of spheres. SIAM Journal on Mathematical Analysis, 44(3): 2120–2145, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henao:2012:SUG

[HM12b] Duvan Henao and Apala Majumdar. Symmetry of uniaxial global Landau–de Gennes minimizers in the theory of nematic liquid crystals. SIAM Journal on Mathematical Analysis, 44(5):3217– 3241, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigendum [HM13].

Henao:2013:CSU

[HM13] Duvan Henao and Apala Majumdar. Corrigendum: Symmetry of Uniaxial Global Landau–de Gennes Minimizers in the Theory of Nematic Liquid Crystals. SIAM Journal on Mathematical Analysis, 45(6):3872–3874, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [HM12b].

Hoell:2014:CDI

[HMN14] Nicholas Hoell, Amir Moradifam, and Adrian Nachman.

Current density impedance imaging of an anisotropic conductivity in a known conformal class. SIAM Journal on Mathematical Analysis, 46(3): 1820–1842, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:2014:CVH

[HMS14] Weimin Han, Stanisław Migórski, and Mircea Sofonea. A class of variational-hemivariational inequalities with applications to frictional contact problems. SIAM Journal on Mathematical Analysis, 46(6): 3891–3912, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herrmann:2013:SPT

[HMSZ13] Michael Herrmann, Karsten Matthies, Hartmut Schwetlick, and Johannes Zimmer. Subsonic phase transition waves in bistable lattice models with small spinodal region. SIAM Journal on Mathematical Analysis, 45(5):2625– 2645, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2011:LTB

[HMW11] Feimin Huang, Ming Mei, and Yong Wang. Large time behavior of solutions to ndimensional bipolar hydrodynamic models for semiconductors. SIAM Journal on Mathematical Analysis, 43(4): 1595–1630, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1595_s1.

Huang:2011:ACS

[HMWY11] Feimin Huang, Ming Mei, Yong Wang, and Huimin Yu. Asymptotic convergence to stationary waves for unipolar hydrodynamic model of semiconductors. SIAM Journal on Mathematical Analysis, 43(1): 411–429, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i1/p411_s1.

Huang:2012:LTB

[HMWY12] Feimin Huang, Ming Mei, Yong Wang, and Tong Yang. Long-time behavior of solutions to the bipolar hydrodynamic model of semiconductors with boundary effect. SIAM Journal on Mathematical Analysis, 44(2):1134– 1164, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:2015:CST

[HMZ15] Bin Han, Qun Mo, and Zhenpeng Zhao. Compactly supported tensor product complex tight framelets with directionality. SIAM Journal on Mathematical Analysis, 47(3): 2464–2494, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hulshof:2013:ESD

[HNP13] J. Hulshof, R. Nolet, and G. Prokert. Existence of solutions to the diffusive VSC model. SIAM Journal on Mathematical Analysis, 45(2): 700–727, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoang:2015:GEG

[HNP15] Luan T. Hoang, Truyen V.

Nguyen, and Tuoc V. Phan. Gradient estimates and global existence of smooth solutions to a cross-diffusion system. SIAM Journal on Mathematical Analysis, 47(3):2122– 2177, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hadzic:2017:LWP

[HNS17] Mahir Hadzić, Gustavo Navarro, and Steve Shkoller. Local well-posedness and global stability of the two-phase Stefan problem. SIAM Journal on Mathematical Analysis, 49(6): 4942–5006, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hangelbroek:2011:KAM

[HNSW11] T. Hangelbroek, F. J. Narcowich, X. Sun, and J. D. Ward. Kernel approximation on manifolds II: The L_{∞} norm of the L_{2} projector. SIAM Journal on Mathematical Analysis, 43(2): 662–684, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i2/p662_s1.

Hangelbroek:2010:KAM

[HNW10] T. Hangelbroek, F. J. Narcowich, and J. D. Ward. Kernel approximation on manifolds I: Bounding the Lebesgue constant. SIAM Journal on Mathematical Analysis, 42(4):1732–1760, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hudson:2015:ASS

[HO15] T. Hudson and C. Ortner. Analysis of stable screw dislocation configurations in an antiplane lattice model. SIAM Journal on Mathematical Analysis, 47(1): 291–320, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoff:2012:LSC

[Hof12]  David Hoff.  Local solu-

tions of a compressible flow problem with Navier boundary conditions in general three-dimensional domains. SIAM Journal on Mathematical Analysis, 44(2):633– 650, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i2/p633_s1.

Hofer:2018:ILP

[Höf18]  Richard M. Höfer.  The

inertialess limit of particle sedimentation modeled by the Vlasov–Stokes equations. SIAM Journal on Mathematical Analysis, 50(5):5446– 5476, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hyvonen:2012:PMN

[HPS12] Nuutti Hyvönen, Petteri Piiroinen, and Otto Seiskari. Point measurements for a Neumann-to–Dirichlet map and the Calderón problem in the plane. SIAM Journal on Mathematical Analysis, 44(5): 3526–3536, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrach:2019:DBM

[HPS19] Bastian Harrach, Valter Pohjola, and Mikko Salo. Dimension bounds in monotonicity methods for the Helmholtz equation. SIAM Journal on Mathematical Analysis, 51(4): 2995–3019, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herrmann:2010:HTW

[HR10] Michael Herrmann and Jens D. M. Rademacher. Heteroclinic travelling waves in convex FPU-type chains. SIAM Journal on Mathematical Analysis, 42(4):1483– 1504, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hur:2012:HPV

[HR12] Youngmi Hur and Amos Ron. High-performance very local Riesz wavelet bases of L_{2}(R^{n}). SIAM Journal on Mathematical Analysis, 44(4): 2237–2265, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hamel:2015:AST

[HR15] François Hamel and Luca Rossi. Admissible speeds of transition fronts for nonautonomous monostable equations. SIAM Journal on Mathematical Analysis, 47(5): 3342–3392, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holden:2019:MDM

[HR19] Helge Holden and Nils Henrik Risebro. Models for dense multilane vehicular traffic. SIAM Journal on Mathematical Analysis, 51(5):3694– 3713, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hamel:2010:SSS

[HS10a] François Hamel and Yannick Sire. Spreading speeds for some reaction–diffusion equations with general initial conditions. SIAM Journal on Mathematical Analysis, 42(6): 2872–2911, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p2872_s1.

Harrach:2010:ESR

[HS10b] Bastian Harrach and Jin Keun Seo. Exact shape-reconstruction by one-step linearization in electrical impedance tomography. SIAM Journal on Mathematical Analysis, 42(4): 1505–1518, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrell:2010:UBS

[HS10c] Evans M. Harrell II and Joachim Stubbe. Universal bounds and semiclassical estimates for eigenvalues of abstract Schrödinger operators. SIAM Journal on Mathematical Analysis, 42(5): 2261–2274, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoang:2013:STG

[HS13] Viet Ha Hoang and Christoph Schwab. Sparse tensor Galerkin discretization of parametric and random parabolic PDEs — analytic regularity and generalized polynomial chaos approximation. SIAM Journal on Mathematical Analysis, 45(5):3050– 3083, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holzer:2014:AFT

[HS14] Matt Holzer and Arnd Scheel. Accelerated fronts in a twostage invasion process. SIAM Journal on Mathematical Analysis, 46(1):397–427, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heinemann:2016:SOC

[HS16] Christian Heinemann and Kevin Sturm. Shape optimization for a class of semilinear variational inequalities with applications to damage models. SIAM Journal on Mathematical Analysis, 48(5): 3579–3617, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holland:2018:PVE

[HS18] Eleanor Holland and R. E.

Showalter. Poro-visco-elastic compaction in sedimentary basins. SIAM Journal on Mathematical Analysis, 50(2): 2295–2316, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hening:2019:PPS

[HS19] Alexandru Hening and Edouard Strickler. On a predator-prey system with random switching that never converges to its equilibrium. SIAM Journal on Mathematical Analysis, 51(5): 3625–3640, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hakkaev:2017:PTW

[HSS17] Sevdzhan Hakkaev, Milena Stanislavova, and Atanas Stefanov. Periodic traveling waves of the regularized short pulse and Ostrovsky equations: Existence and stability. SIAM Journal on Mathematical Analysis, 49(1):674– 698, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2016:SII

[HSV16] Guanghui Hu, Mikko Salo, and Esa V. Vesalainen.

Shape identification in inverse medium scattering problems with a single far-field pattern. SIAM Journal on Mathematical Analysis, 48(1): 152–165, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hardin:2017:GPC

[HSV17] D. P. Hardin, E. B. Saff, and O. V. Vlasiuk. Generating point configurations via hypersingular Riesz energy with an external field. SIAM Journal on Mathematical Analysis, 49(1):646–673, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2017:SSG

[HT17] Feimin Huang and Wenke Tan. On the strong solution of the ghost effect system. SIAM Journal on Mathematical Analysis, 49(5):3496– 3526, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hur:2018:WBS

[HT18] Vera Mikyoung Hur and Lizheng Tao. Wave breaking in a shallow water model. SIAM Journal on Mathematical Analysis, 50(1):354– 380, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howell:2018:CEB

[HTW18] Jason S. Howell, Daniel Toundykov, and Justin T. Webster. A cantilevered extensible beam in axial flow: Semigroup well-posedness and postflutter regimes. SIAM Journal on Mathematical Analysis, 50(2):2048–2085, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2015:UUC

[HTX15] Younghun Hong, Kenneth Taliaferro, and Zhihui Xie. Unconditional uniqueness of the cubic Gross–Pitaevskii hierarchy with low regularity. SIAM Journal on Mathematical Analysis, 47(5):3314– 3341, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrach:2013:MBS

[HU13] Bastian Harrach and Marcel Ullrich. Monotonicitybased shape reconstruction in electrical impedance tomography. SIAM Journal on Mathematical Analysis, 45(6): 3382–3403, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hupkes:2013:NDT

[HV13] H. J. Hupkes and E. S.

Van Vleck. Negative diffusion and traveling waves in high dimensional lattice systems. SIAM Journal on Mathematical Analysis, 45(3): 1068–1135, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoefer:2011:DMH

[HW11] M. A. Hoefer and M. I. Weinstein. Defect modes and homogenization of periodic Schrödinger operators. SIAM Journal on Mathematical Analysis, 43(2):971– 996, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p971_s1.

Hu:2013:GEO

[HW13a] Xianpeng Hu and Guochun Wu. Global existence and optimal decay rates for threedimensional compressible viscoelastic flows. SIAM Journal on Mathematical Analysis, 45(5):2815–2833, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2013:GST

[HW13b] Xianpeng Hu and Hao Wu. Global solution to the threedimensional compressible flow of liquid crystals. SIAM Journal on Mathematical Analysis, 45(5):2678–2699, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2014:GSS

[HW14] Xiangdi Huang and Yun Wang. Global strong solution with vacuum to the two dimensional densitydependent Navier–Stokes system. SIAM Journal on Mathematical Analysis, 46(3):1771– 1788, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2017:SSS

[HW17] Hakho Hong and Teng Wang.

Stability of stationary solutions to the inflow problem for full compressible Navier–Stokes equations with a large initial perturbation. SIAM Journal on Mathematical Analysis, 49(3):2138– 2166, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:2013:LBE

[HWWY13] Feimin Huang, Yi Wang, Yong Wang, and Tong Yang. The limit of the Boltzmann equation to the Euler equations for Riemann problems. SIAM Journal on Mathematical Analysis, 45(3):1741– 1811, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2012:EPG

[HWZ12] Qingwen Hu, Jianhong Wu, and Xingfu Zou. Estimates of periods and global continua of periodic solutions for state-dependent delay equations. SIAM Journal on Mathematical Analysis, 44(4): 2401–2427, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2010:GWP

[HX10] Lingbing He and Li Xu. Global well-posedness for viscoelastic fluid system in bounded domains. SIAM Journal on Mathematical Analysis, 42(6):2610–2625, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ha:2014:USP

[HX14] Seung-Yeal Ha and Qinghua Xiao. Uniform stability and the propagation of regularity for the relativistic Boltzmann equation. SIAM Journal on Mathematical Analysis, 46(1): 165–191, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2013:RSF

[HY13] Jiaxing Hong and Ge Yang. On the regularity of solutions to FENE models. SIAM Journal on Mathematical Analysis, 45(4):2228–2252, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:2014:WPA

[HY14] Lingbing He and Xiongfeng Yang. Well-posedness and asymptotics of grazing collisions limit of Boltzmann equation with Coulomb interaction. SIAM Journal on Mathematical Analysis, 46(6): 4104–4165, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hong:2019:SCD

[HY19] Younghun Hong and Changhun Yang. Strong convergence for discrete nonlinear Schrödinger equations in the continuum limit. SIAM Journal on Mathematical Analysis, 51(2): 1297–1320, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hynd:2013:PRW

[Hyn13] Ryan Hynd. Partial regularity of weak solutions of the viscoelastic Navier– Stokes equations with damping. SIAM Journal on Mathematical Analysis, 45(2):495– 517, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hynd:2019:LCS

[Hyn19] Ryan Hynd. Lagrangian coordinates for the sticky particle system. SIAM Journal on Mathematical Analysis, 51(5): 3769–3795, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:2010:MES

[HZ10] Bin Han and Xiaosheng Zhuang. Matrix extension with symmetry and its application to symmetric orthonormal multiwavelets. SIAM Journal on Mathematical Analysis, 42(5):2297– 2317, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:2019:GCS

[HZ19] Dian Hu and Yongqian Zhang. Global conic shock wave for the steady supersonic flow past a curved cone. SIAM Journal on Mathematical Analysis, 51(3):2372– 2389, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hou:2013:AGS

[HZFQ13] Yu Hou, Peng Zhao, Engui Fan, and Zhijun Qiao. Algebro-geometric solutions for the Degasperis–Procesi hierarchy. SIAM Journal on Mathematical Analysis, 45(3):

1216–1266, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Igbida:2012:PIE

[Igb12] Noureddine Igbida. A partial integrodifferential equation in granular matter and its connection with a stochastic model. SIAM Journal on Mathematical Analysis, 44(3): 1950–1975, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Igbida:2017:MCS

[Igb17]  Noureddine Igbida.  Met-

ric character for the sub-Hamilton–Jacobi obstacle equation. SIAM Journal on Mathematical Analysis, 49(4): 3143–3160, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ignat:2010:SES

[Ign10] Liviu I. Ignat. Strichartz estimates for the Schrödinger equation on a tree and applications. SIAM Journal on Mathematical Analysis, 42(5): 2041–2057, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ignat:2015:CTA

[IISD15] Liviu I. Ignat, Tatiana I.

Ignat, and Denisa Stancu-Dumitru. A compactness tool for the analysis of nonlocal evolution equations. SIAM Journal on Mathematical Analysis, 47(2):1330– 1354, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ibrahim:2011:GSS

[IK11] Slim Ibrahim and Sahbi Keraani. Global small solutions for the Navier–Stokes– Maxwell system. SIAM Journal on Mathematical Analysis, 43(5):2275–2295, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2275_s1.

Ishige:2017:AES

[IKM17] Kazuhiro Ishige, Tatsuki Kawakami, and Hironori Michihisa. Asymptotic expansions of solutions of fractional diffusion equations. SIAM Journal on Mathematical Analysis, 49(3):2167– 2190, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ioannidis:2012:WPM

[IKS12] Andreas D. Ioannidis, Gerhard Kristensson, and Ioannis G. Stratis. On the wellposedness of the Maxwell system for linear bianisotropic media. SIAM Journal on Mathematical Analysis, 44(4): 2459–2473, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ilmavirta:2016:CQT

[Ilm16] Joonas Ilmavirta. Coherent quantum tomography. SIAM Journal on Mathematical Analysis, 48(5):3039– 3064, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iida:2011:SFC

[ILN11] Masato Iida, Roger Lui, and Hirokazu Ninomiya. Stacked fronts for cooperative systems with equal diffusion coefficients. SIAM Journal on Mathematical Analysis, 43(3): 1369–1389, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1369_s1.

Isaza:2016:PRS

[ILP16] Pedro Isaza, Felipe Linares, and Gustavo Ponce. On the propagation of regularity of solutions of the Kadomtsev– Petviashvili equation. SIAM Journal on Mathematical Analysis, 48(2):1006–1024, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ianni:2017:CPB

[ILR17] Isabella Ianni, Stefan Le Coz, and Julien Royer. On the Cauchy problem and the black solitons of a singularly perturbed Gross–Pitaevskii equation. SIAM Journal on Mathematical Analysis, 49(2): 1060–1099, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Isakov:2016:ISC

[ILW16] Victor Isakov, Ru-Yu Lai, and Jenn-Nan Wang. Increasing stability for the conductivity and attenuation coefficients. SIAM Journal on Mathematical Analysis, 48(1): 569–594, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ibrahim:2010:RCP

[IM10] H. Ibrahim and R. Monneau. On the rate of convergence in periodic homogenization of scalar first-order ordinary differential equations. SIAM Journal on Mathematical Analysis, 42(5):2155– 2176, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iglesias:2018:LTE

[IM18] Susely Figueroa Iglesias and Sepideh Mirrahimi. Long time evolutionary dynamics of phenotypically structured populations in time-periodic environments. SIAM Journal on Mathematical Analysis, 50(5): 5537–5568, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Isaev:2013:NGS

[IN13] M. I. Isaev and R. G.

Novikov. New global stability estimates for monochromatic inverse acoustic scattering. SIAM Journal on Mathematical Analysis, 45(3): 1495–1504, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iyer:2010:ETD

[INRZ10] Gautam Iyer, Alexei Novikov, Lenya Ryzhik, and Andrej Zlatoš. Exit times of diffusions with incompressible drift. SIAM Journal on Mathematical Analysis, 42(6): 2484–2498, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ignat:2014:URO

[INSZ14] Radu Ignat, Luc Nguyen, Valeriy Slastikov, and Arghir Zarnescu. Uniqueness results for an ODE related to a generalized Ginzburg– Landau model for liquid crystals. SIAM Journal on Mathematical Analysis, 46(5):3390– 3425, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iwaniec:2016:SDW

[IO16] Tadeusz Iwaniec and Jani Onninen. Smoothing defected welds and hairline cracks. SIAM Journal on Mathematical Analysis, 48(1): 281–301, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ichihara:2013:LTB

[IS13] Naoyuki Ichihara and Shuenn-Jyi Sheu. Large time behavior of solutions of Hamilton– Jacobi–Bellman equations with quadratic nonlinearity in gradients. SIAM Journal on Mathematical Analysis, 45(1): 279–306, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Inglis:2015:MFL

[IT15] J. Inglis and D. Talay. Meanfield limit of a stochastic particle system smoothly interacting through threshold hitting-times and applications to neural networks with dendritic component. SIAM Journal on Mathematical Analysis, 47(5):3884– 3916, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ito:2018:AMT

[Ito18] Ryo Ito. Analysis of the minimal traveling wave speed via the methods of Young measures. SIAM Journal on Mathematical Analysis, 50(4): 3478–3534, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Imanuvilov:2012:IBV

[IY12] O. Yu. Imanuvilov and M. Yamamoto. Inverse boundary value problem for Schrödinger equation in two dimensions. SIAM Journal on Mathematical Analysis, 44(3):1333– 1339, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iyer:2019:SPL

[Iye19] Sameer Iyer. Steady Prandtl layers over a moving boundary: Nonshear Euler flows. SIAM Journal on Mathematical Analysis, 51(3):1657– 1695, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiaqi:2018:SNN

[JH18] Yang Jiaqi and Yin Huicheng.

On the steady non-Newtonian fluids in domains with noncompact boundaries. SIAM Journal on Mathematical Analysis, 50(1):283–338, ???? 2018. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Jerez-Hanckes:2012:EVF

[JHN12] Carlos Jerez-Hanckes and Jean-Claude Nédélec. Explicit variational forms for the inverses of integral logarithmic operators over an interval. SIAM Journal on Mathematical Analysis, 44(4): 2666–2694, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2012:SPG

[Jia12] Jin-Cheng Jiang. Smoothing property of the gain term of the Boltzmann collision operator. SIAM Journal on Mathematical Analysis, 44(3): 1522–1543, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2019:ESE

[Jia19] Jie Jiang. Eventual smoothness and exponential stabilization of global weak solutions to some chemotaxis systems. SIAM Journal on Mathematical Analysis, 51(6): 4604–4644, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2018:SEM

[JJ18] Fei Jiang and Song Jiang. On the stabilizing effect of the magnetic fields in the magnetic Rayleigh–Taylor problem. SIAM Journal on Mathematical Analysis, 50(1):491– 540, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2010:ILC

[JJL10] Song Jiang, Qiangchang Ju, and Fucai Li. Incompressible limit of the compressible magnetohydrodynamic equations with vanishing viscosity coefficients. SIAM Journal on Mathematical Analysis, 42(6): 2539–2553, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2016:ILN

[JJL16] Song Jiang, Qiangchang Ju, and Fucai Li. Incompressible limit of the nonisentropic ideal magnetohydrodynamic equations. SIAM Journal on Mathematical Analysis, 48(1): 302–319, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jessle:2013:NSF

[JJN13] Didier Jesslé, Bum Ja Jin, and Antonín Novotný. Navier– Stokes–Fourier system on unbounded domains: Weak solutions, relative entropies, weakstrong uniqueness. SIAM Journal on Mathematical Analysis, 45(3):1907–1951,

????  2013.  CODEN

SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jenssen:2010:ODC

[JK10] Helge Kristian Jenssen and Trygve Klovning Karper. One-dimensional compressible flow with temperature dependent transport coefficients. SIAM Journal on Mathematical Analysis, 42(2): 904–930, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2019:WPE

[JL19] Ning Jiang and Yi-Long Luo. On well-posedness of Ericksen–Leslie’s hyperbolic incompressible liquid crystal model. SIAM Journal on Mathematical Analysis, 51(1): 403–434, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ju:2013:ALF

[JLL13] Qiangchang Ju, Fucai Li, and Yong Li. Asymptotic limits of the full compressible magnetohydrodynamic equations. SIAM Journal on Mathematical Analysis, 45(5): 2597–2624, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2015:NLC

[JLX15] Song Jiang, Fucai Li, and Feng Xie. Nonrelativistic limit of the compressible Navier–Stokes–Fourier– P1 approximation model arising in radiation hydrodynamics. SIAM Journal on Mathematical Analysis, 47(5):3726– 3746, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2018:GCS

[JLZ18] Ning Jiang, Yanan Liu, and Teng-Fei Zhang. Global classical solutions to a compressible model for micro-macro polymeric fluids near equilibrium. SIAM Journal on Mathematical Analysis, 50(4): 4149–4179, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jang:2012:DOL

[JM12] Juhi Jang and Nader Masmoudi. Derivation of Ohm’s Law from the kinetic equations. SIAM Journal on Mathematical Analysis, 44(5): 3649–3669, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jager:2011:HLM

[JMNR11] Willi Jäger, Andro Mikelić, and Maria Neuss-Radu. Homogenization limit of a model system for interaction of flow, chemical reactions, and mechanics in cell tissues. SIAM Journal on Mathematical Analysis, 43(3):1390– 1435, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1390_s1.

Jiu:2014:TDI

[JMWZ14] Quansen Jiu, Changxing Miao, Jiahong Wu, and Zhifei Zhang. The two-dimensional incompressible Boussinesq equations with general critical dissipation. SIAM Journal on Mathematical Analysis, 46(5): 3426–3454, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2018:EAS

[JMZ18] Ansgar Jüngel, Jirí Mikyska, and Nicola Zamponi. Existence analysis of a singlephase flow mixture with van der Waals pressure. SIAM Journal on Mathematical Analysis, 50(1):1367– 1395, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:2013:SSP

[Joh13] Mathew A. Johnson. Stability of small periodic waves in fractional KdV-type equations. SIAM Journal on Mathematical Analysis, 45(5): 3168–3193, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jurak:2019:TPT

[JRK19] Mladen Jurak, Ivana Radisić, and Ana Zgaljić Keko. Twophase two-component flow in porous media in low solubility regime. SIAM Journal on Mathematical Analysis, 51(3): 2019–2052, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jia:2013:MID

[JS13a] Hao Jia and Vladimír Sverák.

Minimal L^{3}-initial data for potential Navier–Stokes singularities. SIAM Journal on Mathematical Analysis, 45(3): 1448–1459, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2013:EAM

[JS13b] Ansgar Jüngel and Ines Viktoria Stelzer. Existence analysis of Maxwell–Stefan systems for multicomponent mixtures. SIAM Journal on Mathematical Analysis, 45(4): 2421–2440, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jesenko:2014:CCR

[JS14] Martin Jesenko and Bernd Schmidt. Closure and commutability results for Γ-limits and the geometric linearization and homogenization of multiwell energy functionals. SIAM Journal on Mathematical Analysis, 46(4):2525– 2553, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jung:2013:SPB

[JT13] Chang-Yeol Jung and Roger Temam. Singular perturbations and boundary layer theory for convection-diffusion equations in a circle: The generic noncompatible case. SIAM Journal on Mathematical Analysis, 44(6):4274– 4296, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jang:2016:CVS

[JTW16] Juhi Jang, Ian Tice, and Yanjin Wang. The compressible viscous surface-internal wave problem: Local wellposedness. SIAM Journal on Mathematical Analysis, 48(4): 2602–2673, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jungel:2010:GWS

[Jün10] Ansgar Jüngel. Global weak solutions to compressible Navier–Stokes equations for quantum fluids. SIAM Journal on Mathematical Analysis, 42(3):1025–1045, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Junca:2014:HFW

[Jun14] Stéphane Junca. High frequency waves and the maximal smoothing effect for nonlinear scalar conservation laws. SIAM Journal on Mathematical Analysis, 46(3): 2160–2184, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiu:2013:VBA

[JWX13] Quansen Jiu, Yi Wang, and Zhouping Xin. Vacuum behaviors around rarefaction waves to 1D compressible Navier–Stokes equations with density-dependent viscosity. SIAM Journal on Mathematical Analysis, 45(5):3194– 3228, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2015:DLB

[JX15] Ning Jiang and Linjie Xiong.

Diffusive limit of the Boltzmann equation with fluid initial layer in the periodic domain. SIAM Journal on Mathematical Analysis, 47(3): 1747–1777, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2016:HLK

[JXZ16] Ning Jiang, Linjie Xiong, and Teng-Fei Zhang. Hydrodynamic limits of the kinetic self-organized models. SIAM Journal on Mathematical Analysis, 48(5):3383– 3411, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:2010:TIP

[JZ10] Mathew A. Johnson and Kevin Zumbrun. Transverse instability of periodic traveling waves in the generalized Kadomtsev–Petviashvili equation. SIAM Journal on Mathematical Analysis, 42(6): 2681–2702, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jin:2018:HUR

[JZ18] Shi Jin and Yuhua Zhu. Hypocoercivity and uniform regularity for the Vlasov– Poisson–Fokker–Planck system with uncertainty and multiple scales. SIAM Journal on Mathematical Analysis, 50(2):1790–1816, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:2019:BEI

[JZ19] Ning Jiang and Xu Zhang.

The Boltzmann equation with incoming boundary condition: Global solutions and Navier– Stokes limit. SIAM Journal on Mathematical Analysis, 51(3):2504–2534, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:2011:NSV

[JZN11] Mathew A. Johnson, Kevin Zumbrun, and Pascal Noble. Nonlinear stability of viscous roll waves. SIAM Journal on Mathematical Analysis, 43(2): 577–611, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p577_s1.

Kachmar:2014:GLO

[Kac14]  Ayman Kachmar.  The

Ginzburg–Landau order parameter near the second critical field. SIAM Journal on Mathematical Analysis, 46(1): 572–587, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaiser:2010:GDP

[Kai10]  Ralf Kaiser.  The geo-

magnetic direction problem: The two-dimensional and the three-dimensional axisymmetric cases. SIAM Journal on Mathematical Analysis, 42(2): 701–728, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaib:2017:SSA

[Kai17] Gunnar Kaib. Stationary states of an aggregation equation with degenerate diffusion and bounded attractive potential. SIAM Journal on Mathematical Analysis, 49(1): 272–296, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalvin:2012:LAP

[Kal12] Victor Kalvin. Limiting absorption principle and perfectly matched layer method for Dirichlet Laplacians in quasi-cylindrical domains. SIAM Journal on Mathematical Analysis, 44(1):355– 382, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p355_s1.

Karatson:2012:CMI

[Kar12] János Karátson. Characterizing mesh independent quadratic convergence of Newton’s method for a class of elliptic problems. SIAM Journal on Mathematical Analysis, 44(3):1279–1303, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Karakhanyan:2016:IPR

[Kar16] Aram L. Karakhanyan. An inverse problem for the refractive surfaces with parallel lighting. SIAM Journal on Mathematical Analysis, 48(1): 740–784, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Katzourakis:2019:RSI

[Kat19] Nikos Katzourakis. An L^{∞} regularization strategy to the inverse source identification problem for elliptic equations. SIAM Journal on Mathematical Analysis, 51(2): 1349–1370, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kollar:2019:DCS

[KDT19] Richard Kollár, Bernard Deconinck, and Olga Trichtchenko. Direct characterization of spectral stability of smallamplitude periodic waves in scalar Hamiltonian problems via dispersion relation. SIAM Journal on Mathematical Analysis, 51(4):3145– 3169, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Khapalov:2013:MSD

[Kha13] Alexander Y. Khapalov. Micromotions of a swimmer in the 3-D incompressible fluid governed by the nonstationary Stokes equation. SIAM Journal on Mathematical Analysis, 45(6):3360–3381, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kian:2016:RTD

[Kia16] Yavar Kian. Recovery of timedependent damping coefficients and potentials appearing in wave equations from partial data. SIAM Journal on Mathematical Analysis, 48(6):4021–4046, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2009:LFL

[Kim09] Namkwon Kim. Large friction limit and the inviscid limit of 2D Navier–Stokes equations under Navier friction condition. SIAM Journal on Mathematical Analysis, 41(4): 1653–1663, ???? 2009. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigendum [Kim13].

Kim:2013:CLF

[Kim13] Namkwon Kim. Corrigendum: Large Friction Limit and the Inviscid Limit of 2D Navier–Stokes Equations under Navier Friction Condition. SIAM Journal on Mathematical Analysis, 45(3): 1992–1994, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Kim09].

Komech:2010:GAS

[KK10] Alexander Komech and Andrew Komech. Global attraction to solitary waves for a nonlinear Dirac equation with mean field interaction. SIAM Journal on Mathematical Analysis, 42(6): 2944–2964, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p2944_s1.

Kim:2015:WSE

[KK15] Hyunseok Kim and Young-Heon Kim. On weak solutions of elliptic equations with singular drifts. SIAM Journal on Mathematical Analysis, 47(2): 1271–1290, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kreisbeck:2016:HTF

[KK16] Carolin Kreisbeck and Stefan Krömer. Heterogeneous thin films: Combining homogenization and dimension reduction with directors. SIAM Journal on Mathematical Analysis, 48(2): 785–820, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kang:2017:EWS

[KK17a] Kyungkeun Kang and Hwa Kil Kim. Existence of weak solutions in Wasserstein space for a chemotaxis model coupled to fluid equations. SIAM Journal on Mathematical Analysis, 49(4):2965– 3004, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kwon:2017:IJR

[KK17b] Oh Sung Kwon and Jae Ryong Kweon. Interior jump and regularity of compressible viscous Navier–Stokes flows through a cut. SIAM Journal on Mathematical Analysis, 49(3): 1982–2008, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Komech:2018:OSG

[KK18] A. Komech and E. Kopylova. On orbital stability of ground states for finite crystals in fermionic Schrödinger– Poisson model. SIAM Journal on Mathematical Analysis, 50(1):64–85, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klar:2017:TED

[KKT17] Axel Klar, Lisa Kreusser, and Oliver Tse. Trend to equilibrium for a delay Vlasov–Fokker–Planck equation and explicit decay estimates. SIAM Journal on Mathematical Analysis, 49(4): 3277–3298, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kirsch:2018:LAP

[KL18a] Andreas Kirsch and Armin Lechleiter. The limiting absorption principle and a radiation condition for the scattering by a periodic layer. SIAM Journal on Mathematical Analysis, 50(3):2536– 2565, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuwert:2018:AWM

[KL18b] Ernst Kuwert and Yuxiang Li. Asymptotics of Willmore minimizers with prescribed small isoperimetric ratio. SIAM Journal on Mathematical Analysis, 50(4):4407– 4425, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kutyniok:2012:OSA

[KLL12] Gitta Kutyniok, Jakob Lemvig, and Wang-Q Lim. Optimally sparse approximations of 3D functions by compactly supported shearlet frames. SIAM Journal on Mathematical Analysis, 44(4):2962– 3017, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolehmainen:2010:CIP

[KLO10] Ville Kolehmainen, Matti Lassas, and Petri Ola. Calderón’s inverse problem with an imperfectly known boundary and reconstruction up to a conformal deformation. SIAM Journal on Mathematical Analysis, 42(3):1371– 1381, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keller:2016:NPS

[KLO16] Johannes Keller, Caroline Lasser, and Tomoki Ohsawa. A new phase space density for quantum expectations. SIAM Journal on Mathematical Analysis, 48(1): 513–537, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krupchyk:2011:DEH

[KLS11] Katsiaryna Krupchyk, Matti Lassas, and Samuli Siltanen. Determining electrical and heat transfer parameters using coupled boundary measurements. SIAM Journal on Mathematical Analysis, 43(5):

2096–2115, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2096_s1.

Kuan:2015:EMA

[KLS15] Rulin Kuan, Yi-Hsuan Lin, and Mourad Sini. The enclosure method for the anisotropic Maxwell system. SIAM Journal on Mathematical Analysis, 47(5):3488– 3527, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kavallaris:2017:DNP

[KLW17] Nikos I. Kavallaris, Johannes Lankeit, and Michael Winkler. On a degenerate nonlocal parabolic problem describing infinite dimensional replicator dynamics. SIAM Journal on Mathematical Analysis, 49(2): 954–983, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kamotski:2013:LWW

[KM13] I. V. Kamotski and V. G.

Maz’ya. On the linear water wave problem in the presence of a critically submerged body. SIAM Journal on Mathematical Analysis, 44(6): 4222–4249, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kappeler:2017:WPD

[KM17] Thomas Kappeler and Jan-Cornelius Molnar. On the well-posedness of the defocusing mKdV equation below L^{2}. SIAM Journal on Mathematical Analysis, 49(3): 2191–2219, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krishnan:2018:MAR

[KM18] Venkateswaran P. Krishnan and Rohit Kumar Mishra. Microlocal analysis of a restricted ray transform on symmetric m-tensor fields in R^{n}. SIAM Journal on Mathematical Analysis, 50(6):6230– 6254, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurzke:2011:VML

[KMM11] Matthias Kurzke, Christof Melcher, and Roger Moser. Vortex motion for the Landau– Lifshitz–Gilbert equation with spin-transfer torque. SIAM Journal on Mathematical Analysis, 43(3):1099–1121, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1099_s1.

Kroner:2015:TFB

[KMS15] Dietmar Kröner, Thomas Müller, and Lena Maria Strehlau. Traces for functions of bounded variation on manifolds with applications to conservation laws on manifolds with boundary. SIAM Journal on Mathematical Analysis, 47(5):3944–3962, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurzke:2017:GPV

[KMS17] Matthias Kurzke, Jeremy L.

Marzuola, and Daniel Spirn. Gross–Pitaevskii vortex motion with critically scaled inhomogeneities. SIAM Journal on Mathematical Analysis, 49 (1):471–500, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Karper:2013:EWS

[KMT13] Trygve K. Karper, Antoine Mellet, and Konstantina Trivisa. Existence of weak solutions to kinetic flocking models. SIAM Journal on Mathematical Analysis, 45(1):215– 243, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Killip:2018:IVP

[KMV18] Rowan Killip, Jason Murphy, and Monica Visan. The initial-value problem for the cubic-quintic NLS with nonvanishing boundary conditions. SIAM Journal on Mathematical Analysis, 50(3):

2681–2739, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kukavica:2014:LWP

[KMVW14] Igor Kukavica, Nader Masmoudi, Vlad Vicol, and Tak Kwong Wong. On the local well-posedness of the Prandtl and hydrostatic Euler equations with multiple monotonicity regions. SIAM Journal on Mathematical Analysis, 46(6):3865– 3890, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Knupfer:2018:OSI

[KN18] Hans Knüpfer and Florian Nolte. Optimal shape of isolated ferromagnetic domains. SIAM Journal on Mathematical Analysis, 50(6):5857– 5886, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Korzec:2012:GWS

[KNR12] M. D. Korzec, P. Nayar, and P. Rybka. Global weak solutions to a sixth order Cahn–Hilliard type equation. SIAM Journal on Mathematical Analysis, 44(5):3369– 3387, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krahmer:2015:CSR

[KNW15] Felix Krahmer, Deanna Needell, and Rachel Ward. Compressive sensing with redundant dictionaries and structured measurements. SIAM Journal on Mathematical Analysis, 47(6):4606–4629, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koch:2016:VOS

[Koc16] Hans Koch. Vertex order in some large constrained random graphs. SIAM Journal on Mathematical Analysis, 48(4):2588–2601, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kollar:2011:HMN

[Kol11] Richard Kollár. Homotopy method for nonlinear eigenvalue pencils with applications. SIAM Journal on Mathematical Analysis, 43(2): 612–633, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p612_s1.

Komech:2015:CGS

[Kom15]  A. I. Komech.  On the

crystal ground state in the Schrödinger–Poisson model. SIAM Journal on Mathematical Analysis, 47(2):1001– 1021, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kotschote:2012:DCN

[Kot12] Matthias Kotschote. Dynamics of compressible nonisothermal fluids of non-Newtonian Korteweg type. SIAM Journal on Mathematical Analysis, 44(1):74– 101, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p74_s1.

Kroo:2013:SMP

[KP13] András Kroó and Allan Pinkus. On stability of the metric projection operator. SIAM Journal on Mathematical Analysis, 45(2): 639–661, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2018:CBW

[KP18a] Young-Heon Kim and Brendan Pass. A canonical barycenter via Wasserstein regularization. SIAM Journal on Mathematical Analysis, 50(2):1817–1828, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kraitzman:2018:PBS

[KP18b] Noa Kraitzman and Keith Promislow. Pearling bifurcations in the strong functionalized Cahn–Hilliard free energy. SIAM Journal on Mathematical Analysis, 50(3): 3395–3426, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Khusainov:2015:SME

[KPR15] Denys Khusainov, Michael Pokojovy, and Reinhard Racke. Strong and mild extrapolated L^{2}-solutions to the heat equation with constant delay. SIAM Journal on Mathematical Analysis, 47(1): 427–454, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keimer:2018:NSC

[KPS18] Alexander Keimer, Lukas Pflug, and Michele Spinola. Nonlocal scalar conservation laws on bounded domains and applications in traffic flow. SIAM Journal on Mathematical Analysis, 50(6):6271– 6306, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kassmann:2019:HLT

[KPZ19] M. Kassmann, A. Piatnitski, and E. Zhizhina. Homogenization of Lévy-type operators with oscillating coefficients. SIAM Journal on Mathematical Analysis, 51(5): 3641–3665, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kloucek:2010:TFB

[KR10] Petr Klouček and Michel V. Romerio. Tracking free boundaries in fluids using a variational principle. SIAM Journal on Mathematical Analysis, 42(4):1519– 1538, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krejcirik:2019:CMF

[Kre19] David Krejcirík. Complex magnetic fields: an improved Hardy–Laptev–Weidl inequality and quasi-self-adjointness. SIAM Journal on Mathematical Analysis, 51(2):790– 807, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koumatos:2015:DIY

[KRW15] Konstantinos Koumatos, Filip Rindler, and Emil Wiedemann. Differential inclusions and Young measures involving prescribed Jacobians. SIAM Journal on Mathematical Analysis, 47(2):1169– 1195, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:2010:DFS

[Kry10] N. V. Krylov. On divergence form SPDEs with growing coefficients in W_{2}^{1} spaces without weights. SIAM Journal on Mathematical Analysis, 42 (2):609–633, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:2013:EET

[Kry13] N. V. Krylov. An ersatz existence theorem for fully nonlinear parabolic equations without convexity assumptions. SIAM Journal on Mathematical Analysis, 45(6): 3331–3359, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:2014:HTP

[Kry14] N. V. Krylov. Hörmander’s theorem for parabolic equations with coefficients measurable in the time variable. SIAM Journal on Mathematical Analysis, 46(1):854– 870, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kar:2014:RIE

[KS14] Manas Kar and Mourad Sini.

Reconstruction of interfaces from the elastic farfield measurements using CGO solutions. SIAM Journal on Mathematical Analysis, 46(4):

2650–2691, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kian:2019:HSD

[KS19] Yavar Kian and Eric Soccorsi. Hölder stably determining the time-dependent electromagnetic potential of the Schrödinger equation. SIAM Journal on Mathematical Analysis, 51(2):627– 647, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolb:2013:NED

[KSW13] Martin Kolb, Mladen Savov, and Achim Wübker. (Non- )Ergodicity of a degenerate diffusion modeling the fiber lay down process. SIAM Journal on Mathematical Analysis, 45(1):1–13, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kondo:2011:IBV

[KT11] Shintaro Kondo and Atusi Tani. Initial boundary value problem for model equations of resistive drift wave turbulence. SIAM Journal on Mathematical Analysis, 43(2): 925–943, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p925_s1.

Kim:2013:CPF

[KT13] Sungwhan Kim and Alexandru Tamasan. On a Calderón problem in frequency differential electrical impedance tomography. SIAM Journal on Mathematical Analysis, 45(5): 2700–2709, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2017:DSS

[KT17] Chanwoo Kim and Ian Tice.

Dynamics and stability of surface waves with surfactants. SIAM Journal on Mathematical Analysis, 49(2): 1295–1332, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kiselev:2018:GRE

[KT18] Alexander Kiselev and Changhui Tan. Global regularity for 1D Eulerian dynamics with singular interaction forces. SIAM Journal on Mathematical Analysis, 50(6):6208– 6229, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuto:2015:LSS

[Kut15]  Kousuke Kuto.  Limit-

ing structure of shrinking solutions to the stationary Shigesada–Kawasaki– Teramoto model with large cross-diffusion. SIAM Journal on Mathematical Analysis, 47(5):3993–4024, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kondratyev:2016:BVB

[KUV16] Stanislav Kondratyev, José Miguel Urbano, and Dmitry Vorotnikov. On the bulk velocity of Brownian ratchets. SIAM Journal on Mathematical Analysis, 48(2):950– 980, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kondratyev:2019:SHK

[KV19] Stanislav Kondratyev and Dmitry Vorotnikov. Spherical Hellinger–Kantorovich gradient flows. SIAM Journal on Mathematical Analysis, 51(3): 2053–2084, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaliuzhnyi-Verbovetskyi:2017:SHE

[KVM17] Dmitry Kaliuzhnyi-Verbovetskyi and Georgi S. Medvedev. The semilinear heat equation on sparse random graphs. SIAM Journal on Mathematical Analysis, 49(2):1333– 1355, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaliuzhnyi-Verbovetskyi:2018:MFE

[KVM18] Dmitry Kaliuzhnyi-Verbovetskyi and Georgi S. Medvedev. The mean field equation for the Kuramoto model on graph sequences with non-Lipschitz limit. SIAM Journal on Mathematical Analysis, 50(3): 2441–2465, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kimura:2019:PDS

[KvMY19] Masato Kimura, Patrick van Meurs, and Zhenxing Yang. Particle dynamics subject to impenetrable boundaries: Existence and uniqueness of mild solutions. SIAM Journal on Mathematical Analysis, 51(6): 5049–5076, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krahmer:2011:NIJ

[KW11] Felix Krahmer and Rachel Ward. New and improved Johnson–Lindenstrauss embeddings via the restricted isometry property. SIAM Journal on Mathematical Analysis, 43(3):1269–1281, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1269_s1.

Kitagawa:2012:ROT

[KW12] Jun Kitagawa and Micah Warren. Regularity for the optimal transportation problem with Euclidean distance squared cost on the embedded sphere. SIAM Journal on Mathematical Analysis, 44(4): 2871–2887, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kweon:2012:CSD

[Kwe12] Jae Ryong Kweon. Corner singularity dynamics and regularity of compressible viscous Navier–Stokes flows. SIAM Journal on Mathematical Analysis, 44(5):3127– 3161, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2012:BEN

[KY12a] Chanwoo Kim and Seok-Bae Yun. The Boltzmann equation near a rotational local Maxwellian. SIAM Journal on Mathematical Analysis, 44(4): 2560–2598, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:2012:PKS

[KY12b] Inwon Kim and Yao Yao. The Patlak–Keller–Segel model and its variations: Properties of solutions via maximum principle. SIAM Journal on Mathematical Analysis, 44(2): 568–602, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.

siam.org/sima/resource/
1/sjmaah/v44/i2/p568_s1.

Kim:2015:CII

[KY15] Seonghak Kim and Baisheng Yan. Convex integration and infinitely many weak solutions to the Perona–Malik equation in all dimensions. SIAM Journal on Mathematical Analysis, 47(4):2770–2794, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kawano:2011:UNR

[KZ11] A. Kawano and A. Zine.

Uniqueness and nonuniqueness results for a certain class of almost periodic distributions. SIAM Journal on Mathematical Analysis, 43(1): 135–152, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p135_s1.

Kim:2018:RPD

[KZ18] Inwon Kim and Yuming Paul Zhang. Regularity properties of degenerate diffusion equations with drifts. SIAM Journal on Mathematical Analysis, 50(4):4371–4406, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Loy:2014:IRC

[LA14] R. J. Loy and R. S. Anderssen. Interconversion relationships for completely monotone functions. SIAM Journal on Mathematical Analysis, 46(3): 2008–2032, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lacave:2015:UTD

[Lac15] Christophe Lacave. Uniqueness for two-dimensional incompressible ideal flow on singular domains. SIAM Journal on Mathematical Analysis, 47(2):1615–1664, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2014:RPS

[Lai14] Geng Lai. On the Riemann problem for a scalar Zeldovich–von Neumann– Döring combustion model. SIAM Journal on Mathematical Analysis, 46(4):2404– 2443, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2018:ICW

[Lai18] Geng Lai. Interactions of composite waves of the twodimensional full Euler equations for van der Waals gases. SIAM Journal on Mathematical Analysis, 50(4):3535– 3597, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lam:2012:LPS

[Lam12] King-Yeung Lam. Limiting profiles of semilinear elliptic equations with large advection in population dynamics II. SIAM Journal on Mathematical Analysis, 44(3):1808– 1830, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurent:2010:GCS

[Lau10] Camille Laurent. Global controllability and stabilization for the nonlinear Schrödinger equation on some compact manifolds of dimension 3. SIAM Journal on Mathematical Analysis, 42(2):785– 832, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurain:2018:ASS

[Lau18] Antoine Laurain. Analyzing smooth and singular domain perturbations in level set methods. SIAM Journal on Mathematical Analysis, 50(4): 4327–4370, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LeBlanc:2013:PGR

[Le 13]  Valérie Le Blanc.  Per-

sistence of generalized rollwaves under viscous perturbation. SIAM Journal on Mathematical Analysis, 45(2):

572–599, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Le:2010:COK

[Le10] Nam Q. Le. On the convergence of the Ohta–Kawasaki equation to motion by nonlocal Mullins–Sekerka law. SIAM Journal on Mathematical Analysis, 42(4):1602– 1638, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Le:2019:EIS

[Le19]  Hung Le.  On the exis-

tence and instability of solitary water waves with a finite dipole. SIAM Journal on Mathematical Analysis, 51(5): 4074–4104, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lee:2010:NCN

[Lee10]  Paul W. Y. Lee.  New

computable necessary conditions for the regularity theory of optimal transportation. SIAM Journal on Mathematical Analysis, 42(6): 3054–3075, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3054_s1.

Lee:2016:DCP

[Lee16] Minjae Lee. Dirac cones for point scatterers on a honeycomb lattice. SIAM Journal on Mathematical Analysis, 48(2):1459–1488, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lee:2017:UEV

[Lee17] Donghyun Lee. Uniform estimate of viscous free-boundary magnetohydrodynamics with zero vacuum magnetic field. SIAM Journal on Mathematical Analysis, 49(4):2710– 2789, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lei:2010:CEC

[Lei10] Guang-Tsai Lei. Characterization of electromagnetic cavity resonators by integral identities. SIAM Journal on Mathematical Analysis, 42(2): 634–645, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lei:2013:QAS

[Lei13] Yutian Lei. Qualitative analysis for the static Hartree-type equations. SIAM Journal on Mathematical Analysis, 45(1): 388–406, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leitmeyer:2016:ECP

[Lei16] Keith Leitmeyer. Enstrophy cascade in physical scales for the three-dimensional Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 48(1):166–173, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lengeler:2014:WSI

[Len14] Daniel Lengeler. Weak solutions for an incompressible, generalized Newtonian fluid interacting with a linearly elastic Koiter type Shell. SIAM Journal on Mathematical Analysis, 46(4):2614– 2649, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lenells:2016:NSD

[Len16]  Jonatan Lenells.  The

nonlinear steepest descent method: Asymptotics for initial-boundary value problems. SIAM Journal on Mathematical Analysis, 48(3): 2076–2118, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lequeurre:2011:ESS

[Leq11]  Julien Lequeurre.  Exis-

tence of strong solutions to a fluid-structure system. SIAM Journal on Mathematical Analysis, 43(1):389–

410, ????  2011.  CO-

DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p389_s1.

Li:2009:MEF

[Li09] Bo Li. Minimization of electrostatic free energy and the Poisson–Boltzmann equation for molecular solvation with implicit solvent. SIAM Journal on Mathematical Analysis, 40(6):2536–2566, ???? 2009. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Li11].

Li:2011:EME

[Li11] Bo Li. Erratum: “Minimization of Electrostatic Free Energy and the Poisson– Boltzmann Equation for Molecular Solvation with Implicit Solvent”. SIAM Journal on Mathematical Analysis, 43(6):2776–2777, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2776_s1.

See [Li09].

Li:2019:GWP

[Li19]  Jinkai Li.  Global well-

posedness of the one-dimensional compressible Navier–Stokes equations with constant heat conductivity and nonnegative density. SIAM Journal on Mathematical Analysis, 51(5): 3666–3693, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Littmann:2013:QEB

[Lit13] Friedrich Littmann. Quadrature and extremal bandlimited functions. SIAM Journal on Mathematical Analysis, 45 (2):732–747, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2018:GIL

[Liu18a]  Boya Liu.  Global iden-

tifiability of low regularity fluid parameters in acoustic tomography of moving fluid. SIAM Journal on Mathematical Analysis, 50(6): 6348–6364, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2018:MRG

[Liu18b]  Xin Liu.  A model of

radiational gaseous stars. SIAM Journal on Mathematical Analysis, 50(6):6100– 6155, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2019:MSC

[Liu19] Grace Liu. Modified scattering for the cubic Schrödinger equation small data solution on product space. SIAM Journal on Mathematical Analysis, 51(5):4023–4073, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lessard:2017:CAF

[LJ17] Jean-Philippe Lessard and J. D. Mireles James. Computer assisted Fourier analysis in sequence spaces of varying regularity. SIAM Journal on Mathematical Analysis, 49 (1):530–561, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2011:RSP

[LL11] Y. Charles Li and Zhiwu Lin. A resolution of the Sommerfeld paradox. SIAM Journal on Mathematical Analysis, 43(4):1923–1954, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1923_s1.

Lee:2012:NES

[LL12] Paul W. Y. Lee and Jiayong Li. New examples satisfying Ma–Trudinger–Wang conditions. SIAM Journal on Mathematical Analysis, 44 (1):61–73, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p61_s1.

Lakshtanov:2016:DFM

[LL16a] Evgeny Lakshtanov and Armin Lechleiter. Difference factorizations and monotonicity in inverse medium scattering for contrasts with fixed sign on the boundary. SIAM Journal on Mathematical Analysis, 48(6):3688–3707, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2016:ALR

[LL16b] Hongjie Li and Hongyu Liu. On anomalous localized resonance for the elastostatic system. SIAM Journal on Mathematical Analysis, 48(5):3322– 3344, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2018:GDC

[LL18a] Lei Li and Jian-Guo Liu. A generalized definition of Caputo derivatives and its application to fractional ODEs. SIAM Journal on Mathematical Analysis, 50(3):2867– 2900, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2018:SCC

[LL18b] Lei Li and Jian-Guo Liu. Some compactness criteria for weak solutions of time fractional PDEs. SIAM Journal on Mathematical Analysis, 50(4):3963–3995, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2019:IES

[LL19] Jianliang Li and Peijun Li. Inverse elastic scattering for a random source. SIAM Journal on Mathematical Analysis, 51(6):4570–4603, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2014:ESN

[LLLM14] Chi-Kun Lin, Chi-Tien Lin, Yanping Lin, and Ming Mei. Exponential stability of nonmonotone traveling waves for Nicholson’s blowflies equation. SIAM Journal on Mathematical Analysis, 46(2): 1053–1084, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2019:DRS

[LLM19] Jingzhi Li, Hongyu Liu, and Shiqi Ma. Determining a random Schrödinger equation with unknown source and potential. SIAM Journal on Mathematical Analysis, 51(4): 3465–3491, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2019:SLS

[LLN19] Jianfeng Lu, Yulong Lu, and James Nolen. Scaling limit of the Stein variational gradient descent: The mean field regime. SIAM Journal on Mathematical Analysis, 51(2): 648–671, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levermore:2016:GDB

[LLP16] C. David Levermore, Hailiang Liu, and Robert L. Pego. Global dynamics of Bose– Einstein condensation for a model of the Kompaneets equation. SIAM Journal on Mathematical Analysis, 48(4): 2454–2494, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2015:GFE

[LLW15] Junyu Lin, Baishun Lai, and Changyou Wang. Global finite energy weak solutions to the compressible nematic liquid crystal flow in dimension three. SIAM Journal on Mathematical Analysis, 47(4): 2952–2983, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lai:2017:FSS

[LLW17] Baishun Lai, Junyu Lin, and Changyou Wang. Forward self-similar solutions to the viscoelastic Navier– Stokes equation with damping. SIAM Journal on Mathematical Analysis, 49(1):501– 529, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2018:SSV

[lLWW18] Hai liang Li, Teng Wang, and Yi Wang. Stability of the superposition of a viscous contact wave with two rarefaction waves to the bipolar Vlasov– Poisson–Boltzmann system. SIAM Journal on Mathematical Analysis, 50(2):1829– 1876, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lecureux-Mercier:2011:GSS

[LM11]  Magali  Lécureux-Mercier.

Global smooth solutions of Euler equations for van der Waals gases. SIAM Journal on Mathematical Analysis, 43 (2):877–903, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p877_s1.

Lecureux-Mercier:2014:PEA

[LM14] Magali Lécureux-Mercier. A priori estimates and analytical construction of radially symmetric solutions in the gas dynamics. SIAM Journal on Mathematical Analysis, 46(4):

2853–2883, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurencot:2017:SST

[LM17] Philippe Laurençot and Bogdan-Vasile Matioc. Self-similarity in a thin film muskat problem. SIAM Journal on Mathematical Analysis, 49(4):2790– 2842, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lukacova-Medvidova:2017:GER

[LMMNR17] Mária Lukácová-Medvidová, Hana Mizerová, Sárka Necasová, and Michael Renardy. Global existence result for the generalized Peterlin viscoelastic model. SIAM Journal on Mathematical Analysis, 49(4): 2950–2964, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lattanzio:2010:SSR

[LMN^{+}10] Corrado Lattanzio, Corrado Mascia, Toan Nguyen, Ramón G. Plaza, and Kevin Zumbrun. Stability of scalar radiative shock profiles. SIAM Journal on Mathematical Analysis, 41(6):2165– 2206, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lattanzio:2011:MBC

[LMP11] Corrado Lattanzio, Amelio Maurizi, and Benedetto Piccoli. Moving bottlenecks in car traffic flow: a PDE–ODE coupled model. SIAM Journal on Mathematical Analysis, 43 (1):50–67, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p50_s1.

Lemou:2013:SGS

[LMR13] Mohammed Lemou, Florian Méhats, and Cyril Rigault. Stable ground states and self-similar blow-up solutions for the gravitational Vlasov– Manev system. SIAM Journal on Mathematical Analysis, 44(6):3928–3968, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lawley:2015:SSI

[LMR15] Sean D. Lawley, Jonathan C.

Mattingly, and Michael C. Reed. Stochastic switching in infinite dimensions with applications to random parabolic PDE. SIAM Journal on Mathematical Analysis, 47(4): 3035–3063, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liero:2016:OTC

[LMS16] Matthias Liero, Alexander Mielke, and Giuseppe Savaré.

Optimal transport in competition with reaction: The Hellinger–Kantorovich distance and geodesic curves. SIAM Journal on Mathematical Analysis, 48(4):2869– 2911, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laul:2015:LEE

[LMTT15] Parul Laul, Jason Metcalfe, Shreyas Tikare, and Mihai Tohaneanu. Localized energy estimates for wave equations on (1 + 4)-dimensional Myers–Perry space-times. SIAM Journal on Mathematical Analysis, 47 (3):1933–1957, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2017:SSC

[LMW17] Fucai Li, Yanmin Mu, and Dehua Wang. Strong solutions to the compressible Navier–Stokes–Vlasov– Fokker–Planck equations: Global existence near the equilibrium and large time behavior. SIAM Journal on Mathematical Analysis, 49(2):984– 1026, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2017:SHM

[LMZZ17] Jingyu Li, Ming Mei, Guojing Zhang, and Kaijun Zhang.

Steady hydrodynamic model of semiconductors with sonic boundary: (I) subsonic doping profile. SIAM Journal on Mathematical Analysis, 49(6): 4767–4811, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2018:SHM

[LMZZ18] Jingyu Li, Ming Mei, Guojing Zhang, and Kaijun Zhang. Steady hydrodynamic model of semiconductors with sonic boundary: (II) supersonic doping profile. SIAM Journal on Mathematical Analysis, 50 (1):718–734, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2010:UCP

[LN10] Ching-Lung Lin and Gen Nakamura. Unique continuation property for a coupled second-fourth order dynamical system and its application. SIAM Journal on Mathematical Analysis, 42(5): 2318–2336, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2014:ESS

[LNZ14] Nan Lu, Andrea R. Nahmod, and Chongchun Zeng. Equivariante and self-similar standing waves for a Hamiltonian hyperbolic-hyperbolic spin-field system. SIAM Journal on Mathematical Analysis, 46(6):3913–3956, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lobus:2018:BSS

[Löb18] Jörg-Uwe Löbus. Boundedness of the stationary solution to the Boltzmann equation with spatial smearing, diffusive boundary conditions, and Lions’ collision kernel. SIAM Journal on Mathematical Analysis, 50(6):5761– 5782, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lopez:2012:BCW

[Lóp12] Rafael López. Bifurcation of cylinders for wetting and dewetting models with striped geometry. SIAM Journal on Mathematical Analysis, 44(2): 946–965, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorenzi:2014:DPP

[LP14] Alfredo Lorenzi and Viatcheslav Priimenko. Direct problems for poroelastic waves with fractional derivatives. SIAM Journal on Mathematical Analysis, 46(3):1874– 1892, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2016:LWL

[LP16] Huimin Liu and Xueke Pu.

Long wavelength limit for the quantum Euler–Poisson equation. SIAM Journal on Mathematical Analysis, 48(4): 2345–2381, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Le:2019:CPM

[LP19] Uyen Le and Dmitry E. Pelinovsky. Convergence of Petviashvili’s method near periodic waves in the fractional Korteweg–de Vries equation. SIAM Journal on Mathematical Analysis, 51(4):2850– 2883, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lepoutre:2012:GWP

[LPR12] Thomas Lepoutre, Michel Pierre, and Guillaume Rolland. Global well-posedness of a conservative relaxed cross diffusion system. SIAM Journal on Mathematical Analysis, 44(3):1674–1693, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2010:WBO

[LPS10] Yue Liu, Dmitry Pelinovsky, and Anton Sakovich. Wave breaking in the Ostrovsky– Hunter equation. SIAM Journal on Mathematical Analysis, 42(5):1967–1985, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Linares:2013:WPS

[LPS13] Felipe Linares, Didier Pilod, and Jean-Claude Saut. Wellposedness of strongly dispersive two-dimensional surface wave Boussinesq systems. SIAM Journal on Mathematical Analysis, 44(6): 4195–4221, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Linares:2018:CPF

[LPS18] Felipe Linares, Didier Pilod, and Jean-Claude Saut. The Cauchy problem for the fractional Kadomtsev– Petviashvili equations. SIAM Journal on Mathematical Analysis, 50(3):3172–3209, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2011:ODN

[LR11] Dong Li and José L. Rodrigo. On a one-dimensional nonlocal flux with fractional dissipation. SIAM Journal on Mathematical Analysis, 43(1): 507–526, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p507_s1.

Lewin:2013:DPP

[LR13] Mathieu Lewin and Nicolas Rougerie. Derivation of Pekar’s polarons from a microscopic model of quantum crystal. SIAM Journal on Mathematical Analysis, 45(3): 1267–1301, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lechleiter:2015:IOD

[LR15a] Armin Lechleiter and Marcel Rennoch. Inside-outside duality and the determination of electromagnetic interior transmission eigenvalues. SIAM Journal on Mathematical Analysis, 47(1):684– 705, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lemarie–Rieusset:2015:SCT

[LR15b] Pierre Gilles Lemarié-Rieusset.

On some classes of timeperiodic solutions for the Navier–Stokes equations in the whole space. SIAM Journal on Mathematical Analysis, 47(2):1022–1043, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lindsey:2017:OTM

[LR17] Michael Lindsey and Yanir A.

Rubinstein. Optimal transport via a Monge–Ampère optimization problem. SIAM Journal on Mathematical Analysis, 49(4):3073–3124, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2018:QLS

[LRdS18] Wei Liu, Michael Röckner, and José Luís da Silva. Quasi-linear (stochastic) partial differential equations with time-fractional derivatives. SIAM Journal on Mathematical Analysis, 50(3):2588– 2607, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2010:GEW

[LS10] Buyang Li and Weiwei Sun. Global existence of weak solution for nonisothermal multicomponent flow in porous textile media. SIAM Journal on Mathematical Analysis, 42(6): 3076–3102, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3076_s1.

Li:2012:GWS

[LS12a] Buyang Li and Weiwei Sun. Global weak solution for a heat and sweat transport system in three-dimensional fibrous porous media with Condensation/Evaporation and absorption. SIAM Journal on Mathematical Analysis, 44(3):

1448–1473, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2012:CDP

[LS12b] Hailiang Liu and Jaemin Shin.

The Cauchy–Dirichlet problem for the finitely extensible nonlinear elastic dumbbell model of polymeric fluids. SIAM Journal on Mathematical Analysis, 44(5):3617– 3648, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lamacz:2013:EME

[LS13a] Agnes Lamacz and Ben Schweizer. Effective Maxwell equations in a geometry with flat rings of arbitrary shape. SIAM Journal on Mathematical Analysis, 45(3):1460– 1494, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LeCrone:2013:WPS

[LS13b] Jeremy LeCrone and Gieri Simonett. On well-posedness, stability, and bifurcation for the axisymmetric surface diffusion flow. SIAM Journal on Mathematical Analysis, 45(5): 2834–2869, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2013:PMA

[LS13c] Xin Yang Lu and Dejan Slepcev. Properties of minimizers of average-distance problem via discrete approximation of measures. SIAM Journal on Mathematical Analysis, 45(5):3114– 3131, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2015:SRT

[LS15] Junshan Lin and Fadil Santosa. Scattering resonances for a two-dimensional potential well with a thick barrier. SIAM Journal on Mathematical Analysis, 47(2):1458– 1488, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lamacz:2016:NIM

[LS16] A. Lamacz and B. Schweizer. A negative index metamaterial for Maxwell’s equations. SIAM Journal on Mathematical Analysis, 48(6): 4155–4174, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lopez:2017:STW

[LS17] José Luis López and Juan Soler. A space-time Wigner function approach to long time Schrödinger–Poisson dynamics. SIAM Journal on Mathematical Analysis, 49(6): 4915–4941, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leslie:2018:CIE

[LS18] Trevor M. Leslie and Roman Shvydkoy. Conditions implying energy equality for weak solutions of the Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 50(1): 870–890, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levermore:2012:LMN

[LST12] C. David Levermore, Weiran Sun, and Konstantina Trivisa. A low Mach number limit of a dispersive Navier–Stokes system. SIAM Journal on Mathematical Analysis, 44(3): 1760–1807, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2017:GAT

[LSW17] Yulong Lu, Andrew Stuart, and Hendrik Weber. Gaussian approximations for transition paths in Brownian dynamics. SIAM Journal on Mathematical Analysis, 49(4): 3005–3047, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2011:GUS

[LT11] Shitao Liu and Roberto Triggiani. Global uniqueness and stability in determining the damping coefficient of an inverse hyperbolic problem with NonHomogeneous Neumann B.C. through an additional Dirichlet boundary trace. SIAM Journal on Mathematical Analysis, 43(4): 1631–1666, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1631_s1.

Lattanzio:2013:RED

[LT13] Corrado Lattanzio and Athanasios E. Tzavaras. Relative entropy in diffusive relaxation. SIAM Journal on Mathematical Analysis, 45(3): 1563–1584, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2017:EUW

[LT17] Jinkai Li and Edriss S. Titi. Existence and uniqueness of weak solutions to viscous primitive equations for a certain class of discontinuous initial data. SIAM Journal on Mathematical Analysis, 49 (1):1–28, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2019:GEW

[LT19] Xin Liu and Edriss S. Titi.

Global existence of weak solutions to the compressible primitive equations of atmospheric dynamics with degenerate viscosities. SIAM Journal on Mathematical Analysis, 51(3):1913–1964, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lakshtanov:2017:UIC

[LTV17] Evgeny Lakshtanov, Jorge Tejero, and Boris Vainberg. Uniqueness in the inverse conductivity problem for complex-valued Lipschitz conductivities in the plane. SIAM Journal on Mathematical Analysis, 49(5):3766– 3775, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lou:2014:AIF

[LTW14] Yuan Lou, Youshan Tao, and Michael Winkler. Approaching the ideal free distribution in two-species competition models with fitness-dependent dispersal. SIAM Journal on Mathematical Analysis, 46(2): 1228–1262, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2013:EGW

[Lu13] Yun-Guang Lu. Existence of global weak entropy solutions to some nonstrictly hyperbolic systems. SIAM Journal on Mathematical Analysis, 45(6): 3592–3610, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lancia:2010:IHF

[LV10] Maria Rosaria Lancia and Paola Vernole. Irregular heat flow problems. SIAM Journal on Mathematical Analysis, 42(4):1539–1567, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lakshtanov:2012:EIT

[LV12] E. Lakshtanov and B. Vainberg. Ellipticity in the interior transmission problem in anisotropic media. SIAM Journal on Mathematical Analysis, 44(2):1165– 1174, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2013:SSR

[LV13] Xiaodong Li and Vladislav Voroninski. Sparse signal recovery from quadratic measurements via convex programming. SIAM Journal on Mathematical Analysis, 45(5): 3019–3033, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lakshtanov:2015:SWL

[LV15] E. Lakshtanov and B. Vainberg. Sharp Weyl law for signed counting function of positive interior transmission eigenvalues. SIAM Journal on Mathematical Analysis, 47(4): 3212–3234, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laurencot:2015:AGS

[LvR15] Philippe Laurençot and Henry van Roessel. Absence of gelation and self-similar behavior for a coagulationfragmentation equation. SIAM Journal on Mathematical Analysis, 47(3):2355–2374, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2012:STF

[LW12] Yi Li and Yaping Wu. Stability of traveling front solutions with algebraic spatial decay for some autocatalytic chemical reaction systems. SIAM Journal on Mathematical Analysis, 44(3): 1474–1521, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2014:TEW

[LW14a] Xiaobiao Lin and Martin Wechselberger. Transonic evaporation waves in a spherically symmetric nozzle. SIAM Journal on Mathematical Analysis, 46(2):1472– 1504, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2014:SBL

[LW14b] Cheng-Jie Liu and Ya-Guang Wang. Stability of boundary layers for the nonisentropic compressible circularly symmetric 2D flow. SIAM Journal on Mathematical Analysis, 46 (1):256–309, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lipshutz:2015:EUS

[LW15] David Lipshutz and Ruth J. Williams. Existence, uniqueness, and stability of slowly oscillating periodic solutions for delay differential equations with nonnegativity constraints. SIAM Journal on Mathematical Analysis, 47(6): 4467–4535, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lasiecka:2016:FSF

[LW16a] Irena Lasiecka and Justin T. Webster. Feedback stabilization of a fluttering panel in an inviscid subsonic potential flow. SIAM Journal on Mathematical Analysis, 48(3): 1848–1891, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liechty:2016:TLS

[LW16b] Karl Liechty and Dong Wang.

Two Lax systems for the Painlevé II equation, and two related kernels in random matrix theory. SIAM Journal on Mathematical Analysis, 48(5): 3618–3666, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2017:VSR

[LW17] Tai-Ping Liu and Haitao Wang. Viscous scalar rarefaction waves. SIAM Journal on Mathematical Analysis, 49(3): 2061–2100, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2011:LRG

[LWX11] Yongsheng Li, Yifei Wu, and Guixiang Xu. Low regularity global solutions for the focusing mass-critical NLS in R. SIAM Journal on Mathematical Analysis, 43(1): 322–340, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p322_s1.

Li:2016:GCS

[LWX16] Wei-Xi Li, Di Wu, and Chao-Jiang Xu. Gevrey class smoothing effect for the Prandtl equation. SIAM Journal on Mathematical Analysis, 48(3):1672–1726, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lichtner:2011:SDD

[LWY11] M. Lichtner, M. Wolfrum, and S. Yanchuk. The spectrum of delay differential equations with large delay. SIAM Journal on Mathematical Analysis, 43(2):788– 802, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p788_s1.

Li:2018:GMS

[LWY18] Jun Li, Ingo Witt, and Huicheng Yin. Global multidimensional shock waves of 2-dimensional and 3dimensional unsteady potential flow equations. SIAM Journal on Mathematical Analysis, 50(1):933–1009, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2011:ESU

[LWZ11] Peijun Li, Haijun Wu, and Weiying Zheng. Electromagnetic scattering by unbounded rough surfaces. SIAM Journal on Mathematical Analysis, 43(3):1205– 1231, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1205_s1.

Lafitte:2015:HFS

[LWZ15] Olivier Lafitte, Mark Williams, and Kevin Zumbrun. Highfrequency stability of detonations and turning points at infinity. SIAM Journal on Mathematical Analysis, 47(3): 1800–1878, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lafitte:2016:BDO

[LWZ16] Olivier Lafitte, Mark Williams, and Kevin Zumbrun. Blockdiagonalization of ODEs in the semiclassical limit and C^{ω} versus C^{∞} stationary phase. SIAM Journal on Mathematical Analysis, 48(3): 1773–1797, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lam:2018:TWC

[LWZ18] King-Yeung Lam, Xueying Wang, and Tianran Zhang. Traveling waves for a class of diffusive disease-transmission models with network structures. SIAM Journal on Mathematical Analysis, 50(6): 5719–5748, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2016:ETM

[LX16] Jian-Guo Liu and Xiangsheng Xu. Existence theorems for a multidimensional crystal surface model. SIAM Journal on Mathematical Analysis, 48(6): 3667–3687, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2017:OEP

[LX17a] Haigang Li and Longjuan Xu. Optimal estimates for the perfect conductivity problem with inclusions close to the boundary. SIAM Journal on Mathematical Analysis, 49(4): 3125–3142, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2017:ESP

[LX17b] Hongyu Liu and Jingni Xiao. On electromagnetic scattering from a penetrable corner. SIAM Journal on Mathematical Analysis, 49(6):5207– 5241, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2017:AVC

[LX17c] Jian-Guo Liu and Xiangsheng Xu. Analytical validation of a continuum model for the evolution of a crystal surface in multiple space dimensions. SIAM Journal on Mathematical Analysis, 49(3): 2220–2245, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2019:CPS

[LX19] Hao-Guang Li and Chao-Jiang Xu. Cauchy problem for the spatially homogeneous Landau equation with Shubin class initial datum and Gelfand–Shilov smoothing effect. SIAM Journal on Mathematical Analysis, 51(1):532– 564, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2019:JPA

[LXY19] Cheng-Jie Liu, Feng Xie, and Tong Yang. Justification of Prandtl ansatz for MHD boundary layer. SIAM Journal on Mathematical Analysis, 51(3):2748–2791, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2013:GCS

[LXZ13] Hai-Liang Li, Xinying Xu, and Jianwen Zhang. Global classical solutions to 3D compressible magnetohydrodynamic equations with large oscillations and vacuum. SIAM Journal on Mathematical Analysis, 45(3):1356– 1387, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2019:LEU

[LY19] Xin Liu and Yuan Yuan. Local existence and uniqueness of strong solutions to the free boundary problem of the full compressible Navier– Stokes equations in three dimensions. SIAM Journal on Mathematical Analysis, 51(2): 748–789, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2016:SSB

[LYZ16] Hai-Liang Li, Tong Yang, and Mingying Zhong. Spectrum structure and behaviors of the Vlasov–Maxwell–Boltzmann systems. SIAM Journal on Mathematical Analysis, 48(1): 595–669, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:2014:ODC

[LYZZ14] Hongxia Liu, Tong Yang, Huijiang Zhao, and Qingyang Zou. One-dimensional compressible Navier–Stokes equations with temperature dependent transport coefficients and large data. SIAM Journal on Mathematical Analysis, 46(3): 2185–2228, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2017:GRC

[LZ17] Mingjie Li and Qingtian Zhang. Generic regularity of conservative solutions to Camassa–Holm type equations. SIAM Journal on Mathematical Analysis, 49(4): 2920–2949, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:2018:REW

[LZ18] Yong Lu and Zhifei Zhang.

Relative entropy, weak-strong uniqueness, and conditional regularity for a compressible Oldroyd–B model. SIAM Journal on Mathematical Analysis, 50(1):557–590, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:2019:MAS

[LZ19] Junshan Lin and Hai Zhang.

Mathematical analysis of surface plasmon resonance by a nano-gap in the plasmonic metal. SIAM Journal on Mathematical Analysis, 51(6): 4448–4489, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:2015:GMV

[LZZ15] Jing Li, Jianwen Zhang, and Junning Zhao. On the global motion of viscous compressible barotropic flows subject to large external potential forces and vacuum.

SIAM Journal on Mathematical Analysis, 47(2):1121– 1153, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liang:2017:PED

[LZZ17] Xing Liang, Lei Zhang, and Xiao-Qiang Zhao. The principal eigenvalue for degenerate periodic reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 49(5): 3603–3636, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maeda:2017:EAS

[Mae17] Masaya Maeda. Existence and asymptotic stability of quasi-periodic solutions of discrete NLS with potential. SIAM Journal on Mathematical Analysis, 49(5):3396– 3426, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marzuola:2010:CSP

[Mar10]  Jeremy L. Marzuola.  A

class of stable perturbations for a minimal mass soliton in three-dimensional saturated nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 42(3): 1382–1403, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marinelli:2018:WPS

[Mar18] Carlo Marinelli. On wellposedness of semilinear stochastic evolution equations on L_{p} spaces. SIAM Journal on Mathematical Analysis, 50(2): 2111–2143, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masaki:2011:ESS

[Mas11]  Satoshi Masaki.  Energy

solution to a Schrödinger– Poisson system in the twodimensional whole space. SIAM Journal on Mathematical Analysis, 43(6):2719– 2731, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2719_s1.

Mohammed:2016:FDL

[MB16] Wael W. Mohammed and Dirk Blömker. Fast diffusion limit for reaction–diffusion systems with stochastic Neumann boundary conditions. SIAM Journal on Mathematical Analysis, 48(5):3547– 3578, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mohammed:2013:MES

[MBK13] Wael W. Mohammed, Dirk Blömker, and Konrad Klepel. Modulation equation for stochastic Swift–Hohenberg equation. SIAM Journal on Mathematical Analysis, 45(1): 14–30, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mirrahimi:2013:SHJ

[MBPS13] Sepideh Mirrahimi, Guy Barles, Benoît Perthame, and Panagiotis E. Souganidis. A singular Hamilton–Jacobi equation modeling the tail problem. SIAM Journal on Mathematical Analysis, 44(6): 4297–4319, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mora-Corral:2014:QEC

[MC14] Carlos Mora-Corral. Quasistatic evolution of cavities in nonlinear elasticity. SIAM Journal on Mathematical Analysis, 46(1):532– 571, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Medvedev:2014:NHE

[Med14] Georgi S. Medvedev. The nonlinear heat equation on dense graphs and graph limits. SIAM Journal on Mathematical Analysis, 46(4):2743– 2766, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mei:2010:BAP

[Mei10] Ming Mei. Best asymptotic profile for hyperbolic p-system with damping. SIAM Journal on Mathematical Analysis, 42(1):1–23, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Melcher:2010:TFL

[Mel10] Christof Melcher. Thin-film limits for Landau–Lifshitz– Gilbert equations. SIAM Journal on Mathematical Analysis, 42(1):519–537, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Melenk:2012:MPC

[Mel12] Jens Markus Melenk. Mapping properties of combined field Helmholtz boundary integral operators. SIAM Journal on Mathematical Analysis, 44(4):2599–2636, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Merono:2018:FAS

[Mer18] Cristóbal J. Meroño. Fixed angle scattering: Recovery of singularities and its limitations. SIAM Journal on Mathematical Analysis, 50(5): 5616–5636, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Milton:2018:AET

[Mil18] Graeme W. Milton. Approximating the effective tensor as a function of the component tensors in two-dimensional composites of two anisotropic phases. SIAM Journal on Mathematical Analysis, 50(3): 3327–3364, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mitchener:2010:MFM

[Mit10] W. Garrett Mitchener. Meanfield and measure-valued differential equation models for language variation and change in a spatially distributed population. SIAM Journal on Mathematical Analysis, 42(5): 1899–1933, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mizumachi:2011:SSF

[Miz11]  Tetsu Mizumachi.  N-

soliton states of the Fermi– Pasta–Ulam lattices. SIAM Journal on Mathematical Analysis, 43(5):2170–2210, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2170_s1.

Manjunath:2014:DRD

[MJ14] G. Manjunath and H. Jaeger. The dynamics of random difference equations is remodeled by closed relations. SIAM Journal on Mathematical Analysis, 46(1):459– 483, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mindt:2019:EPC

[MLD19] Pascal Mindt, Jens Lang, and Pia Domschke. Entropypreserving coupling of hierarchical gas models. SIAM Journal on Mathematical Analysis, 51(6):4754–4775, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mahadevan:2011:HSC

[MM11] Rajesh Mahadevan and T. Muthukumar. Homogenization of some cheap control problems. SIAM Journal on Mathematical Analysis, 43(5):2211– 2229, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2211_s1.

Ma:2017:SLL

[MM17] To Fu Ma and Rodrigo Nunes Monteiro. Singular limit and long-time dynamics of Bresse systems. SIAM Journal on Mathematical Analysis, 49(4): 2468–2495, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maekawa:2018:ISS

[MM18a] Yasunori Maekawa and Hideyuki Miura. On isomorphism for the space of solenoidal vector fields and its application to the incompressible flows. SIAM Journal on Mathematical Analysis, 50(1): 339–353, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masaki:2018:LRS

[MM18b] Satoshi Masaki and Hayato Miyazaki. Long range scattering for nonlinear Schrödinger equations with critical homogeneous nonlinearity. SIAM Journal on Mathematical Analysis, 50(3):3251–3270, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mailybaev:2011:RLT

[MMB11] A. A. Mailybaev, D. Marchesin, and J. Bruining. Resonance in low-temperature oxidation waves for porous media. SIAM Journal on Mathematical Analysis, 43(5): 2230–2252, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i5/p2230_s1.

Mitsoudis:2013:HEA

[MMP13] D. A. Mitsoudis, Ch. Makridakis, and M. Plexousakis. Helmholtz equation with artificial boundary conditions in a two-dimensional waveguide. SIAM Journal on Mathematical Analysis, 44(6): 4320–4344, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marchese:2019:MTP

[MMT19] Andrea Marchese, Annalisa Massaccesi, and Riccardo Tione. A multimaterial transport problem and its convex relaxation via rectifiable Gcurrents. SIAM Journal on Mathematical Analysis, 51(3): 1965–1998, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muratov:2012:GEC

[MN12] C. B. Muratov and M. Novaga. Global exponential convergence to variational traveling waves in cylinders. SIAM Journal on Mathematical Analysis, 44(1):293– 315, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p293_s1.

Mucci:2016:LGE

[MN16] Domenico Mucci and Lorenzo Nicolodi. On the Landau– de Gennes elastic energy of constrained biaxial nematics. SIAM Journal on Mathematical Analysis, 48(3):1954– 1987, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McKinley:2018:ADG

[MN18] Scott A. McKinley and Hung D. Nguyen. Anomalous diffusion and the generalized Langevin equation. SIAM Journal on Mathematical Analysis, 50(5):5119– 5160, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miyazaki:2011:DFC

[MNS11] Rinko Miyazaki, Toshiki Naito, and Jong Son Shin. Delayed feedback control by commutative gain matrices. SIAM Journal on Mathematical Analysis, 43(3):1122– 1144, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i3/p1122_s1.

Moradifam:2013:CIO

[MNT13] Amir Moradifam, Adrian Nachman, and Alexandru Tamasan. Conductivity imaging from one interior measurement in the presence of perfectly conducting and insulating inclusions. SIAM Journal on Mathematical Analysis, 44(6):3969–3990, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muller:2014:ACD

[MO14] Stefan Müller and Heiner Olbermann. Almost conical deformations of thin sheets with rotational symmetry. SIAM Journal on Mathematical Analysis, 46(1): 25–44, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Milisic:2015:SML

[MO15] Vuk Milisić and Dietmar Oelz.

On a structured model for load-dependent reaction kinetics of transient elastic linkages mediating nonlinear friction. SIAM Journal on Mathematical Analysis, 47(3): 2104–2121, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ma:2019:LNS

[MO19] Manjun Ma and Chunhua Ou.

Linear and nonlinear speed selection for mono-stable wave propagations. SIAM Journal on Mathematical Analysis, 51 (1):321–345, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moameni:2011:SUC

[Moa11] Abbas Moameni. Stability under Γ-convergence of least energy solutions for semilinear problems in the whole R^{N}. SIAM Journal on Mathematical Analysis, 43(4):1759– 1786, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1759_s1.

Monard:2016:IAG

[Mon16] François Monard. Inversion of the attenuated geodesic Xray transform over functions and vector fields on simple surfaces. SIAM Journal on Mathematical Analysis, 48(2): 1155–1177, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moon:2016:DFC

[Moo16] Sunghwan Moon. On the determination of a function from its conical Radon transform with a fixed central axis. SIAM Journal on Mathematical Analysis, 48(3): 1833–1847, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCormick:2016:LBB

[MOR^{+}16] David S. McCormick, Eric J. Olson, James C. Robinson, Jose L. Rodrigo, Alejandro Vidal-López, and Yi Zhou. Lower bounds on blowingup solutions of the threedimensional Navier–Stokes equations in Ḣ^{3/2}, Ḣ^{5/2}, and Ḃ_{2,1}^{5/2}. SIAM Journal on Mathematical Analysis, 48(3): 2119–2132, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mori:2019:VFA

[Mor19]  Ryunosuke Mori.  Valid-

ity of formal asymptotic expansions for singularly perturbed competition-diffusion systems. SIAM Journal on Mathematical Analysis, 51(2): 820–853, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mielke:2014:ANV

[MOS14] Alexander Mielke, Christoph Ortner, and Yasemin Sengül. An approach to nonlinear viscoelasticity via metric gradient flows. SIAM Journal on Mathematical Analysis, 46(2): 1317–1347, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mosco:2018:FTS

[Mos18]  Umberto Mosco.  Finite-

time self-organized-criticality on synchronized infinite grids. SIAM Journal on Mathematical Analysis, 50(3):2409– 2440, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mei:2010:GSM

[MOZ10] Ming Mei, Chunhua Ou, and Xiao-Qiang Zhao. Global stability of monostable traveling waves for nonlocal timedelayed reaction–diffusion equations. SIAM Journal on Mathematical Analysis, 42(6): 2762–2790, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [MOZ12].

Mei:2012:EGS

[MOZ12] Ming Mei, Chunhua Ou, and Xiao-Qiang Zhao. Erratum: “Global Stability of Monostable Traveling Waves For Nonlocal Timedelayed Reaction-diffusion Equations”. SIAM Journal on Mathematical Analysis, 44 (1):538–540, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p538_s1.

See [MOZ10].

Mikulevicius:2012:ESS

[MP12] R. Mikulevicius and H. Pragarauskas. On L_{p}-estimates of some singular integrals related to jump processes. SIAM Journal on Mathematical Analysis, 44(4):2305– 2328, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Melcher:2013:LLS

[MP13] Christof Melcher and Mariya Ptashnyk. Landau–Lifshitz– Slonczewski equations: Global weak and classical solutions. SIAM Journal on Mathematical Analysis, 45(1):407– 429, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muller:2014:DIR

[MP14] Stefan Müller and Mariapia Palombaro. On a differential inclusion related to the Born– Infeld equations. SIAM Journal on Mathematical Analysis, 46(4):2385–2403, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Murray:2016:ADE

[MP16] Ryan W. Murray and Robert L.

Pego. Algebraic decay to equilibrium for the Becker–Döring equations. SIAM Journal on Mathematical Analysis, 48(4): 2819–2842, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mallet-Paret:2014:ANS

[MPN14] John Mallet-Paret and Roger D.

Nussbaum. Analyticity and nonanalyticity of solutions of delay-differential equations.

SIAM Journal on Mathematical Analysis, 46(4):2468– 2500, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Manfredi:2010:AMV

[MPR10] Juan J. Manfredi, Mikko Parviainen, and Julio D. Rossi. An asymptotic mean value characterization for a class of nonlinear parabolic equations related to tug-of-war games. SIAM Journal on Mathematical Analysis, 42(5): 2058–2081, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mora:2017:CID

[MPS17] Maria Giovanna Mora, Mark A. Peletier, and Lucia Scardia. Convergence of interactiondriven evolutions of dislocations with Wasserstein dissipation and slip-plane confinement. SIAM Journal on Mathematical Analysis, 49(5): 4149–4205, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mazzone:2019:MFF

[MPS19] Giusy Mazzone, Jan Prüss, and Gieri Simonett. On the motion of a fluid-filled rigid body with Navier boundary conditions. SIAM Journal on Mathematical Analysis, 51(3): 1582–1606, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maddalena:2018:VPF

[MPT18] Francesco Maddalena, Danilo Percivale, and Franco Tomarelli. Variational problems for Föppl–von Kármán plates. SIAM Journal on Mathematical Analysis, 50(1):251– 282, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Menoukeu-Pamen:2019:SSS

[MPT19] Olivier Menoukeu-Pamen and Ludovic Tangpi. Strong solutions of some one-dimensional SDEs with random and unbounded drifts. SIAM Journal on Mathematical Analysis, 51(5):4105–4141, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mucha:2015:HCC

[MPZ15] P. B. Mucha, M. Pokorný, and E. Zatorska. Heatconducting, compressible mixtures with multicomponent diffusion: Construction of a weak solution. SIAM Journal on Mathematical Analysis, 47(5):3747–3797, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marinelli:2012:EWS

[MQS12] Carlo Marinelli and Lluís Quer-Sardanyons. Existence of weak solutions for a class of semilinear stochastic wave equations. SIAM Journal on Mathematical Analysis, 44(2): 906–925, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Munnier:2015:AAN

[MR15] Alexandre Munnier and Karim Ramdani. Asymptotic analysis of a Neumann problem in a domain with cusp. application to the collision problem of rigid bodies in a perfect fluid. SIAM Journal on Mathematical Analysis, 47(6): 4360–4403, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercuri:2016:RRL

[MRS16] Carlo Mercuri, Giuseppe Riey, and Berardino Sciunzi. A regularity result for the p-Laplacian near uniform ellipticity. SIAM Journal on Mathematical Analysis, 48(3): 2059–2075, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mazon:2014:OMP

[MRT14] José M. Mazón, Julio D. Rossi, and Julián Toledo. An optimal matching problem for the Euclidean distance. SIAM Journal on Mathematical Analysis, 46(1): 233–255, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ming:2015:MSR

[MRT15] Mei Ming, Frederic Rousset, and Nikolay Tzvetkov. Multi-solitons and related solutions for the water-waves system. SIAM Journal on Mathematical Analysis, 47(1): 897–954, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morassi:2012:SDR

[MRV12] Antonino Morassi, Edi Rosset, and Sergio Vessella. Stable determination of a rigid inclusion in an anisotropic elastic plate. SIAM Journal on Mathematical Analysis, 44(3): 2204–2235, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morassi:2019:OSI

[MRV19] Antonino Morassi, Edi Rosset, and Sergio Vessella. Optimal stability in the identification of a rigid inclusion in an isotropic Kirchhoff–Love plate. SIAM Journal on Mathematical Analysis, 51(2): 731–747, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCann:2011:HCO

[MS11] R. J. McCann and M. Sosio. Hölder continuity for optimal multivalued mappings. SIAM Journal on Mathematical Analysis, 43(4):1855– 1871, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1855_s1.

Mascia:2013:MNP

[MS13a] Corrado Mascia and Marta Strani. Metastability for nonlinear parabolic equations with application to scalar viscous conservation laws. SIAM Journal on Mathematical Analysis, 45(5):3084– 3113, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mugnai:2013:CRA

[MS13b] Luca Mugnai and Christian Seis. On the coarsening rates for attachment-limited kinetics. SIAM Journal on Mathematical Analysis, 45(1): 324–344, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morini:2014:CMN

[MS14] Massimiliano Morini and Peter Sternberg. Cascade of minimizers for a nonlocal isoperimetric problem in thin domains. SIAM Journal on Mathematical Analysis, 46(3): 2033–2051, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Montaru:2016:SSR

[MS16] Alexandre Montaru and Boyan Sirakov. Stationary states of reaction–diffusion and Schrödinger systems with inhomogeneous or controlled diffusion. SIAM Journal on Mathematical Analysis, 48(4): 2561–2587, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masaki:2018:RSE

[MS18a] Satoshi Masaki and Jun-Ichi Segata. Refinement of Strichartz estimates for Airy equation in nondiagonal case and its application. SIAM Journal on Mathematical Analysis, 50(3):2839– 2866, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meszaros:2018:VFS

[MS18b] Alpár Richárd Mészáros and Francisco J. Silva. On the variational formulation of some stationary secondorder mean field games systems. SIAM Journal on Mathematical Analysis, 50(1):

1255–1277, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mitake:2016:LAW

[MSTY16] H. Mitake, A. Siconolfi, H. V. Tran, and N. Yamada. A Lagrangian approach to weakly coupled Hamilton–Jacobi systems. SIAM Journal on Mathematical Analysis, 48(2): 821–846, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ming:2013:LTE

[MSZ13] Mei Ming, Jean Claude Saut, and Ping Zhang. Long-time existence of solutions to Boussinesq systems. SIAM Journal on Mathematical Analysis, 44(6): 4078–4100, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Migorski:2019:WPH

[MSZ19] Stanisław Migórski, Mircea Sofonea, and Shengda Zeng. Well-posedness of historydependent sweeping processes. SIAM Journal on Mathematical Analysis, 51(2): 1082–1107, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Matthies:2013:SAJ

[MT13] Karsten Matthies and Florian Theil. A semigroup approach to the justification of kinetic theory. SIAM Journal on Mathematical Analysis, 44(6): 4345–4379, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mizuno:2015:CAC

[MT15] Masashi Mizuno and Yoshihiro Tonegawa. Convergence of the Allen–Cahn equation with Neumann boundary conditions. SIAM Journal on Mathematical Analysis, 47(3): 1906–1932, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [MT16].

Mizuno:2016:ET

[MT16] Masashi Mizuno and Yoshihiro Tonegawa. Erratum to “Convergence of the Allen– Cahn Equation with Neumann Boundary Conditions”. SIAM Journal on Mathematical Analysis, 48(4):3035– 3036, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [MT15].

Matthies:2019:ROS

[MT19] Karsten Matthies and Florian Theil. Rescaled objective solutions of Fokker– Planck and Boltzmann equations. SIAM Journal on Mathematical Analysis, 51(2): 1321–1348, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Munoz:2012:DSL

[Muñ12] Claudio Muñoz. Dynamics of soliton-like solutions for slowly varying, generalized KdV equations: Refraction versus reflection. SIAM Journal on Mathematical Analysis, 44(1):1–60, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p1_s1.

Murphy:2014:INC

[Mur14] Jason Murphy. Intercritical NLS: Critical Ḣ^{s}-bounds imply scattering. SIAM Journal on Mathematical Analysis, 46 (1):939–997, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Molchanov:2019:PDM

[MV19] Stanislav Molchanov and Boris Vainberg. Population dynamics with moderate tails of the underlying random walk. SIAM Journal on Mathematical Analysis, 51(3): 1824–1835, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Metcalfe:2017:SCA

[MW17] Jason Metcalfe and Chengbo Wang. The Strauss conjecture on asymptotically flat spacetimes. SIAM Journal on Mathematical Analysis, 49(6): 4579–4594, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miao:2019:RIC

[MX19] Changxing Miao and Liutang Xue. On the regularity issues of a class of drift-diffusion equations with nonlocal diffusion. SIAM Journal on Mathematical Analysis, 51(4): 2927–2970, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Matsumura:2012:ABS

[MY12] Akitaka Matsumura and Natsumi Yoshida. Asymptotic behavior of solutions to the Cauchy problem for the scalar viscous conservation law with partially linearly degenerate flux. SIAM Journal on Mathematical Analysis, 44(4):2526– 2544, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miroshnikov:2017:WSI

[MY17] Alexey Miroshnikov and Robin Young. Weak* solutions II: The vacuum in Lagrangian gas dynamics.

SIAM Journal on Mathematical Analysis, 49(3):1810– 1843, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mohammed:2013:SBE

[MZ13] Salah-Eldin A. Mohammed and Tusheng Zhang. Stochastic Burgers equation with random initial velocities: a Malliavin calculus approach. SIAM Journal on Mathematical Analysis, 45(4):2396– 2420, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Michel:2018:SAK

[MZ18] Laurent Michel and Maciej Zworski. A semiclassical approach to the Kramers–Smoluchowski equation. SIAM Journal on Mathematical Analysis, 50(5): 5362–5379, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ming:2012:LWA

[MZZ12] Mei Ming, Ping Zhang, and Zhifei Zhang. Long-wave approximation to the 3-D capillary-gravity waves. SIAM Journal on Mathematical Analysis, 44(4):2920– 2948, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nadin:2010:ESR

[Nad10] Grégoire Nadin. The effect of the Schwarz rearrangement on the periodic principal eigenvalue of a nonsymmetric operator. SIAM Journal on Mathematical Analysis, 41(6): 2388–2406, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nazarov:2012:CZH

[Naz12] A. I. Nazarov. A centennial of the Zaremba–Hopf–Oleinik Lemma. SIAM Journal on Mathematical Analysis, 44(1): 437–453, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p437_s1.

Nesensohn:2014:GVF

[Nes14] Manuel Nesensohn. Generalized viscoelastic fluids with a free boundary without surface tension. SIAM Journal on Mathematical Analysis, 46 (1):428–458, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neumayer:2016:SFQ

[Neu16] Robin Neumayer. A strong form of the quantitative Wulff inequality. SIAM Journal on Mathematical Analysis, 48(3): 1727–1772, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2010:ASN

[Ngu10] Toan Nguyen. On asymptotic stability of noncharacteristic viscous boundary layers. SIAM Journal on Mathematical Analysis, 42(3):1156– 1178, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2013:RSA

[Ngu13] Hoai-Minh Nguyen. On a regularized scheme for approximate acoustic cloaking using transformation optics. SIAM Journal on Mathematical Analysis, 45(5):3034– 3049, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2015:ALD

[Ngu15] Linh V. Nguyen. On artifacts in limited data spherical Radon transform: Flat observation surfaces. SIAM Journal on Mathematical Analysis, 47(4):2984–3004, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2016:PLP

[Ngu16]  Quang-Huy Nguyen.  A

pseudo-local property of gravity water waves system. SIAM Journal on Mathematical Analysis, 48(3):1988– 2027, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2017:CAO

[Ngu17] Hoai-Minh Nguyen. Cloaking an arbitrary object via anomalous localized resonance: The cloak is independent of the object. SIAM Journal on Mathematical Analysis, 49(4):3208– 3232, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Niikuni:2012:DSG

[Nii12] Hiroaki Niikuni. On the degenerate spectral gaps of the one-dimensional Schrödinger operators with periodic point interactions. SIAM Journal on Mathematical Analysis, 44 (4):2847–2870, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2019:AQR

[NKV19] Huy Tuan Nguyen, Vo Anh Khoa, and Van Au Vo. Analysis of a quasi-reversibility method for a terminal value quasi-linear parabolic problem with measurements. SIAM Journal on Mathematical Analysis, 51(1):60– 85, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nesenenko:2012:WPD

[NN12] Sergiy Nesenenko and Patrizio Neff. Well-posedness for dislocation based gradient viscoplasticity I: Subdifferential case. SIAM Journal on Mathematical Analysis, 44(3): 1694–1712, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2019:ILN

[NN19] Toan T. Nguyen and Trinh T. Nguyen. The inviscid limit of Navier–Stokes equations for vortex-wave data on R^{2}. SIAM Journal on Mathematical Analysis, 51(3):2575– 2598, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naidenov:2018:LCV

[NNS18] Nikola Naidenov, Geno Nikolov, and Alexei Shadrin. On the largest critical value of T_{n}^{(}k). SIAM Journal on Mathematical Analysis, 50(3): 2389–2408, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nolen:2011:IPR

[Nol11] James Nolen. An invariance principle for random traveling waves in one dimension. SIAM Journal on Mathematical Analysis, 43(1): 153–188, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p153_s1.

Nishibata:2012:ASB

[NOS12] Shinya Nishibata, Masashi Ohnawa, and Masahiro Suzuki. Asymptotic stability of boundary layers to the Euler– Poisson equations arising in plasma physics. SIAM Journal on Mathematical Analysis, 44(2):761–790, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novack:2018:DRL

[Nov18]  Michael R. Novack.  Di-

mension reduction for the Landau–de Gennes model: The vanishing nematic correlation length limit. SIAM Journal on Mathematical Analysis, 50(6):6007–6048, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novack:2019:WST

[Nov19] Matthew D. Novack. On the weak solutions to the threedimensional inviscid quasigeostrophic system. SIAM Journal on Mathematical Analysis, 51(3):2686–2712, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novotny:2011:WVS

[NP11] Antonín Novotný and Milan Pokorný. Weak and variational solutions to steady equations for compressible heat conducting fluids. SIAM Journal on Mathematical Analysis, 43(3):1158–1188, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1158_s1.

Nguyen:2016:WYM

[NP16] Hong Thai Nguyen and Dariusz Paczka. Weak and Young measure solutions for hyperbolic initial boundary value problems of elastodynamics in the Orlicz–Sobolev space setting. SIAM Journal on Mathematical Analysis, 48(2): 1297–1331, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nahmod:2013:ASE

[NPS13] Andrea R. Nahmod, Natasa Pavlović, and Gigliola Staffilani. Almost sure existence of global weak solutions for supercritical Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 45(6):3431–3452, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neves:2018:STC

[NPS18] Wladimir Neves, Evgeniy Panov, and Jean Silva. Strong traces for conservation laws with general nonautonomous flux. SIAM Journal on Mathematical Analysis, 50(6):6049– 6081, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nijholt:2017:CMC

[NRS17] Eddie Nijholt, Bob Rink, and Jan Sanders. Center manifolds of coupled cell networks. SIAM Journal on Mathematical Analysis, 49(5): 4117–4148, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nakanishi:2012:IMA

[NS12] K. Nakanishi and W. Schlag.

Invariant manifolds around soliton manifolds for the nonlinear Klein–Gordon equation. SIAM Journal on Mathematical Analysis, 44(2): 1175–1210, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2013:SAC

[NS13] Toan T. Nguyen and Walter A. Strauss. Stability analysis of collisionless plasmas with specularly reflecting boundary. SIAM Journal on Mathematical Analysis, 45(2):

777–808, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Neukamm:2017:SHN

[NSS17] Stefan Neukamm, Mathias Schäffner, and Anja Schlömerkemper. Stochastic homogenization of nonconvex discrete energies with degenerate growth. SIAM Journal on Mathematical Analysis, 49(3):1761–1809, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nazarov:2013:LEE

[NT13] Sergey A. Nazarov and Jari Taskinen. Localization estimates for eigenfrequencies of waves trapped by a freely floating body in a channel. SIAM Journal on Mathematical Analysis, 45(4):2523– 2545, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Negreanu:2014:TSC

[NT14] Mihaela Negreanu and J. Ignacio Tello. On a two species chemotaxis model with slow chemical diffusion. SIAM Journal on Mathematical Analysis, 46(6):3761– 3781, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2018:KEZ

[NT18] Toan T. Nguyen and Minh-Binh Tran. On the kinetic equation in Zakharov’s wave turbulence theory for capillary waves. SIAM Journal on Mathematical Analysis, 50(2): 2020–2047, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguyen:2019:ACT

[NT19] Hoai-Minh Nguyen and Loc X.

Tran. Approximate cloaking for time-dependent Maxwell equations via transformation optics. SIAM Journal on Mathematical Analysis, 51(5): 4142–4171, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nochetto:2019:DFG

[NTW19] Ricardo H. Nochetto, Konstantina Trivisa, and Franziska Weber. On the dynamics of ferrofluids: Global weak solutions to the rosensweig system and rigorous convergence to equilibrium. SIAM Journal on Mathematical Analysis, 51(6): 4245–4286, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nagayasu:2011:RPO

[NUW11] Sei Nagayasu, Gunther Uhlmann, and Jenn-Nan Wang. Reconstruction of penetrable obstacles in acoustic scattering. SIAM Journal on Mathematical Analysis, 43(1): 189–211, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p189_s1.

Nguyen:2012:ACF

[NV12] Hoai-Minh Nguyen and Michael S. Vogelius. Approximate cloaking for the full wave equation via change of variables. SIAM Journal on Mathematical Analysis, 44(3):1894– 1924, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ninomiya:2017:TCW

[NW17] Hirokazu Ninomiya and Chang-Hong Wu. Traveling curved waves in two-dimensional excitable media. SIAM Journal on Mathematical Analysis, 49 (2):777–817, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Oh:2010:IGM

[Oh10]  Tadahiro Oh.  Invariance

of the Gibbs measure for the Schrödinger–Benjamin– Ono system. SIAM Journal on Mathematical Analysis, 41(6):2207–2225, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Oh:2015:NSE

[Oh15]  Tadahiro Oh.  On non-

linear Schrödinger equations with almost periodic initial data. SIAM Journal on Mathematical Analysis, 47(2): 1253–1270, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohnawa:2014:SCS

[Ohn14]  Masashi Ohnawa.  L^{∞}-

stability of continuous shock waves in a radiating gas model. SIAM Journal on Mathematical Analysis, 46(3): 2136–2159, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohnawa:2015:ASP

[Ohn15] Masashi Ohnawa. Asymptotic stability of plasma boundary layers to the Euler–Poisson equations with fluid-boundary interaction. SIAM Journal on Mathematical Analysis, 47(4): 2795–2831, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohnawa:2016:SDT

[Ohn16]  Masashi Ohnawa.  L^{∞}-

stability of discontinuous traveling waves in a hyperbolicelliptic coupled system. SIAM Journal on Mathematical Analysis, 48(6):3820–3839,

????  2016.  CODEN

SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olbermann:2019:LWF

[Olb19] Heiner Olbermann. On a Γlimit of Willmore functionals with additional curvature penalization term. SIAM Journal on Mathematical Analysis, 51(3):2599–2632, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Onodera:2011:SIH

[Ono11] Michiaki Onodera. Stability of the interface of a Hele– Shaw flow with two injection points. SIAM Journal on Mathematical Analysis, 43(4): 1810–1834, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1810_s1.

Osting:2017:CDP

[OR17] Braxton Osting and Todd Harry Reeb. Consistency of Dirichlet partitions. SIAM Journal on Mathematical Analysis, 49(5): 4251–4274, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ozanski:2019:PRS

[OR19] Wojciech S. Ozański and James C. Robinson. Partial regularity for a surface growth model. SIAM Journal on Mathematical Analysis, 51 (1):228–255, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Otto:2019:OTR

[OSW19] Felix Otto, Sebastian Scholtes, and Maria G. Westdickenberg. Optimal L^{1}-type relaxation rates for the Cahn– Hilliard equation on the line. SIAM Journal on Mathematical Analysis, 51(6):4645– 4682, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Otway:2010:USB

[Otw10] Thomas H. Otway. Unique solutions to boundary value problems in the cold plasma model. SIAM Journal on Mathematical Analysis, 42(6): 3045–3053, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p3045_s1.

Ovcharov:2011:SEK

[Ovc11] Evgeni Y. Ovcharov. Strichartz estimates for the kinetic transport equation. SIAM Journal on Mathematical Analysis, 43(3):1282–1310, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1282_s1.

Otto:2014:REO

[OW14] Felix Otto and Maria G. Westdickenberg. Relaxation to equilibrium in the onedimensional Cahn–Hilliard equation. SIAM Journal on Mathematical Analysis, 46(1): 720–756, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pallard:2014:SMV

[Pal14] Christophe Pallard. Space moments of the Vlasov– Poisson system: Propagation and regularity. SIAM Journal on Mathematical Analysis, 46(3):1754–1770, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Panov:2012:WCS

[Pan12a] Evgeniy Panov. On weak completeness of the set of entropy solutions to a degenerate nonlinear parabolic equation. SIAM Journal on Mathematical Analysis, 44(1): 513–535, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p513_s1.

Pantea:2012:PGS

[Pan12b] Casian Pantea. On the persistence and global stability of mass-action systems. SIAM Journal on Mathematical Analysis, 44(3):1636– 1673, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paoli:2015:VIP

[Pao15] Laetitia Paoli. Vibro-impact problems with dry friction — Part I: Existence result. SIAM Journal on Mathematical Analysis, 47(5):3285– 3313, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paoli:2016:VIP

[Pao16] Laetitia Paoli. Vibro-impact problems with dry friction — Part II: Tangential contacts and frictional catastrophes. SIAM Journal on Mathematical Analysis, 48(2): 1272–1296, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pass:2011:UMS

[Pas11]  Brendan Pass.  Unique-

ness and Monge solutions in the multimarginal optimal transportation problem. SIAM Journal on Mathematical Analysis, 43(6):2758– 2775, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i6/p2758_s1.

Pass:2013:COT

[Pas13] Brendan Pass. On a class of optimal transportation problems with infinitely many marginals. SIAM Journal on Mathematical Analysis, 45(4): 2557–2575, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinto:2017:CFR

[PD17] M. Campos Pinto and B. Després. Constructive formulations of resonant Maxwell’s equations. SIAM Journal on Mathematical Analysis, 49(5): 3637–3670, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pedregal:2015:WLN

[Ped15] Pablo Pedregal. Weak limits in nonlinear conductivity. SIAM Journal on Mathematical Analysis, 47(2):1154– 1168, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2015:UGS

[Pen15] Yue-Jun Peng. Uniformly global smooth solutions and convergence of Euler–Poisson systems with small parameters. SIAM Journal on Mathematical Analysis, 47(2):1355– 1376, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2017:CLD

[Pen17] Guanying Peng. Convergence of the Lawrence–Doniach energy for layered superconductors with magnetic fields near H_{c_{1}}. SIAM Journal on Mathematical Analysis, 49(2):1225– 1266, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perepelitsa:2010:ATR

[Per10] Misha Perepelitsa. Asymptotics toward rarefaction waves and vacuum for 1-D compressible Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 42(3): 1404–1412, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perepelitsa:2015:KFE

[Per15] Misha Perepelitsa. On a kinetic formulation of the Euler equations. SIAM Journal on Mathematical Analysis, 47(3): 2074–2083, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peszek:2015:DCS

[Pes15] Jan Peszek. Discrete Cucker– Smale flocking model with a weakly singular weight.

SIAM Journal on Mathematical Analysis, 47(5):3671– 3686, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pimentel:2016:USG

[Pim16] Juliana F. S. Pimentel. Unbounded Sturm global attractors for semilinear parabolic equations on the circle. SIAM Journal on Mathematical Analysis, 48(6):3860– 3882, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinezich:2019:PSN

[Pin19] John D. Pinezich. Propagation of singularities in nonconvex Hamilton–Jacobi problems: Local structure in two dimensions. SIAM Journal on Mathematical Analysis, 51(5): 3796–3818, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Plakhov:2014:PMR

[Pla14]  Alexander Plakhov.  The

problem of minimal resistance for functions and domains. SIAM Journal on Mathematical Analysis, 46(4): 2730–2742, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perez-Llanos:2018:OFM

[PLPSS18] Mayte Pérez-Llanos, Juan P.

Pinasco, Nicolas Saintier, and Analía Silva. Opinion formation models with heterogeneous persuasion and zealotry. SIAM Journal on Mathematical Analysis, 50(5):4812– 4837, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pohjola:2015:URI

[Poh15] Valter Pohjola. A uniqueness result for an inverse problem of the steady state convection-diffusion equation. SIAM Journal on Mathematical Analysis, 47(3):2084– 2103, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Polacik:2017:PPT

[Pol17] P. Polácik. Planar propagating terraces and the asymptotic one-dimensional symmetry of solutions of semilinear parabolic equations. SIAM Journal on Mathematical Analysis, 49(5):3716– 3740, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Poon:2019:MSS

[PP19] Clarice Poon and Gabriel Peyré. Multidimensional sparse super-resolution. SIAM Journal on Mathematical Analysis, 51(1):1–44, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piat:2013:LES

[PPP13] V. Chiadò Piat, I. Pankratova, and A. Piatnitski. Localization effect for a spectral problem in a perforated domain with Fourier boundary conditions. SIAM Journal on Mathematical Analysis, 45(3): 1302–1327, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pagani:2016:CAG

[PPPV16] Carlo D. Pagani, Dario Pierotti, Angela Pistoia, and Giusi Vaira. Concentration along geodesics for a nonlinear Steklov problem arising in corrosion modeling. SIAM Journal on Mathematical Analysis, 48(2):1085– 1108, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Popoff:2013:WML

[PR13] Nicolas Popoff and Nicolas Raymond. When the 3D magnetic Laplacian meets a curved edge in the semiclassical limit. SIAM Journal on Mathematical Analysis, 45(4): 2354–2395, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Prange:2013:AAB

[Pra13] Christophe Prange. Asymptotic analysis of boundary layer correctors in periodic homogenization. SIAM Journal on Mathematical Analysis, 45 (1):345–387, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piccoli:2015:CFK

[PRT15] Benedetto Piccoli, Francesco Rossi, and Emmanuel Trélat. Control to flocking of the kinetic Cucker–Smale model. SIAM Journal on Mathematical Analysis, 47(6):4685– 4719, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinski:2015:KLA

[PSSW15] F. J. Pinski, G. Simpson, A. M. Stuart, and H. Weber. Kullback–Leibler approximation for probability measures on infinite dimensional spaces. SIAM Journal on Mathematical Analysis, 47(6): 4091–4122, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peletier:2010:DRC

[PSV10] Mark A. Peletier, Giuseppe Savaré, and Marco Veneroni. From diffusion to reaction via Γ-Convergence. SIAM Journal on Mathematical Analysis, 42(4):1805–1825, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pellet:2019:HHC

[PSZ19a] Xavier Pellet, Lucia Scardia, and Caterina Ida Zeppieri. Homogenization of highcontrast Mumford–Shah energies. SIAM Journal on Mathematical Analysis, 51(3): 1696–1729, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piasecki:2019:SDC

[PSZ19b] Tomasz Piasecki, Yoshihiro Shibata, and Ewelina Zatorska. On strong dynamics of compressible two-component mixture flow. SIAM Journal on Mathematical Analysis, 51(4):2793–2849, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Privault:2011:DEF

[PT11] Nicolas Privault and Giovanni Luca Torrisi. Density estimation of functionals of spatial point processes with application to wireless networks. SIAM Journal on Mathematical Analysis, 43(3): 1311–1344, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

siam.org/sima/resource/
1/sjmaah/v43/i3/p1311_s1.

Piccoli:2018:GBE

[PT18] Benedetto Piccoli and Magali Tournus. A general BV existence result for conservation laws with spatial heterogeneities. SIAM Journal on Mathematical Analysis, 50(3): 2901–2927, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pu:2013:DLE

[Pu13] Xueke Pu. Dispersive limit of the Euler–Poisson system in higher dimensions. SIAM Journal on Mathematical Analysis, 45(2):834– 878, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Palacios:2018:QAM

[PUW18] Benjamin Palacios, Gunther Uhlmann, and Yiran Wang. Quantitative analysis of metal artifacts in X-ray tomography. SIAM Journal on Mathematical Analysis, 50(5):4914– 4936, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Plotnikov:2015:INS

[PW15] P. I. Plotnikov and W. Weigant.

Isothermal Navier–Stokes equations and Radon transform.

SIAM Journal on Mathematical Analysis, 47(1):626– 653, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perrin:2018:ODG

[PW18] C. Perrin and M. Westdickenberg. One-dimensional granular system with memory effects. SIAM Journal on Mathematical Analysis, 50(6): 5921–5946, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2011:RLG

[PWG11] Yue-Jun Peng, Shu Wang, and Qilong Gu. Relaxation limit and global existence of smooth solutions of compressible Euler–Maxwell equations. SIAM Journal on Mathematical Analysis, 43(2): 944–970, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p944_s1.

Pu:2017:LLE

[PWW17] Xueke Pu, Meng Wang, and Wendong Wang. The Landau–Lifshitz equation of the ferromagnetic spin chain and Oseen–Frank flow. SIAM Journal on Mathematical Analysis, 49(6):5134–5157, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Procesi:2013:QTF

[PX13] Michela Procesi and Xindong Xu. Quasi-Töplitz functions in KAM Theorem. SIAM Journal on Mathematical Analysis, 45(4):2148– 2181, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pedregal:2010:DMM

[PY10] Pablo Pedregal and Baisheng Yan. A duality method for micromagnetics. SIAM Journal on Mathematical Analysis, 41(6):2431–2452, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Polacik:2014:LSS

[PY14] Peter Polácik and Eiji Yanagida.

Localized solutions of a semilinear parabolic equation with a recurrent nonstationary asymptotics. SIAM Journal on Mathematical Analysis, 46 (5):3481–3496, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paicu:2011:GER

[PZ11] Marius Paicu and Arghir Zarnescu. Global existence and regularity for the full coupled Navier– Stokes and Q-tensor system.

SIAM Journal on Mathematical Analysis, 43(5):2009– 2049, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i5/p2009_s1.

Pawlow:2013:GRS

[PZ13] Irena Pawlow and Wojciech M. Zajaczkowski. Global regular solutions to a Kelvin– Voigt type thermoviscoelastic system. SIAM Journal on Mathematical Analysis, 45(4): 1997–2045, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piatnitski:2017:PHN

[PZ17] A. Piatnitski and E. Zhizhina. Periodic homogenization of nonlocal operators with a convolution-type kernel. SIAM Journal on Mathematical Analysis, 49(1):64–81, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:2019:ABP

[PZZ19] Rui Peng, Guanghui Zhang, and Maolin Zhou. Asymptotic behavior of the principal eigenvalue of a linear second order elliptic operator with small/large diffusion coefficient. SIAM Journal on Mathematical Analysis, 51(6): 4724–4753, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qin:2015:WPD

[Qin15] Wen-Xin Qin. Wave propagation in diatomic lattices. SIAM Journal on Mathematical Analysis, 47(1):477– 497, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Quittner:2012:SCS

[QS12] Pavol Quittner and Philippe Souplet. Symmetry of components for semilinear elliptic systems. SIAM Journal on Mathematical Analysis, 44(4): 2545–2559, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qin:2011:LTB

[QW11] Xiaohong Qin and Yi Wang.

Large-time behavior of solutions to the inflow problem of full compressible Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 43(1):341–366, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sima/resource/
1/sjmaah/v43/i1/p341_s1.

Qiao:2019:VTP

[QWE19] Yangyang Qiao, Huanyao Wen, and Steinar Evje. Viscous two-phase flow in porous media driven by source terms: Analysis and numerics. SIAM Journal on Mathematical Analysis, 51(6):5103– 5140, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rein:2018:ABS

[Rei18] Gerhard Rein. The asymptotic behavior of solutions to the repulsive n-body problem. SIAM Journal on Mathematical Analysis, 50(1):1–4, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rey:2012:BAA

[Rey12] Thomas Rey. Blow up analysis for anomalous granular gases. SIAM Journal on Mathematical Analysis, 44(3): 1544–1561, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Riaza:2010:SLQ

[Ria10] Ricardo Riaza. Stability loss in quasilinear DAEs by divergence of a pencil eigenvalue. SIAM Journal on Mathematical Analysis, 41(6): 2226–2245, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rodiac:2016:RPS

[Rod16] Rémy Rodiac. Regularity properties of stationary harmonic functions whose Laplacian is a Radon measure. SIAM Journal on Mathematical Analysis, 48(4):2495– 2531, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roubicek:2010:TRI

[Rou10] Tomáš Roubíček. Thermodynamics of rate-independent processes in viscous solids at small strains. SIAM Journal on Mathematical Analysis, 42 (1):256–297, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roubicek:2013:ACV

[Rou13]  Tomás Roubícek.  Adhe-

sive contact of visco-elastic bodies and defect measures arising by vanishing viscosity. SIAM Journal on Mathematical Analysis, 45(1):101– 126, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Riva:2018:VTM

[RP18] Matteo Dalla Riva and Luigi Provenzano. On vibrating thin membranes with mass concentrated near the boundary: an asymptotic analysis. SIAM Journal on Mathematical Analysis, 50(3):2928– 2967, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ramming:2013:SSE

[RR13] Tobias Ramming and Gerhard Rein. Spherically symmetric equilibria for self-gravitating kinetic or fluid models in the nonrelativistic and relativistic case — a simple proof for finite extension. SIAM Journal on Mathematical Analysis, 45 (2):900–914, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rocca:2015:E

[RR15] Elisabetta Rocca and Riccarda Rossi. “Entropic” solutions to a thermodynamically consistent PDE system for phase transitions and damage. SIAM Journal on Mathematical Analysis, 47(4):2519– 2586, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rivera:2017:TPT

[RR17] Jaime E. Muñoz Rivera and Reinhard Racke. Transmission problems in (thermo)Viscoelasticity with Kelvin–Voigt damping: Nonexponential, strong, and polynomial stability. SIAM Journal on Mathematical Analysis, 49(5):3741– 3765, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roth:2017:RPE

[RSS17] Gregory Roth, Paul L. Salceanu, and Sebastian J. Schreiber. Robust permanence for ecological maps. SIAM Journal on Mathematical Analysis, 49(5):3527– 3549, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ran:2018:NBV

[RSZ18] Yu Ran, Shu-Ming Sun, and Bing-Yu Zhang. Nonhomogeneous boundary value problems of nonlinear Schrödinger equations in a half plane. SIAM Journal on Mathematical Analysis, 50(3):2773– 2806, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rossi:2017:CRI

[RT17] Riccarda Rossi and Marita Thomas. Coupling rateindependent and rate-dependent processes: Existence results. SIAM Journal on Mathematical Analysis, 49(2):1419– 1494, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Remond-Tiedrez:2019:VSW

[RTT19] Antoine Remond-Tiedrez and Ian Tice. The viscous surface wave problem with generalized surface energies.

SIAM Journal on Mathematical Analysis, 51(6):4894– 4952, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rossi:2017:EFK

[RTV17] Luca Rossi, Andrea Tellini, and Enrico Valdinoci. The effect on Fisher–KPP propagation in a cylinder with fast diffusion on the boundary. SIAM Journal on Mathematical Analysis, 49(6):4595– 4624, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Radu:2016:GDP

[RTY16] Petronela Radu, Grozdena Todorova, and Borislav Yordanov. The generalized diffusion phenomenon and applications. SIAM Journal on Mathematical Analysis, 48(1): 174–203, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2017:CVS

[RTZ17] Zhenjie Ren, Nizar Touzi, and Jianfeng Zhang. Comparison of viscosity solutions of fully nonlinear degenerate parabolic path-dependent PDEs. SIAM Journal on Mathematical Analysis, 49(5): 4093–4116, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ribaud:2012:WPR

[RV12] Francis Ribaud and Stéphane Vento. Well-posedness results for the three-dimensional Zakharov–Kuznetsov equation. SIAM Journal on Mathematical Analysis, 44(4): 2289–2304, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2014:ASD

[RW14] Xiaofeng Ren and Juncheng Wei. Asymmetric and symmetric double bubbles in a ternary inhibitory system. SIAM Journal on Mathematical Analysis, 46(4):2798– 2852, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2011:LTS

[RZ11] Jie Ren and Xicheng Zhang.

Limit theorems for stochastic differential equations with discontinuous coefficients. SIAM Journal on Mathematical Analysis, 43(1):302– 321, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p302_s1.

Renclawowicz:2014:GNN

[RZ14] Joanna Renclawowicz and Wojciech M. Zajaczkowski. Global nonstationary Navier– Stokes motion with large flux. SIAM Journal on Mathematical Analysis, 46(4):2581– 2613, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rodrigues:2016:PCD

[RZ16] L. Miguel Rodrigues and Kevin Zumbrun. Periodiccoefficient damping estimates, and stability of largeamplitude roll waves in inclined thin film flow. SIAM Journal on Mathematical Analysis, 48(1):268–280, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:2017:LDA

[RZ17] Siqi Ren and Weiren Zhao. Linear damping of Alfvén waves by phase mixing. SIAM Journal on Mathematical Analysis, 49(3):2101– 2137, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruland:2018:HSR

[RZZ18] Angkana Rüland, Christian Zillinger, and Barbara Zwicknagl. Higher Sobolev regularity of convex integration solutions in elasticity: The Dirichlet problem with affine data in int(K^{lc}). SIAM Journal on Mathematical Analysis, 50(4): 3791–3841, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sabin:2013:CRS

[Sab13] Julien Sabin. Charge renormalization and static electron/positron pair production for a nonlinear Dirac model with weak interactions. SIAM Journal on Mathematical Analysis, 45(4): 2099–2147, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sacchetti:2018:NSW

[Sac18] Andrea Sacchetti. Nonlinear Stark–Wannier equation. SIAM Journal on Mathematical Analysis, 50(6):5783– 5810, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Salakhudinov:2012:RIE

[Sal12] R. G. Salakhudinov. Refined inequalities for Euclidean moments of a domain with respect to its boundary. SIAM Journal on Mathematical Analysis, 44(4):2949– 2961, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sprengel:2017:TIT

[SCB17] M. Sprengel, G. Ciaramella, and A. Borzì. A theoretical investigation of timedependent Kohn–Sham equations. SIAM Journal on Mathematical Analysis, 49(3): 1681–1704, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [SCB20].

Sprengel:2020:ETI

[SCB20] M. Sprengel, G. Ciaramella, and A. Borzì. Erratum: A Theoretical Investigation of Time-dependent Kohn–Sham Equations. SIAM Journal on Mathematical Analysis, 52(1): 1006–1008, ???? 2020. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [SCB17].

Schwab:2010:PHN

[Sch10] Russell W. Schwab. Periodic homogenization for nonlinear integro-differential equations. SIAM Journal on Mathematical Analysis, 42(6): 2652–2680, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schaeffer:2014:RSC

[Sch14a] Jack Schaeffer. A restriction on shocks in collisionless plasma. SIAM Journal on Mathematical Analysis, 46(4): 2767–2797, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schonbek:2014:HSD

[Sch14b]  Tomas Schonbek.  On a

Helmholtz style decomposition for an exterior domain. SIAM Journal on Mathematical Analysis, 46(5):3497– 3517, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schaeffer:2017:NHS

[Sch17] Jack Schaeffer. On nearly homogeneous states in collisionless plasma. SIAM Journal on Mathematical Analysis, 49(3): 2269–2286, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schachter:2018:NCF

[Sch18a]  Benjamin Schachter.  A

new class of first order displacement convex functionals. SIAM Journal on Mathematical Analysis, 50(2):1779– 1789, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schwenker:2018:GSS

[Sch18b] Sören Schwenker. Generic steady state bifurcations in monoid equivariant dynamics with applications in homogeneous coupled cell systems. SIAM Journal on Mathematical Analysis, 50(3):

2466–2485, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Su:2013:KTQ

[SdlL13] Xifeng Su and Rafael de la Llave. KAM theory for quasiperiodic equilibria in onedimensional quasi-periodic media. SIAM Journal on Mathematical Analysis, 44(6): 3901–3927, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Seiskari:2014:PEP

[Sei14] Otto Seiskari. Point electrode problems in piecewise smooth plane domains. SIAM Journal on Mathematical Analysis, 46(2):1204–1227, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Senik:2017:HNS

[Sen17] Nikita N. Senik. Homogenization for non-self-adjoint periodic elliptic operators on an infinite cylinder. SIAM Journal on Mathematical Analysis, 49(2):874–898, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:2015:SER

[She15] Wen Shen. Slow erosion with rough geological layers. SIAM Journal on Mathematical Analysis, 47(4):3116– 3150, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:2010:NDE

[Sho10] R. E. Showalter. Nonlinear degenerate evolution equations in mixed formulation. SIAM Journal on Mathematical Analysis, 42(5):2114– 2131, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simanek:2016:EIZ

[Sim16] Brian Simanek. An electrostatic interpretation of the zeros of paraorthogonal polynomials on the unit circle. SIAM Journal on Mathematical Analysis, 48(3):2250– 2268, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sincich:2010:SDU

[Sin10] Eva Sincich. Stability for the determination of unknown boundary and impedance with a Robin boundary condition. SIAM Journal on Mathematical Analysis, 42(6):2922– 2943, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v42/i6/p2922_s1.

Sharma:2016:GES

[SM16] Vandana Sharma and Jeff Morgan. Global existence of solutions to reaction–diffusion systems with mass transport type boundary conditions. SIAM Journal on Mathematical Analysis, 48(6): 4202–4240, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Scott:2019:PSE

[SM19] James Scott and Tadele Mengesha. A potential space estimate for solutions of systems of nonlocal equations in peridynamics. SIAM Journal on Mathematical Analysis, 51 (1):86–109, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:2017:RPV

[Smi17] Scott A. Smith. Random perturbations of viscous, compressible fluids: Global existence of weak solutions. SIAM Journal on Mathematical Analysis, 49(6):4521– 4578, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sofonea:2018:CRO

[Sof18] Mircea Sofonea. Convergence results and optimal control for a class of hemivariational inequalities. SIAM Journal on Mathematical Analysis, 50(4): 4066–4086, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Souplet:2019:SAR

[Sou19] Philippe Souplet. A simplified approach to the refined blowup behavior for the nonlinear heat equation. SIAM Journal on Mathematical Analysis, 51(2):991– 1013, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schimperna:2013:CCH

[SP13] Giulio Schimperna and Irena Pawlow. On a class of Cahn– Hilliard models with nonlinear diffusion. SIAM Journal on Mathematical Analysis, 45 (1):31–63, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Spence:2014:WEB

[Spe14] E. A. Spence. Wavenumberexplicit bounds in timeharmonic acoustic scattering. SIAM Journal on Mathematical Analysis, 46(4):2987– 3024, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sisti:2014:ECM

[SR14] Francesco Sisti and Costantino Ricciuti. Effects of concavity on the motion of a body immersed in a Vlasov gas. SIAM Journal on Mathematical Analysis, 46(6):3579– 3611, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Srinivasan:2011:RCS

[Sri11]  Ravi Srinivasan.  Rates

of convergence for Smoluchowski’s coagulation equations. SIAM Journal on Mathematical Analysis, 43(4): 1835–1854, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1835_s1.

Schade:2015:SDC

[SS15] Katharina Schade and Yoshihiro Shibata. On strong dynamics of compressible nematic liquid crystals. SIAM Journal on Mathematical Analysis, 47(5):3963–3992, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schwab:2017:FST

[SS17] Christoph Schwab and Rob Stevenson. Fractional spacetime variational formulations of (Navier–) Stokes equations. SIAM Journal on Mathematical Analysis, 49(4): 2442–2467, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schrecker:2019:VVL

[SS19] Matthew R. I. Schrecker and Simon Schulz. Vanishing viscosity limit of the compressible Navier–Stokes equations with general pressure law. SIAM Journal on Mathematical Analysis, 51(3):2168– 2205, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schouten-Straatman:2019:TWS

[SSH19] Willem M. Schouten-Straatman and Hermen Jan Hupkes. Traveling waves for spatially discrete systems of FitzHugh– Nagumo type with periodic coefficients. SIAM Journal on Mathematical Analysis, 51(4): 3492–3532, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saff:2015:OPA

[SSST15] E. B. Saff, H. Stahl, N. Stylianopoulos, and V. Totik. Orthogonal polynomials for areatype measures and image recovery. SIAM Journal on Mathematical Analysis, 47(3): 2442–2463, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stinner:2014:GWS

[SSW14] Christian Stinner, Christina Surulescu, and Michael Winkler. Global weak solutions in a PDE–ODE system modeling multiscale cancer cell invasion. SIAM Journal on Mathematical Analysis, 46(3): 1969–2007, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Song:2019:NFK

[SSZ19] Jian Song, Xiaoming Song, and Qi Zhang. Nonlinear Feynman–Kac formulas for stochastic partial differential equations with spacetime noise. SIAM Journal on Mathematical Analysis, 51(2): 955–990, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smarrazzo:2010:LTB

[ST10] Flavia Smarrazzo and Alberto Tesei. Long-time behavior of solutions to a class of forward-backward parabolic equations. SIAM Journal on Mathematical Analysis, 42(3): 1046–1093, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shih:2011:GSA

[ST11] Chih-Wen Shih and Jui-Pin Tseng. Global synchronization and asymptotic phases for a ring of identical cells with delayed coupling. SIAM Journal on Mathematical Analysis, 43(4): 1667–1697, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1667_s1.

Sadiq:2015:RCT

[ST15a] Kamran Sadiq and Alexandru Tamasan. On the range characterization of the two-dimensional attenuated Doppler transform. SIAM Journal on Mathematical Analysis, 47(3):2001– 2021, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Strani:2015:TDI

[ST15b] Marta Strani and Benjamin Texier. Time-delayed instabilities in complex Burgers equations. SIAM Journal on Mathematical Analysis, 47(4): 2495–2518, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stinga:2017:RTE

[ST17] Pablo Raúl Stinga and José L.

Torrea. Regularity theory and extension problem for fractional nonlocal parabolic equations and the master equation. SIAM Journal on Mathematical Analysis, 49(5):

3893–3924, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Suzuki:2018:TLS

[ST18] Masahiro Suzuki and Atusi Tani. Time-local solvability of the Degond–Lucquin-Desreux–Morrow model for gas discharge. SIAM Journal on Mathematical Analysis, 50(5):5096–5118, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Slepcev:2019:ALR

[ST19] Dejan Slepcev and Matthew Thorpe. Analysis of p-Laplacian regularization in semisupervised learning. SIAM Journal on Mathematical Analysis, 51(3):2085–2120, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stokols:2019:HCF

[Sto19] Logan F. Stokols. Hölder continuity for a family of nonlocal hypoelliptic kinetic equations. SIAM Journal on Mathematical Analysis, 51(6): 4815–4847, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Strain:2010:GNL

[Str10] Robert M. Strain. Global Newtonian limit for the relativistic Boltzmann equation near vacuum. SIAM Journal on Mathematical Analysis, 42(4):1568–1601, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sugiyama:2016:DFT

[Sug16]  Yuusuke Sugiyama.  De-

generacy in finite time of 1D quasilinear wave equations. SIAM Journal on Mathematical Analysis, 48(2): 847–860, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Suslina:2013:HNP

[Sus13] Tatiana Suslina. Homogenization of the Neumann problem for elliptic systems with periodic coefficients. SIAM Journal on Mathematical Analysis, 45(6):3453–3493, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Savin:2011:DEN

[SV11] Ovidiu Savin and Enrico Valdinoci. Density estimates for a nonlocal variational model via the Sobolev inequality. SIAM Journal on Mathematical Analysis, 43(6): 2675–2687, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2675_s1.

Stan:2014:FKE

[SV14] Diana Stan and Juan Luis Vázquez. The Fisher–KPP equation with nonlinear fractional diffusion. SIAM Journal on Mathematical Analysis, 46(5):3241–3276, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sweers:2018:PHC

[SV18] Guido Sweers and Katerina Vassi. Positivity for a hinged convex plate with stress. SIAM Journal on Mathematical Analysis, 50(1): 1163–1174, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Scala:2019:VAS

[SV19] Riccardo Scala and Nicolas Van Goethem. A variational approach to single crystals with dislocations. SIAM Journal on Mathematical Analysis, 51(1):489–531, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Segatti:2011:FDR

[SW11a] Antonio Segatti and Hao Wu. Finite dimensional reduction and convergence to equilibrium for incompressible smectic-A liquid crystal flows. SIAM Journal on Mathematical Analysis, 43(6): 2445–2481, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i6/p2445_s1.

Shapiro:2011:RDB

[SW11b] A. M. Shapiro and M. I. Weinstein. Radiative decay of bubble oscillations in a compressible fluid. SIAM Journal on Mathematical Analysis, 43 (2):828–876, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p828_s1.

Strauss:2017:SSR

[SW17] Walter A. Strauss and Yilun Wu. Steady states of rotating stars and galaxies. SIAM Journal on Mathematical Analysis, 49(6):4865– 4914, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soave:2018:UCP

[SW18] Nicola Soave and Tobias Weth. The unique continuation property of sublinear equations. SIAM Journal on Mathematical Analysis, 50(4): 3919–3938, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saut:2017:CPL

[SWX17] Jean-Claude Saut, Chao Wang, and Li Xu. The Cauchy problem on large time for surface-waves-type Boussinesq systems II. SIAM Journal on Mathematical Analysis, 49(4):2321–2386, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Song:2015:RSP

[SWZ15] Kyungwoo Song, Qin Wang, and Yuxi Zheng. The regularity of semihyperbolic patches near sonic lines for the 2-D Euler system in gas dynamics. SIAM Journal on Mathematical Analysis, 47(3): 2200–2219, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:2013:BSF

[SX13] Zuowei Shen and Zhiqiang Xu. On B-spline framelets derived from the unitary extension principle. SIAM Journal on Mathematical Analysis, 45 (1):127–151, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Strain:2014:SHB

[SY14] Robert M. Strain and Seok-Bae Yun. Spatially homogeneous Boltzmann equation for relativistic particles. SIAM Journal on Mathematical Analysis, 46(1):917– 938, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stefanov:2017:TPT

[SY17] Plamen Stefanov and Yang Yang. Thermo- and photoacoustic tomography with variable speed and planar detectors. SIAM Journal on Mathematical Analysis, 49(1): 297–310, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sylvester:2012:DTE

[Syl12] John Sylvester. Discreteness of transmission eigenvalues via upper triangular compact operators. SIAM Journal on Mathematical Analysis, 44(1): 341–354, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v44/i1/p341_s1.

Sebert:2011:FPM

[SZ11] Florian M. Sebert and Yi Ming Zou. Factoring pseudoidentity matrix pairs. SIAM Journal on Mathematical Analysis, 43 (2):565–576, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p565_s1.

Salo:2012:IPL

[SZ12a] Mikko Salo and Xiao Zhong. An inverse problem for the p-Laplacian: Boundary determination. SIAM Journal on Mathematical Analysis, 44(4): 2474–2495, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Scardia:2012:LTM

[SZ12b] Lucia Scardia and Caterina Ida Zeppieri. Linetension model for plasticity as the Γ-limit of a nonlinear dislocation energy. SIAM Journal on Mathematical Analysis, 44(4):2372–2400, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taskovic:2018:MLM

[TAGP18] Maja Tasković, Ricardo J. Alonso, Irene M. Gamba, and Natasa Pavlović. On Mittag-Leffler moments for the Boltzmann equation for hard potentials without cutoff. SIAM Journal on Mathematical Analysis, 50(1):834– 869, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Takada:2010:CCE

[Tak10] Ryo Takada. Counterexamples of commutator estimates in the Besov and the Triebel– Lizorkin spaces related to the Euler equations. SIAM Journal on Mathematical Analysis, 42(6):2473–2483, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Takatsu:2013:BED

[Tak13] Asuka Takatsu. Behaviors of φ-exponential distributions in Wasserstein geometry and an evolution equation. SIAM Journal on Mathematical Analysis, 45(4): 2546–2556, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taniguchi:2015:DCC

[Tan15]  Masaharu Taniguchi.  An

(N − 1)-dimensional convex compact set gives an Ndimensional traveling front in the Allen–Cahn equation. SIAM Journal on Mathematical Analysis, 47(1):455– 476, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tang:2018:PSC

[Tan18] Hao Tang. On the pathwise solutions to the Camassa– Holm equation with multiplicative noise. SIAM Journal on Mathematical Analysis, 50(1):1322–1366, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taylor:2018:AED

[Tay18] Jamie M. Taylor. An analysis of equilibria in dense nematic liquid crystals. SIAM Journal on Mathematical Analysis, 50(2):1918–1957, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tian:2017:TTS

[TD17] Xiaochuan Tian and Qiang Du. Trace theorems for some nonlocal function spaces with heterogeneous localization. SIAM Journal on Mathematical Analysis, 49(2): 1621–1644, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tejero:2017:RSP

[Tej17] Jorge Tejero. Reconstruction and stability for piecewise smooth potentials in the plane. SIAM Journal on Mathematical Analysis, 49(1): 398–420, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terracina:2011:QBT

[Ter11] Andrea Terracina. Qualitative behavior of the twophase entropy solution of a forward-backward parabolic problem. SIAM Journal on Mathematical Analysis, 43(1): 228–252, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p228_s1.

Thaller:2019:ESS

[Tha19] Maximilian Thaller. Existence of static solutions of the Einstein–Vlasov–Maxwell system and the thin shell limit. SIAM Journal on Mathematical Analysis, 51(3): 2231–2260, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thirouin:2019:AQS

[Thi19]  Joseph Thirouin.  About

the quadratic Szegő hierarchy. SIAM Journal on Mathematical Analysis, 51(2):1454– 1495, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Treinen:2013:SSS

[Tre13]  Ray Treinen.  On the

symmetry of solutions to some floating drop problems. SIAM Journal on Mathematical Analysis, 44(6):3834– 3847, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Triay:2018:DDG

[Tri18] Arnaud Triay. Derivation of the dipolar Gross–Pitaevskii energy. SIAM Journal on Mathematical Analysis, 50(1):

33–63, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Troy:2017:PSF

[Tro17] William C. Troy. Phaselocked solutions of the finite size Kuramoto coupled oscillator model. SIAM Journal on Mathematical Analysis, 49(3): 1912–1931, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trillos:2018:CLP

[TSA18] Nicolás García Trillos and Daniel Sanz-Alonso. Continuum limits of posteriors in graph Bayesian inverse problems. SIAM Journal on Mathematical Analysis, 50(4): 4020–4040, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tsugawa:2012:LWP

[Tsu12] Kotaro Tsugawa. Local wellposedness of the KdV equation with quasi-periodic initial data. SIAM Journal on Mathematical Analysis, 44(5): 3412–3428, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trofimchuk:2018:TWB

[TV18] Sergei Trofimchuk and Vitaly Volpert. Traveling waves for a bistable reaction– diffusion equation with delay. SIAM Journal on Mathematical Analysis, 50(1):1175– 1199, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temam:2010:SSM

[TW10] R. Temam and D. Wirosoetisno.

Stability of the slow manifold in the primitive equations. SIAM Journal on Mathematical Analysis, 42(1): 427–458, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tao:2011:CHM

[TW11a] Youshan Tao and Michael Winkler. A chemotaxishaptotaxis model: The roles of nonlinear diffusion and logistic source. SIAM Journal on Mathematical Analysis, 43 (2):685–704, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i2/p685_s1.

Thakur:2011:SBR

[TW11b] Gaurav Thakur and Hau-Tieng Wu. Synchrosqueezingbased recovery of instantaneous frequency from nonuniform samples. SIAM Journal on Mathematical Analysis, 43(5):2078–2095, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sima/resource/
                                        1/sjmaah/v43/i5/p2078_s1.

Tao:2015:LTB

[TW15] Youshan Tao and Michael Winkler. Large time behavior in a multidimensional chemotaxis-haptotaxis model with slow signal diffusion. SIAM Journal on Mathematical Analysis, 47(6):4229– 4250, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tan:2018:GWP

[TW18a] Zhong Tan and Yanjin Wang. Global well-posedness of an initial-boundary value problem for viscous non-resistive MHD systems. SIAM Journal on Mathematical Analysis, 50(1):1432–1470, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tsai:2018:ENC

[TW18b] Dong-Ho Tsai and Xiao-Liu Wang. The evolution of nonlocal curvature flow arising in a Hele–Shaw problem. SIAM Journal on Mathematical Analysis, 50(1):1396– 1431, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tan:2015:SSS

[TWW15] Zhong Tan, Yanjin Wang, and Yong Wang. Stability of steady states of the Navier–Stokes–Poisson equations with non-flat doping profile. SIAM Journal on Mathematical Analysis, 47(1): 179–209, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temple:2011:TPL

[TY11] Blake Temple and Robin Young. Time-periodic linearized solutions of the compressible Euler equations and a problem of small divisors. SIAM Journal on Mathematical Analysis, 43(1): 1–49, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p1_s1.

Tan:2013:GSO

[TYZZ13] Zhong Tan, Tong Yang, Huijiang Zhao, and Qingyang Zou. Global solutions to the one-dimensional compressible Navier–Stokes–Poisson equations with large data. SIAM Journal on Mathematical Analysis, 45(2):547–571, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tilli:2013:AFL

[TZ13] Paolo Tilli and Davide Zucco. Asymptotics of the first Laplace eigenvalue with Dirichlet regions of prescribed length. SIAM Journal on Mathematical Analysis, 45(6): 3266–3282, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tilli:2015:WBP

[TZ15] Paolo Tilli and Davide Zucco. Where best to place a Dirichlet condition in an anisotropic membrane? SIAM Journal on Mathematical Analysis, 47(4): 2699–2721, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tao:2018:CPW

[TZ18] Tao Tao and Liqun Zhang. On the continuous periodic weak solutions of Boussinesq equations. SIAM Journal on Mathematical Analysis, 50(1): 1120–1162, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ueda:2012:DSR

[UWK12] Yoshihiro Ueda, Shu Wang, and Shuichi Kawashima. Dissipative structure of the regularity-loss type and time asymptotic decay of solutions for the Euler–Maxwell system. SIAM Journal on Mathematical Analysis, 44(3):2002– 2017, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Valkonen:2015:JSU

[Val15] Tuomo Valkonen. The jump set under geometric regularization. Part 1: Basic technique and first-order denoising. SIAM Journal on Mathematical Analysis, 47(4):2587– 2629, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vonBrecht:2014:NST

[vBM14] James H. von Brecht and Scott G. McCalla. Nonlinear stability through algebraically decaying point spectrum: Applications to nonlocal interaction equations. SIAM Journal on Mathematical Analysis, 46(6):3727–3760, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanBaalen:2011:TPS

[vBW11] Guillaume van Baalen and Peter Wittwer. Time periodic solutions of the Navier–Stokes equations with nonzero constant boundary conditions at infinity. SIAM Journal on Mathematical Analysis, 43(4): 1787–1809, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.

                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1787_s1.

vandenBerg:2011:RNS

[vdBMJLM11] Jan Bouwe van den Berg, Jason D. Mireles-James, Jean-Philippe Lessard, and Konstantin Mischaikow. Rigorous numerics for symmetric connecting orbits: Even homoclinics of the Gray–Scott equation. SIAM Journal on Mathematical Analysis, 43(4): 1557–1594, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i4/p1557_s1.

vandenBerg:2019:RCS

[vdBW19] Jan Bouwe van den Berg and J. F. Williams. Rigorously computing symmetric stationary states of the Ohta– Kawasaki problem in three dimensions. SIAM Journal on Mathematical Analysis, 51(1): 131–158, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vessella:2015:SEI

[Ves15]  Sergio Vessella.  Stabil-

ity estimates for an inverse hyperbolic initial boundary value problem with unknown boundaries. SIAM Journal on Mathematical Analysis, 47(2): 1419–1457, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Veltz:2013:CMR

[VF13] Romain Veltz and Olivier Faugeras. A center manifold result for delayed neural fields equations. SIAM Journal on Mathematical Analysis, 45(3): 1527–1562, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [VF15].

Veltz:2015:ECM

[VF15] Romain Veltz and Olivier Faugeras. Erratum: A center manifold result for delayed neural fields equations. SIAM Journal on Mathematical Analysis, 47(2): 1665–1670, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [VF13].

Visintin:2018:SCS

[Vis18]  Augusto Visintin.  Struc-

tural compactness and stability of semi-monotone flows. SIAM Journal on Mathematical Analysis, 50(3):2628– 2663, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vo:2018:CIS

[VK18] Hoang-Hung Vo and Seonghak Kim. Convex integration for scalar conservation laws in one space dimension. SIAM Journal on Mathematical Analysis, 50(3): 3122–3146, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanNeerven:2012:MRS

[vNVW12] Jan van Neerven, Mark Veraar, and Lutz Weis. Maximal L^{p}-regularity for stochastic evolution equations. SIAM Journal on Mathematical Analysis, 44(3): 1372–1414, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vazquez:2011:ESF

[VW11] Juan Luis Vázquez and Michael Winkler. The evolution of singularities in fast diffusion equations: Infinitetime blow-down. SIAM Journal on Mathematical Analysis, 43(4):1499–1535, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i4/p1499_s1.

Vasseur:2015:ILC

[VW15a] Alexis Vasseur and Yi Wang. The inviscid limit to a contact discontinuity for the compressible Navier–Stokes– Fourier system using the relative entropy method. SIAM Journal on Mathematical Analysis, 47(6):4350– 4359, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vo:2015:CFS

[VW15b] Theodore Vo and Martin Wechselberger. Canards of folded saddle-node type I. SIAM Journal on Mathematical Analysis, 47(4):3235– 3283, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vasseur:2016:GWS

[VY16] Alexis F. Vasseur and Cheng Yu. Global weak solutions to the compressible quantum Navier–Stokes equations with damping. SIAM Journal on Mathematical Analysis, 48(2): 1489–1511, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vergara:2015:ODE

[VZ15] Vicente Vergara and Rico Zacher. Optimal decay estimates for time-fractional and other nonlocal subdiffusion equations via energy methods. SIAM Journal on Mathematical Analysis, 47(1):210– 239, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walker:2014:LBJ

[Wal14] Stephen G. Walker. On a lower bound for the Jensen inequality. SIAM Journal on Mathematical Analysis, 46(5): 3151–3157, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2011:DLV

[Wan11] Yanjin Wang. The diffusive limit of the Vlasov– Boltzmann system for binary fluids. SIAM Journal on Mathematical Analysis, 43(1): 253–301, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p253_s1.

Wang:2012:GST

[Wan12] Yanjin Wang. Global solution and time decay of the Vlasov–Poisson–Landau system in R^{3}. SIAM Journal on Mathematical Analysis, 44(5): 3281–3323, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2013:PNS

[Wan13] Yuzhao Wang. Periodic nonlinear Schrödinger equation in critical H^{s}(T^{n}) spaces. SIAM Journal on Mathematical Analysis, 45(3):1691– 1703, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2018:GWP

[Wan18] Yong Wang. Global wellposedness of the relativistic Boltzmann equation. SIAM Journal on Mathematical Analysis, 50(5):5637– 5694, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2019:FBP

[Wan19] Chunpeng Wang. A free boundary problem of a degenerate elliptic equation and subsonic-sonic flows with general sonic curves. SIAM Journal on Mathematical Analysis, 51(6):4977–5010, ???? 2019. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walsh:2013:SWW

[WBS13] Samuel Walsh, Oliver Bühler, and Jalal Shatah. Steady water waves in the presence of wind. SIAM Journal on Mathematical Analysis, 45(4): 2182–2227, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wei:2018:KSN

[WDL18] Jinlong Wei, Jinqiao Duan, and Guangying Lv. Kinetic solutions for nonlocal scalar conservation laws. SIAM Journal on Mathematical Analysis, 50(2):1521– 1543, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinmann:2012:IMR

[Wei12] Andreas Weinmann. Interpolatory multiscale representation for functions between manifolds. SIAM Journal on Mathematical Analysis, 44(1): 162–191, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i1/p162_s1.

Weng:2014:SSF

[Wen14] Shangkun Weng. On steady subsonic flows for Euler– Poisson models. SIAM Journal on Mathematical Analysis, 46(1):757–779, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2012:ABG

[WFL12] Shu Wang, Yuehong Feng, and Xin Li. The asymptotic behavior of globally smooth solutions of bipolar nonisentropic compressible Euler–Maxwell system for plasma. SIAM Journal on Mathematical Analysis, 44(5): 3429–3457, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wheeler:2013:LAS

[Whe13] Miles H. Wheeler. Largeamplitude solitary water waves with vorticity. SIAM Journal on Mathematical Analysis, 45(5):2937–2994, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Winkler:2015:LDG

[Win15]  Michael Winkler.  Large-

data global generalized solutions in a chemotaxis system with tensor-valued sensitivities. SIAM Journal on Mathematical Analysis, 47(4): 3092–3115, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wroblewska-Kaminska:2017:AAC

[WK17] Aneta Wróblewska-Kamińska.

The asymptotic analysis of the complete fluid system on a varying domain: From the compressible to the incompressible flow. SIAM Journal on Mathematical Analysis, 49(5):3299–3334, ???? 2017. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2016:GPN

[WLT16] Yong Wang, Chun Liu, and Zhong Tan. A generalized Poisson–Nernst–Planck– Navier–Stokes model on the fluid with the crowded charged particles: Derivation and its well-posedness. SIAM Journal on Mathematical Analysis, 48(5):3191–3235, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weller:2013:AFB

[WNRJ13] Frederic Frank Weller, Maria Neuss-Radu, and Willi Jäger. Analysis of a free boundary problem modeling thrombus growth. SIAM Journal on Mathematical Analysis, 45(2): 809–833, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolansky:2019:CAL

[Wol19]  Gershon Wolansky.  Con-

tact angles of liquid drops subjected to a rough boundary. SIAM Journal on Mathematical Analysis, 51(3):2286– 2305, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ward:2014:APS

[WU14a] John Paul Ward and Michael Unser. Approximation properties of Sobolev splines and the construction of compactly supported equivalents. SIAM Journal on Mathematical Analysis, 46(3):1843– 1858, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2014:PBL

[Wu14b]  Kung-Chien Wu.  Point-

wise behavior of the linearized Boltzmann equation on a torus. SIAM Journal on Mathematical Analysis, 46 (1):639–656, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2014:WPD

[Wu14c] Lei Wu. Well-posedness and decay of the viscous surface wave. SIAM Journal on Mathematical Analysis, 46(3): 2084–2135, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2016:SFD

[Wu16] Yilun Wu. Simplicity and finiteness of discrete spectrum of the Benjamin–Ono scattering operator. SIAM Journal on Mathematical Analysis, 48(2):1348–1367, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2017:JSD

[Wu17] Yilun Wu. Jost solutions and the direct scattering problem of the Benjamin–Ono equation. SIAM Journal on Mathematical Analysis, 49(6): 5158–5206, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wunsch:2010:GHS

[Wun10] Marcus Wunsch. The generalized Hunter–Saxton system. SIAM Journal on Mathematical Analysis, 42(3):1286– 1304, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wei:2010:SSS

[WW10] Juncheng Wei and Matthias Winter. Stability of spiky solutions in a reaction–diffusion system with four morphogens on the real line. SIAM Journal on Mathematical Analysis, 42(6):2818–2841, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2012:MLP

[WW12] Shu Wang and Ke Wang. The mixed layer problem and quasi-neutral limit of the drift-diffusion model for semiconductors. SIAM Journal on Mathematical Analysis, 44(2): 699–717, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v44/i2/p699_s1.

Wang:2015:SST

[WW15] Teng Wang and Yi Wang. Stability of superposition of two viscous shock waves for the Boltzmann equation. SIAM Journal on Mathematical Analysis, 47(2):1070– 1120, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wheeler:2018:GOR

[WW18] Aric Wheeler and Mark Williams. Geometric optics for Rayleigh wavetrains in ddimensional. SIAM Journal on Mathematical Analysis, 50 (4):4563–4615, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2012:GRV

[WWW12] Zhi-An Wang, Michael Winkler, and Dariusz Wrzosek. Global regularity versus infinitetime singularity formation in a chemotaxis model with volume-filling effect and degenerate diffusion. SIAM Journal on Mathematical Analysis, 44(5):3502–3525, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2015:GSS

[WWX15] Jiahong Wu, Yifei Wu, and Xiaojing Xu. Global small solution to the 2D MHD system with a velocity damping term. SIAM Journal on Mathematical Analysis, 47(4): 2630–2656, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2011:SLS

[WX11] Jing Wang and Feng Xie. Singular limit to strong contact discontinuity for a 1D compressible radiation hydrodynamics model. SIAM Journal on Mathematical Analysis, 43(3):1189–1204, ???? 2011. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i3/p1189_s1.

Wang:2012:PPQ

[WX12] Yang Wang and Zhiqiang Xu. The performance of PCM quantization under tight frame representations. SIAM Journal on Mathematical Analysis, 44(4):2802– 2823, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:2013:SSG

[WX13] Hao Wu and Xiang Xu. Strong solutions, global regularity, and stability of a hydrodynamic system modeling vesicle and fluid interactions. SIAM Journal on Mathematical Analysis, 45(1): 181–214, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2015:GSS

[WX15] Chunpeng Wang and Zhouping Xin. Global smooth supersonic flows in infinite expanding nozzles. SIAM Journal on Mathematical Analysis, 47(4): 3151–3211, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2016:SCS

[WX16] Chunpeng Wang and Zhouping Xin. On sonic curves of smooth subsonic-sonic and transonic flows. SIAM Journal on Mathematical Analysis, 48(4):2414–2453, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2019:EFB

[WX19] Chunpeng Wang and Zhouping Xin. On an elliptic free boundary problem and subsonic jet flows for a given surrounding pressure. SIAM Journal on Mathematical Analysis, 51(2): 1014–1045, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2015:LWP

[WXY15a] Ya-Guang Wang, Feng Xie, and Tong Yang. Local wellposedness of Prandtl equations for compressible flow in two space variables. SIAM Journal on Mathematical Analysis, 47(1):321–346, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2015:URV

[WXY15b] Yong Wang, Zhouping Xin, and Yan Yong. Uniform regularity and vanishing viscosity limit for the compressible Navier–Stokes with general Navier-slip boundary conditions in three-dimensional domains. SIAM Journal on Mathematical Analysis, 47(6): 4123–4191, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wei:2013:VRP

[WY13] Juncheng Wei and Jun Yang. Vortex ring pinning for the Gross–Pitaevskii equation in three-dimensional space. SIAM Journal on Mathematical Analysis, 44(6):3991– 4047, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2015:SSS

[WY15] Ya-Guang Wang and Fang Yu. Structural stability of supersonic contact discontinuities in three-dimensional compressible steady flows. SIAM Journal on Mathematical Analysis, 47(2):1291– 1329, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2013:IRC

[WZ13a] Wendong Wang and Zhifei Zhang. On the interior regularity criteria for suitable weak solutions of the magnetohydrodynamics equations. SIAM Journal on Mathematical Analysis, 45(5): 2666–2677, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wen:2013:GCL

[WZ13b] Huanyao Wen and Changjiang Zhu. Global classical large solutions to Navier–Stokes equations for viscous compressible and heat-conducting fluids with vacuum. SIAM Journal on Mathematical Analysis, 45(2):431–468, ???? 2013. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2016:DSH

[WZ16] Feng-Yu Wang and Xicheng Zhang. Degenerate SDE with Hölder–Dini drift and non-Lipschitz noise coefficient. SIAM Journal on Mathematical Analysis, 48(3):2189– 2226, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wen:2017:GST

[WZ17] Huanyao Wen and Changjiang Zhu. Global solutions to the three-dimensional full compressible Navier–Stokes equations with vacuum at infinity in some classes of large data. SIAM Journal on Mathematical Analysis, 49(1): 162–221, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:2015:RDL

[WZZ15] Wei Wang, Pingwen Zhang, and Zhifei Zhang. Rigorous derivation from Landau– de Gennes theory to Ericksen– Leslie theory. SIAM Journal on Mathematical Analysis, 47 (1):127–158, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2011:GCS

[Xu11] Jiang Xu. Global classical solutions to the compressible Euler–Maxwell equations. SIAM Journal on Mathematical Analysis, 43(6): 2688–2718, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i6/p2688_s1.

Xu:2016:CRG

[Xu16] Qiang Xu. Convergence rates for general elliptic homogenization problems in Lipschitz domains. SIAM Journal on Mathematical Analysis, 48(6): 3742–3788, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2018:ETC

[Xu18] Xiangsheng Xu. Existence theorems for a crystal surface model involving the p-Laplace operator. SIAM Journal on Mathematical Analysis, 50(4): 4261–4281, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xia:2010:TDM

[XV10] Qinglan Xia and Anna Vershynina. On the transport dimension of measures. SIAM Journal on Mathematical Analysis, 41(6):2407– 2430, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xie:2010:EGS

[XX10] Chunjing Xie and Zhouping Xin. Existence of global steady subsonic Euler flows through infinitely long nozzles. SIAM Journal on Mathematical Analysis, 42(2):751– 784, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2013:GWP

[XXK13] Jiang Xu, Jun Xiong, and Shuichi Kawashima. Global well-posedness in critical Besov spaces for two-fluid Euler–Maxwell equations. SIAM Journal on Mathematical Analysis, 45(3):1422– 1447, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xin:2014:AGR

[XY14] Jack Xin and Yifeng Yu. Asymptotic growth rates and strong bending of turbulent flame speeds of G-equation in steady two-dimensional incompressible periodic flows. SIAM Journal on Mathematical Analysis, 46(4):2444– 2467, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xie:2018:GTS

[XY18] Feng Xie and Tong Yang. Global-in-time stability of 2D MHD boundary layer in the Prandtl–Hartmann regime. SIAM Journal on Mathematical Analysis, 50(6):5749– 5760, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2018:RDB

[XYD18] Jieren Xu, Haizhao Yang, and Ingrid Daubechies. Recursive diffeomorphism-based regression for shape functions. SIAM Journal on Mathematical Analysis, 50(1):5–32, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xin:2019:ASS

[XYY19] Zhouping Xin, Qian Yuan, and Yuan Yuan. Asymptotic stability of shock waves and rarefaction waves under periodic perturbations for 1-D convex scalar conservation laws. SIAM Journal on Mathematical Analysis, 51(4): 2971–2994, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xin:2016:POA

[XYZ16] Jack Xin, Yifeng Yu, and Andrej Zlatos. Periodic orbits of the ABC flow with A = B = C = 1. SIAM Journal on Mathematical Analysis, 48(6): 4087–4093, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2015:GSS

[XZ15] Li Xu and Ping Zhang. Global small solutions to threedimensional incompressible magnetohydrodynamical system. SIAM Journal on Mathematical Analysis, 47(1): 26–65, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:2010:ASC

[XZL10] Xiang Xu, Liyun Zhao, and Chun Liu. Axisymmetric solutions to coupled Navier–Stokes/Allen– Cahn equations. SIAM Journal on Mathematical Analysis, 41(6):2246–2282, ???? 2010. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yamamoto:2013:AES

[Yam13] Masakazu Yamamoto. Asymptotic expansion of solutions to the dissipative equation with fractional Laplacian. SIAM Journal on Mathematical Analysis, 44(6):3786– 3805, ???? 2013. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Yam16].

Yamamoto:2016:EAE

[Yam16] Masakazu Yamamoto. Erratum: Asymptotic Expansion of Solutions to the Dissipative Equation with Fractional Laplacian. SIAM Journal on Mathematical Analysis, 48(4): 3037–3038, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Yam13].

Yao:2019:LST

[Yao19] Peng-Fei Yao. Linear strain tensors on hyperbolic surfaces and asymptotic theories for thin shells. SIAM Journal on Mathematical Analysis, 51(2): 1387–1435, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yi:2010:PSG

[YCW10] Taishan Yi, Yuming Chen, and Jianhong Wu. Periodic solutions and the global attractor in a system of delay differential equations. SIAM Journal on Mathematical Analysis, 42(1):24– 63, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yannacopoulos:2011:WCS

[YFK11] Athanasios N. Yannacopoulos, Nikolaos E. Frangos, and Ioannis Karatzas. Wiener chaos solutions for linear backward stochastic evolution equations. SIAM Journal on Mathematical Analysis, 43(1): 68–113, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

tronic). URL http://epubs.
siam.org/sima/resource/
1/sjmaah/v43/i1/p68_s1.

Yin:2010:AMH

[YMYC10] G. Yin, Xuerong Mao, Chenggui Yuan, and Dingzhou Cao. Approximation methods for hybrid diffusion systems with state-dependent switching processes: Numerical algorithms and existence and uniqueness of solutions. SIAM Journal on Mathematical Analysis, 41(6):2335– 2352, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshida:2017:ABS

[Yos17] Natsumi Yoshida. Asymptotic behavior of solutions toward a multiwave pattern to the Cauchy problem for the scalar conservation law with the Ostwald–de Waele-type viscosity. SIAM Journal on Mathematical Analysis, 49(3): 2009–2036, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshida:2018:ABS

[Yos18] Natsumi Yoshida. Asymptotic behavior of solutions toward the viscous shock waves to the Cauchy problem for the scalar conservation law with nonlinear flux and viscosity. SIAM Journal on Mathematical Analysis, 50(1): 891–932, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ye:2011:GET

[YT11] J. Ye and S. Tanveer. Global existence for a translating near-circular Hele–Shaw bubble with surface tension. SIAM Journal on Mathematical Analysis, 43(1):457– 506, ???? 2011. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sima/resource/
                                        1/sjmaah/v43/i1/p457_s1.

Yun:2015:EBM

[Yun15] Seok-Bae Yun. Ellipsoidal BGK model near a global Maxwellian. SIAM Journal on Mathematical Analysis, 47(3): 2324–2354, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2010:GCS

[YY10] Tong Yang and Hongjun Yu.

Global classical solutions for the Vlasov–Maxwell–Fokker– Planck system. SIAM Journal on Mathematical Analysis, 42 (1):459–488, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2014:SCT

[YY14] Haizhao Yang and Lexing Ying. Synchrosqueezed curvelet transform for twodimensional mode decomposition. SIAM Journal on Mathematical Analysis, 46(3): 2052–2083, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2018:GSS

[YY18] Tong Yang and Hongjun Yu. Global solution for the spatially inhomogeneous non-cutoff Kac equation. SIAM Journal on Mathematical Analysis, 50(4):4503– 4562, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2014:EOV

[YZ14a] Yisong Yang and Ruifeng Zhang. Existence of optical vortices. SIAM Journal on Mathematical Analysis, 46(1): 484–498, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yuan:2014:LTS

[YZ14b] Xiaoping Yuan and Jing Zhang. Long time stability of Hamiltonian partial differential equations. SIAM Journal on Mathematical Analysis, 46(5):3176–3222, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yi:2015:ABS

[YZ15] Taishan Yi and Xingfu Zou. Asymptotic behavior, spreading speeds, and traveling waves of nonmonotone dynamical systems. SIAM Journal on Mathematical Analysis, 47(4):3005–3034, ???? 2015. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yu:2016:SCE

[YZ16] Xiang Yu and Shiqing Zhang. Saari’s conjecture for elliptical motions and minimizing solutions of the N-body problem. SIAM Journal on Mathematical Analysis, 48(1):

709–724, ???? 2016. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:2018:GWP

[YZ18] Kailong Yang and Lifeng Zhao. Global well-posedness and scattering for masscritical, defocusing, infinite dimensional vector-valued resonant nonlinear Schrödinger system. SIAM Journal on Mathematical Analysis, 50(2): 1593–1655, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yao:2010:EAB

[YZZ10] Lei Yao, Ting Zhang, and Changjiang Zhu. Existence and asymptotic behavior of global weak solutions to a 2D viscous liquid-gas two-phase flow model. SIAM Journal on Mathematical Analysis, 42(4): 1874–1897, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yao:2012:ILV

[YZZ12] Lei Yao, Changjiang Zhu, and Ruizhao Zi. Incompressible limit of viscous liquid-gas twophase flow model. SIAM Journal on Mathematical Analysis, 44(5):3324–3345, ???? 2012. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhu:2019:MIE

[ZBL19] Jiahui Zhu, Zdzisław Brzeźniak, and Wei Liu. Maximal inequalities and exponential estimates for stochastic convolutions driven by Lévy-type processes in Banach spaces with application to stochastic quasi-geostrophic equations. SIAM Journal on Mathematical Analysis, 51(3): 2121–2167, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2015:CCM

[ZCO15] Kewei Zhang, Elaine Crooks, and Antonio Orlando. Compensated convexity, multiscale medial axis maps and sharp regularity of the squared-distance function. SIAM Journal on Mathematical Analysis, 47(6):4289– 4331, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2016:CCM

[ZCO16] Kewei Zhang, Elaine Crooks, and Antonio Orlando. Compensated convexity methods for approximations and interpolations of sampled functions in Euclidean spaces: Theoretical foundations. SIAM Journal on Mathematical Analysis, 48(6):4126–4154, ???? 2016. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2010:CFD

[ZF10] Ting Zhang and Daoyuan Fang. Compressible flows with a density-dependent viscosity coefficient. SIAM Journal on Mathematical Analysis, 41(6): 2453–2488, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2012:GES

[ZF12] Ting Zhang and Daoyuan Fang. Global existence of strong solution for equations related to the incompressible viscoelastic fluids in the critical L^{p} framework. SIAM Journal on Mathematical Analysis, 44(4):2266– 2288, ???? 2012. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhan:2010:LSS

[ZH10] Wang Zhan and Yin Huicheng.

Local structural stability of a multidimensional centered rarefaction wave for the threedimensional steady supersonic Euler flow around a sharp corner. SIAM Journal on Mathematical Analysis, 42(4): 1639–1687, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2010:UCL

[Zha10] Kewei Zhang. On universal coercivity in linear elasticity. SIAM Journal on Mathematical Analysis, 42(1): 298–322, ???? 2010. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2014:FSK

[Zha14]  Xicheng Zhang.  Funda-

mental solution of kinetic Fokker–Planck operator with anisotropic nonlocal dissipativity. SIAM Journal on Mathematical Analysis, 46(3): 2254–2280, ???? 2014. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2019:LSA

[Zha19] Katherine Zhiyuan Zhang. Linear stability analysis of the relativistic Vlasov–Maxwell system in an axisymmetric domain. SIAM Journal on Mathematical Analysis, 51(6): 4683–4723, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhigun:2019:GGS

[Zhi19] Anna Zhigun. Generalized global supersolutions with mass control for systems with taxis. SIAM Journal on Mathematical Analysis, 51(3): 2425–2443, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhou:2015:IRF

[Zho15] Wei Zhou. Interior regularity of fully nonlinear degenerate elliptic equations I: Bellman equations with constant coefficients. SIAM Journal on Mathematical Analysis, 47(3): 2375–2415, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhou:2018:LMR

[Zho18] Hanming Zhou. The local magnetic ray transform of tensor fields. SIAM Journal on Mathematical Analysis, 50(2): 1753–1778, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhu:2015:CSC

[Zhu15] Shengguo Zhu. On classical solutions of the compressible magnetohydrodynamic equations with vacuum. SIAM Journal on Mathematical Analysis, 47(4): 2722–2753, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zelati:2015:MPS

[ZK15] Michele Coti Zelati and Piotr Kalita. Minimality properties of set-valued processes and their pullback attractors. SIAM Journal on Mathematical Analysis, 47(2): 1530–1561, ???? 2015. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhu:2018:GCP

[ZLMZ18] Neng Zhu, Zhengrong Liu, Vincent R. Martinez, and Kun Zhao. Global Cauchy problem of a system of parabolic conservation laws arising from a Keller–Segel type chemotaxis model. SIAM Journal on Mathematical Analysis, 50(5): 5380–5425, ???? 2018. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zelati:2019:GSR

[ZN19] Vittorio Coti Zelati and Margherita Nolasco. Ground state for the relativistic one electron atom in a selfgenerated electromagnetic field. SIAM Journal on Mathematical Analysis, 51(3): 2206–2230, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zheng:2017:LWP

[ZT17] Yunrui Zheng and Ian Tice. Local well posedness of the near-equilibrium contact line problem in 2-dimensional Stokes flow. SIAM Journal on Mathematical Analysis, 49 (2):899–953, ???? 2017. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:2014:GWP

[ZZ14] Qian Zhang and Xiaoxin Zheng. Global well-posedness for the two-dimensional incompressible chemotaxis-Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 46(4):3078–3105, ???? 2014. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhao:2019:IIP

[ZZ19] Hongkai Zhao and Yimin Zhong. Instability of an inverse problem for the stationary radiative transport near the diffusion limit. SIAM Journal on Mathematical Analysis, 51(5):3750– 3768, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).